### v1

# On the contraction properties of Sinkhorn semigroups

Abstract
We develop a novel semigroup contraction analysis based on Lyapunov techniques to prove the exponential convergence of Sinkhorn equations on weighted Banach spaces. This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences as well as in weighted Banach spaces. To the best of our knowledge, these are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm. We also illustrate the impact of these results in the context of multivariate linear Gaussian models as well as statistical finite mixture models including Gaussian-kernel density estimation of complex data distributions arising in generative models.
Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, Lyapunov functions.
Mathematics Subject Classification: 37M25, 49Q22, 47H09, 60J20; secondary: 37M25, 60J05, 94A17.
1 Introduction
1.1 Background
The Schrödinger bridge problem is to find a random evolution with minimal relative entropy w.r.t. some reference physical model and with prescribed distributions of the initial and final time random states. This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [56, 66, 68] for recent surveys with numerous references on this subject.
These optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [5, 31, 54, 68], statistical barycenter problems [1, 4, 10, 24, 29, 32], economy [13], computer vision [42, 76], control theory [18, 19], and many others.
Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure. For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [39, 46, 55, 74, 75] as well as the more recent studies [18, 28, 57].
Sinkhorn iterations are based on nonlinear integral transformations over the whole state space. Exact finite-dimensional solutions cannot be computed for general models. However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [39, 46, 55, 74, 75]. For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [2, 24, 25, 51].
Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [14, 45, 74, 75, 77] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [18, 33, 47, 40]. The recent article [2] also presents a refined convergence analysis for Gaussian models based on matrix Riccati difference equations. It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [68] as well as the more recent articles [2, 20, 23] and the references therein.
Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [3, 17, 44]. Linear rates on non-necessarily compact spaces were first obtained by Léger in [58] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [31, 52]). Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [48] under some exponential moments estimates on Schrödinger potentials.
There are very few articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of two recent significant contributions [23, 20]. In [23], the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic cost functions associated with a sufficiently large regularization parameter. These exponential entropic decays were recently refined in [20] to apply to all values of the regularization parameter.
To the best of our knowledge, the exponential convergence of Sinkhorn iterates in the context of statistical mixture models remains an important open research question.
As underlined in [2], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step. These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [6, 34, 35, 36, 70, 71, 72] and references therein.
The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject. The articles [43, 70, 71] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities. Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging. There is a need for new tools in order to handle these models.
1.2 Contribution
In this paper, we introduce a novel approach to Sinkhorn exponential convergence analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
To be specific, we investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups. These semigroup techniques are basic quantitative methods for studying the stability of nonlinear and time-varying stochastic models, see for instance [36] as well as Section 2 in [6]. In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [49] and Meyn and Tweedie [60], see also the book [61] and references therein. These powerful off-the-shelf semigroup techniques are mainly based on two conditions:
  • 1)
    the existence of a Lyapunov function satisfying a drift-type condition, and
  • 2)
    a local minorization condition on the sub-level sets of the Lyapunov function.
Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function. In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator.
Based on these techniques, the analysis developed in this article yields novel contraction estimates with respect to general divergence criteria. It also applies to multivariate Gaussian models as well as statistical finite mixture models.
1.3 Outline of the paper
After presenting some basic notation in Section 1.4, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [2] for a more detailed discussion on these models within our framework). Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C. Section 3.1 recalls some basics facts on the Hilbert projective metric, operator norms, entropy and related $\phi$-divergences. The operator contraction coefficients associated with these criteria are introduced in Section 3.2.
With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups. Section 4.1 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups. The main contraction theorems introduced in this article are presented in Section 4.2 and Section 4.3.
We illustrate the impact of these results in Section 5, which provides several examples, including strong mixing models, Gaussian models as well as Schödinger bridges between finite mixture distributions.
Strong mixing Sinkhorn semigroups are discussed in Section 5.1 in terms of Dobrushin contraction coefficients. In this context, we show that Sinkhorn semigroups contract for any $\phi$-divergence criteria, including the relative entropy, the total variation distance, as well as $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, Jeffreys and Rényi divergences.
Section 5.2 discusses the contraction properties of Gaussian Sinkhorn models. Last but not least, the convergence analysis of Sinkhorn semigroups in the context of statistical finite mixture models, including Gaussian mixture, is carried out in Section 5.3.
Most of the technical proofs, as well as some extended analyses, are provided in Appendices A through E.
1.4 Some basic notation
We let ${\cal B}(\mathbb{X})$ be the algebra of locally bounded measurable functions on a locally compact Polish space $\mathbb{X}$. We denote by ${\cal B}_{b}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ the sub-algebra of bounded measurable functions $f$ on $\mathbb{X}$ endowed with the supremum norm $\|f\|:=\sup_{x\in\mathbb{X}}|f(x)|$. For a given uniformly positive function $\varphi\in{\cal B}(\mathbb{X})$, we let ${\cal B}_{\varphi}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the Banach space of functions $f\in{\cal B}(\mathbb{X})$ equipped with the norm $\|f\|_{\varphi}:=\|f/\varphi\|$.
Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures. Let ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the convex subset of probability measures on $\mathbb{X}$. The Lebesgue integral of a integrable function $f\in{\cal B}(\mathbb{X})$ with respect to (w.r.t.) some $\nu\in{\cal M}(\mathbb{X})$ is denoted by
$\nu(f):=\int_{\mathbb{X}}f(x)~{}\nu(dx)$
For indicator functions $f=1_{A}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
Consider a bounded integral operator ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\cal M}(\mathbb{X})$ and a function $f\in{\cal B}(\mathbb{Y})$. We denote by $\mu{\cal K}\in{\cal M}(\mathbb{Y})$ and ${\cal K}(f)\in{\cal B}(\mathbb{X})$ the measure and the function defined by
$(\mu{\cal K})(dy):=\int~{}\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y)$ (1)
For indicator functions $f=1_{A}$ sometimes we write ${\cal K}(x,A)$ instead of ${\cal K}(1_{A})(x)$. For bounded integral operator ${\cal K}_{1}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and ${\cal K}_{2}(y,dx)$ from $\mathbb{Y}$ into $\mathbb{X}$, we also denote by $({\cal K}_{1}{\cal K}_{2})$ the bounded integral operator from $\mathbb{X}$ into $\mathbb{X}$ defined by the integral composition
$({\cal K}_{1}{\cal K}_{2})(x_{1},dx_{2}):=\int{\cal K}_{1}(x_{1},dy){\cal K}_{2}(y,dx_{2}).$
Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we see that
$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(x,y))=\mu(dy)\,{\cal K}(y,dx).$
A measure $\nu_{1}\in{\cal M}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\cal M}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$. Whenever $\nu_{1}\ll\nu_{2}$, we denote by ${d\nu_{1}}/{d\nu_{2}}$ the Radon-Nikodym derivative of $\nu_{1}$ w.r.t. $\nu_{2}$. We also consider the partial order, denoted $\nu_{1}\leq\nu_{2}$, when $\nu_{1}(A)\leq\nu_{2}(A)$ for any measurable subset $A\subset\mathbb{X}$. The measures are said comparable when we have $c~{}\nu_{1}\leq\nu_{2}\leq c^{-1}~{}\nu_{1}$ for some scalar constant $c>0$.
2 Description of the models
2.1 Entropic optimal transport
Consider locally compact Polish spaces $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ endowed with some locally bounded positive measures $\lambda$ and $\nu$. Consider some probability measures of the form
$\lambda_{U}(dx):=e^{-U(x)}~{}\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~{}\nu(dy)$
for some normalized potential functions $U:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$. Let $\Pi_{U,V}$ be the set of coupling probability measures ${\cal P}$ on the product space $(\mathbb{X}\times\mathbb{Y})$, with prescribed first and second coordinate marginals $(\lambda_{U},\nu_{V})$. Denote ${\cal Q}$ a Markov transition from $\mathbb{X}$ into $\mathbb{Y}$ of the form
${\cal Q}(x,dy)=e^{-W(x,y)}~{}\nu(dy)\quad\mbox{\rm and set}\quad{\cal R}(y,dx):=e^{-W(x,y)}~{}\lambda(dx).$ (2)
In the latter expression, $W(x,y)$ stands for some potential function on $(\mathbb{X}\times\mathbb{Y})$. In optimal transport literature $W$ is sometimes called the cost function.
Consider the reference probability measure
${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{}{\cal Q}(x,dy).$
The entropic transport problem (a.k.a. dynamic Schrödinger bridge problem) w.r.t. the reference measure ${\cal P}_{0}$ is formulated as
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0}),$ (3)
where $\text{Ent}({\cal P}~{}|~{}{\cal P}_{0})=\int\log\frac{\delta{\cal P}}{d{\cal P}_{0}}d{\cal P}$ is the relative entropy (or Kullback-Leibler divergence) between ${\cal P}$ and ${\cal P}_{0}$. It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0})<\infty$. This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (3) (cf. the seminal article by Csiszár [27], as well as Section 6 in the Lecture Notes by Nutz [66], see also the survey article by Léonard [56] and references therein).
2.2 Schrödinger bridges
The Schrödinger bridge distribution $\mathbb{P}$ defined in (3) can rarely be computed analytically. However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [28, 74, 75]. This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions
${\cal P}_{2n}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad{\cal P}_{2n+1}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}$ (4)
starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dx)$ and a transition operator ${\cal S}(x,dy)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
$(\mu\times{\cal S})^{\flat}(d(x,y)):=\mu(dy){\cal S}(y,dx).$
The Markov transitions ${\cal S}_{n}$ in (4) are defined by the Sinkhorn iterations
$\left\{\begin{array}[]{l}(\pi_{2n}\times{\cal S}_{2n+1})^{\flat}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\cal S}_{2(n+1)}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\cal S}_{2n+1}.\end{array}\right.$ (5)
In the dual formulation, the bridge distributions ${\cal P}_{n}$ are often written as
${\cal P}_{n}(d(x,y))=e^{-W(x,y)}~{}\lambda_{U_{n}}(dx)~{}\nu_{V_{n}}(dy),$ (6)
with the initial potential functions $(U_{0},V_{0})=(U,0)$ and potential functions $(U_{n},V_{n})$ defined sequentially by the integral recursions
$U_{2n+1}$ $=$ $~{}U_{2n}~{}:=U+\log{{\cal Q}(e^{-V_{2n}})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2(n+1)}}$ $=$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2n+1}}:=V+\log{{\cal R}(e^{-U_{2n}})}$ (7)
When $n\to\infty$, the sequence of distributions ${\cal P}_{2n}$ converges towards the Schrödinger bridge (3) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t. the reference measure ${\cal P}_{0}$) while the sequence ${\cal P}_{2n+1}$ converges towards the Schrödinger bridge from $\nu_{V}$ to $\lambda_{U}$ (w.r.t. the reference measure ${\cal P}_{1}$).
As shown in Section 6 of [2], under mild integrability assumptions the Schrödinger bridge distribution $\mathbb{P}$ solving (3) can also be computed from any Sinkhorn bridge distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{2n}).$ (8)
The Sinkhorn transition operators can also be written in terms of the potential functions $(U_{n},V_{n})$, namely,
${\cal S}_{2n}(x,dy)=\frac{{\cal Q}(x,dy)e^{-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\cal S}_{2n+1}(y,dx)=\frac{{\cal R}(y,dx)~{}e^{-U_{2n+1}(x)}}{{\cal R}(e^{-U_{2n+1}})(y)},$ (9)
while Proposition 6.3 in [2] yields the formulae
$V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$ (10)
for the potentials $(U_{n},V_{n})$, $n\geq 0$, expressed in terms of the distribution flow of the internal states of the Gibbs-loop process.
In addition, for any $n\geq 0$ we also have the commutation formulae
${\cal S}_{2n}\left(\frac{d\nu_{V}}{d\pi_{2n}}\right)=\frac{d\pi_{2n+1}}{d\lambda_{U}}\quad\mbox{and}\quad{\cal S}_{2(n+1)}\left(\frac{d\pi_{2n}}{d\nu_{V}}\right)=\frac{d\lambda_{U}}{d\pi_{2n+1}}.$ (11)
In the same vein, for any $n\geq 1$ we have
${\cal S}_{2n-1}\left(\frac{d\lambda_{U}}{d\pi_{2n-1}}\right)=\frac{d\pi_{2n}}{d\nu_{V}}\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)=\frac{d\nu_{V}}{d\pi_{2n}}.$ (12)
A proof for (11) and (12) is provided in Appendix E on page E.
2.3 Sinkhorn semigroups
For any $n\geq 0$ we have the fixed point equations
$\lambda_{U}{\cal S}^{\circ}_{2n+1}=\lambda_{U}\quad\mbox{\rm and}\quad\nu_{V}{\cal S}^{\circ}_{2(n+1)}=\nu_{V}$ (13)
with the forward-backward Gibbs-type Markov transitions
${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}.$
The above fixed point equations yield, for any $n\geq 1$, the time inhomogeneous Markov evolution equations
$\pi_{2n}=\pi_{2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2n-1}{\cal S}^{\circ}_{2n+1}.$ (14)
For any $0\leq p\leq n$ we have the semigroup equations
$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{2n+1}}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
The Sinkhorn transitions ${\cal S}^{\circ}_{2p,2n}$ and ${\cal S}^{\circ}_{2p+1,2n+1}$ are defined inductively by the time inhomogeneous semigroup equations
${\cal S}^{\circ}_{2p,2n}={\cal S}^{\circ}_{2p,2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2p+1,2n+1}={\cal S}^{\circ}_{2p+1,2n-1}{\cal S}^{\circ}_{2n+1},$
with ${\cal S}^{\circ}_{2n,2n}=I={\cal S}^{\circ}_{2n+1,2n+1}$ the identity integral operator. The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (14) sharing the same invariant measures (13) at each individual step. Using the commutation properties (11) and (12) for any $n\geq p$ we readily check that
$\frac{d\pi_{2n}}{d\nu_{V}}$ $=$ ${\cal S}^{\circ}_{2n}\left(\frac{d\pi_{2(n-1)}}{d\nu_{V}}\right)={\cal S}^{\circ}_{2n,2p}\left(\frac{d\pi_{2p}}{d\nu_{V}}\right)\quad\text{and}$
$\frac{d\pi_{2n+1}}{d\lambda_{U}}$ $=$ ${\cal S}_{2n+1}^{\circ}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)={\cal S}^{\circ}_{2n+1,2p-1}\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right),$ (15)
with the backward semigroups
${\cal S}^{\circ}_{2n,2p}:={\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1,2p-1}:={\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1}.$
By (11) and (12) we also obtain
$\frac{d\lambda_{U}}{d\pi_{2n+1}}$ $=$ $\left({\cal S}_{2(n+1)}\,{\cal S}^{\circ}_{2n,2p}\,{\cal S}_{2p-1}\right)\left(\frac{d\lambda_{U}}{d\pi_{2p-1}}\right)\quad\text{and}$
$\frac{d\nu_{V}}{d\pi_{2n}}$ $=$ $\left({\cal S}_{2n+1}\,{\cal S}^{\circ}_{2n-1,2p-1}\,{\cal S}_{2(p-1)}\right)\left(\frac{d\nu_{V}}{d\pi_{2(p-1)}}\right).$ (16)
3 Contraction of Markov operators
We recall some definitions of contraction coefficients for Markov operators, both in terms of the Hilbert projective metric and $\Phi$-entropies. These coefficients will be key to the analysis of contraction properties of Sinkhorn semigroups in Section 4.
3.1 Metrics and divergences
3.1.1 Hilbert projective metric
The Hilbert (projective) metric is defined, for any comparable measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$, by the formula
$H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~{}\|\nu_{1}-\nu_{2}\|_{\sf tv},$ (17)
[…]
3.1.2 Integral operator norms
Given some functions $\varphi\in{\cal B}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\cal M}_{\varphi}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\cal M}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}$ ::=:= $\sup\{|\nu(f)|~{}:~{}\|f\|_{\varphi}\leq 1\}=\sup\{|\nu(f)|~{}:~{}\mbox{\rm osc}_{\varphi}(f)\leq 1\}$ (18)
$=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation}\quad|\nu|:=\nu_{+}+\nu_{-}.$
In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\cal B}(\mathbb{X})$, which is given by
$\mbox{\rm osc}_{\varphi}(f):=\sup_{x_{1},x_{2}}\frac{|f(x_{1})-f(x_{2})|}{\varphi(x_{1})+\varphi(x_{2})}\leq\|f\|_{\varphi}.$
For a detailed proof of the equivalent formulations in the latter definition (18), we refer to Proposition 8.2.16 in [38].
Whenever $\varphi\geq 1$, for any $\kappa>0$, we can construct the function $\phi_{\kappa}=1+\kappa\phi$ and obtain the norm equivalence formulae
$\kappa~{}\|\nu\|_{\sf tv}\leq\kappa~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi_{\kappa}}\leq(2+\kappa){\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}.$ (19)
The choice of condition $\varphi\geq 1/2$ in the definition of the $\varphi$-norm given in (18) is imposed only to recover the conventional total variation dual distance ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\mu_{1}-\mu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{1/2}=\|\mu_{1}-\mu_{2}\|_{\sf tv}$ between probability measures $\mu_{1},\mu_{2}\in{\cal M}_{1}(\mathbb{X})={\cal M}_{1,\frac{1}{2}}(\mathbb{X})$, where ${\cal M}_{1,\varphi}(\mathbb{X})$ denotes the subset of probability measures $\nu$ s.t. $\nu(\varphi)<\infty$. Recall that for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
$\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\cal M}_{1}(S)~{}:~{}\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$ (20)
3.1.3 Entropy and divergences
Consider a convex function
$\begin{array}[]{cccc}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\\ \end{array}$
such that for any $\alpha,v,w\in\mathbb{R}_{+}$ we have
$\Phi(\alpha w,\alpha v)=\alpha~{}\Phi(u,v)\quad\mbox{\rm and}\quad\Phi(1,1)=0.$ (21)
The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
$D_{\Phi}(\nu_{1},\nu_{2}):=\int~{}\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~{}\gamma(dx).$
By homogeneity arguments, the above definition does not depend on the choice of the dominating measure $\gamma$. Note that
$\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~{}|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$ (22)
Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [26], i.e., we also have
$D_{\Phi}(\nu_{1},\nu_{2})=\int~{}\phi\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~{}\nu_{2}(dx)\quad\text{where}\quad\phi(u):=\Phi(u,1)$
and we have used the convention $D_{\Phi}(\nu_{1},\nu_{2})=\infty$ when $\nu_{1}\not\ll\nu_{2}$. Therefore, the $\Phi$-entropy generalizes the classical relative entropy (a.k.a. Kullback-Leibler divergence or $I$-divergence). More precisely, choosing the functions
$\Phi(u,v)=\Phi_{1}(u,v):=u\log{(u/v)}\quad\mbox{\rm and}\quad\phi(u)=u\log{u}$
we recover the classical formulae
$D_{\Phi_{1}}(\nu_{1},\nu_{2})=\mbox{\rm Ent}(\nu_{1}~{}|~{}\nu_{2}):=\int~{}\log{\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)}~{}\nu_{1}(dx)=\int~{}\phi_{1}\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~{}\nu_{2}(dx),$
with the convex function with unique minimum $\phi_{1}(1)=0=\partial\phi_{1}(1)$ defined by
$\phi_{1}(u)=u\log{u}-(u-1).$
Other examples of $\phi$-divergences includes $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [26, 37, 63, 65].
3.2 Some contraction coefficients
Let $\delta_{x}$ denote the Dirac delta measure located at $x$. Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$, we denote by $\imath_{x_{1},x_{2}}({\cal K})(y)$ the Radon-Nikodym function
$\imath_{x_{1},x_{2}}({\cal K})(y):=\frac{d\delta_{x_{1}}{\cal K}}{d\delta_{x_{2}}{\cal K}}(y).$ (23)
We also consider the $[0,1]$-valued parameters
$\jmath({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~{}\inf_{y\in\mathbb{Y}}~{}\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
$\hbar({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (24)
and we note that
$\delta_{x_{1}}{\cal K}(dy_{1})~{}\delta_{x_{2}}{\cal K}(dy_{2})~{}\geq\hbar({\cal K})~{}\delta_{x_{2}}{\cal K}(dy_{1})~{}\delta_{x_{1}}{\cal K}(dy_{2}).$
Integrating the above inequality w.r.t. the variable $y_{2}$ we readily see that
$\delta_{x_{1}}{\cal K}(dy_{1})~{}\geq\hbar({\cal K})~{}\delta_{x_{2}}{\cal K}(dy_{1}),\quad\text{which readily yields}\quad\jmath({\cal K})\geq\hbar({\cal K})\geq\jmath({\cal K})^{2}.$ (25)
3.2.1 Birkhoff contraction coefficient
The Birkhoff contraction coefficient of a Markov transition ${\cal K}$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
[…]
where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B. Combining (25) and (26) we readily obtain the equivalent inequalities
[…]
As remarked in Section 3 of [22], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t. the Hilbert metric.
3.2.2 Dobrushin contraction coefficients
[…]
$\chi_{\varphi,\psi}({\cal K})=\sup_{\nu_{1},\nu_{2}\in{\cal M}_{1,\varphi}(S)}\frac{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}{\cal K}-\nu_{2}{\cal K}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}-\nu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}}$
When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\cal K})$ instead of $\chi_{\varphi,\varphi}({\cal K})$. Given a Markov transition ${\cal K}_{0}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a Markov transition ${\cal K}_{1}(x,dy)$ from $\mathbb{Y}$ into $\mathbb{X}$ we have
$\chi_{\varphi}({\cal K}_{0}{\cal K}_{1})=\chi_{\varphi,\varphi}({\cal K}_{0}{\cal K}_{1})\leq\chi_{\varphi,\psi}({\cal K}_{0})~{}\chi_{\psi,\varphi}({\cal K}_{1}).$ (27)
[…]
with the parameter
$\varepsilon({\cal K}):=1-\sup_{(x_{1},x_{2})\,\in\mathbb{X}^{2}}{\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}}.$ (29)
Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$ using the estimate (25) as well as (20) we readily find that
$\varepsilon({\cal K})\geq\jmath({\cal K})\geq\hbar({\cal K})\quad\mbox{\rm and}\quad\raisebox{2.0pt}{$\chi$}({\cal K})\leq 1-\jmath({\cal K})\leq 1-\hbar({\cal K}).$ (30)
3.2.3 Entropy contraction coefficient
[…]
For any convex function $\Phi$ satisfying (21) we recall that
[…]
The estimate (32) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [21] for finite state spaces and [37] for general measurable spaces). For further details on the Dobrushin contraction coefficient we refer the reader to the article [41], see also [35, 37] and Chapter 4 in [34].
4 Sinkhorn contraction properties
4.1 Some regularity conditions
The $r$-sub-level set of a positive function $\varphi\in{\cal B}(\mathbb{X})$ is denoted as
$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~{}:~{}\varphi(x)\leq r\right\}.$
When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$. Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [6] and references therein). In the context of Gibbs-loop semigroups (14), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
Condition (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
There exist some parameters $\epsilon\in]0,1[$ some constant $c>0$ and some integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and for any $n\geq n_{0}$ we have the estimates
${\cal S}_{2n}^{\circ}(\psi)\leq\epsilon~{}\psi+c\quad\mbox{and}\quad{\cal S}_{2n+1}^{\circ}(\varphi)\leq\epsilon~{}\varphi+c.$ (33)
There also exists some $r_{0}\geq 1$ and some function
$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{2n+1}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}^{\circ}_{2n}\|_{\sf tv}\leq 1-\varepsilon(r).$ (34)
The inequality (34) is satisfied if, and only if, there exists some ${\cal X}\in{\cal M}_{1}(\mathbb{X})$ (respectively, ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$) that may depend on $(n,r,(x_{1},x_{2}))$ (respectively, on $(n,r,(y_{1},y_{2}))$) satisfying, for any $i\in\{1,2\}$, the minorization conditions
${\cal S}^{\circ}_{2n}(y_{i},dy)\geq\varepsilon(r)~{}{\cal Y}(dy)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}(x_{i},dx)\geq\varepsilon(r)~{}{\cal X}(dx).$ (35)
The above assertion is rather well known and it can be checked using (20). Note that the drift condition (33) is trivially met for the unit functions $\varphi=1=\psi$ but in this case for any $r\geq 1$ we have $C_{\varphi}(r)=\mathbb{X}$ and $C_{\psi}(r)=\mathbb{Y}$. In that case, the inequalities in (35) are uniform minorization conditions over $\mathbb{X}$ and $\mathbb{Y}$.
Let ${\cal B}_{\infty}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the subalgebra of locally bounded and uniformly positive functions $\varphi$ that grow at infinity; that is, $\sup_{C}\varphi<\infty$ for any compact set $C\subset\mathbb{X}$, and for any
$r\geq\varphi_{\star}:=\inf_{\mathbb{X}}\varphi>0$ (36)
the $r$-sub-level set $C_{\varphi}(r)$ is a non empty compact subset. We denote by ${\cal B}_{0}(\mathbb{X}):=\{1/\varphi~{}:~{}\varphi\in{\cal B}_{\infty}(\mathbb{X})\}\subset{\cal B}_{b}(\mathbb{X})$ the sub-algebra of (bounded) positive functions, locally lower bounded and that vanish at infinity. When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ the $r$-sub-level sets $C_{\varphi}(r)$ and $C_{\psi}(r)$ are compact and the inequality (35) yields local minorization conditions.
The assumption $\varphi\wedge\psi\geq 1$ is only made to simplify the presentation. Indeed, for any function $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ satisfying (33), the rescaled functions $\varphi/\varphi_{\star}\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi/\psi_{\star}\in{\cal B}_{\infty}(\mathbb{Y})$ also satisfy (33).
The drift conditions in (33) ensure that for any $n\geq n_{0}$ we have
${\cal S}^{\circ}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\varphi}(\mathbb{X}).$
Whenever Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, it can be easily shown that
$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.$ (37)
These estimates are direct consequences of the drift conditions (33). A detailed proof is provided in Section E on page E.
4.2 Norm contraction theorems
The drift conditions (33) combined with the local minorization conditions (34) provide a way to control the contraction coefficients of Sinkhorn transitions. More precisely, we have the following estimate for the contraction coefficients of the forward-backward Markov transitions ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$.
Theorem 4.1.
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, then there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that
$\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\cal S}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (38)
Theorem 4.1 is a straightforward consequence of Theorem C.1 in Appendix C.
The contraction estimate (38) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $0\leq p\leq n$ we readily obtain
$\|(\mu_{1}-\mu_{2}){\cal S}^{\circ}_{2p,2n}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\mu_{1}-\mu_{2}\|_{\psi_{\kappa}}.$
Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\pi_{0},\nu_{V})$, the above estimate immediately yields an upper bound on the total variation distance, i.e.,
$\|\pi_{2n}-\nu_{V}\|_{\sf tv}\leq\|\pi_{2n}-\nu_{V}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\pi_{0}-\nu_{V}\|_{\psi_{\kappa}}.$
Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [30, 70, 71]
$\|{\cal S}^{\circ}_{0,2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{0,2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{n}~{}\left(\psi(y_{1})+\psi(y_{2})\right).$
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may make it hard to test –except in the case of Gaussian and linear models. Several Lyapunov functions in this context are designed in Section 5.2. A stronger (but easier to test) condition can be provided in terms of the Sinkhorn transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
Condition (${\cal D}_{\varphi,\psi}({\cal S})$).
There exists some parameter $\epsilon\in]0,1[$, some constant $c>0$ and some integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$ we have the estimates
${\cal S}_{2n}(\psi)\leq\epsilon~{}\varphi+c\quad\mbox{and}\quad{\cal S}_{2n+1}(\varphi)\leq\epsilon~{}\psi+c.$ (39)
There also exists some $r_{0}\geq 1$ and some function
$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (40)
When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, we readily see that
${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y})$
for any $n\geq n_{0}$.
Theorem 4.2.
If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (33) and (34) also hold for any $n>n_{0}$. In addition, there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that, for any $n\geq n_{0}$,
[…]
A proof of Theorem 4.2 is provided in Appendix C (on page C).
4.3 A Lyapunov stability theorem
Finally, we introduce a more tractable sufficient condition, in terms of the potential functions $(U,V,W)$. It enables us to obtain geometric rates for the convergence of the Sinkhorn iterations, as explicitly shown for the case of linear Gaussian models and finite mixture models in Sections 5.2 and 5.3, respectively. Let us assume that $W$ is locally bounded, $(U,V)\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y}))$ and $e^{-W}\in\mathbb{L}_{\infty}(\lambda\otimes\nu)$. In addition, assume that $U$ and $V$ have compact sub-levels sets with non empty interior and the measures $\lambda$ and $\nu$ have full support, i.e., they are finite on compact sets and strictly positive on non-empty open sets. Let us also denote
$W_{U}(y):=\int~{}\lambda_{U}(dx)~{}W(x,y)\quad\text{and}\quad W^{V}(x):=\int~{}\nu_{V}(dy)~{}W(x,y).$ (41)
Condition (${\cal H}_{\delta}(U,V)$).
There is some $\delta\in]0,1[$ such that ${\cal U}_{\delta}:=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$ and ${\cal V}_{\delta}:=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$. In addition, the functions $W_{U},W^{V}$ are locally bounded and
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$
Theorem 4.3.
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$, where $U_{\star}=\inf_{\mathbb{X}}U(x)$ and $V_{\star}=\inf_{\mathbb{Y}}V(y)$. In addition, there exist some parameter $\rho\in[0,1[$ and some constant $c>0$ such that
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c~{}\rho^{2n}\quad\text{for any $n\geq 1$.}$ (42)
A proof of Theorem 4.3 is provided in Appendix D.3.
Remark 4.4.
Using Proposition D.3 in Appendix D and following exactly the same argument as in the proof of Theorem 4.3, the estimate (42) remains valid when the weaker condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, provided that
${}\left(\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\right)\vee\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right)<\infty.$
Let us discuss some fairly straightforward consequences of the estimate (42). First note that (42) ensures the convergence of the series
$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ ::=:= $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ ::=:= $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (43)
where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (7) converge as $n\rightarrow\infty$ to a system of integral equations
$\mathbb{U}=U+\log{{\cal Q}(e^{-\mathbb{V}})}\quad\mbox{\rm and}\quad\mathbb{V}=V+\log{{\cal R}(e^{-\mathbb{U}})}.$
More precisely, using the estimates in (42) it is straightforward to prove the following corollary.
Corollary 4.5.
Under the assumptions of Theorem 4.3 for any $n\geq 1$ we have
$\|\mathbb{U}-U_{2n}\|_{\psi}\vee\|\mathbb{V}-V_{2n}\|_{\varphi}\leq\frac{c}{1-\rho^{2}}~{}\rho^{2n}$ (44)
with the constant $c>0$ as in (42). In addition, we have the relative entropy estimates
$\mbox{\rm Ent}\left(\lambda_{U}~{}|~{}\pi_{2n+1}\right)\leq c~{}\lambda_{U}(\varphi)~{}\rho^{2n}~{}\quad\mbox{and}\quad\mbox{\rm Ent}\left(\nu_{V}~{}|~{}\pi_{2n}\right)\leq c~{}\nu_{V}(\psi)~{}\rho^{2n}.$ (45)
With this notation at hand, the bridge distribution (3) has the form
$\lim_{n\to\infty}{\cal P}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~{}\lambda_{\mathbb{U}}(dx)~{}\nu_{\mathbb{V}}(dy).$ (46)
In addition, we have
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right)$
The exponential estimate (42) also allows the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (45).
Corollary 4.6.
Under the assumptions of Theorem 4.3 we obtain
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})$ $=$ $\sum_{p\geq n}\mbox{\rm Ent}\left(\lambda_{U}~{}|~{}\pi_{2p+1}\right)+\sum_{p\geq n}\mbox{\rm Ent}\left(\nu_{V}~{}|~{}\pi_{2p}\right)\quad\text{for any $n\geq 1$}.$
In addition, there exists some $c>0$ such that, for any $n\geq 1$, we have the entropy exponential decay estimates
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})\leq c~{}\rho^{2n}~{}(\lambda_{U}(\varphi)+\nu_{V}(\psi)).$
5 Some illustrations
5.1 Strong mixing models
As underlined in [45, 62] in the context of finite state space models, Sinkhorn iterations of the form (5) preserve the cross product Radon-Nikodym ratio in (24) and, hence, we have
$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}),\quad\forall n\geq 0.$ (47)
This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [14, 18, 33, 47]. Recall that the Hilbert contraction coefficient of a Markov transition ${\cal Q}$ can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq. (26)), hence the exponential convergence of Sinkhorn iterations w.r.t. the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$. An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
From (25), let us observe that
$\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})\geq\jmath({\cal Q})^{2}.$
Thus, for any given a convex function $\Phi$ satisfying (21), we can use (26) and (32) to show that
$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
while
$\jmath({\cal Q})>0\quad\text{implies}\quad\left\{\begin{array}[]{l}\raisebox{2.0pt}{$\chi$}_{H}({\cal S}_{n})=\raisebox{2.0pt}{$\chi$}_{H}({\cal Q})<1,\\ \text{${\cal D}_{1,1}({\cal S}^{\circ})$ is satisfied with $\varepsilon(r)=\hbar({\cal Q})>0$ for any $r\geq 1$, and}\\ \sup_{n\geq 0}\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal S}_{n})\leq\raisebox{2.0pt}{$\chi$}({\cal S}):=\sup_{n\geq 0}\raisebox{2.0pt}{$\chi$}({\cal S}_{n})\leq 1-\hbar({\cal Q})<1.\end{array}\right.$
We can also obtain the sandwich-type contraction inequalities described below.
Theorem 5.1.
For any convex function $\Phi$ satisfying (21) we have the contraction estimates
$\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~{}D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~{}D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~{}D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~{}D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$ (48)
[…]
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~{}\raisebox{2.0pt}{$\chi$}({\cal S})^{2n}\quad\text{for any $n\geq 0$.}$ (49)
Proof.
The estimates (48) are a direct consequence of (32) combined with the Markov transport formulae
$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\cal S}_{2(n+1)},\pi_{2n+1}{\cal S}_{2(n+1)}),\quad\text{and}$
$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
which are themselves easily obtained from expressions (5) and (13). The proof of (49) follows exactly the same arguments as the proof of the inequality (42) in Theorem 4.3, hence it is skipped.
The following result can be obtained from the series expansion given by Corollary 4.6.
Corollary 5.2.
Under the assumptions of Theorem 5.1, we have the entropy exponential decay estimates
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})\leq c~{}\raisebox{2.0pt}{$\chi$}({\cal S})^{2n},\quad\text{for any $n\geq 0$}.$
Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [35]. For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself. The choice ${\cal Q}(x,dy)~{}\propto~{}e^{-\delta\,\|y-b(x)\|}~{}dx$ yields
$\jmath({\cal Q})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [7], Nash [64] and Varopoulos [78] on Gaussian estimates for heat kernels on manifolds. Nevertheless, the above condition is not met for linear Gaussian transitions.
5.2 Linear-Gaussian models
[…]
Let ${\cal G}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\cal S}^{+}_{d}\subset{\cal G}l_{d}$ the subset of positive definite matrices. We sometimes use the Löwner partial ordering notation $v_{1}\geq v_{2}$ to mean that a symmetric matrix $v_{1}-v_{2}$ is positive semi-definite (equivalently, $v_{2}-v_{1}$ is negative semi-definite), and $v_{1}>v_{2}$ when $v_{1}-v_{2}$ is positive definite (equivalently, $v_{2}-v_{1}$ is negative definite). Given $v\in{\cal S}_{d}^{+}$ we denote by $v^{1/2}$ the principal (unique) symmetric square root.
Denote by $\nu_{m,\sigma}$ the Gaussian distribution on $\mathbb{R}^{d}$ with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\sigma\in{\cal S}^{+}_{d}$. In addition, let $g_{\sigma}$ denote the probability density function of the distribution $\nu_{0,\sigma}$, with covariance matrix $\sigma\in{\cal S}^{+}_{d}$. Consider the target probability measures $(\lambda_{U},\nu_{V})=(\nu_{m,\sigma},\nu_{\overline{m},\overline{\sigma}})$ for some given parameters
$(m,\overline{m})\in(\mathbb{R}^{d}\times\mathbb{R}^{d})\quad\text{and}\quad(\sigma,\overline{\sigma})\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d}).$
Equivalently, the log-density functions $(U,V)$ are defined by
$U(x)=-\log{g_{\sigma}(x-m)}\quad\mbox{\rm and}\quad V(y):=-\log{g_{\overline{\sigma}}(y-\overline{m})}.$ (50)
The linear Gaussian transition ${\cal Q}$ is defined by (2), with the potential
$W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$ (51)
for some given $(\alpha,\beta,\tau)\in\left(\mathbb{R}^{d}\times{\cal G}l_{d}\times{\cal S}^{+}_{d}\right)$. In this context, we have $\jmath({\cal Q})=0$ and Hilbert projective techniques cannot be applied. However, a refined exponential stability analysis of the Sinkhorn algorithm for Gaussian models has been developed in [2]. For instance, following Remark 5.4 in [2], there exists some $n_{0}\geq 1$ such that for any $n\geq n_{0}$ we have the estimate
$\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~{}\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$ (52)
The matrix $\Delta_{\varpi}$ in expression (52) above is defined as
$\Delta_{\varpi}:=\frac{\varpi}{2}+\left(\varpi+\left(\frac{\varpi}{2}\right)^{2}\right)^{1/2}\geq\varpi+(I+\varpi^{-1})^{-1}$ (53)
with
$\varpi:=(\gamma\,\gamma^{\prime})^{-1}\quad\mbox{\rm and}\quad\gamma:=\overline{\sigma}^{1/2}~{}\tau^{-1}\beta~{}\sigma^{1/2}.$
The drift-minorization techniques developed in this paper also enable us to obtain exponential convergence on weighted Banach spaces.
Proposition 5.3.
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied with the Lyapunov functions
$\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$ (54)
A proof is provided in Appendix A (on page A).
The choice of the Lyapunov functions is not unique. For instance, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
$\varphi(x)$ ::=:= $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
$\psi(y)$ ::=:= $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (55)
There also exists some $\delta_{0}>0$ such that ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
$\varphi(x)$ ::=:= $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
$\psi(y)$ ::=:= $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (56)
A detailed proof of the above assertions is provided in Appendix A on page A.
We end this section with an equivalent formulation of condition ${\cal H}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (50) and (51).
Proposition 5.4.
Condition ${\cal H}_{\delta}(U,V)$ is satisfied for the Lyapunov functions in (56) if, and only if,
$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~{}\sigma^{-1}\quad\text{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~{}\overline{\sigma}^{-1}.$ (57)
A proof is provided in Appendix E (on page E).
5.3 Finite mixture models
Consider some (fully supported) discrete probability measures $\xi_{0}$ and $\xi_{1}$ on some some finite states $E_{0}$ and $E_{1}$ and some collection of probability measures $(\lambda_{U_{i}})_{i\in E_{0}}$ and $(\nu_{V_{j}})_{j\in E_{1}}$ for some continuous potential functions $U_{i}:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V_{j}:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$ indexed by $i\in E_{0}$ and $j\in E_{1}$. The mixture distributions $\lambda_{U}$ and $\nu_{V}$ associated with these potential functions take the form
$U=-\log{\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}}\quad\mbox{\rm and}\quad V=-\log{\sum_{j\in E_{1}}\xi_{1}(j)~{}e^{-V_{j}}}.$ (58)
Proposition 5.5.
Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1[$ if conditions ${\cal H}_{\delta}(U_{i},V_{j})$ are met for the same $\delta\in]0,1[$ and almost every $(i,j)\in(E_{0}\times E_{1})$.
A proof is provided in Appendix E (on page E).
Let us now consider the mixture model (58) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (51) and the log-densities
$-U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (59)
where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters. Combining Proposition 5.5 with Proposition 5.4 we readily obtain the following result.
Proposition 5.6.
The Gaussian mixture model defined in (58) with the Gaussian log-densities of (59) defined on $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$ satisfies Condition ${\cal H}_{\delta}(U,V)$ for some $\delta\in]0,1[$ provided that
$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~{}\sigma_{i}^{-1}\quad\mbox{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~{}\overline{\sigma}_{j}^{-1}$ (60)
for almost every $(i,j)\in(E_{0}\times E_{1})$.
Let $(X_{i})_{1\leq i\leq N}$ be $N$ independent and identically distributed samples from some (arbitrary) data distribution $\eta(dx)$ on $\mathbb{X}=\mathbb{R}^{d}$. The Gaussian-kernel estimation model defined by
$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~{}dx.$ (61)
is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$). The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing. Note that (61) coincides with (58) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$. For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [67, 69] as well as the books [16, 73, 79].
The Gaussian mixture distribution (61) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [31] and references therein. Choosing $(\alpha,\beta)=(0,I)$ and $\tau=tI$; for some $t>0$ in (51), the Markov transition ${\cal Q}(x,dy)$ coincides with the the transition semigroup of the $d$-dimensional Brownian motion, namely
${\cal Q}(x,dy)=(2\pi t)^{-d/2}~{}\exp{\left(-\frac{1}{2t}~{}\|y-x\|^{2}\right)}~{}dy.$
In this context, the l.h.s. condition in (60) applied to (61) becomes
$\frac{1}{t}~{}I<(\delta\wedge(1-\delta))~{}\sigma^{-1}.$ (62)
[…]
$e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~{}\sigma^{-1}.$ (63)
Both transitions gradually transform the data distribution $\lambda_{U}$ into a simpler distribution $\lambda_{U}{\cal Q}$ by injecting noise into data. Also note that (62) and (63) are met when the time horizon is chosen sufficiently large.
The choice of the prior distribution $\nu_{V}$ is rather flexible. For instance, we can choose a Gaussian distribution, say $\nu_{0,\overline{\sigma}}$, satisfying the r.h.s. condition in (57). Starting from such Gaussian random states, the backward Sinkhorn transitions (indexed by odd indices) eventually generates samples back into the initial distribution by slowly removing the noise distortion. Conditions (62) and (63) reflect the fact that the forward noise injection as well as the reverse denoising procedure in diffusion-based generative models often requires very large time horizons.
Acknowledgements
This work is supported by the Innovation and Talent Base for Digital Technology and Finance (B21038) and ‘the Fundamental Research Funds for the Central Universities’, Zhongnan University of Economics and Law (2722023EJ002).
J. Miguez acknowledges the financial support of the Office of Naval Research (award no. N00014-22-1-2647) and MICIU/AEI/10.13039/501100011033/FEDER UE (grant no. PID2021-125159NB-I00 TYCHE).
Appendix A The Gaussian Sinkhorn algorithm
The Frobenius matrix norm of a given matrix $v\in\mathbb{R}^{d_{1}\times d_{2}}$ is defined by $\left\|v\right\|_{F}^{2}=\mbox{\rm Tr}(v^{\prime}v)$, with the trace operator $\mbox{\rm Tr}(\cdot)$ and $v^{\prime}$ the transpose of the matrix $v$. Note that, for $x\in\mathbb{R}^{d\times 1}=\mathbb{R}^{d}$, the Frobenius norm $\|x\|_{F}=\sqrt{x^{\prime}x}$ coincides with the Euclidean norm. The spectral norm is defined by $\|v\|_{2}=\sqrt{\lambda_{\sf max}(v^{\prime}v)}$.
Riccati difference equations
Let us associate with some $\varpi\in{\cal S}^{+}_{d}$ the increasing map $\mbox{\rm Ricc}_{\varpi}$ from ${\cal S}^{0}_{d}$ into ${\cal S}^{+}_{d}$ defined by
$\begin{array}[]{lccl}\mbox{\rm Ricc}_{\varpi}:&{\cal S}^{0}_{d}&\mapsto&{\cal S}^{+}_{d}\\ &v&\leadsto&\mbox{\rm Ricc}_{\varpi}(v):=(I+(\varpi+v)^{-1})^{-1}\\ \end{array}.$ (64)
If we define the matrices
$\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\cal S}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~{}\tau^{-1}\beta~{}\sigma^{1/2},$
and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~{}\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [2]) that the recursions
$\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~{}\upsilon_{2n}~{}\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~{}\upsilon_{2n+1}~{}\gamma^{\prime},$ (65)
satisfy the Riccati difference equations
$\upsilon_{2(n+1)}=\mbox{\rm Ricc}_{\varpi}\left(\upsilon_{2n}\right)\quad\mbox{and}\quad\upsilon_{2n+1}=\mbox{\rm Ricc}_{\overline{\varpi}}\left(\upsilon_{2n-1}\right),\quad\text{for $n\geq 0$}.$ (66)
Let us also define
$\tau_{2n}:=\overline{\sigma}^{1/2}~{}\upsilon_{2n}~{}\overline{\sigma}^{1/2}\quad\mbox{and}\quad\tau_{2n+1}:=\sigma^{1/2}~{}\upsilon_{2n+1}~{}\sigma^{1/2},$
as well as the gain matrices
$\beta_{2n}:=\tau_{2n}~{}\tau^{-1}\beta\quad\mbox{and}\quad\beta_{2n+1}:=\tau_{2n+1}~{}\beta^{\prime}\tau^{-1}.$
The recursions for $\tau_{n}$ and $\beta_{n}$ are initialized with $(\beta_{0},\tau_{0})=(\beta,\tau)$, where $(\beta,\tau)$ are the parameters of the reference Gaussian measure.
Sinkhorn transitions
We also denote by $G$ a $d$-dimensional centred Gaussian random variable (r.v.) with unit covariance. The Sinkhorn transitions in (5) can be expressed as
${\cal S}_{2n}(x,dy)=\mathbb{P}\left({\cal Z}_{2n}(x)\in dy\right)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}(y,dx)=\mathbb{P}\left({\cal Z}_{2n+1}(y)\in dx\right)$
with random maps defined, for any $n\geq 0$, by
${\cal Z}_{2n}(x)$ $=$ $m_{2n}+\beta_{2n}(x-m)+\tau_{2n}^{1/2}~{}G\quad\text{and}$
${\cal Z}_{2n+1}(y)$ $=$ $m_{2n+1}+\beta_{2n+1}(y-\overline{m})+\tau_{2n+1}^{1/2}~{}G.$
The parameters $m_{n}$ in the latter expressions are defined, for any $n\geq 0$, by the recursion
$m_{2n+1}$ $=$ $m+\beta_{2n+1}(\overline{m}-m_{2n}),$
$m_{2(n+1)}$ $=$ $\overline{m}+\beta_{2(n+1)}~{}(m-m_{2n+1})$
with initial condition $m_{0}=\alpha_{0}+\beta_{0}m$, where, again, $(\alpha_{0},\beta_{0})=(\alpha,\beta)$ are parameters of the reference Gaussian measure. Note that $\pi_{n}=\nu_{m_{n},\sigma_{n}}$ with
$\sigma_{2n}:=\beta_{2n}~{}\sigma~{}\beta_{2n}^{\prime}+\tau_{2n}\quad\mbox{and}\quad\sigma_{2n+1}:=\beta_{2n+1}~{}\overline{\sigma}~{}\beta_{2n+1}^{\prime}+\tau_{2n+1}.$
Gibbs-loop process
The transitions of the Gibbs-loop process introduced in Section 2.3 can also be rewritten as
${\cal S}^{\circ}_{2n}(y_{1},dy_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n}(y_{1})\in dy_{2}\right)\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1}(x_{1},dx_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n+1}(x_{1})\in dx_{2}\right)$
with the random maps
${\cal Z}^{\circ}_{2n}(y)$ ::=:= $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G$
${\cal Z}^{\circ}_{2n+1}(x)$ ::=:= $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~{}G$
defined in terms of the parameters
$\begin{array}[]{rclcrcl}\beta_{2n+1}^{\circ}&:=&\beta_{2n+1}~{}\beta_{2n},&&\tau_{2n+1}^{\circ}&:=&\tau_{2n+1}+\beta_{2n+1}~{}\tau_{2n}~{}\beta_{2n+1}^{\prime},\\ &&&&&&\\ \beta^{\circ}_{2n}&:=&\beta_{2n}~{}\beta_{2n-1},&&\text{and}~{}~{}\tau^{\circ}_{2n}&:=&\tau_{2n}+\beta_{2n}~{}\tau_{2n-1}~{}\beta_{2n}^{\prime}.\end{array}$
By Lemma 4.9 in [2], we have
$0\leq\overline{\sigma}^{-1/2}~{}\beta^{\circ}_{2(n+1)}~{}\overline{\sigma}^{1/2}=I-\upsilon_{2(n+1)}\leq(I+\varpi)^{-1}$
as well as
$0\leq\sigma^{-1/2}~{}\beta_{2n+1}^{\circ}~{}\sigma^{1/2}=I-\upsilon_{2n+1}\leq(I+\overline{\varpi})^{-1}.$
Moreover, Corollary 4.4 in [2] yields the uniform estimates
$\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}\leq\sigma,~{}~{}\forall n\geq 1.$
Lyapunov functions
From the fixed point equations (13) we have
$\overline{\sigma}=\beta^{\circ}_{2n}\overline{\sigma}\left(\beta^{\circ}_{2n}\right)^{\prime}+\tau^{\circ}_{2n}\geq\tau^{\circ}_{2n}\quad\mbox{and}\quad\sigma=\beta^{\circ}_{2n+1}\sigma\left(\beta^{\circ}_{2n+1}\right)^{\prime}+\tau^{\circ}_{2n+1}\geq\tau^{\circ}_{2n+1},$ (67)
which, in turn, imply that
$\overline{\sigma}_{\varpi}:=\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}^{\circ}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma_{\overline{\varpi}}:=\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}^{\circ}\leq\sigma.$ (68)
Also note that if $\overline{\sigma}^{-1}\leq(\tau_{2n}^{\circ})^{-1}$ then $(\tau_{2n}^{\circ})^{1/2}\overline{\sigma}^{-1}(\tau_{2n}^{\circ})^{1/2}\leq I$. Moreover, we also readily see that
$\begin{array}[]{l}\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\\ \\ =(y-\overline{m})^{\prime}\left(\beta^{\circ}_{2n}\right)^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)\\ \\ =(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)^{\prime}\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)\right]\overline{\sigma}^{-1/2}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right),\end{array}$
which implies
$\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]$ $=$ $(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[I-\upsilon_{2n}\right]^{2}\overline{\sigma}^{-1/2}(y-\overline{m})$ (69)
$+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)$
$\leq$ $(1+\lambda_{\sf min}(\varpi))^{-2}~{}(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})+d.$
We are now in position to prove Proposition 5.3.
Proof of Proposition 5.3
Consider the function
$\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$
The estimate in (69) ensures that
${\cal S}_{2n}^{\circ}(\psi)(y)=\mathbb{E}\left[\psi({\cal Z}^{\circ}_{2n}(y))\right]\leq(1+\lambda_{\sf min}(\varpi))^{-2}~{}\psi(y)+\left(\left(1-(1+\lambda_{\sf min}(\varpi))^{-2}\right)+d\right).$
Next, let us observe that
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})=\frac{1}{\sqrt{\mbox{\rm det}(2\pi\tau_{2n}^{\circ})}}\\ \\ \displaystyle\times~{}\exp{\left(-\frac{1}{2}~{}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\right)}~{}dy_{2}\end{array}$
and
$\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left(y_{2}-\overline{m}\right)+2\left(\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left(\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left((y_{2}-\overline{m})\right)+2~{}(1+\lambda_{\sf min}(\varpi^{-1}))~{}\left(y_{1}-\overline{m}\right)^{\prime}(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\left(y_{1}-\overline{m}\right),\end{array}$
where the second inequality follows from (68). Moreover, arguing again as above we arrive at
$\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)+2~{}(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~{}\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\end{array}$ (70)
The inequality (70) above together with (67) readily yields
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\frac{1}{\sqrt{\mbox{\rm det}(2\pi\overline{\sigma})}}{}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~{}\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\right)}\\ \\ \displaystyle\times~{}\exp{\left(-\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)\right)}~{}dy_{2}.\end{array}$
Thus for any $y_{1}$ such that
$\psi(y_{1})=1+(y_{1}-\overline{m})^{\prime}\overline{\sigma}^{-1}(y_{1}-\overline{m})\leq r$
we have
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\varepsilon(r)~{}\nu_{\overline{m},\overline{\sigma}_{\varpi}/2}(dy_{2})\end{array}$
where
$\varepsilon(r)=\frac{1}{\sqrt{\mbox{\rm det}(2\overline{\sigma}\overline{\sigma}_{\varpi}^{-1})}}~{}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~{}(r-1)\right)}.$
This shows that the function $\psi$, and similarly the function $\varphi$ defined in (54) satisfy the drift-minorization conditions stated in (33) and (34).
Lyapunov functions in (55) and (56)
Using the triangle inequality for Mahalanobis distances we check that
$\begin{array}[]{l}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\\ \\ \leq\left[(\beta^{\circ}_{2n}(y-\overline{m}))^{\prime}\overline{\sigma}^{-1}(\beta^{\circ}_{2n}(y-\overline{m}))\right]^{1/2}+\left[(\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G)^{\prime}\overline{\sigma}^{-1}(\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G)\right]^{1/2}\\ \\ \leq(1+\lambda_{\sf min}(\varpi))^{-1}~{}((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2},\end{array}$
which yields the almost sure estimate
$\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\leq(1+\lambda_{\sf min}(\varpi))^{-1}~{}((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2}.$
This implies that the function $\psi$, and similarly the function $\varphi$ defined in (55) satisfy the drift-minorization conditions stated in (33) and (34).
The analysis of the Lyapunov functions (56) is based on the following technical lemma.
Lemma A.1.
For any matrices $\sigma>\tau$ and any $z\in\mathbb{R}^{d}$ and $0\leq\delta\leq 1$ we have
$\mathbb{E}\left[\exp{\left(\frac{\delta}{2}(z+\tau^{1/2}G)^{\prime}\sigma^{-1}(z+\tau^{1/2}G)\right)}\right]=\displaystyle\sqrt{\frac{\mbox{\rm det}(\sigma)}{\mbox{\rm det}(\sigma-\delta\tau)}}~{}\exp{\left(\frac{\delta}{2}z^{\prime}(\sigma-\delta\tau)^{-1}z\right)}.$
Proof.
For any $g\in\mathbb{R}^{d}$ we have
$\frac{\delta}{2}~{}(z+g)^{\prime}\sigma^{-1}(z+g)-\frac{1}{2}~{}g^{\prime}\tau^{-1}g=\frac{\delta}{2}~{}z^{\prime}\sigma^{-1}z+\frac{\delta^{2}}{2}~{}z^{\prime}\sigma^{-1}\raisebox{2.0pt}{$\chi$}\sigma^{-1}z-\frac{1}{2}~{}(g-\delta\raisebox{2.0pt}{$\chi$}\sigma^{-1}z)^{\prime}\raisebox{2.0pt}{$\chi$}^{-1}(g-\delta\raisebox{2.0pt}{$\chi$}\sigma^{-1}z),$
with
$0<\raisebox{2.0pt}{$\chi$}:=(\tau^{-1}-\delta\sigma^{-1})^{-1}=-\frac{1}{\delta}~{}\sigma+\frac{1}{\delta}\sigma\tau^{-1}(\tau^{-1}-\delta\sigma^{-1})^{-1}.$
We note that
$\sigma^{-1}+\delta\sigma^{-1}\raisebox{2.0pt}{$\chi$}\sigma^{-1}$ $=$ $\sigma^{-1}+\delta\sigma^{-1}\left(-\frac{1}{\delta}~{}\sigma+\frac{1}{\delta}\sigma\tau^{-1}(\tau^{-1}-\delta\sigma^{-1})^{-1}\right)\sigma^{-1}$
$=$ $\tau^{-1}(\tau^{-1}-\delta\sigma^{-1})^{-1}\sigma^{-1}=(\sigma-\delta\tau)^{-1}$
and finally observe that
$\frac{1}{\mbox{\rm det}(\tau\raisebox{2.0pt}{$\chi$}^{-1})}=\frac{\mbox{\rm det}(\sigma)}{\mbox{\rm det}(\sigma-\delta\tau)}.$
Using Lemma A.1 we check that
$\begin{array}[]{l}\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\right)}\right]\\ \\ =\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G\right)^{\prime}\overline{\sigma}^{-1}\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G\right)\right]\right)}\right]\\ \\ =\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~{}\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2})}}~{}\exp{\left(\frac{\delta}{2}~{}(y-\overline{m})^{\prime}~{}(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~{}\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}~{}(y-\overline{m})\right)}.\end{array}$
On the other hand, we have
$\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\leq I$
and arguing as above we conclude that
$(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~{}\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}$ $=$ $(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n})^{\prime}\left(I-\delta~{}\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\right)^{-1}\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}$ (71)
$\leq$ $\frac{1}{1-\delta}~{}(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\leq\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}~{}\overline{\sigma}^{-1}.$ (72)
If we now consider the function
$\psi(y):=\exp{\left(\frac{\delta}{2}(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}$
then we have
$\begin{array}[]{l}\displaystyle\frac{1}{\psi(y)}~{}\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)\\ \\ \leq\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~{}I)}}~{}\exp{\left(-\frac{\delta}{2}~{}\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~{}(y-\overline{m})^{\prime}~{}\overline{\sigma}^{-1}~{}(y-\overline{m})\right)}.\end{array}$
Choosing
$0<\delta<1-\frac{1}{(1+\lambda_{\sf min}(\varpi))^{2}}$
we check that
$\frac{1}{(1-\delta)(1+\lambda_{\sf min}(\varpi))^{2}}<1\quad\text{and}\quad\lim_{|y|\to\infty}\frac{1}{\psi(y)}~{}{\cal S}^{\circ}_{2n}(\psi)(y)=0.$
This implies that the function $\psi$, and similarly the function $\varphi$, defined in (56) satisfy the drift-minorization conditions stated in (33) and (34).
Appendix B Hilbert metric contractions
We recall the Hilbert contraction coefficient formula ([9, 11, 12, 50], see also [8, 22, 59])
[…]
Also note that
$H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})=\sup_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)$
with the Radon-Nikodym function $\imath_{x_{1},x_{2}}({\cal K})(y)$ defined in (23). We also have
$\sup_{(x_{1},x_{2})\in S^{2}}H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})=1/\hbar({\cal K})$
with the parameter $\hbar({\cal K})$ defined in (24). Using the fact that
$\tanh\left(\frac{1}{4}~{}\log{u}\right)=\frac{u^{1/4}-u^{-1/4}}{u^{1/4}-u^{-1/4}}=\frac{u^{1/2}-1}{u^{1/2}+1}=\frac{1-u^{-1/2}}{1+u^{-1/2}}$
we conclude that
[…]
Proposition B.1.
Consider the collection of Sinkhorn transitions ${\cal S}_{n}$ defined in (4). For any $n\geq 0$,
$\hbar({\cal S}_{n})=\hbar({\cal Q})\quad\mbox{and}\quad\raisebox{2.0pt}{$\chi$}_{H}({\cal S}_{n})=\frac{1-\sqrt{\hbar({\cal Q})}}{1+\sqrt{\hbar({\cal Q})}}.$
Proof.
For any $n\geq 0$, we have
$\imath_{y_{1},y_{2}}({\cal S}_{2n+1})(x_{1})~{}\imath_{y_{2},y_{1}}({\cal S}_{2n+1})(x_{2})$ $=$ $\frac{(d\delta_{x_{1}}{\cal S}_{2n}/d\pi_{2n})(y_{1})}{(d\delta_{x_{1}}{\cal S}_{2n}/d\pi_{2n})(y_{2})}\frac{(d\delta_{x_{2}}{\cal S}_{2n}/d\pi_{2n})(y_{2})}{(d\delta_{x_{2}}{\cal S}_{2n}/d\pi_{2n})(y_{1})}$
$=$ $\imath_{x_{2},x_{1}}({\cal S}_{2n})(y_{2})~{}\imath_{x_{1},x_{2}}({\cal S}_{2n})(y_{1})$
which, together with (24), yields
$\hbar({\cal S}_{2n+1})=\hbar({\cal S}_{2n}).$
In the same vein, for any $n\geq 1$ we have
$\imath_{x_{1},x_{2}}({\cal S}_{2n})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal S}_{2n})(y_{2})$ $=$ $\frac{(d\delta_{y_{1}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{1})}{(d\delta_{y_{1}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{2})}~{}\frac{(d\delta_{y_{2}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{2})}{(d\delta_{y_{2}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{1})}$
$=$ $\imath_{y_{2},y_{1}}({\cal S}_{2n-1})(x_{2})~{}\imath_{y_{1},y_{2}}({\cal S}_{2n-1})(x_{1}),$
hence
$\hbar({\cal S}_{2n})=\hbar({\cal S}_{2n-1}).$
Recalling that ${\cal S}_{0}={\cal Q}$, this ends the proof of the proposition.
Appendix C Weighted norm contractions
We associate with some $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ the following condition.
Condition (${\cal L}_{\varphi,\psi}({\cal K})$).
There exists some parameters $\epsilon\in]0,1[$ and $c<\infty$ such that
${\cal K}(\psi)\leq\epsilon~{}\varphi+c.$ (75)
In addition, there exist some constant $r_{0}\geq 1$ and some function
$\begin{array}[]{ccccc}\varepsilon&:&[r_{0},\infty[&\mapsto&]0,1]\\ &&r&\leadsto&\varepsilon(r)\\ \end{array}$
such that, for any $r\geq r_{0}$ and $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$,
$\chi_{\varphi}(r,{\cal K}):=\sup_{(x_{1},x_{2})\in C_{\varphi}(r)^{2}}\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}<1-\varepsilon(r).$ (76)
By (20), condition (76) is equivalent to the existence of some ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$ (that may depends on $(r,x_{1},x_{2})$) satisfying for $i\in\{1,2\}$ the local minorization condition
${\cal K}(x_{i},dy)\geq\varepsilon(r)~{}{\cal Y}(dy).$ (77)
For instance, assume that $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $x_{1},x_{2}\in\mathbb{X}$ and set
$\jmath_{\varphi,\psi}(r,{\cal K}):=\inf_{(x_{1},x_{2})\in C_{\varphi}(r)}~{}\inf_{y\in C_{\psi}(r)}\imath_{x_{1},x_{2}}({\cal K})(y).$ (78)
In this notation, we see that if $\jmath_{\varphi,\psi}(r,{\cal K})\geq\varepsilon(r)$ then (77) holds.
Note that the strength of conditions (75), (76) depends on the strength of the functions $(\varphi,\psi)$. When the functions are bounded, the geometric drift condition (75) is trivially met, but in this case (77) is a uniform contraction condition on the space $\mathbb{X}$.
When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact condition (77) is met as soon as
${\cal K}(x,dy)\geq q(x,y)~{}\nu(dy)$ (79)
for some Radon positive measure $\nu$ on $\mathbb{Y}$ and some density function $q$, satisfying the local minorization condition
$\inf_{C_{\varphi}(r)\times C_{\psi}(r)}q>0\quad\mbox{\rm and}\quad 0<\nu(C_{\psi}(r))<\infty$ (80)
for any $r\geq r_{0}$. For locally compact Polish spaces condition $0<\nu(C_{\psi}(r))<\infty$ is satisfied if $\psi$ has compact sub-levels sets $C_{\psi}(r)$ with non empty interior and $\nu$ is a Radon measure of full support. Also note that the l.h.s. minorization condition (80) is satisfied when $(x,y)\in(\mathbb{X}\times\mathbb{Y})^{\circ}\mapsto q(x,y)$ is a continuous positive function on the interior $(\mathbb{X}\times\mathbb{Y})^{\circ}:=(\mathbb{X}^{\circ}\times\mathbb{Y}^{\circ})$ of $(\mathbb{X}\times\mathbb{Y})$. Also note that
[…]
Indeed, if the inequality above holds then for any $0<\epsilon<\|\Theta\|$, the set $A_{\epsilon}:=\{\Theta\geq\epsilon\}=C_{1/\Theta}(1/\epsilon)$ is a non empty compact subset and we have
${\cal K}(\psi)\leq\epsilon~{}1_{S-A_{\epsilon}}\varphi+1_{A_{\epsilon}}(\Theta~{}\varphi)\leq\epsilon~{}\varphi+c_{\epsilon}\quad\mbox{\rm with}\quad c_{\epsilon}:=\|\Theta\|\sup_{A_{\epsilon}}\varphi.$ (81)
Theorem C.1.
Assume Condition ${\cal L}_{\varphi,\psi}({\cal K})$ is satisfied for some uniformly positive functions $(\varphi,\psi)\in({\cal B}(\mathbb{X})\times{\cal B}(\mathbb{Y}))$. Then, there exist constants $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (75) and on the function $\varepsilon(r)$ in (76) and such that
[…]
Theorem C.1 is a consequence of Lemma 2.3 in [36], see also Theorem 8.2.21 in [38] and Lemma 2.1 in [6]. Several examples satisfying the Lyapunov condition (75) and the local minorization conditions (82) are discussed in [6, 38].
Proof of Theorem 4.2
Assume that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$. From (39), for any $n>n_{0}$ we have the estimates
${\cal S}^{\circ}_{2n}(\psi)$ $=$ ${\cal S}_{2n-1}({\cal S}_{2n}(\psi))\leq\epsilon~{}{\cal S}_{2n-1}(\varphi)+c\leq\epsilon^{2}~{}\psi+(1+\epsilon)~{}c\quad\text{and}$
${\cal S}^{\circ}_{2n+1}(\varphi)$ $=$ ${\cal S}_{2n}({\cal S}_{2n+1}(\varphi))\leq\epsilon~{}{\cal S}_{2n}(\psi)+c\leq\epsilon^{2}~{}\varphi+(1+\epsilon)~{}c.$ (83)
Also note that for any $n>n_{0}$ the inequality (40) in Condition ${\cal D}_{\varphi,\psi}({\cal S})$ implies that
$\|((\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n-1}){\cal S}_{2n}\|_{\sf tv}\vee\|((\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (84)
The estimates in (83) imply that (33) holds, while the inequality (84) implies that (34) is satisfied.
Moreover, combining the fixed point equations in (13) with (83) we arrive at
$\sup_{n\geq n_{0}}\left(\|{\cal S}_{2n}(\psi)\|_{\varphi}\vee\|{\cal S}_{2n+1}(\varphi)\|_{\psi}\right)<\infty\quad\mbox{\rm and}\quad(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))\leq\frac{c}{1-\epsilon},$ (85)
and (39) together with (85) yields
$\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))\leq c_{1}~{}(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))$
for some $c_{1}>0$. Therefore, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied.
Finally, we observe that (39) ensures that inequality (75) is satisfied for the Markov transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$, while (40) yields
$\chi_{\varphi}(r,{\cal S}_{2n})\vee\chi_{\psi}(r,{\cal S}_{2n+1})\leq 1-\varepsilon(r)$
and, hence, the inequality (76) also holds for ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$. As a consequence, Theorem C.1 can be applied to show that there are parameters $\kappa\geq 0$ and $\rho\in[0,1[$ such that
[…]
Appendix D Schrödinger potential functions
D.1 Some quantitative estimates
Proposition D.1.
When $\jmath({\cal Q})>0$, for any $n\geq 0$ we have the uniform estimates
$\|\frac{d\pi_{2n}}{d\nu_{V}}\|\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|\leq\hbar({\cal Q})^{-2}\quad\mbox{and}\quad\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|\leq\hbar({\cal Q})^{-2}.$ (86)
Proof.
Recalling that $\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})$, for any given $x\in\mathbb{X}$ we have
$\hbar({\cal Q})~{}{\cal S}_{2n}(x,dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})=({\cal S}_{2n-1}{\cal S}_{2n})(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})}~{}{\cal S}_{2n}(x,dy_{2}).$
This yields, for any $(y_{1},\overline{y}_{1})\in\mathbb{Y}^{2}$, the estimate
$\hbar({\cal Q})^{2}~{}{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{}{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})$
and integrating w.r.t. $\nu_{V}(d\overline{y}_{1})$ and $\pi_{2(n-1)}(dy_{1})$ we find that
$\hbar({\cal Q})^{2}~{}\nu_{V}(dy_{2})\leq\pi_{2n}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{}\nu_{V}(dy_{2}).$
Similarly, integrating w.r.t. $\nu_{V}(dy_{1})$ and $\pi_{2(n-1)}(d\overline{y}_{1})$ we also obtain
$\hbar({\cal Q})^{2}~{}\pi_{2n}(dy_{2})\leq\nu_{V}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{}\pi_{2n}(dy_{2}).$
From the inequalities above we conclude that
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq\hbar({\cal Q})^{-2}\quad\text{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq\hbar({\cal Q})^{-2}.$
More generally these Radon Nikodym functions are unbounded and we need to carry a more refined analysis.
Proposition D.2.
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds for some $n_{0}\geq 1$ then there exists $c_{\epsilon}>0$ such that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$ (87)
Proof.
Using (15) and (33), we see that, for any $n\geq n_{0}$,
$\frac{d\pi_{2n}}{d\nu_{V}}$$\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~{}\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2n_{0}}\right)\left(\psi\right)$
$\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\left(\epsilon^{n-n_{0}}\psi+\frac{c}{1-\epsilon}\right)\leq c_{\epsilon}~{}\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~{}\psi,\quad\text{with}~{}c_{\epsilon}:=\left(1+\frac{c}{1-\epsilon}\right).$
This implies that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\text{and, similarly,}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$
Proposition D.3.
In addition, when Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, there exists $c_{\epsilon}>0$ such that
$\sup_{n\geq n_{0}}\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)\leq c_{\epsilon}~{}\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right),$
as well as
$\sup_{n\geq n_{0}}\left(\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\right)\leq c_{\epsilon}~{}\left(\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\right).$
Proof.
If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$ then, by Theorem 4.2, the drift condition (33) also holds for any $n>n_{0}$, which implies that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\quad\text{and, similarly,}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}.$
Arguing as in the proof of Proposition D.2, now combining the drift conditions (39) with (16), we find that
$\sup_{n\geq n_{0}}\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}.$
Under the stronger Condition ${\cal H}_{\delta}(U,V)$, more refined estimates can be obtained. With that end, consider the functions
${\cal W}_{V,U}(x):=\log{{\cal Q}_{V}(e^{W_{U}})(x)}\quad\mbox{\rm and}\quad{\cal W}^{U,V}(y):=\log{{\cal R}_{U}(e^{W^{V}})(y)},$ (88)
with the integrated cost functions $(W_{U},W^{V})$ defined in (41) and the integral operators
${\cal Q}_{V}(x,dy)=e^{-W(x,y)}~{}\nu_{V}(dy)\quad\mbox{\rm and}\quad{\cal R}_{U}(y,dx)=e^{-W(x,y)}~{}\lambda_{U}(dx).$
Next technical lemma is pivotal.
Lemma D.4.
For any $n\geq 0$ we have
$\begin{array}[]{rcl}\log{{\cal Q}(\exp{(-V_{2n})})}&\geq&-W^{V}+\nu_{V}(V)\quad\text{and}\\ &&\\ -\nu_{V}(V)+{\cal W}^{U,V}\geq\log{{\cal R}(\exp{(-U_{2n})})}&\geq&-W_{U},\end{array}$ (89)
as well as
$U_{2n}\geq U-W^{V}+\nu_{V}(V).$
In addition, for any $n\geq 1$ we have
${\cal W}_{V,U}\geq\log{{\cal Q}(\exp{(-V_{2n})})}\geq-W^{V}+\nu_{V}(V)$ (90)
as well as
$\begin{array}[]{rcccl}U-W^{V}+\nu_{V}(V)&\leq&U_{2n}&\leq&U+{\cal W}_{V,U}\quad\text{and}\\ &&&&\\ V-W_{U}&\leq&V_{2n}&\leq&V+{\cal W}^{U,V}-\nu_{V}(V).\end{array}$ (91)
Proof.
Formula (10) yields the following monotone properties
$\begin{array}[]{rcccccl}\nu_{V}(V_{2(n+1)})&\leq&\nu_{V}(V_{2n})&=&\nu_{V}(V_{2(n+1)})+\mbox{\rm Ent}(\nu_{V}~{}|~{}\pi_{2n})&\leq&\nu_{V}(V_{0})=0,\\ &&&&&&\\ \lambda_{U}(U_{2(n+1)})&\leq&\lambda_{U}(U_{2n})&=&\lambda_{U}(U_{2(n+1)})+\mbox{\rm Ent}(\lambda_{U}~{}|~{}\pi_{2n+1})&\leq&\lambda_{U}(U_{0}).\end{array}$ (92)
Applying Jensen’s inequality, for any $n\geq 0$ we have
$\log{{\cal Q}(\exp{(-V_{2n})})(x)}$ $=$ $\log{\int~{}\nu_{V}(dy)~{}\exp{(-W(x,y)+V(y)-V_{2n}(y))}}$
$\geq$ $-W^{V}(x)+\nu_{V}(V)-\nu_{V}(V_{2n})\geq-W^{V}(x)+\nu_{V}(V).$
By (92) we also have $\lambda_{U}(U)\geq\lambda_{U}(U_{2n})$ and, therefore, for any $n\geq 0$ we obtain
$\log{{\cal R}(\exp{(-U_{2n})})(y)}$ $=$ $\log{\int~{}\lambda_{U}(dx)~{}\exp{(-W(x,y)+U(x)-U_{2n}(x))}}$
$\geq$ $-W_{U}(y)+\lambda_{U}(U)-\lambda_{U}(U_{2n})\geq-W_{U}(y)$
using Jensen’s inequality again. This ends the proof of (89).
The integral equations in (7) imply that for any $n\geq 0$ we have
$U_{2n}\geq U-W^{V}+\nu_{V}(V)\quad\mbox{\rm and}\quad V_{2(n+1)}\geq V-W_{U}.$
Using again (7) we see that
$U_{2(n+1)}-U=\log{{\cal Q}(e^{-V_{2(n+1)}})}$$\leq$ $\log{{\cal Q}_{V}(e^{W_{U}})}={\cal W}_{V,U}\quad\text{and}$
$V_{2(n+1)}-V=\log{{\cal R}(e^{-U_{2n}})}$$\leq$ $-\nu_{V}(V)+\log{{\cal R}_{U}(e^{W^{V}})}=-\nu_{V}(V)+{\cal W}^{U,V}.$
Proposition D.5.
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then
$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}\vee\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty.$ (93)
In addition, the exists some constant $c>0$ such that, for any $n\geq 1$,
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq c~{}e^{W_{U}}\quad\mbox{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq c~{}e^{W^{V}}.$ (94)
Proof.
Observe that
$\text{if}\quad\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty\quad\text{then}\quad e^{\delta U}=\frac{1}{e^{-\delta U}}\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}>0$
and this yields
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)=\lambda_{U}\left(e^{W^{V}}~{}e^{\delta U}\right)\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~{}\lambda_{U}\left(e^{W^{V}}\right).$
In the same vein, we obtain
$\nu_{(1-\delta)V}\left(e^{W_{U}}\right)=\nu_{V}\left(e^{W_{U}}~{}e^{\delta V}\right)\geq\|e^{-\delta V}\|_{\mathbb{L}_{\infty}(\nu)}^{-1}~{}\nu_{V}\left(e^{W_{U}}\right)$
and conclude that
$\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}$ $=$ $\|{\cal Q}_{V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{V}\left(e^{W_{U}}\right)$
$\leq$ $\|e^{-\delta V}\|_{\mathbb{L}_{\infty}(\nu)}~{}\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
as well as
$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{U}(e^{W^{V}})\|\leq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~{}\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$ by Condition ${\cal H}_{\delta}(U,V)$. This ends the proof of (93).
Using (91) we see that, for any $n\geq 1$,
$-{\cal W}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\cal W}_{V,U}+W^{V}-\nu_{V}(V)$
as well as
$-{\cal W}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\cal W}^{U,V}+W_{U}-\nu_{V}(V).$
Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (99) and (100) in the proof of the commutation formulae (11), we readily obtain the estimates
$e^{\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~{}e^{-W_{U}}\leq\frac{d\pi_{2n}}{d\nu_{V}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}~{}e^{W_{U}}$
as well as
$e^{\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~{}e^{-W^{V}}\leq\frac{d\pi_{2n+1}}{d\lambda_{U}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}~{}e^{W^{V}}$
Remark D.6.
The functions $({\cal W}_{V,U},{\cal W}^{U,V})$ defined in (88) can be computed explicitly for the the Gaussian model defined in (50) and (51).
We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [2]) we have
${\cal W}_{V,U}(x)=\frac{1}{2}~{}\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~{}(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~{}(\beta_{\tau}~{}(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
and
${\cal W}^{U,V}(y)=\frac{1}{2}~{}\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\frac{1}{2}(y-\overline{m})^{\prime}\left(\beta_{\tau}\,\sigma\,\beta_{\tau}^{\prime}-\tau^{-1}\right)(y-\overline{m})-(y-\overline{m})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$ (95)
Note that one readily obtain the relationships
[…]
Using (94), Condition ${\cal H}_{\delta}(U,V)$ ensures that for any $n\geq 1$ we have
$e^{-\delta V}~{}\left(\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\right)\leq c~{}e^{-{\cal V}_{\delta}}\quad\mbox{\rm and}\quad e^{-\delta U}~{}~{}\left(\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\right)\leq c~{}e^{-{\cal U}_{\delta}}$
for some constant $c<\infty$. These estimates immediately yield the following result.
Proposition D.7.
Under the assumptions of Proposition D.5, for any $n\geq 1$ we have
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\in{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\in{\cal B}_{\varphi}(\mathbb{X}).$
In addition, we have the uniform estimates
$\sup_{n\geq 1}~{}\left(\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\right)<\infty\quad\mbox{and}\quad\sup_{n\geq 1}~{}\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)<\infty.$
D.2 Lyapunov inequalities
The result below is a fairly direct consequence of the estimates presented in Lemma D.4.
Lemma D.8.
For any $n\geq 1$ we have the estimates
$e^{\nu_{V}(V)}~{}e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{x}{\cal S}_{2n}}{d\delta_{x}{\cal Q}_{V}}(y)\leq e^{-\nu_{V}(V)}~{}e^{W^{V}(x)+W_{U}(y)}$
and
$e^{\nu_{V}(V)}~{}e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{y}{\cal S}_{2n+1}}{d\delta_{y}{\cal R}_{U}}(x)\leq e^{-\nu_{V}(V)}~{}e^{W^{V}(x)+W_{U}(y)}.$
Proof.
Using the Sinkhorn transition operators as expressed in (9) together with (90) and (91), we find that, for any $n\geq 1$,
${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{(V-V_{2n})(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\geq e^{\nu_{V}(V)}e^{-({\cal W}_{V,U}(x)+{\cal W}^{V,U}(y))}{\cal Q}_{V}(x,dy)~{}$
and, in the same vein,
${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~{}e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\geq e^{\nu_{V}(V)}~{}e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{}{\cal R}_{U}(y,dx).$
Similarly, one can resort to (90) and (91) again to obtain the lower bounds
${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{V-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\leq e^{-\nu_{V}(V)}~{}e^{W^{V}(x)}~{}{\cal Q}_{V}(x,dy)~{}e^{W_{U}(y)}$
and, in the same way,
${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~{}e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\leq e^{-\nu_{V}(V)}~{}e^{W_{U}(y)}~{}{\cal R}_{U}(y,dx)~{}e^{W^{V}(x)}.$
With these estimates at hand, we can now state and prove the following Lyapunov inequalities.
Proposition D.9.
Assume that Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~{}e^{-{\cal U}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\varphi\right)/\psi\leq c_{\delta}~{}e^{-{\cal V}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{Y})$ (96)
for every $n\geq 1$ and some finite constant $c_{\delta}<\infty$.
Proof.
By Lemma D.8 we have
${\cal S}_{2n}(x,dy)~{}\psi(y)/\varphi(x)$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal U}_{\delta}(x)}~{}{\cal Q}_{(1-\delta)V}(x,dy)~{}e^{W_{U}(y)},\quad\text{and}$
${\cal S}_{2n+1}(y,dx)~{}\varphi(x)/\psi(y)$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal V}_{\delta}(y)}~{}{\cal R}_{(1-\delta)U}(y,dx)~{}e^{W^{V}(x)},$
which (via (88)) imply the inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal U}_{\delta}}~{}e^{{\cal W}_{(1-\delta)V,U}},\quad\text{and}$
${\cal S}_{2n+1}\left(\varphi\right)/\psi$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal V}_{\delta}}~{}e^{{\cal W}^{(1-\delta)U,V}}.$
On the other hand, we have the estimates
$\|e^{{\cal W}_{(1-\delta)V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}=\|{\cal Q}_{(1-\delta)V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
and
$\|e^{{\cal W}^{(1-\delta)U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{(1-\delta)U}(e^{W^{V}})\|_{\mathbb{L}_{\infty}(\nu)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ when Condition ${\cal H}_{\delta}(U,V)$ holds. This yields (96) with the constant
$c_{\delta}:=e^{-\nu_{V}(V)}~{}\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\left(\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\right).$
D.3 Proof of Theorem 4.3
Following exactly the same argument leading to (81) in Appendix C, the Lyapunov inequalities (96) in Proposition D.9 ensure that the drift conditions stated in (39) hold with with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
For $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
$e^{{\cal W}_{V,U}(x)}={\cal Q}_{V}(e^{W_{U}})(x)$ $=$ $\int~{}e^{W_{U}(y)-W(x,y)}~{}\nu_{V}(dy)$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\int~{}e^{W_{U}(y)}~{}\nu_{V}(dy)\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
and we conclude that
$w_{V,U}(r):=\sup_{C_{\varphi}(r)}{\cal W}_{V,U}<\infty.$
Similarly, for $\nu$-almost any $y\in C_{\psi}(r)$ we have
$e^{{\cal W}^{U,V}(y)}={\cal R}_{U}(e^{W^{V}})(y)$ $=$ $\int~{}e^{-W(x,y)+W^{V}(x)}~{}\lambda_{U}(dx)$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{U}(e^{W^{V}})\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$
which yields
$w^{U,V}(r):=\sup_{C_{\psi}(r)}{\cal W}^{U,V}<\infty.$
On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ we arrive at
${\cal S}_{2n}(x,dy)$$\geq$ $e^{\nu_{V}(V)}~{}e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{}\nu_{V}(dy)~{}1_{C_{\psi}(r)}(y)$
$\geq$ $e^{-w(r)}~{}\nu_{V}\left(C_{\psi}(r)\right)~{}{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad{\cal Y}_{r}(dy):=\frac{\nu_{V}(dy)~{}1_{C_{\psi}(r)}(y)}{\nu_{V}\left(C_{\psi}(r)\right)}$
and the parameter
$w(r):=\sup_{C_{\varphi}(r)\times C_{\psi}(r)}W+w_{V,U}(r)+w^{U,V}(r)-\nu_{V}(V).$
Note that for any $r>1$ we have
$\nu_{V}\left(C_{\psi}(r)\right)$ $=$ $e^{-V_{\star}}~{}\nu\left(e^{-(V-V_{\star})}~{}1_{(V-V_{\star})\leq\delta^{-1}\log{r}}\right)$
$\geq$ $w_{0}(r):=e^{-V_{\star}}~{}e^{-(\delta^{-1}\log{r})}~{}\nu\left(C_{\psi}(r)\right)>0$
and this yields, for any $x\in C_{\varphi}(r)$, the estimate
${\cal S}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{}{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~{}e^{-w(r)}.$
In the same vein, for any $y\in C_{\psi}(r)$ we have
${\cal S}_{2n+1}(y,dx)$$\geq$ $e^{\nu_{V}(V)}~{}e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{}\lambda_{U}(dx)~{}1_{C_{\varphi}(r)}(x)$
and, arguing as above, we arrive at
$\lambda_{U}\left(C_{\varphi}(r)\right)\geq w_{1}(r):=e^{-U_{\star}}~{}e^{-(\delta^{-1}\log{r})}~{}\lambda\left(C_{\varphi}(r)\right)>0.$
Thus, for any $y\in C_{\psi}(r)$ we have
${\cal S}_{2n+1}(y,dx)\geq\varepsilon_{1}(r)~{}{\cal X}_{r}(dx)$
where
$\varepsilon_{1}(r):=w_{1}(r)~{}e^{-w(r)}\quad\mbox{\rm and}\quad{\cal X}_{r}(dx):=\frac{\lambda_{U}(dx)~{}1_{C_{\varphi}(r)}(x)}{\lambda_{U}\left(C_{\varphi}(r)\right)}.$
The above estimates ensure that the local minorization condition (40) is met with $n_{0}=1$ and the function
$\varepsilon(r):=\varepsilon_{0}(r)\wedge\varepsilon_{1}(r)>0.$
Hence, we conclude that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied with $n_{0}=1$.
Using (13) and (15), for any $n\geq p\geq 1$ we verify that
$\frac{d\pi_{2n}}{d\nu_{V}}(y)-1=\left[(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right]\left(\frac{d\pi_{2p}}{d\nu_{V}}\right),$
which implies the inequality
$\left|\frac{d\pi_{2n}}{d\nu_{V}}(y)-1\right|\leq\|\frac{d\pi_{2p}}{d\nu_{V}}\|_{\psi}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}.$
On the other hand, combining Proposition D.7 with the norm equivalence (19) and Theorem 4.2, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
$~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}$$\leq$ $c_{\kappa,1}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi_{\kappa}}$
$\leq$ $c_{\kappa,1}~{}\rho^{2(n-p)}~{}|\delta_{y}-\nu_{V}|_{\psi_{\kappa}}\leq c_{\kappa,2}~{}\rho^{2(n-p)}~{}(\psi(y)+\nu_{V}(\psi)).$
Therefore, for any $n\geq 1$, we obtain
$\left\|\frac{d\pi_{2n}}{d\nu_{V}}-1\right\|_{\psi}\leq c_{\kappa,3}~{}\rho^{2n}~{}\quad\text{and, similarly,}\quad\left\|\frac{d\nu_{V}}{d\pi_{2n}}-1\right\|_{\psi}\leq c_{\kappa,4}~{}\rho^{2n},$ (97)
for some constants $c_{\kappa,i}$, $i=3,4$. By the same argument, using (13) we have, for any $n\geq p\geq 1$,
$\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)-1=\left[(\delta_{x}-\lambda_{U})({\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1})\right]\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right)$
and arrive at
$\left\|\frac{d\pi_{2n+1}}{d\lambda_{U}}-1\right\|_{\varphi}\vee\left\|\frac{d\lambda_{U}}{d\pi_{2n+1}}-1\right\|_{\varphi}\leq c_{\kappa,5}~{}\rho^{2n}~{}$ (98)
for any $n\geq 1$ and some constant $c_{\kappa,5}$.
Finally, note that for any function $f(x)>0$ we have
$\left|\log{f(x)}\right|$ $=$ $\log{\left(1+(f(x)-1)\right)}~{}1_{f(x)\geq 1}+\log{\left(1+\frac{1}{f(x)}-1\right)}~{}1_{{1}/{f(x)}\geq 1}$
$\leq$ $\left(f(x)-1\right)~{}1_{f(x)\geq 1}+\left(\frac{1}{f(x)}-1\right)~{}1_{{1}/{f(x)}\geq 1},$
where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $01+u>$ 0. This implies that
$\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}$
The inequality above combined with (97) and (98) yields the estimates
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c_{6,\kappa}~{}\rho^{2n},$
for some finite constant $c_{\kappa,6}$. This completes the proof of (42).
Appendix E Some technical proofs
Proof of (11) and (12)
We first note, using (5) and (6), that
$\frac{d{\cal P}_{2n}}{d{\cal P}_{2n+1}}(x,y)=\frac{d\pi_{2n}}{d\nu_{V}}(y)=\exp{\left(V_{2n+1}(y)-V_{2n}(y)\right)}$ (99)
and, in the same vein,
$\frac{d{\cal P}_{2n+1}}{d{\cal P}_{2(n+1)}}(x,y)=\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)=\exp{\left(U_{2(n+1)}(x)-U_{2n+1}(x)\right)}.$ (100)
If we additionally observe that
${\cal P}_{2(n+1)}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{}{\cal P}_{2n}(d(x,y))~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\cal P}_{2n+1}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n-1}}(x)~{}{\cal P}_{2n-1}(d(x,y))~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)$
then, using (4), we obtain the formulae
${\cal S}_{2(n+1)}(x,dy)$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{}{\cal S}_{2n}(x,dy)~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\cal S}_{2n+1}(y,dx)$ $=$ $~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)~{}{\cal S}_{2n-1}(y,dx)~{}\frac{d\lambda_{U}}{d\pi_{2n-1}}(x),$
and the commutation properties now follow from elementary manipulations.
Proof of (37)
Assume Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied. Combining (33) with (13) we arrive at
$\nu_{V}(\psi)=\nu_{V}({\cal S}_{2n}^{\circ}(\psi))\leq\epsilon~{}\nu_{V}(\psi)+c\quad\text{which implies}\quad\nu_{V}(\psi)\leq\frac{c}{1-\epsilon}.$
In a similar way, one can also show that
$\lambda_{U}(\varphi)\leq\frac{c}{1-\epsilon}.$
Using the Markov evolution equations in (14) together with (33) we find that, for any $n\geq n_{0}$,
$\pi_{2n}(\psi)$ $=$ $\pi_{2(n-1)}({\cal S}_{2n}^{\circ}(\psi))$
$\leq$ $\epsilon~{}\pi_{2(n-1)}(\psi)+c\leq\epsilon^{n-n_{0}}~{}\pi_{2n_{0}}(\psi)+\frac{c}{1-\epsilon}$
and, by the same argument,
$\pi_{2n+1}(\varphi)$ $=$ $\pi_{2n-1}({\cal S}_{2n+1}^{\circ}(\varphi))$
$\leq$ $\epsilon~{}\pi_{2n-1}(\varphi)+c\leq\epsilon^{n-n_{0}}~{}\pi_{2n_{0}+1}(\varphi)+\frac{c}{1-\epsilon}.$
Proof of Proposition 5.4
It is straightforward to show that
$2W_{U}(y)$ $=$ $\left\|\tau^{-1/2}(y-m_{0})\right\|_{F}^{2}+\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\log{\mbox{\rm det}(2\pi\tau)}\quad\text{and}$
$2W^{V}(x)$ $=$ $\left\|\tau^{-1/2}(\beta(x-m)+(m_{0}-\overline{m}))\right\|_{F}^{2}+~{}\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\log{\mbox{\rm det}(2\pi\tau)},$
with $m_{0}:=(\alpha+\beta m)$ and $\sigma_{\beta}:=(\beta\sigma\beta^{\prime})$, hence $W_{U}$ and $W^{V}$ are locally bounded. Moreover, for any $\delta\geq 0$ we obtain
$2(\delta U-W^{V})$ $=$ $-\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\delta~{}\log{\mbox{\rm det}(2\pi\sigma)}-\log{\mbox{\rm det}(2\pi\tau)}$
$+(x-m)^{\prime}\left(\delta~{}\sigma^{-1}-\beta^{\prime}\tau^{-1}\beta\right)(x-m)$
$-2(m_{0}-\overline{m})^{\prime}\tau^{-1}\beta(x-m)-(m_{0}-\overline{m})^{\prime}\tau^{-1}(m_{0}-\overline{m})$
and, in the same way,
$2(\delta V-W_{U})$ $=$ $-\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\delta~{}\log{\mbox{\rm det}(2\pi\overline{\sigma})}-\log{\mbox{\rm det}(2\pi\tau)}$
$+(y-\overline{m})^{\prime}\left(\delta~{}\overline{\sigma}^{-1}-\tau^{-1}\right)(y-\overline{m})$
$-2(\overline{m}-m_{0})^{\prime}\tau^{-1}(y-\overline{m})-(\overline{m}-m_{0})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$
Therefore,
${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{R}^{d})\quad\mbox{\rm and}\quad{\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{R}^{d})$
if, and only if,
$\beta^{\prime}\tau^{-1}\beta<\delta~{}\sigma^{-1}\quad\mbox{\rm and}\quad\tau^{-1}<\delta~{}\overline{\sigma}^{-1}.$
On the other hand, straightforward manipulations yield
$2(W^{V}-(1-\delta)U)$ $=$ $\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})-(1-\delta)~{}\log{\mbox{\rm det}(2\pi\sigma)}+\log{\mbox{\rm det}(2\pi\tau)}$
$-(x-m)^{\prime}\left((1-\delta)~{}\sigma^{-1}-\beta^{\prime}\tau^{-1}\beta\right)(x-m)$
$+2(m_{0}-\overline{m})^{\prime}\tau^{-1}\beta(x-m)+(m_{0}-\overline{m})^{\prime}\tau^{-1}(m_{0}-\overline{m}),$
which shows that
$\lambda_{(1-\delta)U}(e^{W^{V}})<\infty\Longleftrightarrow\beta^{\prime}\tau^{-1}\beta<(1-\delta)~{}\sigma^{-1}.$
In the same vein, we have
$2(W_{U}-(1-\delta)V)$ $=$ $\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})-(1-\delta)~{}\log{\mbox{\rm det}(2\pi\overline{\sigma})}-\log{\mbox{\rm det}(2\pi\tau)}$
$-(y-\overline{m})^{\prime}\left((1-\delta)~{}\overline{\sigma}^{-1}-\tau^{-1}\right)(y-\overline{m})$
$+2(\overline{m}-m_{0})^{\prime}\tau^{-1}(y-\overline{m})+(\overline{m}-m_{0})^{\prime}\tau^{-1}(\overline{m}-m_{0}),$
which shows that
$\nu_{(1-\delta)V}(e^{W_{U}})<\infty\Longleftrightarrow\tau^{-1}<(1-\delta)~{}\overline{\sigma}^{-1}.$
Proof of Proposition 5.5
Let us first recall that
$\lambda_{U}(dx)=\sum_{i\in E_{0}}\xi_{0}(i)~{}\lambda_{U_{i}}(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy)=\sum_{j\in E_{1}}\xi_{1}(j)~{}\nu_{V_{j}}(dx),$
which implies that
$W_{U}=\sum_{i\in E_{0}}~{}\xi_{0}(i)~{}W_{U_{i}}\quad\mbox{\rm and}\quad W^{V}=\sum_{j\in E_{1}}~{}\xi_{1}(j)~{}W^{V_{j}}$ (101)
and ensures that the functions $W_{U}$ and $W^{V}$ are locally bounded.
Moreover, we note that
$\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}\leq e^{-\inf_{i\in E_{0}}U_{i}},$
which implies that
$-U=\log{\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}}\leq-\inf_{i\in E_{0}}U_{i}$
and, hence,
$U\geq\inf_{i\in E_{0}}U_{i}.$
As a consequence, we readily obtain the inequalities
$\delta U-W^{V}\geq\inf_{i\in E_{0}}\sum_{j\in E_{1}}~{}\xi_{1}(j)~{}(\delta~{}U_{i}-~{}W^{V_{j}})\geq\inf_{(i,j)\in(E_{0}\times E_{1})}~{}(\delta~{}U_{i}-~{}W^{V_{j}})$
which, in turn, enable us to show that
$\left\{x~{}:~{}\delta U(x)-W^{V}(x)\leq r\right\}$$\subset$ $\left\{x~{}:~{}\inf_{(i,j)\in(E_{0}\times E_{1})}~{}(\delta~{}U_{i}(x)-~{}W^{V_{j}}(x))\leq r\right\}$
$=$ $\bigcup_{(i,j)\in(E_{0}\times E_{1})}\left\{x~{}:~{}(\delta~{}U_{i}(x)-~{}W^{V_{j}}(x))\leq r\right\}$
and this ensures that ${\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$. In the same vein, one can also prove that ${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$.
On the other hand, using the fact that $(a+b)^{1-\delta}\leq a^{1-\delta}+b^{1-\delta}$ for any $a,b\geq 0$, we have
$\lambda_{(1-\delta)U}(dx)$ $=$ $\left(\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}(x)}\right)^{1-\delta}~{}\lambda(dx)$ (102)
$\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}e^{-(1-\delta)U_{i}(x)}~{}\lambda(dx)=\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}\lambda_{(1-\delta)U_{i}}(dx)$
and, by way of Hölder’s inequality, we also have
$\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~{}e^{\xi_{1}(j)~{}W^{V_{j}}}\right)\leq\prod_{j\in E_{1}}~{}\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}.$ (103)
Combining (101), (102) and (103) we obtain the estimates
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$$\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~{}e^{\xi_{1}(j)~{}W^{V_{j}}}\right)$
$\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}\prod_{j\in E_{1}}~{}\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}$
from which we conclude that $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ (because $\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)<\infty$ for almost every $(i,j)$). Using the same argument, we can also prove that $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$.

### v2

@@
 Abstract
-We develop a novel semigroup contraction analysis based on Lyapunov techniques to prove the exponential convergence of Sinkhorn equations on weighted Banach spaces.
-This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences as well as in weighted Banach spaces.
-To the best of our knowledge, these are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm.
-We also illustrate the impact of these results in the context of multivariate linear Gaussian models as well as statistical finite mixture models including Gaussian-kernel density estimation of complex data distributions arising in generative models.
-Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, Lyapunov functions.
+We develop a novel semigroup stability analysis based on Lyapunov techniques and contraction coefficients to prove exponential convergence of Sinkhorn equations on weighted Banach spaces.
+This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences and Kantorovich-type criteria, including the relative entropy, squared Hellinger integrals, $\alpha$-divergences as well as weighted total variation norms and Wasserstein distances.
+To the best of our knowledge, these contraction inequalities are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm.
+We also provide Lyapunov contractions principles under minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of Sinkhorn semigroups.
+We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models.
+Last but not least, our approach also allows to consider statistical finite mixture of the above models, including kernel-type density estimators of complex data distributions arising in generative modeling.
+Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, weighted total variation norms, Lyapunov functions, Kantorovich and entropy criteria, Wasserstein semi-distances.
 Mathematics Subject Classification: 37M25, 49Q22, 47H09, 60J20; secondary: 37M25, 60J05, 94A17.
 1 Introduction
+1.1 Entropic optimal transport
+Consider locally compact Polish spaces $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ endowed with some locally bounded positive measures $\lambda$ and $\nu$.
+Consider some probability measures of the form
+$\lambda_{U}(dx):=e^{-U(x)}~\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~\nu(dy)$ (1)
+for some normalized potential functions $U:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$.
+Let $\Pi_{U,V}$ be the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{Y})$, with prescribed first and second coordinate marginals $(\lambda_{U},\nu_{V})$.
+Denote $\cal Q$} a Markov transition from $\mathbb{X}$ into $\mathbb{Y}$ of the form
+${\cal Q}(x,dy)=e^{-W(x,y)}~\nu(dy)\quad\mbox{\rm and set}\quad{\cal R}(y,dx):=e^{-W(x,y)}~\lambda(dx).$ (2)
+In the latter expression, the transition potential $W(x,y)$ stands for some real valued function on $(\mathbb{X}\times\mathbb{Y})$. Consider the reference probability measure constructed as
+${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{\cal Q}(x,dy).$ (3)
+The entropic transport problem (or Schrödinger bridge problem) with respect to (w.r.t.) the reference measure ${\cal P}_{0}$ is formulated as
+$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0}),\quad\mbox{\rm with the relative entropy}\quad\text{Ent}({\cal P}~|~{\cal P}_{0}):=\int\log\frac{d{\cal P}}{d{\cal P}_{0}}~d{\cal P}.$ (4)
+It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})<\infty$.
+This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (4) (cf.
+the seminal article by Csiszár [csiszar-2], as well as Section 6 in the Lecture Notes by Nutz [nutz], and also the survey article by Léonard [leonard] and references therein).
+Let us note that the product distribution
+$(\lambda_{U}\otimes\nu_{V})(d(x,y)):=\lambda_{U}(dx)~\nu_{V}(dy)$
+trivially belongs to $\Pi_{U,V}$. In addition, choosing ${\cal P}=(\lambda_{U}\times{\cal K})\,\in\,\Pi_{U,V}$ for some Markov transition ${\cal K}(x,dy)$ we have
+$\frac{d{\cal P}}{d{\cal P}_{0}}(x,y)=\frac{d(\lambda_{U}\otimes\nu_{V})}{d{\cal P}_{0}}(x,y)\times\frac{d{\cal P}}{d(\lambda_{U}\otimes\nu_{V})}(x,y)$
+which yields
+$\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=\int~W(x,y)~{\cal P}(d(x,y))+\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})-\int V(y)\nu_{V}(dy).$
+Choosing a potential transition of the form $W(x,y)={\cal W}(x,y)/t$, for some parameter $t>0$, problem (4) is equivalent to the optimal entropic transport problem
+$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\left(\int~{\cal W}(x,y)~{\cal P}(d(x,y))+t~\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})\right).$ (5)
+In the optimal transport literature, function $\cal W$} is sometimes called the cost function and the parameter $t>0$ is interpreted as a regularization parameter.
+In this context, the non-regularized problem corresponding to $t=0$ coincides with the Kantorovich optimal transport problem with cost function $\cal W$}.
+For a more detailed discussion on this subject we refer to [dobrushin2006perturbation, rachev1985monge, villani2008optimal] and the numerous references therein.
 1.2 Background
-The Schrödinger bridge problem is to find a random evolution with minimal relative entropy w.r.t.
+The Schrödinger bridge problem (4) is to find a random evolution with minimal relative entropy w.r.t.
 some reference physical model and with prescribed distributions of the initial and final time random states.
-This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [56, 66, 68] for recent surveys with numerous references on this subject.
-These optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [5, 31, 54, 68], statistical barycenter problems [1, 4, 10, 24, 29, 32], economy [13], computer vision [42, 76], control theory [18, 19], and many others.
+This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [leonard, nutz, peyre] for recent surveys with numerous references on this subject.
+Optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [arjovsky, doucet-bortoli, kolouri, peyre, rioux2023semi, rioux2024entropic], statistical barycenter problems [agueh, andersen, bigot, cramer-2, cuturi-w, loubes], economy [bojitov], computer vision [dominitz, solomon], control theory [chen, chen-2], and many others.
 Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure.
-For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [39, 46, 55, 74, 75] as well as the more recent studies [18, 28, 57].
-Sinkhorn iterations are based on nonlinear integral transformations over the whole state space.
-Exact finite-dimensional solutions cannot be computed for general models.
-However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [39, 46, 55, 74, 75].
-For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [2, 24, 25, 51].
-Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [14, 45, 74, 75, 77] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [18, 33, 47, 40].
-The recent article [2] also presents a refined convergence analysis for Gaussian models based on matrix Riccati difference equations.
-It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [68] as well as the more recent articles [2, 20, 23] and the references therein.
-Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [3, 17, 44].
-Linear rates on non-necessarily compact spaces were first obtained by Léger in [58] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [31, 52]).
-Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [48] under some exponential moments estimates on Schrödinger potentials.
-There are very few articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of two recent significant contributions [23, 20].
-In [23], the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic cost functions associated with a sufficiently large regularization parameter.
-These exponential entropic decays were recently refined in [20] to apply to all values of the regularization parameter.
-To the best of our knowledge, the exponential convergence of Sinkhorn iterates in the context of statistical mixture models remains an important open research question.
-As underlined in [2], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step.
-These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [6, 34, 35, 36, 70, 71, 72] and references therein.
-The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject.
-The articles [43, 70, 71] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities.
+For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3] as well as the more recent studies [chen, cuturi, leonard-2].
+Sinkhorn bridges are based on nonlinear integral transformations over the whole state space (cf.
+Section 2) and exact finite-dimensional solutions cannot be computed for general models.
+However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3].
+For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [adm-24, cramer-2, cramer-3, janadi].
+Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [borwein, fienberg, sinkhorn-2, sinkhorn-3, soules] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [chen, deligiannidis, franklin, marino].
+See [greco2023non] for a refined analysis of Sinkhorn iterates and their gradient on the $d$-dimensional torus as well as the more recent article [chizat2025sharper] on compact spaces and bounded cost functions.
+As remarked in Section 3 of [cohen], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t.
+the Hilbert metric, see for instance [borwein, chen, deligiannidis, franklin] and Section 4.1 in the present article.
+Contraction estimates w.r.t.
+the Hilbert metric yield exponential decays w.r.t.
+$\phi$-divergences, including the total variation norm, the relative entropy and Wasserstein distances of any order.
+Nevertheless they can rarely be used to derive contraction estimates with respect to some $\phi$-divergence criteria.
+At this point, we underline that even on totally bounded state spaces such as the open unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ the cost function $W$ as well as the potential functions $(U,V)$ may be unbounded and the reference Markov transition may have a unit Birkhoff contraction coefficient (i.e.
+equal to $1$), see for instance Section 3.2.1 and the series of illustrations discussed in Section 5.3 and Section 5.4.
+The landmark example of quadratic costs on Euclidian spaces is another class of unbounded costs functions.
+In this context, the recent article [adm-24] also presents a refined stability analysis of Sinkhorn bridges for Gaussian models based on matrix Riccati difference equations.
+It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [peyre] as well as the more recent articles [adm-24, chiarini, durmus] and the references therein.
+Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [alts-2017, chak-2018, dvu-2017].
+Linear rates on non-necessarily compact spaces were first obtained by Léger in [leger] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [doucet-bortoli, karimi-2024]).
+Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [promit-2022] under some exponential moments estimates on Schrödinger potentials.
+There are very few and recent articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of three significant contributions [chiarini, durmus, eckstein2025hilbert].
+The first exponential stability analysis of unbounded models and non necessarily compactly supported marginals is presented in the article [durmus].
+In this study, the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic symmetric cost functions associated with a sufficiently large regularization parameter.
+In our framework this particular class of cost functions corresponds to symmetric Gaussian reference transitions.
+The approach taken in [durmus] relies on well known links between gradient and Hessian of Sinkhorn and Schrödinger potentials along Hamilton-Jacobi-Bellman equations with conditional expectations and covariances [chewi2023entropic, pooladian2021entropic], see also [del2025entropic].
+The main conditions in [durmus] are expressed in terms of the integrated convexity profiles of marginal potentials.
+Extending ideas from [conforti2024weak] the proofs are based on the propagation of these convexity profiles based on diffusion reflection techniques introduced in [eberle2016reflection].
+The exponential entropic decays presented in [durmus] were further developed and recently refined in [chiarini] to apply to all values of the regularization parameter and some classes of perturbation of log-concave models.
+The article [chiarini] has opened new perspectives in the convergence analysis for entropically regularized transport problems with quadratic symmetric costs.
+In the context of distance-like cost functions, one of the basic tool is a remarkable stability theorem, Theorem 1.1 in [chiarini], that allows to control the entropy of Schrödinger bridges by the entropy of their marginal distributions.
+However this stability approach requires to estimate the semi-concavity properties of generally unknown dual Schrödinger bridge transitions and it only applies to entropically regularized transport problems with symmetric quadratic-type costs.
+Thus, this framework excludes the important cases when the reference transition is associated with linear Gaussian transitions arising in Ornstein-Uhlenbeck-type diffusion generative models and denoising diffusions.
+In this connection, the more recent article [del2025stability] complements the stability approach in  [chiarini] with a novel entropic continuity theorem that applies to more general classes of non necessarily symmetric transition potentials and log-concave type target marginals, including strongly convex at infinity potentials.
+The article [eckstein2025hilbert] is also based on a novel version of Hilbert’s projective metric for spaces of integrable functions of bounded growth.
+Extending results in [chen, deligiannidis] to unbounded cost functions, the recent article  [eckstein2025hilbert] presents exponential decays of these new classes of Hilbert’s projective metrics with respect to some cone of functions in settings where the marginal distributions have sufficiently light tails compared to the growth of the cost function.
+These results also imply the exponential convergence of the total variation distance.
+On normed finite dimensional spaces, the main condition in  [eckstein2025hilbert] is that $W$ grows faster that a polynomial of order $p$ and the tails of both marginals decay faster that $e^{-r^{p+q}}$ for some $q>0$.
+This condition doesn’t cover Gaussian models.
+As underlined in [adm-24], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step.
+These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [dpa, dm-04, dg-01, horton-dp, saloff-zuniga, saloff-zuniga-2, saloff-zuniga-3] and references therein.
+The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject, much more difficult to handle than their time homogeneous counterparts; this is because the operators may drastically change during the semigroup evolution.
+The articles [douc, saloff-zuniga, saloff-zuniga-2] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities.
 Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging.
 There is a need for new tools in order to handle these models.
-1.2 Contribution
-In this paper, we introduce a novel approach to Sinkhorn exponential convergence analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
-To be specific, we investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups.
-These semigroup techniques are basic quantitative methods for studying the stability of nonlinear and time-varying stochastic models, see for instance [36] as well as Section 2 in [6].
-In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [49] and Meyn and Tweedie [60], see also the book [61] and references therein.
-These powerful off-the-shelf semigroup techniques are mainly based on two conditions:
-  • 1)
-    the existence of a Lyapunov function satisfying a drift-type condition, and
-  • 2)
-    a local minorization condition on the sub-level sets of the Lyapunov function.
+In this paper, we introduce a novel approach to Sinkhorn stability analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
+1.3 Contributions
+We investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups.
+These semigroup techniques are basic quantitative methods for studying the stability of nonlinear and time-varying stochastic models, see for instance [horton-dp] as well as Section 2 in [dpa].
+In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [hasminskii] and Meyn and Tweedie [meyn-1], see also the book [meyn-2] and references therein.
+These powerful off-the-shelf semigroup techniques are mainly based on two conditions: 1) the existence of a Lyapunov function satisfying a drift-type condition; 2) a local minorization condition on the sub-level sets of the Lyapunov function.
 Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function.
 In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator.
-Based on these techniques, the analysis developed in this article yields novel contraction estimates with respect to general divergence criteria.
-It also applies to multivariate Gaussian models as well as statistical finite mixture models.
-1.3 Outline of the paper
-After presenting some basic notation in Section 1.4, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [2] for a more detailed discussion on these models within our framework).
+For a more detailed discussion on this contraction coefficients we refer to [dpa, horton-dp, penev].
+This Lyapunov approach and the operator-theoretic framework presented in this article strongly differ from the approaches discussed earlier and taken in the series of recent articles [chiarini, durmus, del2025stability, eckstein2025hilbert].
+Our main contributions can be summarized in the following way:
+  • In the context of bounded transition potentials, we provide novel strict contraction estimates with respect to general $\Phi$-divergence criteria, including $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [csiszar, dm-03, majernik, nielsen] (see, for instance, Theorem 2.2 and Section 4.1).
+  • In the context of locally bounded cost functions and marginal potentials with compact sub-level sets, we develop new weighted total variation norm estimates, including Kantorovich and Wasserstein-type distances (see Section 4.4 and Section 4.5) as well as relative entropy bridge estimates (in Section 4.6).
+    We also emphasize that the choice of the Wasserstein semi-metric and the Lyapunov functions depends on the marginal target measures (see Theorem 2.3 and Section 4.5).
+  • We develop general Lyapunov contractions principles under practical and minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of models.
+    We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models (e.g., semi-discrete marginal models).
+Last but not least, our approach also allows to consider statistical finite mixture of the above models including kernel-type density estimators of complex data distributions arising in generative models.
+To the best of our knowledge, the stability results discussed above are the first results of this type in the literature on entropic optimal transport and Sinkhorn semigroups.
+1.4 Outline of the article
+After presenting some basic notation in Section 1.5, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [adm-24] for a more detailed discussion on these models within our framework).
+Section 2.3, presents a brief summary of the main stability theorems of the article.
+All the methodologies and regularity conditions discussed in the article are illustrated in a variety of examples in Section 5.
+To motivate our results, however, Section 2.4 already includes the discussion of some application models.
+We also give some detailed comments on our regularity conditions and propose an avenue of open research projects.
 Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C.
@@
 With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups.
-Section 4.1 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups.
-The main contraction theorems introduced in this article are presented in Section 4.2 and Section 4.3.
-We illustrate the impact of these results in Section 5, which provides several examples, including strong mixing models, Gaussian models as well as Schödinger bridges between finite mixture distributions.
 Strong mixing Sinkhorn semigroups are discussed in Section 4.1 in terms of Dobrushin contraction coefficients.
 In this context, we show that Sinkhorn semigroups contract for any $\phi$-divergence criteria, including the relative entropy, the total variation distance, as well as $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, Jeffreys and Rényi divergences.
-Section 5.2 discusses the contraction properties of Gaussian Sinkhorn models.
-Last but not least, the convergence analysis of Sinkhorn semigroups in the context of statistical finite mixture models, including Gaussian mixture, is carried out in Section 5.3.
-Most of the technical proofs, as well as some extended analyses, are provided in Appendices A through E.
+Section 4.2 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups.
+The main contraction theorems introduced in this article are presented in Section 4.3.
+Section 4.4 presents a simple and new regularity condition that allows to provide quantitative exponential stability estimates for a large class of models.
+Section 4.5 presents quantitative exponential decays w.r.t.
+Wasserstein semi-metrics and Section 4.6 is dedicated to relative entropy decays and the convergence of Sinkhorn bridges.
+Section 5 is dedicated to a variety of illustrations.
+In particular, polynomial growth models are discussed in Section 5.1, Section 5.2 is dedicated to heavy tailed models, and in Section 5.3 we study bounded Lipschitz domains.
+Beta and Weibull distributions are discussed, respectively, in Section 5.4 and Section 5.5.
+An explicit description of the key regularity conditions and results in the context of multivariate linear Gaussian distributions is provided in Section 5.8.
+Finally, Section 5.9 is dedicated to statistical finite (non necessarily Gaussian) mixture models.
+Detailed proofs and some extended analyses can be found in the appendices to this manuscript.
+Appendix A is devoted to the Gaussian Sinkhorn algorithm, Appendices B and C contain material related to Hilbert metric and weighted norm contractions, respectively.
+Appendix D is devoted to Schrödinger potential functions.
+Finally, some additional technical proofs provided for completeness in Appendix E.
 1.5 Some basic notation
@@
 Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures.
-Let ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the convex subset of probability measures on $\mathbb{X}$.
-The Lebesgue integral of a integrable function $f\in{\cal B}(\mathbb{X})$ with respect to (w.r.t.) some $\nu\in{\cal M}(\mathbb{X})$ is denoted by
-$\nu(f):=\int_{\mathbb{X}}f(x)~{}\nu(dx)$
-For indicator functions $f=1_{A}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
+We denote the convex subset of probability measures on $\mathbb{X}$ as ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$.
+The Lebesgue integral of an integrable function $f\in{\cal B}(\mathbb{X})$ w.r.t. $\nu\in{\cal M}(\mathbb{X})$ is denoted by
+$\nu(f):=\int_{\mathbb{X}}f(x)~\nu(dx).$
+For indicator functions $f=1_{A}$ of a measurable subset $A\subset\mathbb{X}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
 Consider a bounded integral operator ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\cal M}(\mathbb{X})$ and a function $f\in{\cal B}(\mathbb{Y})$.
 We denote by $\mu{\cal K}\in{\cal M}(\mathbb{Y})$ and ${\cal K}(f)\in{\cal B}(\mathbb{X})$ the measure and the function defined by
-$(\mu{\cal K})(dy):=\int~{}\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y)$ (1)
+$(\mu{\cal K})(dy):=\int~\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y).$ (6)
 For indicator functions $f=1_{A}$ sometimes we write ${\cal K}(x,A)$ instead of ${\cal K}(1_{A})(x)$.
@@
 $({\cal K}_{1}{\cal K}_{2})(x_{1},dx_{2}):=\int{\cal K}_{1}(x_{1},dy){\cal K}_{2}(y,dx_{2}).$
-Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we see that
-$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(x,y))=\mu(dy)\,{\cal K}(y,dx).$
+Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we set
+$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(y,x)):=\mu(dx)\,{\cal K}(x,dy).$
 A measure $\nu_{1}\in{\cal M}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\cal M}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$.
@@
 The measures are said comparable when we have $c~\nu_{1}\leq\nu_{2}\leq c^{-1}~\nu_{1}$ for some scalar constant $c>0$.
-2 Description of the models
-2.1 Entropic optimal transport
-Consider locally compact Polish spaces $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ endowed with some locally bounded positive measures $\lambda$ and $\nu$.
-Consider some probability measures of the form
-$\lambda_{U}(dx):=e^{-U(x)}~{}\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~{}\nu(dy)$
-for some normalized potential functions $U:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$.
-Let $\Pi_{U,V}$ be the set of coupling probability measures ${\cal P}$ on the product space $(\mathbb{X}\times\mathbb{Y})$, with prescribed first and second coordinate marginals $(\lambda_{U},\nu_{V})$.
-Denote ${\cal Q}$ a Markov transition from $\mathbb{X}$ into $\mathbb{Y}$ of the form
-${\cal Q}(x,dy)=e^{-W(x,y)}~{}\nu(dy)\quad\mbox{\rm and set}\quad{\cal R}(y,dx):=e^{-W(x,y)}~{}\lambda(dx).$ (2)
-In the latter expression, $W(x,y)$ stands for some potential function on $(\mathbb{X}\times\mathbb{Y})$. In optimal transport literature $W$ is sometimes called the cost function.
-Consider the reference probability measure
-${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{}{\cal Q}(x,dy).$
-The entropic transport problem (a.k.a. dynamic Schrödinger bridge problem) w.r.t. the reference measure ${\cal P}_{0}$ is formulated as
-$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0}),$ (3)
-where $\text{Ent}({\cal P}~{}|~{}{\cal P}_{0})=\int\log\frac{\delta{\cal P}}{d{\cal P}_{0}}d{\cal P}$ is the relative entropy (or Kullback-Leibler divergence) between ${\cal P}$ and ${\cal P}_{0}$.
-It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0})<\infty$.
-This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (3) (cf.
-the seminal article by Csiszár [27], as well as Section 6 in the Lecture Notes by Nutz [66], see also the survey article by Léonard [56] and references therein).
-2.2 Schrödinger bridges
+The positive and negative part of $a\in\mathbb{R}$ are defined respectively by $a_{+}=\max(a,0)$ and $a_{-}=\max(-a,0)$.
+Given $a,b\in\mathbb{R}$ we also set $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$.
+Kantorovich and Wasserstein criteria
+We consider the weighted total variation norm associated with some positive function $\varphi\in{\cal B}(\mathbb{X})$ defined as
+[…]
+where $\Pi(\mu_{1},\mu_{2})$ stands for the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{X})$, with prescribed first and second coordinate marginals $(\mu_{1},\mu_{2})$ such that $\mu_{i}(\varphi)<\infty$.
+For a given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ and $p\geq 1$ we also consider the Wasserstein semi-metric $\mathbb{W}_{p,\varrho}$ defined by
+$\mathbb{W}_{p,\varrho}(\mu_{1},\mu_{2}):=\inf_{{\cal P}\in\Pi(\mu_{1},\mu_{2})}\left(\int~\varrho(x_{1},x_{2})^{p}~{\cal P}(d(x_{1},x_{2}))\right)^{1/p}.$ (8)
+[…]
+When $\varrho(x_{1},x_{2})$ is a metric on $\mathbb{X}$ (possibly different from the metric of the Polish space) we recall (see for instance [edwards2011kantorovich, fernique-81, kellerer1982duality, Szulga-83, villani2008optimal]) the Kantorovich-Rubinstein duality
+$\mathbb{W}_{\varrho}(\mu_{1},\mu_{2})=\sup{\left\{(\mu_{1}-\mu_{2})(f)~;~f~:~\mathbb{X}\mapsto\mathbb{R}\quad s.t.\quad|f(x_{1})-f(x_{2})|\leq\varrho(x_{1},x_{2})\right\}}.$
+Note that
+[…]
+More generally, for any given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ such that
+$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm and}\quad\varphi(x)\geq c>0$ (9)
+we have the estimates
+[…]
+For instance on a given normed vector space $(\mathbb{X},\|\mbox{\LARGE.}\|)$, for any $b,c>0$ and $p,q\geq 1$, if
+$\varphi(x):=c~\exp{\left(b\|x\|^{p}\right)}$
+then
+$\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{\tiny exp}(x_{1},x_{2}):=2c~\left(\exp{\left(\frac{b}{2^{1+(p-1)_{+}}}~\|x_{1}-x_{2}\|^{p}\right)}-1\right).$
+We check this claim using the fact that $u+v\geq 2\sqrt{u}\sqrt{v}$ and $(u^{p}+v^{p})\geq 2^{-(p-1)_{+}}(u+v)^{p}$ for any $p>0$ and $u,v\geq 0$. In this context, we have
+[…]
+The choice of the semi-metric $\varrho$ in (10) is not unique. Indeed, developing the exponential function for any $q\geq 1$ we also have the estimate
+$\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{p,q}(x_{1},x_{2}):=c_{p,q}~\|x_{1}-x_{2}\|^{pq}\quad\mbox{\rm with}\quad c_{p,q}:=2c~\frac{1}{q!}\left(\frac{b}{2^{1+(p-1)_{+}}}\right)^{q},$ (11)
+and
+[…]
+Entropy and divergences
+Consider a convex function
+$\begin{array}[]{cccl}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\quad\mbox{\rm s.t.
+$\forall a\in\mathbb{R}_{+}$~~~ $\Phi(au,av)=a~\Phi(u,v)$ and $\Phi(1,1)=0$.}\end{array}$ (12)
+The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
+$D_{\Phi}(\nu_{1},\nu_{2}):=\int~\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~\gamma(dx).$ (13)
+By homogeneity arguments, the definition (13) does not depend on the choice of the dominating measure $\gamma$. Note that
+$\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$ (14)
+Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [csiszar], i.e., we also have
+$D_{\Phi}(\nu_{1},\nu_{2})=\int~\phi\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~\nu_{2}(dx)\quad\text{where}\quad\phi(u):=\Phi(u,1)$
+and we have used the convention $D_{\Phi}(\nu_{1},\nu_{2})=\infty$ when $\nu_{1}\not\ll\nu_{2}$.
+Therefore, the $\Phi$-entropy generalizes the classical relative entropy (a.k.a.
+Kullback-Leibler divergence or $I$-divergence).
+More precisely, choosing the functions
+$\Phi(u,v)=\Phi_{1}(u,v):=u\log{(u/v)}\quad\mbox{\rm and}\quad\phi(u)=u\log{u}\$
+we recover the classical formulae
+$D_{\Phi_{1}}(\nu_{1},\nu_{2})=\mbox{\rm Ent}(\nu_{1}~|~\nu_{2}):=\int~\log{\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)}~\nu_{1}(dx)=\int~\phi_{1}\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~\nu_{2}(dx),$
+with the convex function with unique minimum $\phi_{1}(1)=0=\partial\phi_{1}(1)$ defined by
+$\phi_{1}(u)=u\log{u}-(u-1).$
+2 Statement of main results
+2.1 Sinkhorn bridges
 The Schrödinger bridge distribution $\mathbb{P}$ defined in (4) can rarely be computed analytically.
-However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [28, 74, 75].
-This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions
+However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [cuturi, sinkhorn-2, sinkhorn-3].
+This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions (${\cal P}_{n}$) and transition operators (${\cal S}_{n}$) given by
 ${\cal P}_{2n}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad{\cal P}_{2n+1}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}$ (15)
-starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dx)$ and a transition operator ${\cal S}(x,dy)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
+starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dy)$ and a transition operator ${\cal S}(y,dx)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
 $(\mu\times{\cal S})^{\flat}(d(x,y)):=\mu(dy){\cal S}(y,dx).$
 The Markov transitions ${\cal S}_{n}$ in (15) are defined by the Sinkhorn iterations
-$\left\{\begin{array}[]{l}(\pi_{2n}\times{\cal S}_{2n+1})^{\flat}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\cal S}_{2(n+1)}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\cal S}_{2n+1}.\end{array}\right.$ (5)
+$\left\{\begin{array}[]{l}(\pi_{2n}\times{\cal S}_{2n+1})^{\flat}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\cal S}_{2(n+1)}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\cal S}_{2n+1}.\end{array}\right$.
+(16)
 In the dual formulation, the bridge distributions ${\cal P}_{n}$ are often written as
@@
 $U_{2n+1}$ $=$ $~U_{2n}~:=U+\log{{\cal Q}(e^{-V_{2n}})}$
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2(n+1)}}$ $=$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2n+1}}:=V+\log{{\cal R}(e^{-U_{2n}})}$ (7)
+$V_{2(n+1)}$ $=$ $V_{2n+1}:=V+\log{{\cal R}(e^{-U_{2n}})}.$ (18)
 When $n\to\infty$, the sequence of distributions ${\cal P}_{2n}$ converges towards the Schrödinger bridge (4) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t.
@@
 the reference measure ${\cal P}_{1}$).
-As shown in Section 6 of [2], under mild integrability assumptions the Schrödinger bridge distribution $\mathbb{P}$ solving (3) can also be computed from any Sinkhorn bridge distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
+Under mild integrability assumptions (see for instance Section 6 in [adm-24]) the Schrödinger bridge distribution $\mathbb{P}$ solving (4) can also be computed from any Sinkhorn bridge reference distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
 $\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{2n}).$ (19)
@@
 ${\cal S}_{2n}(x,dy)=\frac{{\cal Q}(x,dy)e^{-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\cal S}_{2n+1}(y,dx)=\frac{{\cal R}(y,dx)~e^{-U_{2n+1}(x)}}{{\cal R}(e^{-U_{2n+1}})(y)},$ (20)
-while Proposition 6.3 in [2] yields the formulae
+while Proposition 6.3 in [adm-24] yields the formulae
 $V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$ (21)
 for the potentials $(U_{n},V_{n})$, $n\geq 0$, expressed in terms of the distribution flow of the internal states of the Gibbs-loop process.
-In addition, for any $n\geq 0$ we also have the commutation formulae
-${\cal S}_{2n}\left(\frac{d\nu_{V}}{d\pi_{2n}}\right)=\frac{d\pi_{2n+1}}{d\lambda_{U}}\quad\mbox{and}\quad{\cal S}_{2(n+1)}\left(\frac{d\pi_{2n}}{d\nu_{V}}\right)=\frac{d\lambda_{U}}{d\pi_{2n+1}}.$ (11)
-In the same vein, for any $n\geq 1$ we have
-${\cal S}_{2n-1}\left(\frac{d\lambda_{U}}{d\pi_{2n-1}}\right)=\frac{d\pi_{2n}}{d\nu_{V}}\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)=\frac{d\nu_{V}}{d\pi_{2n}}.$ (12)
-A proof for (11) and (12) is provided in Appendix E on page E.
 2.2 Sinkhorn semigroups
@@
 For any $0\leq p\leq n$ we have the semigroup equations
-$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{2n+1}}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
+$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
 The Sinkhorn transitions ${\cal S}^{\circ}_{2p,2n}$ and ${\cal S}^{\circ}_{2p+1,2n+1}$ are defined inductively by the time inhomogeneous semigroup equations
@@
 The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (23) sharing the same invariant measures (22) at each individual step.
-Using the commutation properties (11) and (12) for any $n\geq p$ we readily check that
-$\frac{d\pi_{2n}}{d\nu_{V}}$ $=$ ${\cal S}^{\circ}_{2n}\left(\frac{d\pi_{2(n-1)}}{d\nu_{V}}\right)={\cal S}^{\circ}_{2n,2p}\left(\frac{d\pi_{2p}}{d\nu_{V}}\right)\quad\text{and}$
-$\frac{d\pi_{2n+1}}{d\lambda_{U}}$ $=$ ${\cal S}_{2n+1}^{\circ}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)={\cal S}^{\circ}_{2n+1,2p-1}\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right),$ (15)
-with the backward semigroups
-${\cal S}^{\circ}_{2n,2p}:={\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1,2p-1}:={\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1}.$
-By (11) and (12) we also obtain
-$\frac{d\lambda_{U}}{d\pi_{2n+1}}$ $=$ $\left({\cal S}_{2(n+1)}\,{\cal S}^{\circ}_{2n,2p}\,{\cal S}_{2p-1}\right)\left(\frac{d\lambda_{U}}{d\pi_{2p-1}}\right)\quad\text{and}$
-$\frac{d\nu_{V}}{d\pi_{2n}}$ $=$ $\left({\cal S}_{2n+1}\,{\cal S}^{\circ}_{2n-1,2p-1}\,{\cal S}_{2(p-1)}\right)\left(\frac{d\nu_{V}}{d\pi_{2(p-1)}}\right).$ (16)
+Remark 2.1.
+At every time step $n\geq 0$, the conjugate formulae (16) also show that ${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure
+${\cal P}_{2n}(d(x,y))=\lambda_{U}(dx)~{\cal S}_{2n}(x,dy)=\pi_{2n}(dy)~{\cal S}_{2n+1}(y,dx)$
+For $n=0$, the two-blocks Gibbs sampler with the forward/backward transition ${\cal S}^{\circ}_{1}:={\cal S}_{0}{\cal S}_{1}$ and target product measure ${\cal P}_{0}=\lambda_{U}\times{\cal S}_{0}$ can be interpreted as a sampling technique for $\lambda_{U}$ based on data augmentation or auxiliary variables (a.k.a.
+proximal sampling), see for instance [damlen1999gibbs, rendell2020global, vono2020asymptotically, vono2019split] and references therein.
+For a more detailed discussion on proximal samplers we also refer to the series of articles [mou, chewi-chen, chewi-phd, jiaming, jiaojiao, lee, guan].
+Similarly, ${\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure ${\cal P}_{2n+1}$.
+2.3 Some stability theorems
+In contrast with Kantorovich criteria or related Wasserstein-type semi-distances, even when $\mathbb{X}=\mathbb{Y}$ and the marginals coincide, $\lambda_{U}=\nu_{V}$, the identity-map deterministic coupling has infinite entropy.
+That is, for any reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ of the form (3) we have
+${\cal P}(d(x,y)):=\lambda_{U}(dx)~\delta_{x}(dy)\in\Pi_{U,V}\Longrightarrow\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=+\infty.$
+This fact, together with (5), shows that the Schrödinger bridge $\mathbb{P}$ and the convergence analysis of Sinkhorn bridges ${\cal P}_{2n}$ strongly depend on the form of the reference transition $\cal Q$} and the transition potential function $W$ discussed in (2).
+Markov transitions of the form (2) with a bounded potential $W$ typically arise on finite or compact spaces.
+For instance, Markov processes on compact manifolds, such as diffusions on bounded manifolds reflected at the boundary, have Markov transitions of the form (2) with a uniformly bounded transition potential $W$ that depends on the time parameter [dpa, aronson, nash, varopoulos].
+The convergence analysis of Sinkhorn bridges in this context is often developed using Hilbert Hilbert projective metrics [chen, deligiannidis, franklin, marino].
+In the context of bounded transition potentials, the contraction principles of Markov operators presented in the article [dm-03] yield an alternative stability analysis that also applies to any generalized $\Phi$-entropies introduced in (13).
+Theorem 2.2.
+Assume that the transition potential $W$ is uniformly bounded.
+In this case, for any $\Phi$-entropy criteria $D_{\Phi}$ there is some parameters $\rho\in]0,1[$ and $c>0$ such that for any $n\geq 0$ we have the contraction inequalities
+$D_{\Phi}\left(\pi_{2n},\nu_{V}\right)\leq\rho^{n}~D_{\Phi}\left(\pi_{0},\nu_{V}\right)\quad\mbox{and}\quad\left\|\log{({d\pi_{2n}}/{d\nu_{V}})}\right\|\leq c~\rho^{n}.$ (24)
+The above inequalities remain valid when we replace $(\pi_{2n},\nu_{V})$ by $(\nu_{V},\pi_{2n})$.
+Similar estimates hold for $(\lambda_{U},\pi_{2n+1})$.
+There also exists $c>0$ such that for any $n\geq 0$ we have
+$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\rho^{n}.$ (25)
+The estimates (24) are a consequence of the sandwich-type contraction inequalities stated in Theorem 4.1.
+A more refined version of the entropy estimate (25) in terms of Dobrushin contraction coeffcients is provided in Corollary 4.2
+We underline that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}=]0,1[$ the transition potential function $W$ may be unbounded.
+Indeed, consider the reference Markov transition defined for any $x\in]0,1[$ by
+${\cal Q}(x,dy)=\frac{1}{q(x)}~\left(1-|x-y|\right)~1_{]0,1[}(y)~dy\quad\mbox{\rm with}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.$ (26)
+In this context, we have
+$W(x,y)=-\log{\left((x\wedge y)+(1-(x\vee y))\right)}+\log{q(x)}.$
+The function $W(x,y)$ is locally bounded on $0,1[^{2}$ but tends to $\infty$ on boundaries. For instance, we obtain
+$W\left(\frac{1}{n},1-\frac{1}{n}\right)\geq\log{\frac{n}{2}}-2\log{2}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}+\infty.$
+This rather simple model does not satisfy the boundedness condition of Theorem 2.2. To handle locally bounded transition potentials, consider the functions
+${\cal U}_{\delta}:=\delta U-W^{V}\quad\mbox{\rm and}\quad{\cal V}_{\delta}:=\delta V-W_{U}$ (27)
+with the integrated costs
+$W_{U}(y):=\int~\lambda_{U}(dx)~W(x,y)\quad\text{and}\quad W^{V}(x):=\int~\nu_{V}(dy)~W(x,y).$ (28)
+Intuitively, the next theorem shows that Sinkhorn semigroups are stable as soon as the marginal-potentials $(U,V)$ dominate at infinity the (integrated) transition-potentials $(W_{U},W^{V})$.
+In terms of potential energy conversion, the marginal potentials should be large enough to compensate high energy transition losses.
+A more thorough discussion on these regularity conditions is provided in Section 4.4.
+We formalize this physically intuitive principle with the second main Wasserstein-type stability theorem this article, which takes basically the following form.
+Theorem 2.3.
+Assume that $W$ is locally bounded and bounded below by a real number and $(U,V)$ are locally bounded with compact sub-level sets.
+In addition, there exists some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta/2},{\cal V}_{\delta/2})$ are locally bounded with compact sub-level sets and we have $\lambda_{(1-\delta)U}(1)\wedge\nu_{(1-\delta)V}(1)<\infty$.
+Then, there is some $\rho\in]0,1[$, $n_{0}\geq 1$ and $c>0$ s.t.
+for any $n\geq n_{0}$ we have
+[…]
+When $\psi$ satisfies the lower bound estimate (9) for some semi-metric $\varrho$ we also have the exponential Wasserstein estimates
+$\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\vee\mathbb{W}_{\varrho}(\pi_{2n},\nu_{V})\leq c~\rho^{n}.$ (30)
+[…]
+The estimates (29) are a consequence of Corollary 4.15 and the Wasserstein formulation (7) of weighted norms.
+The estimates (30) are a consequence of (10) and (29).
+More refined strict contraction estimates are provided in Corollary 4.14 and Corollary 4.16.
+2.4 Illustrations and comments
+Some illustrations
+In Section 5, the stability theorems presented in Section 2.3 and Section 4 are illustrated in a variety of situations ranging from polynomial growth potentials (Section 5.1) and heavy tailed marginals on general normed spaces (Section 5.2), to more sophisticated boundary state space models (Section 5.3), semi-compact models including semi-discrete target spaces (Section 5.7) as well as finite mixture of statistical models (Section 5.9).
+Section 5.8 is dedicated to explicit descriptions and equivalent regularity conditions in the context of linear Gaussian models.
+When $W$ is bounded, it does not appear useful to consider weighted total variation norms.
+In this context, the stability of Sinkhorn semigroups w.r.t.
+Hilbert’s metric is well known [borwein, chen, deligiannidis, franklin] and already yields exponential decays w.r.t.
+the total variation norm.
+On bounded subspaces $\mathbb{X}\subset\mathbb{R}^{d}$ the total variation norm also dominates Wasserstein distances w.r.t.
+the Euclidian norm but in general the target space $\mathbb{X}$ is non complete and then not Polish w.r.t.
+the Euclidean distance.
+In this context, it is desirable to find a complete distance on the target space that makes the corresponding Wasserstein distance in distribution space complete [bolley2008separability].
+It is easy to construct examples of marginal models on non complete metric spaces.
+For instance, consider the marginal Beta distributions $\lambda_{U}=\nu_{V}$ on $\mathbb{X}=\mathbb{Y}=]0,1[$ equipped with the Euclidean metric $\varrho_{e}\left(x_{1},x_{2}\right)=|x_{1}-x_{2}|$ and defined by
+$\nu_{V}(dx)=\frac{1}{6}~x(1-x)~\nu(dx)\quad\mbox{\rm with}\quad\nu(dx):=1_{]0,1[}(x)~dx.$
+Applying (24) to the total variation distance (see (14)) we readily check the Wasserstein exponential decays
+$\mathbb{W}_{\varrho_{e}}(\pi_{2n},\nu_{V})\leq\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\leq\rho^{n}~\|\pi_{0}-\nu_{V}\|_{\tiny tv}.$
+Note that the open unit interval $\mathbb{X}=]0,1[$ is not complete for the Euclidean metric but it is complete for the metric defined as
+$\varrho_{c}\left(x_{1},x_{2}\right):=\left|\frac{1}{x_{1}}-\frac{1}{x_{2}}\right|+\left|\frac{1}{1-x_{1}}-\frac{1}{1-x_{2}}\right|\geq 2~{\varrho_{e}}(x_{1},x_{2}).$
+On the other hand, for any parameter $\delta\in]0,1[$, if we have
+$\psi(x):=\exp{(\delta V(x))}\geq\frac{6^{\delta}}{2}~\left(\frac{1}{x^{\delta}}+\frac{1}{(1-x)^{\delta}}\right)$
+then
+$\psi(x_{1})+\psi(x_{2})\geq\varrho_{c,\delta}\left(x_{1},x_{2}\right):=c_{\delta}~\varrho_{c}\left(x_{1},x_{2}\right)^{\delta}$
+for some $c_{\delta}>0$. In this context, the l.h.s. estimate (29) yields the $\varrho_{c,\delta}$-Wasserstein estimates
+[…]
+as well as the weighted total variation norm decays
+[…]
+Theorem 2.3 also applies to Gaussian marginal distributions on $\mathbb{Y}=\mathbb{X}=\mathbb{R}^{d}$ and a transition potential $W$ with at most quadratic growth (cf.
+Section 5.1).
+In this context, applying (11) to $(p,q)=(2,1)$ we have the $2$-Wasserstein estimate (30), which implies
+[…]
+Section 5.8 provides an explicit description of the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ in the context of linear-Gaussian models.
+For instance, for one-dimensional centered Gaussian marginals with unit variance and a symmetric Gaussian reference transition $\cal Q$} with variance $\tau>2$ we have
+${2}/{\tau}<\delta<1\Longrightarrow{\cal U}_{\delta/2}(x)=c_{\delta}+\frac{x^{2}}{2}~\left(\frac{\delta}{2}-\frac{1}{\tau}\right)\stackrel{{\scriptstyle|x|\to\infty}}{{\longrightarrow}}+\infty\quad\mbox{\rm for some $c_{\delta}\in\mathbb{R}$.
+}$
+A symmetric Gaussian reference transition $\cal Q$} with variance $\tau$ can be interpreted as the transition probability of a Brownian motion on a time horizon $\tau$.
+In this context, the exponential stability of Sinkhorn bridges w.r.t.
+the relative entropy for log-concave marginals at infinity, including Gaussian marginals, also holds for any time horizon $\tau>0$, see for instance [del2025stability, chiarini, durmus].
+The article [adm-24] also provides Wasserstein exponential decays for any Gaussian marginals and any linear-Gaussian reference transition.
+For these classes of log-concave-type models, our Lyapunov approach yields new weighted total variation norm estimates for sufficiently large time horizons.
+Energy barriers
+Section 2.3 and Section 5 discuss a series of application models that can be handled using our Lyapunov approach.
+This section is concerned with some limitations of our approach and related important open research problems.
+Firstly, note that Theorem 2.3 applies when $W$ is locally bounded.
+But even well defined absolutely continuous reference transitions on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}=]0,1[$ may have a potential $W$ which is not well defined on the whole space.
+Indeed, consider the reference Markov transition defined for any $x\in]0,1[$ by
+$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{q(x)}~|x-y|~\nu(dy)\\ \\ \displaystyle\mbox{\rm with}\quad\nu(dy):=1_{]0,1[}(y)~dy\quad\mbox{\rm and}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.\end{array}$
+This rather elementary example is a a slight modification of (26) but in this case the function $W$ is not defined on the diagonal and for any $x\not=y$ we have
+$\begin{array}[]{l}\displaystyle W(x,y)=-\log{|x-y|}+\log{q(x)}\Longrightarrow W(x,x)=+\infty.\end{array}$
+To the best of our knowledge, the stability analysis of Sinkhorn bridges in the context of energy-barrier cost functions of this form remains an open research question.
+Absolutely continuous target marginals may also have an infinite potential, for instance
+$\begin{array}[]{l}\displaystyle\lambda_{U}(dx):=4~|x-1/2|~\lambda(dx)\quad\mbox{\rm with}\quad\lambda(dx):=1_{]0,1[}(x)~dx\\ \\ \displaystyle\text{~~yields~~}U(x)=-\log{4}-\log{|x-1/2|}\Longrightarrow U(1/2)=+\infty.\end{array}$
+In this case the potential function $U$ is not well defined on the whole state space, thus this marginal model cannot be written in terms of a Boltzmann-Gibbs measure of the form (1).
+Nevertheless for uniformly bounded transition potential functions, Theorem 2.2 applies to any class of marginal distributions.
+Last but not least, note that the Weibull marginal distribution with shape parameter strictly less than one is locally bounded on $\mathbb{X}=]0,\infty[$ but does not have compact sub-level sets; for instance, we have
+$\lambda_{U}(dx)=\frac{1}{2\sqrt{x}}~e^{-\sqrt{x}}~1_{]0,\infty[}(x)~dx\Longrightarrow U(x)=\sqrt{x}+\log{\sqrt{x}}+\log{2}\stackrel{{\scriptstyle x\to 0^{+}}}{{\longrightarrow}}-\infty.$
+The potential function $U$ above is well defined on the whole state space but does not have compact sub-level sets.
+Even for bounded transition potential functions Theorem 2.3 cannot be used (but as mentioned above, Theorem 2.2 applies to any marginal models).
 3 Contraction of Markov operators
-We recall some definitions of contraction coefficients for Markov operators, both in terms of the Hilbert projective metric and $\Phi$-entropies.
+Section 3.1 is dedicated to a brief description of the Hilbert (projective) metric and integral operator norms.
+Section 3.2 recalls some definitions of contraction coefficients for Markov operators in the context of Hilbert projective metrics, weighted total variation norms and $\Phi$-entropies.
 These coefficients will be key to the analysis of contraction properties of Sinkhorn semigroups in Section 4.
@@
 $H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~\|\nu_{1}-\nu_{2}\|_{\sf tv},$ (31)
-[…]
+where $\|\cdot\|$ and $\|\cdot\|_{\sf tv}$ denote the supremum and total variation norms, respectively, and $H(\nu_{1},\nu_{2})=\infty$ when the measures $\nu_{1},\nu_{2}$ are not comparable.
+The proof of the r.h.s.
+estimate in (31) is provided in [atar] (cf.
+Lemma 1), see also [legland].
+For a more thorough discussion on Hilbert projective metrics we refer the reader to the recent review and self contained article on Hilbert projective metrics [cohen].
 3.1.2 Integral operator norms
 Given some functions $\varphi\in{\cal B}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\cal M}_{\varphi}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\cal M}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
-${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}$ ::=:= $\sup\{|\nu(f)|~{}:~{}\|f\|_{\varphi}\leq 1\}=\sup\{|\nu(f)|~{}:~{}\mbox{\rm osc}_{\varphi}(f)\leq 1\}$ (18)
-$=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation}\quad|\nu|:=\nu_{+}+\nu_{-}.$
+[…]
+$=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation measure}\quad|\nu|:=\nu_{+}+\nu_{-}.$
 In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\cal B}(\mathbb{X})$, which is given by
 $\mbox{\rm osc}_{\varphi}(f):=\sup_{x_{1},x_{2}}\frac{|f(x_{1})-f(x_{2})|}{\varphi(x_{1})+\varphi(x_{2})}\leq\|f\|_{\varphi}.$
-For a detailed proof of the equivalent formulations in the latter definition (18), we refer to Proposition 8.2.16 in [38].
-Whenever $\varphi\geq 1$, for any $\kappa>0$, we can construct the function $\phi_{\kappa}=1+\kappa\phi$ and obtain the norm equivalence formulae
-$\kappa~{}\|\nu\|_{\sf tv}\leq\kappa~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi_{\kappa}}\leq(2+\kappa){\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}.$ (19)
-The choice of condition $\varphi\geq 1/2$ in the definition of the $\varphi$-norm given in (18) is imposed only to recover the conventional total variation dual distance ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\mu_{1}-\mu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{1/2}=\|\mu_{1}-\mu_{2}\|_{\sf tv}$ between probability measures $\mu_{1},\mu_{2}\in{\cal M}_{1}(\mathbb{X})={\cal M}_{1,\frac{1}{2}}(\mathbb{X})$, where ${\cal M}_{1,\varphi}(\mathbb{X})$ denotes the subset of probability measures $\nu$ s.t.
-$\nu(\varphi)<\infty$.
-Recall that for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
+For a detailed proof of the equivalent formulations in the latter definition (32), we refer to Proposition 8.2.16 in [penev].
+The weighted Wasserstein formulation (7) of the $\varphi$-norm is a consequence of Kellerer’s extension of the Kantorovich-Rubinstein theorem [edwards2011kantorovich, kellerer1984duality, kellerer1982duality], see also the recent article [delgerber2025].
+[…]
+Recall that, given an arbitrary Polish space $S$, for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
 $\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\cal M}_{1}(S)~:~\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$ (33)
-3.1.3 Entropy and divergences
-Consider a convex function
-$\begin{array}[]{cccc}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\\ \end{array}$
-such that for any $\alpha,v,w\in\mathbb{R}_{+}$ we have
-$\Phi(\alpha w,\alpha v)=\alpha~{}\Phi(u,v)\quad\mbox{\rm and}\quad\Phi(1,1)=0.$ (21)
-The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
-$D_{\Phi}(\nu_{1},\nu_{2}):=\int~{}\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~{}\gamma(dx).$
-By homogeneity arguments, the above definition does not depend on the choice of the dominating measure $\gamma$. Note that
-$\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~{}|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$ (22)
-Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [26], i.e., we also have
-$D_{\Phi}(\nu_{1},\nu_{2})=\int~{}\phi\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~{}\nu_{2}(dx)\quad\text{where}\quad\phi(u):=\Phi(u,1)$
-and we have used the convention $D_{\Phi}(\nu_{1},\nu_{2})=\infty$ when $\nu_{1}\not\ll\nu_{2}$.
-Therefore, the $\Phi$-entropy generalizes the classical relative entropy (a.k.a.
-Kullback-Leibler divergence or $I$-divergence).
-More precisely, choosing the functions
-$\Phi(u,v)=\Phi_{1}(u,v):=u\log{(u/v)}\quad\mbox{\rm and}\quad\phi(u)=u\log{u}$
-we recover the classical formulae
-$D_{\Phi_{1}}(\nu_{1},\nu_{2})=\mbox{\rm Ent}(\nu_{1}~{}|~{}\nu_{2}):=\int~{}\log{\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)}~{}\nu_{1}(dx)=\int~{}\phi_{1}\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~{}\nu_{2}(dx),$
-with the convex function with unique minimum $\phi_{1}(1)=0=\partial\phi_{1}(1)$ defined by
-$\phi_{1}(u)=u\log{u}-(u-1).$
-Other examples of $\phi$-divergences includes $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [26, 37, 63, 65].
+Choosing $\varphi\geq 1$ and $\kappa>0$, for any $\nu\in{\cal M}_{b}(\mathbb{X})$ we also have the norm equivalence formulae
+[…]
+For exponential functions of the form
+$\varphi(x)=c~e^{g(x)}\mbox{\rm~~with some parameter $c\geq 0$ and~~}g(x_{1})+g(x_{2})\geq\varrho_{g}(x_{1},x_{2})$
+for some semi-metric $\varrho_{g}$ we also have the estimate
+$\varphi(x_{1})+\varphi(x_{2})\geq 2c~e^{g(x_{1})/2}e^{g(x_{2})/2}\geq\varrho_{\varphi}(x_{1},x_{2}):=2c~\left(e^{\frac{1}{2}\varrho_{g}(x_{1},x_{2})}-1\right).$
+Also note that for any $p\geq 1$ we have
+$\varrho_{\varphi}(x_{1},x_{2})\geq c_{p}~\varrho_{g}(x_{1},x_{2})^{p}\quad\mbox{\rm with}\quad c_{p}:=\frac{c}{2^{p-1}p!}$
+and this implies that
+[…]
+Last but not least, on normed vector spaces consider a function of the form
+$\varphi(x)\geq a+\|b(x)\|^{p}$
+for some function $b(x)$ and $a,p>0$. In this case we have
+$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho(x_{1},x_{2}):=\frac{1}{2^{(p-1)_{+}}}~\|b(x_{1})-b(x_{2})\|^{p}$ (36)
+as well as the Wasserstein estimate (10).
 3.2 Some contraction coefficients
@@
 We also consider the $[0,1]$-valued parameters
-$\jmath({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~{}\inf_{y\in\mathbb{Y}}~{}\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
-$\hbar({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (24)
+$\jmath({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~\inf_{y\in\mathbb{Y}}~\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
+$\hbar({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (38)
 and we note that
@@
 The Birkhoff contraction coefficient of a Markov transition $\cal K$} from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
-[…]
+$\raisebox{2.0pt}{$\chi$}_{H}({\cal K}):=\sup\frac{H(\nu_{1}{\cal K},\nu_{2}{\cal K})}{H(\nu_{1},\nu_{2})}=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}},$ (40)
 where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B.
 Combining (39) and (40) we readily obtain the equivalent inequalities
+${\raisebox{2.0pt}{$\chi$}_{H}({\cal K})<1}\Longleftrightarrow\hbar({\cal K})>0\Longleftrightarrow\jmath({\cal K})>0.$
+We also emphasize that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}$ there is an interesting twist because the uniform minorization condition $\jmath({\cal K})>0$ is generally not satisfied.
+We illustrate this assertion with the semi-circle transition on the unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ defined by
+${\cal K}(x,dy)=\frac{1}{q(x)}~\sqrt{1-(x-y)^{2}}~1_{]0,1[}(y)~dy$ (41)
+for some normalizing constant $1/3\leq q(x)\leq 1$. In this case, choosing $x_{1}=1/(2n)=1-y=1-x_{2}$ in (37) we check that
+$0\leq\jmath({\cal K})\leq 3~\sqrt{1-(1-1/n)^{2}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0\Longrightarrow\jmath({\cal K})=0.$
+The entropic transport model (4) on the unit interval ]0,1[]0,1[ with reference transition (41) is discussed in Section 5.4. Note that in that case, the cost function is unbounded.
+3.2.2 Dobrushin contraction coefficients
+Consider functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ for any $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.
+The $(\varphi,\psi)$-Dobrushin coefficient $\raisebox{2.0pt}{$\chi$}_{\varphi,\psi}({\cal K})$ of a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by the norm operator
 […]
-As remarked in Section 3 of [22], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t.
-the Hilbert metric.
-3.2.2 Dobrushin contraction coefficients
-[…]
-$\chi_{\varphi,\psi}({\cal K})=\sup_{\nu_{1},\nu_{2}\in{\cal M}_{1,\varphi}(S)}\frac{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}{\cal K}-\nu_{2}{\cal K}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}-\nu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}}$
 When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\cal K})$ instead of $\chi_{\varphi,\varphi}({\cal K})$.
@@
 $\chi_{\varphi}({\cal K}_{0}{\cal K}_{1})=\chi_{\varphi,\varphi}({\cal K}_{0}{\cal K}_{1})\leq\chi_{\varphi,\psi}({\cal K}_{0})~\chi_{\psi,\varphi}({\cal K}_{1}).$ (42)
-[…]
-[…]
+The terminology $\varphi$-Dobrushin coefficient comes from the fact that we recover the standard Dobrushin contraction coefficient $\raisebox{2.0pt}{$\chi$}({\cal K})$ by choosing constant functions.
+Indeed, if $\varphi=1=\psi$, then
+$\raisebox{2.0pt}{$\chi$}_{1,1}({\cal K})=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{\|\nu_{1}{\cal K}-\nu_{2}{\cal K}\|_{\sf tv}}{\|\nu_{1}-\nu_{2}\|_{\sf tv}}=:\raisebox{2.0pt}{$\chi$}({\cal K})=1-\varepsilon({\cal K}),$ (43)
 with the parameter
@@
 3.2.3 Entropy contraction coefficient
+Given a convex function $\Phi$ satisfying (12), the Dobrushin $\Phi$-contraction coefficient $\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})$ of some Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
+$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K}):=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{D_{\Phi}(\nu_{1}{\cal K},\nu_{2}{\cal K})}{D_{\Phi}(\nu_{1},\nu_{2})},$ (46)
+where the supremum is taken over all probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ such that $D_{\Phi}(\nu_{1},\nu_{2})>0$.
+Here, again, the terminology $\Phi$-Dobrushin coefficient comes from the fact that choosing $\Phi(u,v)=\Phi_{0}(u,v)=\frac{1}{2}|u-v|$ we recover the standard Dobrushin contraction coefficient, i.e., $\raisebox{2.0pt}{$\chi$}_{\Phi_{0}}({\cal K})=\raisebox{2.0pt}{$\chi$}({\cal K})$.
+For any convex function $\Phi$ satisfying (12) we recall that
+$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})\leq\raisebox{2.0pt}{$\chi$}({\cal K}).$ (47)
+The estimate (47) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [cohen-0] for finite state spaces and [dm-03] for general measurable spaces).
+For further details on the Dobrushin contraction coefficient we refer the reader to the article [dob-56], see also [dg-01, dm-03] and Chapter 4 in [dm-04].
+4 Sinkhorn contraction properties
+4.1 Strong mixing models
+As underlined in [fienberg, mosteller] in the context of finite state space models, Sinkhorn iterations of the form (16) preserve the cross product Radon-Nikodym ratio in (38) and, hence, for any $n\geq 0$ we have
+$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}).$ (48)
+This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [borwein, chen, deligiannidis, franklin].
+Recall that the Hilbert contraction coefficient of a Markov transition $\cal Q$} can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq.
+(40)), hence the exponential convergence of Sinkhorn iterations w.r.t.
+the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$.
+An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
+From (39), let us observe that
+$\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})\geq\jmath({\cal Q})^{2}.$
+Thus, for any given convex function $\Phi$ satisfying (12), we can use (40) and (47) to show that
+$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
+while
 […]
-[…]
-[…]
-For any convex function $\Phi$ satisfying (21) we recall that
-[…]
-The estimate (32) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [21] for finite state spaces and [37] for general measurable spaces).
-For further details on the Dobrushin contraction coefficient we refer the reader to the article [41], see also [35, 37] and Chapter 4 in [34].
-4 Sinkhorn contraction properties
-4.1 Some regularity conditions
+We can also obtain the sandwich-type contraction inequalities described below.
+Theorem 4.1.
+For any convex function $\Phi$ satisfying (12) we have the contraction estimates
+$\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$ (50)
+The above inequalities remain valid when we replace $(\nu_{V},\pi_{2n})$ by $(\lambda_{U},\pi_{2n+1})$.
+If we assume that $\jmath({\cal Q})>0$, then $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ and, in addition, there exists some $c>0$ such that
+$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~\raisebox{2.0pt}{$\chi$}({\cal S})^{2n}\quad\text{for any $n\geq 0$.}$ (51)
+Proof.
+The estimates (50) are a direct consequence of (47) combined with the Markov transport formulae
+$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\cal S}_{2(n+1)},\pi_{2n+1}{\cal S}_{2(n+1)}),\quad\text{and}$
+$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
+which are themselves easily obtained from expressions (16) and (22). The proof of (51) follows exactly the same arguments as the proof of the inequality (71) in Theorem 4.13, hence it is skipped.
+The following result can be obtained from the series expansion given by Corollary 4.20.
+Corollary 4.2.
+Under the assumptions of Theorem 4.1, we have the entropy exponential decay estimates
+$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\raisebox{2.0pt}{$\chi$}({\cal S})^{2n},\quad\text{for any $n\geq 0$}.$
+Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [dg-01].
+For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself.
+The choice ${\cal Q}(x,dy)~\propto~e^{-\delta\,\|y-b(x)\|}~dx$ yields
+$\jmath({\cal Q})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
+In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [aronson], Nash [nash] and Varopoulos [varopoulos] on Gaussian estimates for heat kernels on manifolds.
+Nevertheless, the above condition is not met for Gaussian transitions.
+4.2 Gibbs-Loop Sinkhorn semigroups
 The $r$-sub-level set of a positive function $\varphi\in{\cal B}(\mathbb{X})$ is denoted as
-$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~{}:~{}\varphi(x)\leq r\right\}.$
+$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~:~\varphi(x)\leq r\right\}.$ (52)
 When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$.
-Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [6] and references therein).
+Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [dpa] and references therein).
 In the context of Gibbs-loop semigroups (23), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
-Condition (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
+Condition 4.3 (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
 There exist some parameters $\epsilon\in]0,1[$ some constant $c>0$ and some integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and for any $n\geq n_{0}$ we have the estimates
@@
 There also exists some $r_{0}\geq 1$ and some function
-$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
+[…]
 such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
@@
 Whenever Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, it can be easily shown that
-$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.$ (37)
+$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.\$(57)
 These estimates are direct consequences of the drift conditions (53). A detailed proof is provided in Section E on page E.
-4.2 Norm contraction theorems
+4.3 Weighted norm contractions
 The drift conditions (53) combined with the local minorization conditions (54) provide a way to control the contraction coefficients of Sinkhorn transitions.
@@
 Theorem 4.4.
-If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, then there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that
+If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied some integer $n_{0}\geq 1$, then there exist parameters $\kappa>0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (53) and on the function $\varepsilon(r)$ in (54) and such that
 $\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\cal S}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (58)
 Theorem 4.4 is a straightforward consequence of Theorem C.2 in Appendix C.
-The contraction estimate (38) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $0\leq p\leq n$ we readily obtain
-$\|(\mu_{1}-\mu_{2}){\cal S}^{\circ}_{2p,2n}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\mu_{1}-\mu_{2}\|_{\psi_{\kappa}}.$
-Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\pi_{0},\nu_{V})$, the above estimate immediately yields an upper bound on the total variation distance, i.e.,
-$\|\pi_{2n}-\nu_{V}\|_{\sf tv}\leq\|\pi_{2n}-\nu_{V}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\pi_{0}-\nu_{V}\|_{\psi_{\kappa}}.$
-Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [30, 70, 71]
-$\|{\cal S}^{\circ}_{0,2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{0,2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{n}~{}\left(\psi(y_{1})+\psi(y_{2})\right).$
-Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may make it hard to test –except in the case of Gaussian and linear models.
+The contraction estimate (58) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $n_{0}\leq p\leq n$ we readily obtain
+[…]
+Note that
+[…]
+Choosing $p=n_{0}$ and $(\mu_{1},\mu_{2})=(\pi_{2(n_{0}-1)},\nu_{V})$, the above estimate immediately yields, for any $n\geq n_{0}$, an upper bound on the $\psi$-norm.
+We summarize the above discussion with the following corollary.
+Corollary 4.5.
+Under the assumptions of Theorem 4.4 for any $n\geq n_{0}$ we have
+[…]
+[…]
+with the parameters $\rho\in[0,1[$ and $\kappa>0$ as in (58).
+Choosing $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [persi, saloff-zuniga, saloff-zuniga-2]
+$\|{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{1+n-n_{0}}~\left(\psi_{\kappa}(y_{1})+\psi_{\kappa}(y_{2})\right).$
+Remark 4.6.
+The proximal sampler discussed in Remark 2.1 has a Markov transition ${\cal S}^{\circ}_{1}$ with invariant measure $\lambda_{U}$.
+The weighted-norm contraction properties of these models are simply based on the existence of a Lyapunov function $\varphi$ satisfying the r.h.s.
+condition in (53) with $n=0$.
+As in (54) we also require a local minorization condition $\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{1}\|_{\sf tv}\leq 1-\varepsilon(r)$ on the sub-level sets $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$ of the Lyapunov function to ensure that $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{1})<1$, for some parameter $\kappa\geq 0$ (see for instance Theorem C.2).
+Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may not be easy to test – except in the case of Gaussian and linear models.
 Several Lyapunov functions in this context are designed in Section 5.8.
-A stronger (but easier to test) condition can be provided in terms of the Sinkhorn transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
-Condition (${\cal D}_{\varphi,\psi}({\cal S})$).
+A stronger (but easier to test) condition can be provided in terms of the transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
+Condition 4.7 (${\cal D}_{\varphi,\psi}({\cal S})$).
 There exists some parameter $\epsilon\in]0,1[$, some constant $c>0$ and some integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$ we have the estimates
@@
 There also exists some $r_{0}\geq 1$ and some function
-$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
+[…]
 such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
 $\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (61)
-When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, we readily see that
-${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y})$
-for any $n\geq n_{0}$.
+When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, for any $n\geq n_{0}$ we readily see that
+${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y}).$
+In addition, we have
+$\pi_{2n+1}(\varphi)=\nu_{V}{\cal S}_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c$ (62)
+[…]
 Theorem 4.8.
-If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (33) and (34) also hold for any $n>n_{0}$.
-In addition, there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that, for any $n\geq n_{0}$,
+Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds as soon as condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied and the estimates (59) hold for some $n_{0}\geq 1$ and some parameters $\rho\in[0,1[$ and $\kappa>0$.
+When ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied there also exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (60) and on the function $\varepsilon(r)$ in (61) and such that, for any $n\geq n_{0}$,
+$\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa},\psi_{\kappa}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa},\varphi_{\kappa}}({\cal S}_{2n+1})<\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
+A proof of Theorem 4.8 is provided in Appendix C (on page C).
+Let us emphasize that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is stronger than Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$.
+In particular, the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ can be stable even when the Sinkhorn transitions ${\cal S}_{n}$ are unstable.
+Intuitively, Gibbs-loop transitions go forward and backward towards their invariant measures (22) while the Sinkhorn transitions in (16) transport asymptotically a target marginal measure to the other.
+For instance, by Lemma 4.9 in [adm-24] (see also Section A) the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ of the Gaussian Sinkhorn algorithm are always stable.
+However, in the context of regularized Gaussian entropic problems on a time horizon $t$ (cf.
+Section 3.4 in [adm-24]) the Sinkhorn transitions ${\cal S}_{n}$ may be unstable when the time horizon is not sufficiently large (by Remarks 3.9 and 4.6 in [adm-24]).
+4.4 A Lyapunov stability theorem
+Regularity conditions
+This section presents some sufficient conditions for ${\cal D}_{\varphi,\psi}({\cal S})$, in terms of the potential functions $(U,V,W)$.
+These simple conditions enable us to obtain geometric rates for the convergence of the Sinkhorn iterations for a variety of models, including polynomial growth models (Section 5.1) and heavy tailed models (Section 5.2), as well as boundary states models including Beta and Weibull distributions (Section 5.3 and Section 5.4).
+Illustrations in the context of Linear-Gaussian models and finite (non necessarily Gaussian) mixture marginals are also discussed respectively in Section 5.8 and Section 5.9.
+Let us assume that $U,V,W$ are locally bounded and $e^{-W}\in\mathbb{L}_{\infty}(\lambda\otimes\nu)$.
+In addition, $U,V$ have compact sub-level sets.
+These conditions ensure there exists some $W_{\star}\in\mathbb{R}$ such that
+$W(x,y)\geq W_{\star}\quad\mbox{\rm for $(\lambda\otimes\nu)$-almost every $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.}$ (63)
+In addition, $(U,V)$ are bounded below by a real number so that $e^{-U}$ and $e^{-V}$ are uniformly bounded.
+This excludes probability densities that tends to infinity at boundary states, such that some classes of Weibull or the Beta densities for some shape parameters.
+Also note that for any $\delta>0$ we have
+$(e^{\delta U},e^{\delta V})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
+Consider the following conditions:
+Condition 4.9 (${\cal H}_{\delta}(U,V)$).
+There is some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ defined in (27) are locally bounded with compact sub-level sets and we have
+$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$ (64)
+Note that (63) and (64) together imply the inequalities
+$\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)\leq e^{-W_{\star}}~\left(\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\right)<\infty.$
+This yields the exponential moment estimates
+${\cal H}_{\delta}(U,V)\Longrightarrow\lambda_{(1-\delta)U}\left(1\right)=\lambda_{U}\left(\exp{(\delta U)}\right)<\infty\quad\mbox{\rm and}\quad\nu_{(1-\delta)V}(1)=\nu_{V}\left(\exp{(\delta V)}\right).$
+In particular all absolute moments $\lambda_{U}(|U|^{p})\vee\nu_{V}(|V|^{p})<\infty$, with $p\geq 0$ are finite. Given the reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ also have
+${\cal H}_{\delta}(U,V)\Longrightarrow\mbox{\rm Ent}((\lambda_{U}\otimes\nu_{V})~|~{\cal P}_{0})=(\lambda_{U}\otimes\nu_{V})(W)-\nu_{V}(V)<\infty.$ (65)
+To check this claim, note that
+$(\lambda_{U}\otimes\nu_{V})(W)$ $=$ $\lambda_{U}(W^{V})=W_{\star}+\lambda_{U}(W^{V}-W_{\star})$
+$\leq$ $W_{\star}+\lambda_{U}(e^{W^{V}-W_{\star}})=W_{\star}+e^{-W_{\star}}~\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}).$
+On the other hand, for any $r>0$ we have
+$\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}})$ $\leq$ $e^{-r}~\lambda_{(1-\delta)U}(1)+\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}~1_{{\cal U}_{\delta}\leq r})$
 […]
-A proof of Theorem 4.2 is provided in Appendix C (on page C).
-4.3 A Lyapunov stability theorem
-Finally, we introduce a more tractable sufficient condition, in terms of the potential functions $(U,V,W)$.
-It enables us to obtain geometric rates for the convergence of the Sinkhorn iterations, as explicitly shown for the case of linear Gaussian models and finite mixture models in Sections 5.2 and 5.3, respectively.
-Let us assume that $W$ is locally bounded, $(U,V)\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y}))$ and $e^{-W}\in\mathbb{L}_{\infty}(\lambda\otimes\nu)$.
-In addition, assume that $U$ and $V$ have compact sub-levels sets with non empty interior and the measures $\lambda$ and $\nu$ have full support, i.e., they are finite on compact sets and strictly positive on non-empty open sets.
-Let us also denote
-$W_{U}(y):=\int~{}\lambda_{U}(dx)~{}W(x,y)\quad\text{and}\quad W^{V}(x):=\int~{}\nu_{V}(dy)~{}W(x,y).$ (41)
-Condition (${\cal H}_{\delta}(U,V)$).
-There is some $\delta\in]0,1[$ such that ${\cal U}_{\delta}:=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$ and ${\cal V}_{\delta}:=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$.
-In addition, the functions $W_{U},W^{V}$ are locally bounded and
-$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$
+Condition 4.10 (${\cal H}^{\prime}_{\delta}(U,V)$).
+There is some $\delta\in]0,1/2[$ such that ${\cal U}_{\delta}$ and ${\cal V}_{\delta}$ are locally bounded with compact sub-level sets and we have
+$\lambda_{(1-2\delta)U}\left(1\right)\vee\nu_{(1-2\delta)V}\left(1\right)<\infty.$ (66)
+We have
+${\cal H}^{\prime}_{\delta}(U,V)\Longrightarrow{\cal H}_{\delta}(U,V).$ (67)
+To verify this claim, note that the regularity condition on $({\cal U}_{\delta},{\cal V}_{\delta})$ ensures that
+$(e^{{\cal U}_{\delta}},e^{{\cal V}_{\delta}})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
+When ${\cal U}_{\delta}$ is locally bounded with compact sub-level sets $C_{{\cal U}}(r):=\{{\cal U}_{\delta}\leq r\}$ for any $0<\delta<1/2$ we have
+$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $=$ $\lambda_{(1-2\delta)U}\left(e^{-{\cal U}_{\delta}}\right)$
+$\leq$ $\left(e^{-r}+e^{c_{\delta}(r)}\right)\lambda_{(1-2\delta)U}\left(1\right)\quad\mbox{\rm with}\quad c_{\delta}(r):=\sup_{x\in C_{{\cal U}}(r)}|{\cal U}_{\delta}(x)|.$
+In this context, we readily see that
+[…]
+This ends the proof of (67).
+Remark 4.11.
+Whenever $\|W\|<\infty$ we have $\|W_{U}\|\vee\|W^{V}\|<\infty$. In this case ${\cal H}_{\delta}(U,V)$ is satisfied as soon as there is some $\delta\in]0,1[$ such that
+$\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)<\infty.$ (69)
+Remark 4.12.
+Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions and we have $\delta U-W^{V}\geq{\cal U}_{\delta}^{-}$ and $\delta V-W_{U}\geq{\cal V}_{\delta}^{-}$, for some functions $({\cal U}_{\delta}^{-},{\cal V}_{\delta}^{-})$ with compact level sets.
+In this case, the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ are locally bounded with compact level sets.
+A Lyapunov stability theorem
 Theorem 4.13.
-If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$, where $U_{\star}=\inf_{\mathbb{X}}U(x)$ and $V_{\star}=\inf_{\mathbb{Y}}V(y)$.
+If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions
+$(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})\quad\mbox{with}\quad(U_{\star},V_{\star}):=(\inf_{\mathbb{X}}U,\inf_{\mathbb{Y}}V)\in(\mathbb{R}\times\mathbb{R}).$ (70)
 In addition, there exist some parameter $\rho\in[0,1[$ and some constant $c>0$ such that
@@
 A proof of Theorem 4.13 is provided in Appendix D.4.
+The above theorem combined with the estimates (62) ensures that
+$\sup_{n\geq n_{0}}\pi_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c=\epsilon~e^{-\delta V_{\star}}~\nu_{(1-\delta)V}(1)+c$
+$\sup_{n\geq n_{0}}\pi_{2n}(\psi)$ $\leq$ $\epsilon~\lambda_{U}(\varphi)+c=\epsilon~e^{-\delta U_{\star}}~\lambda_{(1-\delta)U}(1)+c$
+with the parameters $(\epsilon,c)$ as in (60).
+The rescaling of the potential functions (70) is only made to have $\varphi\wedge\psi\geq 1$.
+Combining Theorem 4.13 with Theorem 4.8 we obtain the following corollary.
+Corollary 4.14.
+Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
+[…]
+[…]
+with the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$.
+The r.h.s. term in (72) can be estimate with (62). Also note that
+[…]
+This yields the exponential decays.
+Corollary 4.15.
+Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
+[…]
+[…]
+with the functions $(\psi,\varphi)$ as in Corollary 4.14.
+4.5 Wasserstein exponential decays
+Assume that the Lyapunov functions $(\psi,\varphi)$ defined in (70) satisfy the estimates
+$e^{\delta U(x_{1})}+e^{\delta U(x_{2})}\geq\varrho_{\varphi}(x_{1},x_{2})\quad\mbox{\rm and}\quad e^{\delta V(y_{1})}+e^{\delta V(y_{2})}\geq\varrho_{\psi}(y_{1},y_{2})$ (74)
+for some semi-metric $\varrho_{\varphi},\varrho_{\psi}$. In this case, by (10) there exists $c_{u},c_{v}>0$ such that
+[…]
+with the Wasserstein semi-metrics $\mathbb{W}_{\varrho_{\varphi}},\mathbb{W}_{\varrho_{\psi}}$ defined in (8). This yields the following Wasserstein estimates.
+Corollary 4.16.
+Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+In addition, the Lyapunov functions $(e^{\delta U},e^{\delta V})$ satisfy (74) for some semi-metric $(\varrho_{\varphi},\varrho_{\psi})$.
+In this situation, there exists some $n_{0}\geq 0$ and $c_{u},c_{v}\geq 0$ such that for any $n\geq n_{0}$ we have
+[…]
+[…]
+In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
+Assume there exists some semi-metric $\varrho_{U},\varrho_{V}$ and some constants $c_{u},c_{v}\in\mathbb{R}$ such that
+$U(x_{1})+U(x_{2})$ $\geq$ $c_{u}+\varrho_{U}(x_{1},x_{2}),$
+$V(y_{1})+V(y_{2})$ $\geq$ $c_{v}+\varrho_{V}(y_{1},y_{2}).$ (76)
+In this case, by (35), for any $p\geq 1$ there exists some constant $c_{p}>0$ such that
+[…]
+with the $p$-Wasserstein semi-metrics $\mathbb{W}_{p,\varrho_{U}},\mathbb{W}_{p,\varrho_{V}}$ defined in (8). This yields the following Wasserstein estimates.
+Corollary 4.17.
+Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+In addition, the target potential functions $(U,V)$ satisfy (76) for some semi-metric $(\varrho_{U},\varrho_{V})$.
+In this situation, for any $p\geq 1$ there exists some $c_{p}\geq 0$ such that for any $n\geq n_{0}$ we have
+[…]
+[…]
+In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
 Remark 4.18.
@@
 $\left(\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\right)\vee\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right)<\infty.$
+4.6 Entropy exponential decays
 Let us discuss some fairly straightforward consequences of the estimate (71). First note that (71) ensures the convergence of the series
-$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ ::=:= $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
-$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ ::=:= $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (43)
+$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ $:=$ $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
+$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ $:=$ $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (78)
 where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (18) converge as $n\rightarrow\infty$ to a system of integral equations
@@
 $\lim_{n\to\infty}{\cal P}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~\lambda_{\mathbb{U}}(dx)~\nu_{\mathbb{V}}(dy).$ (81)
-In addition, we have
+Moreover, we have
 $\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
-$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right)$
+$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right).$
 The exponential estimate (71) also allows the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (80).
@@
 5 Some illustrations
-5.1 Strong mixing models
-As underlined in [45, 62] in the context of finite state space models, Sinkhorn iterations of the form (5) preserve the cross product Radon-Nikodym ratio in (24) and, hence, we have
-$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}),\quad\forall n\geq 0.$ (47)
-This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [14, 18, 33, 47].
-Recall that the Hilbert contraction coefficient of a Markov transition ${\cal Q}$ can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq.
-(26)), hence the exponential convergence of Sinkhorn iterations w.r.t.
-the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$.
-An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
-From (25), let us observe that
-$\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})\geq\jmath({\cal Q})^{2}.$
-Thus, for any given a convex function $\Phi$ satisfying (21), we can use (26) and (32) to show that
-$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
-while
-$\jmath({\cal Q})>0\quad\text{implies}\quad\left\{\begin{array}[]{l}\raisebox{2.0pt}{$\chi$}_{H}({\cal S}_{n})=\raisebox{2.0pt}{$\chi$}_{H}({\cal Q})<1,\\ \text{${\cal D}_{1,1}({\cal S}^{\circ})$ is satisfied with $\varepsilon(r)=\hbar({\cal Q})>0$ for any $r\geq 1$, and}\\ \sup_{n\geq 0}\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal S}_{n})\leq\raisebox{2.0pt}{$\chi$}({\cal S}):=\sup_{n\geq 0}\raisebox{2.0pt}{$\chi$}({\cal S}_{n})\leq 1-\hbar({\cal Q})<1.\end{array}\right.$
-We can also obtain the sandwich-type contraction inequalities described below.
-Theorem 5.1.
-For any convex function $\Phi$ satisfying (21) we have the contraction estimates
-$\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~{}D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~{}D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~{}D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~{}D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$ (48)
+5.1 Polynomial growth models
+Consider some finite dimensional normed vector space $\mathbb{X}=\mathbb{Y}$ equipped with measure $\lambda=\nu$ such that for any $p>0$ and $\tau>0$ we have
+$\int e^{-\tau\|x\|^{p}}~\lambda(dx)<\infty.$ (82)
+In this context, we have $\lambda_{U}(dx)=e^{-U(x)}~\lambda(dx)$ and $\nu_{V}(dx)=\lambda_{V}(dx)=e^{-V(x)}~\lambda(dx)$.
+Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions.
+In addition, there exists some parameters $c_{w},c_{u},c_{v}\in\mathbb{R}$ as well as $\varsigma_{u},\varsigma_{v},\tau_{u},\tau_{v},p_{u},p_{v}>0$ and $q_{u},q_{v}\geq 0$ such that
+$W(x,y)$ $\leq$ $c_{w}+\varsigma_{u}~\|x\|^{p_{u}}+\varsigma_{v}~\|y\|^{p_{v}},$
+$U(x)$ $\geq$ $c_{u}+\tau_{u}~\|x\|^{p_{u}+q_{u}}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\tau_{v}~\|x\|^{p_{v}+q_{v}}.$ (83)
+Proposition 5.1.
+Condition ${\cal H}^{\prime}_{\delta}(U,V)$ is satisfied for some $0<\delta<1/2$ whenever, for any parameter $\iota\in\{u,v\}$, one of the following assertions is satisfied
+$\left(q_{\iota}>0\right)\quad\mbox{or}\quad\left(q_{\iota}=0\quad\mbox{and}\quad\tau_{\iota}/2>\varsigma_{\iota}\right).$ (84)
+Proof.
+The exponential moment condition (82) ensures that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments
+$m^{p_{u}}_{u}:=\int\|x\|^{p_{u}}\lambda_{U}(dx)<\infty\quad\mbox{\rm and}\quad m^{p_{v}}_{v}:=\int\|x\|^{p_{u}}\lambda_{V}(dx)<\infty.$
+Also note that
+${\cal U}_{\delta}(x)=\delta U(x)-W^{V}(x)$ $\geq$ $(\delta c_{u}-c_{w}-\varsigma_{v}~m^{p_{v}})+\delta\tau_{u}~\|x\|^{p_{u}+q_{u}}-\varsigma_{u}~\|x\|^{p_{u}}.$
+This shows that ${\cal U}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
+$q_{u}>0\quad\mbox{\rm or}\quad\left(q_{u}=0\quad\mbox{\rm and}\quad\delta\tau_{u}>\varsigma_{u}\right).$
+In the same vein, condition (84) with $a=v$ ensures ${\cal V}_{\delta}$ has compact sub-level sets.
+For instance, for bi-Laplace transitions on $\mathbb{X}=\mathbb{R}=\mathbb{Y}$ we have
+$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{\varsigma}{2}~e^{-\varsigma|x-y|}~dy\quad\mbox{\rm with $\varsigma>0$}\\ \\ \Longrightarrow W(x,y)=c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma|x|+\varsigma|y|\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }\\ \\ \Longrightarrow|W(x,z_{1})-W(x,z_{2})|\vee|W(z_{1},y)-W(z_{2},y)|\leq\varsigma~|z_{1}-z_{2}|.\end{array}$
+For linear Gaussian models, the r.h.s.
+condition in (84) is satisfied as soon as the covariance of the linear Gaussian transition $\cal Q$} is defined by (2) is sufficiently large.
+In the same context of linear and Gaussian models, we show in Proposition 5.3 that this condition is in fact equivalent to condition ${\cal H}_{\delta}(U,V)$.
+By (83) there exists some constant $\overline{\tau}_{u}>0$ such that
+$U(x_{1})+U(x_{2})$ $\geq$ $\overline{c}_{u}+\overline{\tau}_{u}~\left(\|x_{1}\|+\|x_{2}\|\right)^{p_{u}+q_{u}}\quad\mbox{\rm with}\quad\overline{c}_{u}:=2c_{u}.$
+This yields the estimate
+$U(x_{1})+U(x_{2})\geq\overline{c}_{u}+\varrho_{U}(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho_{U}(x_{1},x_{2}):=\overline{\tau}_{u}~\|x_{1}-x_{2}\|^{p_{u}+q_{u}}.$
+Similarly, there exist some constants $\overline{c}_{u}\in\mathbb{R}$ and $\overline{\tau}_{v}>0$ such that
+$V(x_{1})+V(x_{2})\geq\overline{c}_{v}+\varrho_{V}(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho_{V}(x_{1},x_{2}):=\overline{\tau}_{v}~\|x_{1}-x_{2}\|^{p_{v}+q_{v}}.$
+We conclude that $(U,V)$ satisfy (76) with the semi-metric $(\varrho_{U},\varrho_{V})$ so that the Wasserstein estimates (77) hold.
+5.2 Heavy tailed transitions
+Consider the normed vector space $\mathbb{X}=\mathbb{Y}$ and the reference measure $\lambda=\nu$ discussed in Section 5.1.
+Assume some parameter $c_{w}\in\mathbb{R}$, $\tau_{p,w},\tau_{q,w}>0$ and $p_{w}\wedge q_{w}>1$ such that
+$W(x,y)\leq c_{w}+\log\left(1+\tau_{p,w}~\|x\|^{p_{w}}+\tau_{q,w}~\|y\|^{q_{w}}\right).$
+For instance, for Cauchy transitions on $\mathbb{X}=\mathbb{R}=\mathbb{Y}$ we have
+$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{\pi a}~\frac{1}{1+((x-y)/a)^{2}}~dy\quad\mbox{\rm with}\quad a>0\\ \\ \displaystyle\Longrightarrow W(x,y)=\log{(a\pi)}+\log{\left(1+((x-y)/a)^{2}\right)}\quad\mbox{\rm and}\quad|\partial_{x}W(x,y)|\vee|\partial_{y}W(x,y)|\leq 1/a\\ \\ \displaystyle\Longrightarrow c_{w}=\log{(a\pi)},\qquad\tau_{p,w}=2/a^{2}=\tau_{q,w}\quad\mbox{\rm and}\quad p_{w}=q_{w}=2.\end{array}$
+Exponential targets
+Consider some continuous target potential functions $U,V$ such that
+$U(x)\geq c_{u}+\tau_{u}~\|x\|^{p_{u}}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\tau_{v}~\|x\|^{q_{v}}\quad\mbox{\rm with $c_{u},c_{v}\in\mathbb{R}$ and $p_{u},q_{v}>0$}.$
+Also assume that $(W_{U},W^{V})$ are continuous functions.
+Following the proof of Proposition 5.1, the exponential moment condition (82) ensures that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\nu_{V}=\lambda_{V}$ have finite $p_{w}$ and $q_{w}$ absolute moments.
+By Jensen’s inequality, we also have
+${\cal U}_{\delta}(x)$ $\geq$ $\delta c_{u}+\delta\tau_{u}~\|x\|^{p_{u}}-c_{w}-\log\left(1+\tau_{p,w}~\|x\|^{p_{w}}+\tau_{q,w}~m_{v}^{q_{w}}\right)\stackrel{{\scriptstyle\|x\|\rightarrow\infty}}{{\longrightarrow}}\infty$
+In the same vein we check that
+$\lim_{\|x\|\rightarrow\infty}(\delta U(x)-W^{V}(x))=\infty=\lim_{\|x\|\rightarrow\infty}(\delta V(x)-W_{U}(x)).$ (85)
+We conclude that ${\cal H}^{\prime}_{\delta}(U,V)$ is met for any $0<\delta<1/2$. Also note that $(U,V)$ satisfy (76) with the semi-metrics
+$\varrho_{U}(x_{1},x_{2}):=\overline{\tau}_{u}~\|x_{1}-x_{2}\|^{p_{u}}\quad\mbox{\rm and}\quad\varrho_{V}(x_{1},x_{2}):=\overline{\tau}_{v}~\|x_{1}-x_{2}\|^{q_{v}}$
+for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (77) hold.
+Heavy tailed targets
+Consider continuous target potential functions $U,V$ such that
+$U(x)\geq c_{u}+\log{\left(1+\tau_{u}~\|x\|^{p_{u}}\right)}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\log{\left(1+\tau_{v}~\|x\|^{q_{v}}\right)}$
+with $c_{u},c_{v}\in\mathbb{R}$ and some parameters
+$p_{u}>1+2p_{w}\quad\mbox{\rm and}\quad q_{v}>1+2q_{w}.$
+Condition (68) is met as soon as
+$(1-2\delta)p_{u}>1\quad\mbox{\rm and}\quad(1-2\delta)q_{v}>1$
+On the other hand, the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{w}$ and $q_{w}$ absolute moments as soon as
+$p_{u}-p_{w}>1\quad\mbox{\rm and}\quad q_{v}-q_{w}>1.$
+Finally, (85) holds when
+$\delta p_{u}-p_{w}>0\quad\mbox{\rm and}\quad\delta q_{v}-q_{w}>0.$
+This shows that ${\cal H}_{\delta}(U,V)$ for any $0<\delta<1/2$ such that
+$\frac{p_{w}}{p_{u}}<\delta<\frac{1}{2}~\left(1-\frac{1}{p_{u}}\right)\quad\mbox{\rm and}\quad\frac{q_{w}}{q_{v}}<\delta<\frac{1}{2}~\left(1-\frac{1}{q_{v}}\right).$
+Also note that the exponential functions $(e^{\delta U},e^{\delta V})$ satisfy (74) with the semi-metrics
+$\varrho_{\varphi}(x_{1},x_{2}):=\overline{\tau}_{u}~\|x_{1}-x_{2}\|^{p_{u}}\quad\mbox{\rm and}\quad\varrho_{\psi}(x_{1},x_{2}):=\overline{\tau}_{v}~\|x_{1}-x_{2}\|^{q_{v}}$
+for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (75) hold.
+5.3 Some boundary states
+Consider a bounded Lipschitz domain $\mathbb{X}=\mathbb{Y}\subset\mathbb{R}^{d}$ equipped with the uniform measure $\lambda=\nu$.
+Denote by $\varrho_{\partial}(x,\partial\mathbb{X})$ the distance from $x\in\mathbb{X}$ to the boundary $\partial\mathbb{X}$.
+We assume that the target distribution $\lambda_{U}$ has a continuous potential function $U$ and satisfies the estimates
+$\lambda_{U}(dx)\leq c_{q,u}~\varrho_{\partial}(x,\partial\mathbb{X})^{q_{u}}~\lambda(dx)$
+for some parameters $c_{p,u},c_{q,u}>0$ and $q_{u}>0$.
+Similarly, $\nu_{V}=\lambda_{V}$ has a continuous potential function $V$ and satisfies the above estimates for some $c_{p,v},c_{q,v}>0$ and $q_{v}>0$.
+Finally, we assume that the integrated functions $(W^{V},W_{U})$ are continuous.
+When $W(x,y)$ is bounded, Sinkhorn transitions are strongly mixing, in the sense that $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ with the uniform contraction parameter $\raisebox{2.0pt}{$\chi$}({\cal S})$ defined in (49).
+The exponential decays of $\Phi$-entropies associated with these strong mixing models are discussed in Section 4.1.
+Note that for any $x\in\mathbb{X}$ we have
+$U(x)\geq-\log{c_{q,u}}+q_{u}~\log{\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})}}\Longleftrightarrow e^{\delta U(x)}\geq\frac{1}{c_{q,u}^{\delta}}~\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})^{\delta q_{u}}}.$
+By Remark 4.11 the above models satisfy ${\cal H}_{\delta}(U,V)$ for any $\delta\in]0,1[$.
+In addition, following the proof of (36) we check that the exponential function $e^{\delta U}$ satisfies (74) with the semi-metric
+$\varrho_{\varphi}(x_{1},x_{2})=\overline{c}_{u}~\left(\frac{1}{\varrho_{\partial}(x_{1},\partial\mathbb{X})}-\frac{1}{\varrho_{\partial}(x_{2},\partial\mathbb{X})}\right)^{\delta q_{u}}~1_{x_{1}\not=x_{2}}\quad\mbox{\rm for some parameter $\overline{c}_{u}>0$.}$
+Similarly, the exponential function $e^{\delta V}$ satisfies (74) with the metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(c_{u},q_{u})$ by $(c_{v},q_{v})$, so that the Wasserstein estimates (75) hold.
+When $W(x,y)=W(y,x)$ is symmetric and non necessarily bounded, assume we have an estimate of the form
+$W(x,y)\leq c_{w}+q_{w}~\log{\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})}}$
+for some $c_{w}\in\mathbb{R}$ and some parameters $0\leq q_{w}<(q_{u}\wedge q_{v})/2$. In this case condition ${\cal H}_{\delta}(U,V)$ holds for any $\frac{q_{w}}{q_{u}\wedge q_{v}}<\delta<1/2$.
+5.4 Beta marginal models
+Next we illustrate the boundary state models discussed in Section 5.3 with the open unit interval $\mathbb{X}=\mathbb{Y}=]0,1[$ with boundary $\partial\mathbb{X}=\{0,1\}$.
+The distance to the boundary is given explicitly by the function
+$x\in\mathbb{X}\mapsto\varrho_{\partial}(x,\partial\mathbb{X})=x~1_{]0,1/2[}+(1-x)~1_{[1/2,1[}(x)\quad\mbox{\rm and}\quad\lambda(dx)=1_{]0,1[}(x)dx.$
+Note that
+$\frac{1}{2}~\varrho_{\partial}(x,\partial\mathbb{X})\leq x(1-x)\leq\varrho_{\partial}(x,\partial\mathbb{X})$
+and consider the reference Markov transition $\cal Q$} defined by the semi-circle distribution (41). Let $(\lambda_{U},\lambda_{V})$ be some Beta marginal distributions given by
+$\lambda_{U}(dx)=\frac{1}{c_{u}}~x^{a_{u}}(1-x)^{b_{u}}~\lambda(dx)\quad\mbox{\rm and}\quad\lambda_{V}(dx)=\frac{1}{c_{v}}~x^{a_{v}}(1-x)^{b_{v}}~\lambda(dx)$
+for some normalizing constants $c_{u},c_{v}>0$ and some parameters $a_{u},a_{v},b_{u},b_{v}>1$. In this context, we have
+$\frac{1}{2^{a_{u}\vee b_{u}}c_{u}}~\varrho_{\partial}(x,\partial\mathbb{X})^{a_{u}\vee b_{u}}~\lambda(dx)\leq\lambda_{U}(dx)\leq\frac{1}{c_{u}}~\varrho_{\partial}(x,\partial\mathbb{X})^{a_{u}\wedge b_{u}}~\lambda(dx).$
+Also observe that
+$1\leq\frac{1}{1-(x-y)^{2}}$ $=$ $\frac{1}{(1-(x-y))(1+(x-y))}$
+$=$ $\frac{1}{y+(1-x)}~\frac{1}{x+(1-y)}\leq\frac{1}{x(1-x)}\wedge\frac{1}{y(1-y)},$
+which yields the estimates
+$W(x,y)$ $=$ $\log{q(x)}+\frac{1}{2}~\log{\left(\frac{1}{(1-(x-y))(1+(x-y))}\right)}$
+$\leq$ $\frac{1}{2}~\log\left(\frac{1}{x(1-x)}\right)\quad\mbox{\rm with}\quad\frac{1}{3}\leq q(x):=\int_{0}^{1}\sqrt{1-(x-y)^{2}}dy\leq 1.$
+Note that for any $(x,y)\in]0,1[^{2}$ we obtain
+$|\partial_{x}\sqrt{1-(x-y)^{2}}|=\frac{|y-x|}{\sqrt{((1-x)+y)((1-y)+x)}}\leq g(y):=\frac{1}{\sqrt{y(1-y)}}\Longrightarrow\lambda(g)<\infty$
+as well as
+$|\partial_{x}\log{(1-(x-y)^{2})}|\leq g(x)^{2}\wedge g(y)^{2}.$
+In this example, the functions $(U,V)$ and $(W_{U},W^{V})$ are differentiable functions. Condition (66) is clearly met for any $0<\delta<1/2$ and we have
+$\delta U(x)-W^{V}(x)\geq\delta\log{(c_{u})}+(\delta a_{u}-1/2)\log{\frac{1}{x}}+(\delta b_{u}-1/2)\log{\frac{1}{1-x}}.$
+This shows that ${\cal H}^{\prime}_{\delta}(U,V)$ is met for any $((a_{u}\wedge b_{u})\wedge(a_{v}\wedge b_{v}))^{-1}<2\delta<1$. We also have
+$U(x_{1})+U(x_{2})\geq\overline{c}_{u}:+\varrho_{U}(x_{1},x_{2})$
+with $\overline{c}_{u}:=2\log{(c_{u})}$ and the metric
+$\varrho_{U}(x_{1},x_{2}):=a_{u}~\left|\log{\frac{1}{x_{1}}}-\log{\frac{1}{x_{2}}}\right|+b_{u}\left|\log{\frac{1}{1-x_{1}}}-\log{\frac{1}{1-x_{2}}}\right|.$
+We conclude that $U$ satisfies (76). Similarly $V$ satisfies (76) with a metric $\varrho_{V}$ defined as $\varrho_{U}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$.
+Recall that $uv\geq(u+v)/2$ and $u^{p}v^{q}\geq(uv)^{p\wedge q}$ for any $u,v\geq 1$ and any parameters $p,q>0$ we see that
+$e^{\delta U(x)}$ $\geq$ $\frac{c_{u}^{\delta}}{2}~\left(\left(\frac{1}{x}\right)^{\delta a_{u}}+\left(\frac{1}{1-x}\right)^{\delta b_{u}}\right),$
+$e^{\delta U(x)}$ $\geq$ $c_{u}^{\delta}~\left(\frac{1}{x(1-x)}\right)^{\delta(a_{u}\wedge b_{u})}\geq\overline{c}_{u}^{\delta}~\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})^{\delta(a_{u}\wedge b_{u})}}\quad\mbox{\rm with}\quad\overline{c}_{u}:=2^{(a_{u}\wedge b_{u})}c_{u}.$
+Following the proof of (36) we check that the exponential function $e^{\delta U}$ satisfies (74) with
+$\varrho_{\varphi}(x_{1},x_{2})=\overline{c}_{u}~\left(\frac{1}{\varrho_{\partial}(x_{1},\partial\mathbb{X})}-\frac{1}{\varrho_{\partial}(x_{2},\partial\mathbb{X})}\right)^{\delta(a_{u}\wedge b_{u})}$
+as well as with the semi-metric
+$\varrho_{\varphi}(x_{1},x_{2})=\overline{c}_{u}~\left(\left|\frac{1}{x_{1}}-\frac{1}{x_{2}}\right|^{\delta a_{u}}+\left|\frac{1}{1-x_{1}}-\frac{1}{1-x_{2}}\right|^{\delta b_{u}}\right)$
+for some parameter $\overline{c}_{u}>0$.
+Similarly, the exponential function $e^{\delta V}$ satisfies (74) with the complete metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$, so that the Wasserstein estimates (75) hold.
+5.5 Weibull marginal models
+We equip the open half-line $\mathbb{X}=\mathbb{Y}=]0,\infty[$ with the Lebesgue measure $\lambda(dx)=\nu(dx)=1_{]0,\infty[}(x)dx$ and consider the Weibull marginal measure
+$\lambda_{U}(dx)=\frac{a_{u}+1}{b_{u}}~\left(\frac{x}{b_{u}}\right)^{a_{u}}e^{-(x/b_{u})^{a_{u}+1}}~\lambda(dx)$
+with potential
+$U(x)=c_{u}+(x/b_{u})^{a_{u}+1}+a_{u}~\log{\frac{1}{x}},\quad\mbox{\rm where}\quad c_{u}:=\log{\frac{b_{u}}{a_{u}+1}}+a_{u}\log{b_{u}}$
+and $a_{u},b_{u}>0$.
+Note that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments $m^{p_{u}}_{u}$ and $m^{p_{v}}_{v}$ for any $p_{u},p_{v}$.
+We define similarly $\lambda_{V}$ with some parameters $a_{v},b_{v}>0$.
+Consider a continuous function $W(x,y)$ satisfying (83) for some parameters $c_{w}\in\mathbb{R}$ as well as $p_{u},p_{v}>0$. In this context, we have
+$W^{V}(x)\leq c_{w}+\varsigma_{u}~x^{p_{u}}+\varsigma_{v}~m_{v}^{p_{v}}\quad\mbox{\rm with}\quad m^{p_{v}}_{v}:=\int_{0}^{\infty}~x^{p_{u}}\lambda_{V}(dx)<\infty.$
+This implies that
+$\delta U(x)-W^{V}(x)\geq\overline{c}_{u}+\delta a_{u}~\log{\frac{1}{x}}+\frac{\delta}{b_{u}^{a_{u}+1}}~x^{a_{u}+1}-\varsigma_{u}~x^{p_{u}}\quad\mbox{\rm with}\quad\overline{c}_{u}:=\delta c_{u}-c_{w}-\varsigma_{v}~m_{v}^{p_{v}}$
+Assume that $(W_{U},W^{V})$ are both continuous functions. In this situation, ${\cal U}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
+$a_{u}+1>p_{u}\quad\mbox{\rm or}\quad\left(p_{u}=a_{u}+1\quad\mbox{\rm and}\quad\varsigma_{u}~b_{u}^{a_{u}+1}<\delta<1/2\right).$
+The r.h.s. condition requires that $\varsigma_{u}<b_{u}^{-(a_{u}+1)}/2$. The analysis of these Weibull models now follows the same line of arguments as the ones we used for the Beta models.
+5.6 Exponential marginal models
+We equip the half-line $\mathbb{X}=\mathbb{Y}=[0,\infty[$ with the Lebesgue measure $\lambda(dx)=1_{[0,\infty[}(x)dx$ and consider the exponential marginal measures
+$\lambda_{U}(dx)=\tau_{u}~e^{-\tau_{u}x}~\lambda(dx)\quad\mbox{\rm and}\quad\lambda_{V}(dx)=\tau_{v}~e^{-\tau_{v}x}~\lambda(dx)\quad\mbox{\rm with}\quad\tau_{u}\wedge\tau_{v}>0\\$
+for which we have the potentials
+$U(x)=c_{u}+\tau_{u}x\quad\mbox{\rm and}\quad\quad V(x)=c_{v}+\tau_{v}x\quad\mbox{\rm with}\quad(c_{u},c_{v}):=(-\log{\tau_{u}},-\log{\tau_{u}}).$
+Note that the functions $x\in[0,\infty[\mapsto\tau_{\iota}x\in[0,\infty[$ with $\iota\in\{u,v\}$ have compact sub-level sets.
+Let ${\cal Q}(x,dy)$ be the exponential transition defined for any $x\in[0,\infty[$ by
 […]
-$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~{}\raisebox{2.0pt}{$\chi$}({\cal S})^{2n}\quad\text{for any $n\geq 0$.}$ (49)
-Proof.
-The estimates (48) are a direct consequence of (32) combined with the Markov transport formulae
-$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\cal S}_{2(n+1)},\pi_{2n+1}{\cal S}_{2(n+1)}),\quad\text{and}$
-$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
-which are themselves easily obtained from expressions (5) and (13). The proof of (49) follows exactly the same arguments as the proof of the inequality (42) in Theorem 4.3, hence it is skipped.
-The following result can be obtained from the series expansion given by Corollary 4.6.
-Corollary 5.2.
-Under the assumptions of Theorem 5.1, we have the entropy exponential decay estimates
-$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})\leq c~{}\raisebox{2.0pt}{$\chi$}({\cal S})^{2n},\quad\text{for any $n\geq 0$}.$
-Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [35].
-For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself.
-The choice ${\cal Q}(x,dy)~{}\propto~{}e^{-\delta\,\|y-b(x)\|}~{}dx$ yields
-$\jmath({\cal Q})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
-In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [7], Nash [64] and Varopoulos [78] on Gaussian estimates for heat kernels on manifolds.
-Nevertheless, the above condition is not met for linear Gaussian transitions.
+[…]
+which yields the cost function
+$W(x,y)\leq c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma~x+\varsigma~y,\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }.$
+In this context, condition ${\cal H}^{\prime}_{\delta}(U,V)$ holds for any $({\varsigma}/{\tau_{u}})\vee({\varsigma}/{\tau_{v}})<\delta<{1}/{2}$.
+5.7 Semi-compact models
+In many practical situations of interest, the target state spaces are different. Next we illustrate this situation with a semi-discrete model and a semi-compact models.
+Semi-discrete model
+Consider a finite or countable subset $\mathbb{X}\subset\mathbb{R}^{d}$ equipped with the counting measure $\lambda$ and the discrete topology.
+In this context, the potential function $U$ is a function
+$U~:~x\in\mathbb{X}\mapsto U(x)\in[0,\infty[\quad\mbox{\rm such that}\quad\sum_{x\in\mathbb{X}}e^{-U(x)}=1.$
+Let $\nu_{V}(dy)=e^{-V(y)}~dy$ be a Boltzmann-Gibbs probability measure on $\mathbb{R}^{d}$ and assume that $(U,V,W)$ are continuous functions satisfying the polynomial growth conditions (83).
+In this context, the sub-level sets of $U$ are finite, thus compact in $\mathbb{X}$.
+Following the proof of Proposition 5.1 these conditions ensure that ${\cal H}^{\prime}_{\delta}(U,V)$ holds for some $0<\delta<1/2$ as soon as condition (84) is satisfied.
+The analysis of heavy tailed transitions follows word-for-word the same lines of arguments as in Section 5.2, this it is skipped.
+Semi-compact model
+Consider the uniform measure $\lambda_{U}$ on a compact space $\mathbb{X}\subset\mathbb{Y}=\mathbb{R}^{d}$ associated with a null potential $U=0$, which yields $\lambda_{U}=\lambda$.
+Let $\nu_{V}$ be a centered Gaussian distribution with unit variance on $\mathbb{Y}$ and assume that the transition potential is chosen so that
+$W(x,y)\leq w(x)+a~\|y\|^{p}\Longrightarrow W_{U}(y)\leq\lambda(w)+b~\|y\|^{p}$
+for some continuous function $w(x)$ and some parameters $a\leq 2$ and $p\leq 1$. These models satisfy ${\cal H}^{\prime}_{\delta}(U,V)$ as soon as
+$(p=2~~\mbox{\rm and}~~b<\delta/2)\quad\mbox{\rm or}\quad(p<2)$ (86)
+We check this claim using the fact that
+[…]
 5.8 Linear-Gaussian models
-[…]
+Consider the state spaces $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, for some integer $d\geq 1$ and the Lebesgue measure $\lambda(dz)=\nu(dz)=dz$.
+Hereafter, points $x$ in the Euclidean space $\mathbb{X}=\mathbb{R}^{d}$ are represented by $d$-dimensional column vectors (or, equivalently, by $d\times 1$ matrices).
+When there is no possible confusion, we use the notation $\|\cdot\|$ for any equivalent matrix or vector norm.
 Let ${\cal G}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\cal S}^{+}_{d}\subset{\cal G}l_{d}$ the subset of positive definite matrices.
@@
 Consider the target probability measures $(\lambda_{U},\nu_{V})=(\nu_{m,\sigma},\nu_{\overline{m},\overline{\sigma}})$ for some given parameters
-$(m,\overline{m})\in(\mathbb{R}^{d}\times\mathbb{R}^{d})\quad\text{and}\quad(\sigma,\overline{\sigma})\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d}).$
+[…]
 Equivalently, the log-density functions $(U,V)$ are defined by
@@
 $W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$ (88)
-for some given $(\alpha,\beta,\tau)\in\left(\mathbb{R}^{d}\times{\cal G}l_{d}\times{\cal S}^{+}_{d}\right)$.
-In this context, we have $\jmath({\cal Q})=0$ and Hilbert projective techniques cannot be applied.
-However, a refined exponential stability analysis of the Sinkhorn algorithm for Gaussian models has been developed in [2].
-For instance, following Remark 5.4 in [2], there exists some $n_{0}\geq 1$ such that for any $n\geq n_{0}$ we have the estimate
+[…]
 $\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$ (89)
@@
 $\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$ (91)
-A proof is provided in Appendix A (on page A).
+A proof is provided in the Appendix A (on page A).
 The choice of the Lyapunov functions is not unique. For instance, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
-$\varphi(x)$ ::=:= $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
-$\psi(y)$ ::=:= $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (55)
+$\varphi(x)$ $:=$ $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
+$\psi(y)$ $:=$ $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (92)
 There also exists some $\delta_{0}>0$ such that ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
-$\varphi(x)$ ::=:= $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
-$\psi(y)$ ::=:= $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (56)
-A detailed proof of the above assertions is provided in Appendix A on page A.
+$\varphi(x)$ $:=$ $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
+$\psi(y)$ $:=$ $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (93)
+A detailed proof of the above assertions is provided in Appendix A (on page A).
 We end this section with an equivalent formulation of condition ${\cal H}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (87) and (88).
@@
 Let us now consider the mixture model (95) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (88) and the log-densities
-$-U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (59)
+$U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (96)
 where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters.
@@
 The Gaussian-kernel estimation model defined by
-$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~{}dx.$ (61)
+$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~dx$ (98)
 is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$).
@@
 Note that (98) coincides with (95) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$.
-For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [67, 69] as well as the books [16, 73, 79].
-The Gaussian mixture distribution (61) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [31] and references therein.
-Choosing $(\alpha,\beta)=(0,I)$ and $\tau=tI$; for some $t>0$ in (51), the Markov transition ${\cal Q}(x,dy)$ coincides with the the transition semigroup of the $d$-dimensional Brownian motion, namely
+For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [parzen, rosenblatt] as well as the books [chacon, simonov, wand].
+The Gaussian mixture distribution (98) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [doucet-bortoli] and references therein.
+Remark 5.6.
+Schrödinger bridges and the Sinkhorn algorithm are often used in practice to solve high dimensional optimal transport problems.
+In our context, choosing $\tau=tI$ and $(\alpha,\beta)=(0,I)$ in (88), the Schrödinger bridge in the limit $t\rightarrow 0$ of vanishing regularization solves the Monge-Kantorovich problem and converge towards the optimal $2$-Wasserstein coupling.
+In this case, ${\cal S}_{0}=Q$ coincides with the heat kernel on a time horizon $t$ (see also Remark 2.3 in [adm-24]), namely,
 ${\cal Q}(x,dy)=(2\pi t)^{-d/2}~\exp{\left(-\frac{1}{2t}~\|y-x\|^{2}\right)}~dy.$
@@
 $\frac{1}{t}~I<(\delta\wedge(1-\delta))~\sigma^{-1}.$ (99)
-[…]
+In many cases, the Markov transition ${\cal Q}(x,dy)$ of interest is generated by a stochastic diffusion between time 0 and a time horizon $t$ –see, e.g., [leonard, stromme2023sampling, kato2024large].
+In our context, if we choose $(\alpha,\beta)=(0,e^{tA})$ and $\tau=\tau_{t}:=\int_{0}^{t}e^{sA}e^{sA^{\prime}}ds$ for some Hurwitz matrix $A$ and some $t>0$, ${\cal Q}(x,dy)$ coincides with the transition semigroup of the Ornstein-Uhlenbeck diffusion (see Remark 2.3 in [adm-24]).
+In this case, the l.h.s.
+condition in (97) applied to (98) is met for any $t\geq t_{0}$ provided that
 $e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~\sigma^{-1}.$ (100)
@@
 $\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\cal S}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~\tau^{-1}\beta~\sigma^{1/2},$
-and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~{}\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [2]) that the recursions
+and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [adm-24]) that the recursions
 $\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~\upsilon_{2n}~\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~\upsilon_{2n+1}~\gamma^{\prime},$ (102)
@@
 with the random maps
-${\cal Z}^{\circ}_{2n}(y)$ ::=:= $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G$
-${\cal Z}^{\circ}_{2n+1}(x)$ ::=:= $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~{}G$
+${\cal Z}^{\circ}_{2n}(y)$ $:=$ $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G$
+${\cal Z}^{\circ}_{2n+1}(x)$ $:=$ $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~G$
 defined in terms of the parameters
 $\begin{array}[]{rclcrcl}\beta_{2n+1}^{\circ}&:=&\beta_{2n+1}~\beta_{2n},&&\tau_{2n+1}^{\circ}&:=&\tau_{2n+1}+\beta_{2n+1}~\tau_{2n}~\beta_{2n+1}^{\prime},\\ &&&&&&\\ \beta^{\circ}_{2n}&:=&\beta_{2n}~\beta_{2n-1},&&\text{and}~~\tau^{\circ}_{2n}&:=&\tau_{2n}+\beta_{2n}~\tau_{2n-1}~\beta_{2n}^{\prime}.\end{array}$
-By Lemma 4.9 in [2], we have
+By Lemma 4.9 in [adm-24], we have
 $0\leq\overline{\sigma}^{-1/2}~\beta^{\circ}_{2(n+1)}~\overline{\sigma}^{1/2}=I-\upsilon_{2(n+1)}\leq(I+\varpi)^{-1}$
@@
 $0\leq\sigma^{-1/2}~\beta_{2n+1}^{\circ}~\sigma^{1/2}=I-\upsilon_{2n+1}\leq(I+\overline{\varpi})^{-1}.$
-Moreover, Corollary 4.4 in [2] yields the uniform estimates
+Moreover, Corollary 4.4 in [adm-24] yields the uniform estimates
 $\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}\leq\sigma,~~\forall n\geq 1.$
@@
 then we have
-$\begin{array}[]{l}\displaystyle\frac{1}{\psi(y)}~{}\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)\\ \\ \leq\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~{}I)}}~{}\exp{\left(-\frac{\delta}{2}~{}\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~{}(y-\overline{m})^{\prime}~{}\overline{\sigma}^{-1}~{}(y-\overline{m})\right)}.\end{array}$
+$\frac{1}{\psi(y)}~\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)$ $\leq$ $\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~I)}}$$\times$
+$\times~\exp{\left(-\frac{\delta}{2}~\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~(y-\overline{m})^{\prime}~\overline{\sigma}^{-1}~(y-\overline{m})\right)}.$
 Choosing
@@
 Appendix B Hilbert metric contractions
-We recall the Hilbert contraction coefficient formula ([9, 11, 12, 50], see also [8, 22, 59])
-[…]
+We recall the Hilbert contraction coefficient formula ([bauer, birkhoff, birkhoff-2, hopf], see also [atar, cohen, legland])
+$\raisebox{2.0pt}{$\chi$}_{H}({\cal K})=\tanh{\left(\frac{1}{4}\log\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})\right)}.$ (110)
 Also note that
@@
 we conclude that
-[…]
+$\raisebox{2.0pt}{$\chi$}_{H}({\cal K})=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}}.$ (111)
 Proposition B.1.
@@
 We associate with some $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ the following condition.
-Condition (${\cal L}_{\varphi,\psi}({\cal K})$).
+Condition C.1 (${\cal L}_{\varphi,\psi}({\cal K})$).
 There exists some parameters $\epsilon\in]0,1[$ and $c<\infty$ such that
@@
 In addition, there exist some constant $r_{0}\geq 1$ and some function
-$\begin{array}[]{ccccc}\varepsilon&:&[r_{0},\infty[&\mapsto&]0,1]\\ &&r&\leadsto&\varepsilon(r)\\ \end{array}$
+[…]
 such that, for any $r\geq r_{0}$ and $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$,
@@
 When the functions are bounded, the geometric drift condition (112) is trivially met, but in this case (114) is a uniform contraction condition on the space $\mathbb{X}$.
-When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact condition (77) is met as soon as
+When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact and condition (114) is met as soon as we have
 ${\cal K}(x,dy)\geq q(x,y)~\nu(dy)$ (116)
@@
 Then, there exist constants $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (112) and on the function $\varepsilon(r)$ in (113) and such that
-[…]
-Theorem C.1 is a consequence of Lemma 2.3 in [36], see also Theorem 8.2.21 in [38] and Lemma 2.1 in [6].
-Several examples satisfying the Lyapunov condition (75) and the local minorization conditions (82) are discussed in [6, 38].
+$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\cal K})\leq\rho\quad\mbox{with}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (119)
+Theorem C.2 is a consequence of Lemma 2.3 in [horton-dp], see also Theorem 8.2.21 in [penev] and Lemma 2.1 in [dpa].
+Several examples satisfying the Lyapunov condition (112) and the local minorization conditions (119) are discussed in [dpa, penev].
 Proof of Theorem 4.8
+We check that if Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (53) and (54) also hold for any $n>n_{0}$.
 Assume that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$. From (60), for any $n>n_{0}$ we have the estimates
@@
 As a consequence, Theorem C.2 can be applied to show that there are parameters $\kappa\geq 0$ and $\rho\in[0,1[$ such that
-[…]
+$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{{\psi_{\kappa},\varphi_{\kappa}}}({\cal S}_{2n+1})\leq\rho.$
 Appendix D Schrödinger potential functions
-D.1 Some quantitative estimates
+D.1 Commutation formulae
+For any $n\geq 0$ we have the commutation formulae
+${\cal S}_{2n}\left(\frac{d\nu_{V}}{d\pi_{2n}}\right)=\frac{d\pi_{2n+1}}{d\lambda_{U}}\quad\mbox{and}\quad{\cal S}_{2(n+1)}\left(\frac{d\pi_{2n}}{d\nu_{V}}\right)=\frac{d\lambda_{U}}{d\pi_{2n+1}}.$ (123)
+In the same vein, for any $n\geq 1$ we have
+${\cal S}_{2n-1}\left(\frac{d\lambda_{U}}{d\pi_{2n-1}}\right)=\frac{d\pi_{2n}}{d\nu_{V}}\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)=\frac{d\nu_{V}}{d\pi_{2n}}.$ (124)
+A proof for (123) and (124) is provided in Appendix E (on page E).
+Using the commutation properties (123) and (124) for any $n\geq p$ we readily check that
+$\frac{d\pi_{2n}}{d\nu_{V}}$ $=$ ${\cal S}^{\circ}_{2n}\left(\frac{d\pi_{2(n-1)}}{d\nu_{V}}\right)={\cal S}^{\circ}_{2n,2p}\left(\frac{d\pi_{2p}}{d\nu_{V}}\right)\quad\text{and}$
+$\frac{d\pi_{2n+1}}{d\lambda_{U}}$ $=$ ${\cal S}_{2n+1}^{\circ}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)={\cal S}^{\circ}_{2n+1,2p-1}\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right),$ (125)
+with the backward semigroups
+${\cal S}^{\circ}_{2n,2p}:={\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1,2p-1}:={\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1}.$
+By (123) and (124) we also obtain
+$\frac{d\lambda_{U}}{d\pi_{2n+1}}$ $=$ $\left({\cal S}_{2(n+1)}\,{\cal S}^{\circ}_{2n,2p}\,{\cal S}_{2p-1}\right)\left(\frac{d\lambda_{U}}{d\pi_{2p-1}}\right)\quad\text{and}$
+$\frac{d\nu_{V}}{d\pi_{2n}}$ $=$ $\left({\cal S}_{2n+1}\,{\cal S}^{\circ}_{2n-1,2p-1}\,{\cal S}_{2(p-1)}\right)\left(\frac{d\nu_{V}}{d\pi_{2(p-1)}}\right).$ (126)
+D.2 Some quantitative estimates
 Proposition D.1.
@@
 Proof.
-Observe that
-$\text{if}\quad\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty\quad\text{then}\quad e^{\delta U}=\frac{1}{e^{-\delta U}}\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}>0$
-and this yields
+Since $U$ is locally bounded with compact sub-level sets for any $\delta\geq 0$ and any sufficiently large $r>0$ we have
+$e^{-\delta U}=e^{-\delta U}~1_{\{U>r\}}+e^{-\delta U}~1_{\{U\leq r\}}\leq\varsigma^{U}_{\delta}(r)\quad\mbox{\rm and}\quad\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}\geq 1/\varsigma^{U}_{\delta}(r)$
+with the parameters
+$\varsigma^{U}_{\delta}(r):=e^{-\delta r}+e^{\delta c_{U}(r)}\quad\mbox{\rm with}\quad c_{U}(r):=\sup_{x\in C_{U}(r)}|U(x)|\quad\mbox{\rm and}\quad C_{U}(r)=\{x~:~U(x)\leq r\}$
+This yields
 $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)=\lambda_{U}\left(e^{W^{V}}~e^{\delta U}\right)\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~\lambda_{U}\left(e^{W^{V}}\right).$
@@
 Using (132) we see that, for any $n\geq 1$,
-$-{\cal W}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\cal W}_{V,U}+W^{V}-\nu_{V}(V)$
+${\cal W}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\cal W}_{V,U}+W^{V}-\nu_{V}(V)$
 as well as
-$-{\cal W}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\cal W}^{U,V}+W_{U}-\nu_{V}(V).$
+${\cal W}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\cal W}^{U,V}+W_{U}-\nu_{V}(V).$
 Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (140) and (141) in the proof of the commutation formulae (123), we readily obtain the estimates
@@
 The functions $({\cal W}_{V,U},{\cal W}^{U,V})$ defined in (129) can be computed explicitly for the the Gaussian model defined in (87) and (88).
-We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [2]) we have
+We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [adm-24]) we have
 ${\cal W}_{V,U}(x)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~(\beta_{\tau}~(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
@@
 $\sup_{n\geq 1}~\left(\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\right)<\infty\quad\mbox{and}\quad\sup_{n\geq 1}~\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)<\infty.$
-D.2 Lyapunov inequalities
+D.3 Lyapunov inequalities
 The result below is a fairly direct consequence of the estimates presented in Lemma D.4.
@@
 Assume that Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
-${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~{}e^{-{\cal U}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\varphi\right)/\psi\leq c_{\delta}~{}e^{-{\cal V}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{Y})$ (96)
+${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~e^{-{\cal U}_{\delta}}~\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n-1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}~\in{\cal B}_{0}(\mathbb{Y})$ (137)
 for every $n\geq 1$ and some finite constant $c_{\delta}<\infty$.
@@
 $c_{\delta}:=e^{-\nu_{V}(V)}~\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\left(\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\right).$
-D.3 Proof of Theorem 4.3
-Following exactly the same argument leading to (81) in Appendix C, the Lyapunov inequalities (96) in Proposition D.9 ensure that the drift conditions stated in (39) hold with with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
-For $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
+By (130) we have
+${\cal S}_{1}(y,dx)$ $=$ $\frac{{\cal R}(y,dx)~e^{-U(x)}}{{\cal R}(e^{-U})(y)}\leq e^{W_{U}(y)}~{\cal R}(y,dx)~e^{-U(x)}$
+This implies that
+$e^{-\delta V(y)}{\cal S}_{1}\left(e^{\delta U}\right)(y)\leq e^{-(\delta V(y)-W_{U}(y))}~\|e^{-W}\|~\lambda_{(1-\delta)U}\left(1\right)$
+Equivalently, some finite constant $c_{\delta}<\infty$ we have
+${\cal S}_{1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}$~
+This ends the proof of the proposition.
+D.4 Proof of Theorem 4.13
+The Lyapunov inequalities (137) in Proposition D.9 ensure that the drift conditions stated in (60) hold with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
+Recall from the proof of Proposition D.5 that $e^{-\delta U}$ and $e^{-\delta V}$ are uniformly bounded for any $\delta\geq 0$.
+Since $\mathbb{X}$ and $\mathbb{Y}$ can be exhausted respectively by the compact sub-levels sets $C_{U}(r)$ and $C_{V}(r)$ of $U$ and $V$, there exists some $r_{1}>1$ such that $\lambda(C_{U}(r))\wedge\nu(C_{V}(r))>0$ for $r=r_{1}$ and thus for any $r\geq r_{1}$.
+This ensure there exists some $r_{0}>0$ such that for any $r\geq r_{0}$ we have
+$\lambda(C_{\varphi}(r))\wedge\nu(C_{\psi}(r))>0.$
+Following the proof of Proposition D.5, for $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
 $e^{{\cal W}_{V,U}(x)}={\cal Q}_{V}(e^{W_{U}})(x)$ $=$ $\int~e^{W_{U}(y)-W(x,y)}~\nu_{V}(dy)$
-$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\int~{}e^{W_{U}(y)}~{}\nu_{V}(dy)\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
+$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{V}(e^{W_{U}})$
+$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\|e^{-\delta V}\|_{\mathbb{L}_{\infty}(\nu)}~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
 and we conclude that
@@
 $e^{{\cal W}^{U,V}(y)}={\cal R}_{U}(e^{W^{V}})(y)$ $=$ $\int~e^{-W(x,y)+W^{V}(x)}~\lambda_{U}(dx)$
-$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{U}(e^{W^{V}})\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$
+$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{U}(e^{W^{V}})$
+$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$
 which yields
 $w^{U,V}(r):=\sup_{C_{\psi}(r)}{\cal W}^{U,V}<\infty.$
-On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ we arrive at
+On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ and $r\geq r_{0}$ we arrive at
 ${\cal S}_{2n}(x,dy)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~\nu_{V}(dy)~1_{C_{\psi}(r)}(y)$
@@
 $w(r):=\sup_{C_{\varphi}(r)\times C_{\psi}(r)}W+w_{V,U}(r)+w^{U,V}(r)-\nu_{V}(V).$
-Note that for any $r>1$ we have
+Note that for any $r\geq r_{0}$ we have
 $\nu_{V}\left(C_{\psi}(r)\right)$ $=$ $e^{-V_{\star}}~\nu\left(e^{-(V-V_{\star})}~1_{(V-V_{\star})\leq\delta^{-1}\log{r}}\right)$
 $\geq$ $w_{0}(r):=e^{-V_{\star}}~e^{-(\delta^{-1}\log{r})}~\nu\left(C_{\psi}(r)\right)>0$
-and this yields, for any $x\in C_{\varphi}(r)$, the estimate
+and this yields, for any $x\in C_{\varphi}(r)$ and $r\geq r_{0}$, the estimate
 ${\cal S}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~e^{-w(r)}.$
@@
 which implies the inequality
-$\left|\frac{d\pi_{2n}}{d\nu_{V}}(y)-1\right|\leq\|\frac{d\pi_{2p}}{d\nu_{V}}\|_{\psi}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}.$
+[…]
 On the other hand, combining Proposition D.7 with the norm equivalence (34) and Theorem 4.8, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
-$~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}$$\leq$ $c_{\kappa,1}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi_{\kappa}}$
+[…]
 $\leq$ $c_{\kappa,1}~\rho^{2(n-p)}~|\delta_{y}-\nu_{V}|_{\psi_{\kappa}}\leq c_{\kappa,2}~\rho^{2(n-p)}~(\psi(y)+\nu_{V}(\psi)).$
@@
 $\leq$ $\left(f(x)-1\right)~1_{f(x)\geq 1}+\left(\frac{1}{f(x)}-1\right)~1_{{1}/{f(x)}\geq 1},$
-where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $01+u>$ 0. This implies that
+where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $1+u>0$. This implies that
 $\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}$
@@
 which implies that
-$-U=\log{\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}}\leq-\inf_{i\in E_{0}}U_{i}$
+$U=\log{\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}}}\leq-\inf_{i\in E_{0}}U_{i}$
 and, hence,
@@
 $=$ $\bigcup_{(i,j)\in(E_{0}\times E_{1})}\left\{x~:~(\delta~U_{i}(x)-~W^{V_{j}}(x))\leq r\right\}$
-and this ensures that ${\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$. In the same vein, one can also prove that ${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$.
+and this ensures that ${\cal U}_{\delta}=\delta U-W^{V}$ have compact sub-level sets. In the same vein, one can also prove that ${\cal V}_{\delta}=\delta V-W_{U}$ have compact sub-level sets.
 On the other hand, using the fact that $(a+b)^{1-\delta}\leq a^{1-\delta}+b^{1-\delta}$ for any $a,b\geq 0$, we have

### v3

@@
 On the contraction properties of Sinkhorn semigroups
-Abstract
-We develop a novel semigroup stability analysis based on Lyapunov techniques and contraction coefficients to prove exponential convergence of Sinkhorn equations on weighted Banach spaces.
-This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences and Kantorovich-type criteria, including the relative entropy, squared Hellinger integrals, $\alpha$-divergences as well as weighted total variation norms and Wasserstein distances.
-To the best of our knowledge, these contraction inequalities are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm.
-We also provide Lyapunov contractions principles under minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of Sinkhorn semigroups.
-We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models.
-Last but not least, our approach also allows to consider statistical finite mixture of the above models, including kernel-type density estimators of complex data distributions arising in generative modeling.
-Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, weighted total variation norms, Lyapunov functions, Kantorovich and entropy criteria, Wasserstein semi-distances.
-Mathematics Subject Classification: 37M25, 49Q22, 47H09, 60J20; secondary: 37M25, 60J05, 94A17.
-1 Introduction
-1.1 Entropic optimal transport
-Consider locally compact Polish spaces $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ endowed with some locally bounded positive measures $\lambda$ and $\nu$.
-Consider some probability measures of the form
-$\lambda_{U}(dx):=e^{-U(x)}~\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~\nu(dy)$ (1)
-for some normalized potential functions $U:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$.
-Let $\Pi_{U,V}$ be the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{Y})$, with prescribed first and second coordinate marginals $(\lambda_{U},\nu_{V})$.
-Denote $\cal Q$} a Markov transition from $\mathbb{X}$ into $\mathbb{Y}$ of the form
-${\cal Q}(x,dy)=e^{-W(x,y)}~\nu(dy)\quad\mbox{\rm and set}\quad{\cal R}(y,dx):=e^{-W(x,y)}~\lambda(dx).$ (2)
-In the latter expression, the transition potential $W(x,y)$ stands for some real valued function on $(\mathbb{X}\times\mathbb{Y})$. Consider the reference probability measure constructed as
-${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{\cal Q}(x,dy).$ (3)
-The entropic transport problem (or Schrödinger bridge problem) with respect to (w.r.t.) the reference measure ${\cal P}_{0}$ is formulated as
-$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0}),\quad\mbox{\rm with the relative entropy}\quad\text{Ent}({\cal P}~|~{\cal P}_{0}):=\int\log\frac{d{\cal P}}{d{\cal P}_{0}}~d{\cal P}.$ (4)
-It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})<\infty$.
-This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (4) (cf.
+Abstract.
+We develop a novel stability theory for Sinkhorn semigroups based on Lyapunov techniques and quantitative contraction coefficients, and establish exponential convergence of Sinkhorn iterations on weighted Banach spaces.
+This operator-theoretic framework yields explicit exponential decay rates of Sinkhorn iterates toward Schrödinger bridges with respect to a broad class of $\phi$-divergences and Kantorovich-type distances, including relative entropy, squared Hellinger integrals, $\alpha$-divergences, weighted total variation norms, and Wasserstein distances.
+To the best of our knowledge, these results provide the first systematic contraction inequalities of this kind for entropic transport and the Sinkhorn algorithm.
+We further introduce Lyapunov contraction principles under minimal regularity assumptions, leading to quantitative exponential stability estimates for a large family of Sinkhorn semigroups.
+The framework applies to models with polynomially growing potentials and heavy-tailed marginals on general normed spaces, as well as to more structured boundary state-space models, including semicircle transitions and Beta, Weibull, and exponential marginals, together with semi-compact settings.
+Finally, our approach extends naturally to statistical finite mixtures of such models, including kernel-based density estimators arising in modern generative modeling.
+Key words and phrases:
+Entropic optimal transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, weighted total variation norms, Lyapunov functions, Kantorovich and entropy criteria, Wasserstein semi-distances.
+2020 Mathematics Subject Classification:
+Primary 37M25, 49Q22, 47H09, 60J20; secondary: 60J05, 94A17.
+1. Introduction
+1.1. Entropic optimal transport
+Let $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ be locally compact Polish spaces equipped with locally bounded positive measures $\lambda$ and $\nu$. We consider probability measures of the form
+(1.1) $\lambda_{U}(dx):=e^{-U(x)}~\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~\nu(dy)$
+where $U:\mathbb{X}\to\mathbb{R}$ and $V:\mathbb{Y}\to\mathbb{R}$ are measurable potential functions chosen so that $\lambda_{U}$ and $\nu_{V}$ are normalized.
+Let $\Pi_{U,V}$ be the set of probability measures $\mathcal{P}$} on $(\mathbb{X}\times\mathbb{Y})$, with prescribed marginals $(\lambda_{U},\nu_{V})$.
+We introduce a reference Markov kernel $\mathcal{Q}$} from $\mathbb{X}$ to $\mathbb{Y}$ of the form
+(1.2) ${\mathcal{Q}}(x,dy)=e^{-W(x,y)}~\nu(dy)\quad\mbox{\rm and set}\quad{\mathcal{R}}(y,dx):=e^{-W(x,y)}~\lambda(dx).$
+where $W:\mathbb{X}\times\mathbb{Y}\to\mathbb{R}$ is a measurable transition potential. The associated reference probability measure on $\mathbb{X}\times\mathbb{Y}$ is defined by
+(1.3) ${\mathcal{P}}_{0}(d(x,y)):=(\lambda_{U}\times{\mathcal{Q}})(d(x,y)):=\lambda_{U}(dx)~{\mathcal{Q}}(x,dy).$
+The entropic optimal transport problem, also known as the Schrödinger bridge problem, with respect to the reference measure ${\mathcal{P}}_{0}$ consists in finding
+(1.4) $\mathbb{P}=\operatorname*{arg\,min}_{{\mathcal{P}}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0}),\quad\mbox{\rm with the relative entropy}\quad\text{Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0}):=\int\log\frac{d{\mathcal{P}}}{d{\mathcal{P}}_{0}}~d{\mathcal{P}},$
+where $d{\mathcal{P}}/d{\mathcal{P}}_{0}$ denotes the Radon-Nikodym derivative of $\mathcal{P}$} with respect to ${\mathcal{P}}_{0}$.
+We implicitly assume that there exists at least one ${\mathcal{P}}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\mathcal{P}}|{\mathcal{P}}_{0})<\infty$.
+This condition ensures the existence of a solution $\mathbb{P}$ to (1.4) (cf.
 the seminal article by Csiszár [csiszar-2], as well as Section 6 in the Lecture Notes by Nutz [nutz], and also the survey article by Léonard [leonard] and references therein).
-Let us note that the product distribution
+Note that the product distribution
 $(\lambda_{U}\otimes\nu_{V})(d(x,y)):=\lambda_{U}(dx)~\nu_{V}(dy)$
-trivially belongs to $\Pi_{U,V}$. In addition, choosing ${\cal P}=(\lambda_{U}\times{\cal K})\,\in\,\Pi_{U,V}$ for some Markov transition ${\cal K}(x,dy)$ we have
-$\frac{d{\cal P}}{d{\cal P}_{0}}(x,y)=\frac{d(\lambda_{U}\otimes\nu_{V})}{d{\cal P}_{0}}(x,y)\times\frac{d{\cal P}}{d(\lambda_{U}\otimes\nu_{V})}(x,y)$
-which yields
-$\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=\int~W(x,y)~{\cal P}(d(x,y))+\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})-\int V(y)\nu_{V}(dy).$
-Choosing a potential transition of the form $W(x,y)={\cal W}(x,y)/t$, for some parameter $t>0$, problem (4) is equivalent to the optimal entropic transport problem
-$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\left(\int~{\cal W}(x,y)~{\cal P}(d(x,y))+t~\mbox{\rm Ent}({\cal P}~|~\lambda_{U}\otimes\nu_{V})\right).$ (5)
-In the optimal transport literature, function $\cal W$} is sometimes called the cost function and the parameter $t>0$ is interpreted as a regularization parameter.
-In this context, the non-regularized problem corresponding to $t=0$ coincides with the Kantorovich optimal transport problem with cost function $\cal W$}.
-For a more detailed discussion on this subject we refer to [dobrushin2006perturbation, rachev1985monge, villani2008optimal] and the numerous references therein.
-1.2 Background
-The Schrödinger bridge problem (4) is to find a random evolution with minimal relative entropy w.r.t.
-some reference physical model and with prescribed distributions of the initial and final time random states.
-This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [leonard, nutz, peyre] for recent surveys with numerous references on this subject.
-Optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [arjovsky, doucet-bortoli, kolouri, peyre, rioux2023semi, rioux2024entropic], statistical barycenter problems [agueh, andersen, bigot, cramer-2, cuturi-w, loubes], economy [bojitov], computer vision [dominitz, solomon], control theory [chen, chen-2], and many others.
-Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure.
-For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3] as well as the more recent studies [chen, cuturi, leonard-2].
-Sinkhorn bridges are based on nonlinear integral transformations over the whole state space (cf.
-Section 2) and exact finite-dimensional solutions cannot be computed for general models.
-However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3].
-For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [adm-24, cramer-2, cramer-3, janadi].
-Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [borwein, fienberg, sinkhorn-2, sinkhorn-3, soules] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [chen, deligiannidis, franklin, marino].
-See [greco2023non] for a refined analysis of Sinkhorn iterates and their gradient on the $d$-dimensional torus as well as the more recent article [chizat2025sharper] on compact spaces and bounded cost functions.
-As remarked in Section 3 of [cohen], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t.
-the Hilbert metric, see for instance [borwein, chen, deligiannidis, franklin] and Section 4.1 in the present article.
-Contraction estimates w.r.t.
-the Hilbert metric yield exponential decays w.r.t.
-$\phi$-divergences, including the total variation norm, the relative entropy and Wasserstein distances of any order.
-Nevertheless they can rarely be used to derive contraction estimates with respect to some $\phi$-divergence criteria.
-At this point, we underline that even on totally bounded state spaces such as the open unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ the cost function $W$ as well as the potential functions $(U,V)$ may be unbounded and the reference Markov transition may have a unit Birkhoff contraction coefficient (i.e.
-equal to $1$), see for instance Section 3.2.1 and the series of illustrations discussed in Section 5.3 and Section 5.4.
-The landmark example of quadratic costs on Euclidian spaces is another class of unbounded costs functions.
-In this context, the recent article [adm-24] also presents a refined stability analysis of Sinkhorn bridges for Gaussian models based on matrix Riccati difference equations.
-It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [peyre] as well as the more recent articles [adm-24, chiarini, durmus] and the references therein.
+trivially belongs to $\Pi_{U,V}$. Moreover, if ${\mathcal{P}}=(\lambda_{U}\times{\mathcal{K}})\,\in\,\Pi_{U,V}$ for some Markov kernel ${\mathcal{K}}(x,dy)$, then
+$\frac{d{\mathcal{P}}}{d{\mathcal{P}}_{0}}(x,y)=\frac{d(\lambda_{U}\otimes\nu_{V})}{d{\mathcal{P}}_{0}}(x,y)\times\frac{d{\mathcal{P}}}{d(\lambda_{U}\otimes\nu_{V})}(x,y)$
+which yields the decomposition
+$\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0})=\int~W(x,y)~{\mathcal{P}}(d(x,y))+\mbox{\rm Ent}({\mathcal{P}}~|~\lambda_{U}\otimes\nu_{V})-\int V(y)\nu_{V}(dy).$
+Choosing a transition potential of the form $W(x,y)={\mathcal{W}}(x,y)/t$, for some parameter $t>0$, problem (1.4) becomes equivalent to the regularized optimal transport problem
+(1.5) $\mathbb{P}=\operatorname*{arg\,min}_{{\mathcal{P}}\,\in\,\Pi_{U,V}}\left(\int~{\mathcal{W}}(x,y)~{\mathcal{P}}(d(x,y))+t~\mbox{\rm Ent}({\mathcal{P}}~|~\lambda_{U}\otimes\nu_{V})\right).$
+In the optimal transport literature, $\mathcal{W}$} is referred to as the cost function and $t>0$ as the entropic regularization parameter.
+The limiting case $t=0$ formally corresponds to the classical Kantorovich optimal transport problem with cost $\mathcal{W}$}.
+For further background and developments, we refer to [dobrushin2006perturbation, rachev1985monge, villani2008optimal] and the references therein.
+1.2. Background
+The Schrödinger bridge problem (1.4) consists in constructing a random evolution that minimizes the relative entropy with respect to a given reference model while matching prescribed initial and final marginal distributions.
+This convex optimization problem can be interpreted as an entropic regularization of the Monge–Kantorovich optimal transport problem; we refer to [leonard, nutz, peyre] for recent surveys and comprehensive bibliographies.
+Entropic optimal transport and Schrödinger bridges have become central tools in a wide range of applications, including generative modeling and machine learning [arjovsky, doucet-bortoli, kolouri, peyre, rioux2023semi, rioux2024entropic], statistical barycenters [agueh, andersen, bigot, cramer-2, cuturi-w, loubes], economics [bojitov], computer vision [dominitz, solomon], and control theory [chen, chen-2].
+Under mild integrability conditions, the Schrödinger bridge problem admits a unique solution that can be characterized as the fixed point of the Sinkhorn equations, also known as the iterative proportional fitting procedure.
+We refer to the pioneering works [deming, fortet, kruithof, sinkhorn-2, sinkhorn-3] and to more recent developments in [chen, cuturi, leonard-2] for historical background and modern perspectives.
+Sinkhorn bridges involve nonlinear integral operators acting on infinite-dimensional function spaces, and explicit solutions are generally unavailable.
+In finite state spaces, the Sinkhorn algorithm reduces to alternating row and column normalizations of matrices.
+For Gaussian models, however, Sinkhorn equations can be solved analytically using sequential Bayesian conjugacy arguments [adm-24, cramer-2, cramer-3, janadi].
 Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [alts-2017, chak-2018, dvu-2017].
@@
 Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [promit-2022] under some exponential moments estimates on Schrödinger potentials.
-There are very few and recent articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of three significant contributions [chiarini, durmus, eckstein2025hilbert].
-The first exponential stability analysis of unbounded models and non necessarily compactly supported marginals is presented in the article [durmus].
-In this study, the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic symmetric cost functions associated with a sufficiently large regularization parameter.
-In our framework this particular class of cost functions corresponds to symmetric Gaussian reference transitions.
-The approach taken in [durmus] relies on well known links between gradient and Hessian of Sinkhorn and Schrödinger potentials along Hamilton-Jacobi-Bellman equations with conditional expectations and covariances [chewi2023entropic, pooladian2021entropic], see also [del2025entropic].
-The main conditions in [durmus] are expressed in terms of the integrated convexity profiles of marginal potentials.
-Extending ideas from [conforti2024weak] the proofs are based on the propagation of these convexity profiles based on diffusion reflection techniques introduced in [eberle2016reflection].
-The exponential entropic decays presented in [durmus] were further developed and recently refined in [chiarini] to apply to all values of the regularization parameter and some classes of perturbation of log-concave models.
-The article [chiarini] has opened new perspectives in the convergence analysis for entropically regularized transport problems with quadratic symmetric costs.
-In the context of distance-like cost functions, one of the basic tool is a remarkable stability theorem, Theorem 1.1 in [chiarini], that allows to control the entropy of Schrödinger bridges by the entropy of their marginal distributions.
-However this stability approach requires to estimate the semi-concavity properties of generally unknown dual Schrödinger bridge transitions and it only applies to entropically regularized transport problems with symmetric quadratic-type costs.
-Thus, this framework excludes the important cases when the reference transition is associated with linear Gaussian transitions arising in Ornstein-Uhlenbeck-type diffusion generative models and denoising diffusions.
-In this connection, the more recent article [del2025stability] complements the stability approach in  [chiarini] with a novel entropic continuity theorem that applies to more general classes of non necessarily symmetric transition potentials and log-concave type target marginals, including strongly convex at infinity potentials.
-The article [eckstein2025hilbert] is also based on a novel version of Hilbert’s projective metric for spaces of integrable functions of bounded growth.
+Most existing results on the exponential convergence of Sinkhorn iterations concern finite or compact state spaces, or bounded cost functions, typically relying on Hilbert’s projective metric and related techniques [borwein, fienberg, sinkhorn-2, sinkhorn-3, soules, chen, deligiannidis, franklin, marino].
+See also [greco2023non] for refined analyses on the torus and [chizat2025sharper] for recent advances in compact settings.
+As emphasized in [cohen], contraction in the Hilbert metric is a very strong property, as this metric dominates all $\phi$-divergences.
+While Hilbert metric contraction implies exponential decay for a broad class of divergences (including total variation, relative entropy, and Wasserstein distances), it is generally difficult to obtain sharp contraction estimates for specific divergence criteria.
+Moreover, even on totally bounded spaces such as $\mathbb{X}=]0,1[=\mathbb{Y}$, the cost function and Schrödinger potentials may be unbounded, and the associated Markov operators may have unit Birkhoff contraction coefficient; see Sections 3.2.1, 5.3, and 5.4.
+Quadratic costs on Euclidean spaces provide another important class of unbounded models.
+In this setting, [adm-24] develops a refined stability analysis for Gaussian Schrödinger bridges based on matrix Riccati equations.
+Beyond Gaussian models, only a few recent works address exponential convergence of Sinkhorn iterations for non-compact spaces and unbounded costs, notably [durmus, chiarini, eckstein2025hilbert].
+The article [durmus] establishes quantitative exponential convergence for symmetric quadratic costs and marginal distributions with asymptotically positive log-concavity, using diffusion reflection couplings and convexity propagation techniques [conforti2024weak, eberle2016reflection].
+These results were further refined in [chiarini], which applies to all regularization parameters and certain perturbations of log-concave models.
+A central tool in [chiarini] is a stability theorem controlling the entropy of Schrödinger bridges by the entropy of their marginals, under uniform semi-concavity assumptions on unknown dual Schrödinger potentials.
+This approach is also limited to symmetric quadratic-type costs and does not cover important classes of nonsymmetric reference transitions, such as those arising in linear Gaussian models for Ornstein–Uhlenbeck processes and denoising diffusion models.
+Recent work [del2025stability] partially overcomes these limitations by establishing entropic continuity results for nonsymmetric transitions and strongly convex-at-infinity potentials.
+An alternative approach based on generalized Hilbert metrics for weighted function spaces is developed in [eckstein2025hilbert].
 Extending results in [chen, deligiannidis] to unbounded cost functions, the recent article  [eckstein2025hilbert] presents exponential decays of these new classes of Hilbert’s projective metrics with respect to some cone of functions in settings where the marginal distributions have sufficiently light tails compared to the growth of the cost function.
@@
 This condition doesn’t cover Gaussian models.
-As underlined in [adm-24], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step.
-These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [dpa, dm-04, dg-01, horton-dp, saloff-zuniga, saloff-zuniga-2, saloff-zuniga-3] and references therein.
-The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject, much more difficult to handle than their time homogeneous counterparts; this is because the operators may drastically change during the semigroup evolution.
-The articles [douc, saloff-zuniga, saloff-zuniga-2] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities.
-Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging.
-There is a need for new tools in order to handle these models.
-In this paper, we introduce a novel approach to Sinkhorn stability analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
-1.3 Contributions
+Finally, Sinkhorn semigroups belong to the broader class of time-inhomogeneous Markov processes that share a common invariant measure at each step [adm-24].
+Stability analysis for such processes is notoriously delicate, as transition operators may vary substantially over time.
+These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [dpa, dm-04, dg-01, horton-dp, saloff-zuniga-3] and references therein.
+While general tools exist—such as coupling methods, spectral techniques, and functional inequalities [douc, saloff-zuniga, saloff-zuniga-2]-their direct application to Sinkhorn semigroups is challenging due to the nonlinear conjugate structure of the updates.
+This motivates the development of new analytical tools.
+In this work, we introduce a Lyapunov-based, operator-theoretic framework for the quantitative stability analysis of Sinkhorn semigroups on weighted Banach spaces.
+1.3. Contributions
 We investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups.
@@
 In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [hasminskii] and Meyn and Tweedie [meyn-1], see also the book [meyn-2] and references therein.
-These powerful off-the-shelf semigroup techniques are mainly based on two conditions: 1) the existence of a Lyapunov function satisfying a drift-type condition; 2) a local minorization condition on the sub-level sets of the Lyapunov function.
+These powerful off-the-shelf semigroup techniques are mainly based on two conditions:
+  1. (1)
+    The existence of a Lyapunov function satisfying a drift-type condition.
+  2. (2)
+    A local minorization condition on the sub-level sets of the Lyapunov function.
 Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function.
 In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator.
-For a more detailed discussion on this contraction coefficients we refer to [dpa, horton-dp, penev].
+For a more detailed discussion on this contraction coefficients we refer to [dpa, delgerber2025, horton-dp, penev].
 This Lyapunov approach and the operator-theoretic framework presented in this article strongly differ from the approaches discussed earlier and taken in the series of recent articles [chiarini, durmus, del2025stability, eckstein2025hilbert].
 Our main contributions can be summarized in the following way:
-  • In the context of bounded transition potentials, we provide novel strict contraction estimates with respect to general $\Phi$-divergence criteria, including $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [csiszar, dm-03, majernik, nielsen] (see, for instance, Theorem 2.2 and Section 4.1).
-  • In the context of locally bounded cost functions and marginal potentials with compact sub-level sets, we develop new weighted total variation norm estimates, including Kantorovich and Wasserstein-type distances (see Section 4.4 and Section 4.5) as well as relative entropy bridge estimates (in Section 4.6).
-    We also emphasize that the choice of the Wasserstein semi-metric and the Lyapunov functions depends on the marginal target measures (see Theorem 2.3 and Section 4.5).
-  • We develop general Lyapunov contractions principles under practical and minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of models.
-    We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models (e.g., semi-discrete marginal models).
-Last but not least, our approach also allows to consider statistical finite mixture of the above models including kernel-type density estimators of complex data distributions arising in generative models.
-To the best of our knowledge, the stability results discussed above are the first results of this type in the literature on entropic optimal transport and Sinkhorn semigroups.
-1.4 Outline of the article
-After presenting some basic notation in Section 1.5, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [adm-24] for a more detailed discussion on these models within our framework).
-Section 2.3, presents a brief summary of the main stability theorems of the article.
-All the methodologies and regularity conditions discussed in the article are illustrated in a variety of examples in Section 5.
-To motivate our results, however, Section 2.4 already includes the discussion of some application models.
-We also give some detailed comments on our regularity conditions and propose an avenue of open research projects.
-Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C.
-Section 3.1 recalls some basics facts on the Hilbert projective metric, operator norms, entropy and related $\phi$-divergences.
-The operator contraction coefficients associated with these criteria are introduced in Section 3.2.
-With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups.
-Strong mixing Sinkhorn semigroups are discussed in Section 4.1 in terms of Dobrushin contraction coefficients.
-In this context, we show that Sinkhorn semigroups contract for any $\phi$-divergence criteria, including the relative entropy, the total variation distance, as well as $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, Jeffreys and Rényi divergences.
-Section 4.2 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups.
-The main contraction theorems introduced in this article are presented in Section 4.3.
-Section 4.4 presents a simple and new regularity condition that allows to provide quantitative exponential stability estimates for a large class of models.
-Section 4.5 presents quantitative exponential decays w.r.t.
-Wasserstein semi-metrics and Section 4.6 is dedicated to relative entropy decays and the convergence of Sinkhorn bridges.
-Section 5 is dedicated to a variety of illustrations.
-In particular, polynomial growth models are discussed in Section 5.1, Section 5.2 is dedicated to heavy tailed models, and in Section 5.3 we study bounded Lipschitz domains.
-Beta and Weibull distributions are discussed, respectively, in Section 5.4 and Section 5.5.
-An explicit description of the key regularity conditions and results in the context of multivariate linear Gaussian distributions is provided in Section 5.8.
-Finally, Section 5.9 is dedicated to statistical finite (non necessarily Gaussian) mixture models.
-Detailed proofs and some extended analyses can be found in the appendices to this manuscript.
-Appendix A is devoted to the Gaussian Sinkhorn algorithm, Appendices B and C contain material related to Hilbert metric and weighted norm contractions, respectively.
-Appendix D is devoted to Schrödinger potential functions.
-Finally, some additional technical proofs provided for completeness in Appendix E.
-1.5 Some basic notation
-We let ${\cal B}(\mathbb{X})$ be the algebra of locally bounded measurable functions on a locally compact Polish space $\mathbb{X}$.
-We denote by ${\cal B}_{b}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ the sub-algebra of bounded measurable functions $f$ on $\mathbb{X}$ endowed with the supremum norm $\|f\|:=\sup_{x\in\mathbb{X}}|f(x)|$.
-For a given uniformly positive function $\varphi\in{\cal B}(\mathbb{X})$, we let ${\cal B}_{\varphi}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the Banach space of functions $f\in{\cal B}(\mathbb{X})$ equipped with the norm $\|f\|_{\varphi}:=\|f/\varphi\|$.
-Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures.
-We denote the convex subset of probability measures on $\mathbb{X}$ as ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$.
-The Lebesgue integral of an integrable function $f\in{\cal B}(\mathbb{X})$ w.r.t. $\nu\in{\cal M}(\mathbb{X})$ is denoted by
+  • Strict contraction for bounded costs: For bounded transition potentials, we establish strict contraction inequalities for a broad class of $\phi$-divergences, including $\mathbb{L}_{p}$ norms, Havrda–Charvat and $\alpha$-divergences, Jensen–Shannon divergence, squared Hellinger and Kakutani–Hellinger distances, as well as Jeffreys and Rényi divergences (see Theorem 2.2 and Section 4.1).
+  • Weighted-norm stability for unbounded models: For locally bounded costs and marginal potentials with compact sub-level sets, we introduce weighted total variation and Kantorovich–Wasserstein contraction estimates, together with relative entropy bounds for Schrödinger bridges (Sections 4.4, 4.5, and 4.6).
+The choice of Lyapunov functions and Wasserstein metrics is explicitly adapted to the marginal distributions.
+  • General Lyapunov contraction principles: We formulate practical drift–minorization conditions yielding quantitative exponential convergence rates for a large class of Sinkhorn semigroups.
+These results apply to models with polynomial growth, heavy-tailed marginals, boundary state spaces, semi-compact domains, and finite mixture models, including kernel-based density estimators used in generative modeling.
+To the best of our knowledge, these results provide the first general exponential stability estimates for Sinkhorn semigroups with respect to weighted norms, $\phi$-divergences, and Wasserstein-type distances in non-necessarily compact and unbounded settings.
+1.4. Outline of the article
+After introducing notation in Section 1.5, Section 2 recalls the Schrödinger bridge problem and the associated Sinkhorn semigroups, emphasizing their interpretation as time-inhomogeneous nonlinear Markov evolutions (see also [adm-24]).
+Section 2.3 presents a concise overview of the main stability and contraction results established in this work.
+Motivating applications are already discussed in Section 2.4, which highlights the relevance of our approach for non-compact state spaces, unbounded costs, and heavy-tailed marginals.
+Sections 33.2 develop the operator-theoretic foundations underlying our analysis.
+We review Lyapunov drift–minorization techniques and weighted Banach space contractions for Markov integral operators (Section 3, with complements in Appendix C), recall the Hilbert projective metric and $\phi$-divergences (Section 3.1), and introduce the associated contraction coefficients (Section 3.2).
+Building on this framework, Section 4 adapts weighted-norm and Lyapunov contraction techniques to Sinkhorn semigroups.
+Strong mixing regimes are analyzed in Section 4.1, where we establish strict contraction properties for a broad class of $\phi$-divergences, including relative entropy, total variation, $\mathbb{L}_{p}$ norms, Hellinger-type divergences, Jensen–Shannon, Jeffreys, and Rényi divergences.
+General drift–minorization conditions ensuring quantitative contraction bounds are developed in Section 4.2.
+The main exponential stability theorems are presented in Section 4.3.
+Section 4.4 introduces a simple and practically verifiable Lyapunov condition yielding exponential convergence on weighted Banach spaces.
+Extensions to Wasserstein-type semi-metrics are derived in Section 4.5, while Section 4.6 is devoted to relative-entropy decay and the convergence of Sinkhorn bridges.
+Section 5 illustrates the scope of the theory through a wide range of examples, including polynomial growth models (Section 5.1), heavy-tailed marginals (Section 5.2), bounded Lipschitz domains (Section 5.3), Beta and Weibull models (Sections 5.4 and 5.5), and multivariate linear Gaussian models (Section 5.8).
+Section 5.9 extends the analysis to finite mixture models, including non-Gaussian and kernel-based constructions relevant to generative modeling.
+Detailed proofs and complementary material are collected in the appendices.
+Appendix A focuses on Gaussian Sinkhorn algorithms, Appendices B and C provide additional results on Hilbert metrics and weighted-norm contractions, Appendix D reviews Schrödinger potentials, and Appendix E contains further technical arguments.
+1.5. Some basic notation
+We let ${\mathcal{B}}(\mathbb{X})$ be the algebra of locally bounded measurable functions on a locally compact Polish space $\mathbb{X}$.
+We denote by ${\mathcal{B}}_{b}(\mathbb{X})\subset{\mathcal{B}}(\mathbb{X})$ the sub-algebra of bounded measurable functions $f$ on $\mathbb{X}$ endowed with the supremum norm $\|f\|:=\sup_{x\in\mathbb{X}}|f(x)|$.
+For a given uniformly positive function $\varphi\in{\mathcal{B}}(\mathbb{X})$, we let ${\mathcal{B}}_{\varphi}(\mathbb{X})\subset{\mathcal{B}}(\mathbb{X})$ be the Banach space of functions $f\in{\mathcal{B}}(\mathbb{X})$ equipped with the norm $\|f\|_{\varphi}:=\|f/\varphi\|$.
+Let ${\mathcal{M}}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\mathcal{M}}_{b}(\mathbb{X})\subset{\mathcal{M}}(\mathbb{X})$ the subset of bounded measures.
+We denote the convex subset of probability measures on $\mathbb{X}$ as ${\mathcal{M}}_{1}(\mathbb{X})\subset{\mathcal{M}}_{b}(\mathbb{X})$.
+The Lebesgue integral of an integrable function $f\in{\mathcal{B}}(\mathbb{X})$ w.r.t. $\nu\in{\mathcal{M}}(\mathbb{X})$ is denoted by
 $\nu(f):=\int_{\mathbb{X}}f(x)~\nu(dx).$
 For indicator functions $f=1_{A}$ of a measurable subset $A\subset\mathbb{X}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
-Consider a bounded integral operator ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\cal M}(\mathbb{X})$ and a function $f\in{\cal B}(\mathbb{Y})$.
-We denote by $\mu{\cal K}\in{\cal M}(\mathbb{Y})$ and ${\cal K}(f)\in{\cal B}(\mathbb{X})$ the measure and the function defined by
-$(\mu{\cal K})(dy):=\int~\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y).$ (6)
-For indicator functions $f=1_{A}$ sometimes we write ${\cal K}(x,A)$ instead of ${\cal K}(1_{A})(x)$.
-For bounded integral operator ${\cal K}_{1}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and ${\cal K}_{2}(y,dx)$ from $\mathbb{Y}$ into $\mathbb{X}$, we also denote by $({\cal K}_{1}{\cal K}_{2})$ the bounded integral operator from $\mathbb{X}$ into $\mathbb{X}$ defined by the integral composition
-$({\cal K}_{1}{\cal K}_{2})(x_{1},dx_{2}):=\int{\cal K}_{1}(x_{1},dy){\cal K}_{2}(y,dx_{2}).$
-Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we set
-$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(y,x)):=\mu(dx)\,{\cal K}(x,dy).$
-A measure $\nu_{1}\in{\cal M}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\cal M}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$.
+Consider a bounded integral operator ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\mathcal{M}}(\mathbb{X})$ and a function $f\in{\mathcal{B}}(\mathbb{Y})$.
+We denote by $\mu{\mathcal{K}}\in{\mathcal{M}}(\mathbb{Y})$ and ${\mathcal{K}}(f)\in{\mathcal{B}}(\mathbb{X})$ the measure and the function defined by
+(1.6) $(\mu{\mathcal{K}})(dy):=\int~\mu(dx){\mathcal{K}}(x,dy)\quad\mbox{\rm and}\quad{\mathcal{K}}(f)(x):=\int{\mathcal{K}}(x,dy)f(y).$
+For indicator functions $f=1_{A}$ sometimes we write ${\mathcal{K}}(x,A)$ instead of ${\mathcal{K}}(1_{A})(x)$.
+For bounded integral operator ${\mathcal{K}}_{1}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and ${\mathcal{K}}_{2}(y,dx)$ from $\mathbb{Y}$ into $\mathbb{X}$, we also denote by $({\mathcal{K}}_{1}{\mathcal{K}}_{2})$ the bounded integral operator from $\mathbb{X}$ into $\mathbb{X}$ defined by the integral composition
+$({\mathcal{K}}_{1}{\mathcal{K}}_{2})(x_{1},dx_{2}):=\int{\mathcal{K}}_{1}(x_{1},dy){\mathcal{K}}_{2}(y,dx_{2}).$
+Given a Markov transition ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\mathcal{M}}(\mathbb{X})$ we set
+$(\mu\times{\mathcal{K}})(d(x,y)):=\mu(dx)\,{\mathcal{K}}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\mathcal{K}})^{\flat}(d(y,x)):=\mu(dx)\,{\mathcal{K}}(x,dy).$
+A measure $\nu_{1}\in{\mathcal{M}}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\mathcal{M}}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$.
 Whenever $\nu_{1}\ll\nu_{2}$, we denote by ${d\nu_{1}}/{d\nu_{2}}$ the Radon-Nikodym derivative of $\nu_{1}$ w.r.t.
@@
 Kantorovich and Wasserstein criteria
-We consider the weighted total variation norm associated with some positive function $\varphi\in{\cal B}(\mathbb{X})$ defined as
+We consider the weighted total variation norm associated with some positive function $\varphi\in{\mathcal{B}}(\mathbb{X})$ defined as
 […]
-where $\Pi(\mu_{1},\mu_{2})$ stands for the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{X})$, with prescribed first and second coordinate marginals $(\mu_{1},\mu_{2})$ such that $\mu_{i}(\varphi)<\infty$.
+where $\Pi(\mu_{1},\mu_{2})$ stands for the set of coupling probability measures $\mathcal{P}$} on the product space $(\mathbb{X}\times\mathbb{X})$, with prescribed first and second coordinate marginals $(\mu_{1},\mu_{2})$ such that $\mu_{i}(\varphi)<\infty$.
 For a given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ and $p\geq 1$ we also consider the Wasserstein semi-metric $\mathbb{W}_{p,\varrho}$ defined by
-$\mathbb{W}_{p,\varrho}(\mu_{1},\mu_{2}):=\inf_{{\cal P}\in\Pi(\mu_{1},\mu_{2})}\left(\int~\varrho(x_{1},x_{2})^{p}~{\cal P}(d(x_{1},x_{2}))\right)^{1/p}.$ (8)
+(1.8) $\mathbb{W}_{p,\varrho}(\mu_{1},\mu_{2}):=\inf_{{\mathcal{P}}\in\Pi(\mu_{1},\mu_{2})}\left(\int~\varrho(x_{1},x_{2})^{p}~{\mathcal{P}}(d(x_{1},x_{2}))\right)^{1/p}.$
 […]
@@
 More generally, for any given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ such that
-$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm and}\quad\varphi(x)\geq c>0$ (9)
+(1.9) $\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm and}\quad\varphi(x)\geq c>0$
 we have the estimates
@@
 […]
-The choice of the semi-metric $\varrho$ in (10) is not unique. Indeed, developing the exponential function for any $q\geq 1$ we also have the estimate
-$\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{p,q}(x_{1},x_{2}):=c_{p,q}~\|x_{1}-x_{2}\|^{pq}\quad\mbox{\rm with}\quad c_{p,q}:=2c~\frac{1}{q!}\left(\frac{b}{2^{1+(p-1)_{+}}}\right)^{q},$ (11)
+The choice of the semi-metric $\varrho$ in (1.10) is not unique. Indeed, developing the exponential function for any $q\geq 1$ we also have the estimate
+(1.11) $\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{p,q}(x_{1},x_{2}):=c_{p,q}~\|x_{1}-x_{2}\|^{pq}\quad\mbox{\rm with}\quad c_{p,q}:=2c~\frac{1}{q!}\left(\frac{b}{2^{1+(p-1)_{+}}}\right)^{q},$
 and
@@
 Consider a convex function
-$\begin{array}[]{cccl}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\quad\mbox{\rm s.t.
-$\forall a\in\mathbb{R}_{+}$~~~ $\Phi(au,av)=a~\Phi(u,v)$ and $\Phi(1,1)=0$.}\end{array}$ (12)
-The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
-$D_{\Phi}(\nu_{1},\nu_{2}):=\int~\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~\gamma(dx).$ (13)
-By homogeneity arguments, the definition (13) does not depend on the choice of the dominating measure $\gamma$. Note that
-$\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$ (14)
+(1.12) $\begin{array}[]{cccl}\Phi:&\mathbb{R}_{+}^{2}&\rightarrow&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\mapsto&\Phi(u,v)\quad\mbox{\rm s.t.
+$\forall a\in\mathbb{R}_{+}$~~~ $\Phi(au,av)=a~\Phi(u,v)$ and $\Phi(1,1)=0$.}\end{array}$
+The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
+(1.13) $D_{\Phi}(\nu_{1},\nu_{2}):=\int~\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~\gamma(dx).$
+By homogeneity arguments, the definition (1.13) does not depend on the choice of the dominating measure $\gamma$. Note that
+(1.14) $\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$
 Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [csiszar], i.e., we also have
@@
 $\phi_{1}(u)=u\log{u}-(u-1).$
-2 Statement of main results
-2.1 Sinkhorn bridges
-The Schrödinger bridge distribution $\mathbb{P}$ defined in (4) can rarely be computed analytically.
+2. Statement of main results
+This section introduces the Sinkhorn bridge construction, the associated time-inhomogeneous Sinkhorn semigroups, and states our main exponential stability results.
+Detailed proofs and refinements are deferred to later sections.
+2.1. Sinkhorn bridges
+The Schrödinger bridge distribution $\mathbb{P}$ defined in (1.4) can rarely be computed analytically.
 However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [cuturi, sinkhorn-2, sinkhorn-3].
-This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions (${\cal P}_{n}$) and transition operators (${\cal S}_{n}$) given by
-${\cal P}_{2n}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad{\cal P}_{2n+1}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}$ (15)
-starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dy)$ and a transition operator ${\cal S}(y,dx)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
-$(\mu\times{\cal S})^{\flat}(d(x,y)):=\mu(dy){\cal S}(y,dx).$
-The Markov transitions ${\cal S}_{n}$ in (15) are defined by the Sinkhorn iterations
-$\left\{\begin{array}[]{l}(\pi_{2n}\times{\cal S}_{2n+1})^{\flat}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\cal S}_{2(n+1)}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\cal S}_{2n+1}.\end{array}\right$.
-(16)
-In the dual formulation, the bridge distributions ${\cal P}_{n}$ are often written as
-${\cal P}_{n}(d(x,y))=e^{-W(x,y)}~\lambda_{U_{n}}(dx)~\nu_{V_{n}}(dy),$ (17)
+This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions (${\mathcal{P}}_{n}$) and transition operators (${\mathcal{S}}_{n}$) given by
+(2.1) ${\mathcal{P}}_{2n}=\lambda_{U}\times{\mathcal{S}}_{2n}\quad\mbox{\rm and}\quad{\mathcal{P}}_{2n+1}=(\nu_{V}\times{\mathcal{S}}_{2n+1})^{\flat}$
+starting from ${\mathcal{P}}_{0}$ with ${\mathcal{S}}_{0}={\mathcal{Q}}$, where, for a measure $\mu(dy)$ and a transition operator ${\mathcal{S}}(y,dx)$, the notation $(\mu\times{\mathcal{S}})^{\flat}$ denotes the ‘reverse’ measure
+$(\mu\times{\mathcal{S}})^{\flat}(d(x,y)):=\mu(dy){\mathcal{S}}(y,dx).$
+The Markov transitions ${\mathcal{S}}_{n}$ in (2.1) are defined by the Sinkhorn iterations
+(2.2) $\left\{\begin{array}[]{l}(\pi_{2n}\times{\mathcal{S}}_{2n+1})^{\flat}=\lambda_{U}\times{\mathcal{S}}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\mathcal{S}}_{2(n+1)}=(\nu_{V}\times{\mathcal{S}}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\mathcal{S}}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\mathcal{S}}_{2n+1}.\end{array}\right$.
+In the dual formulation, the bridge distributions ${\mathcal{P}}_{n}$ are often written as
+(2.3) ${\mathcal{P}}_{n}(d(x,y))=e^{-W(x,y)}~\lambda_{U_{n}}(dx)~\nu_{V_{n}}(dy),$
 with the initial potential functions $(U_{0},V_{0})=(U,0)$ and potential functions $(U_{n},V_{n})$ defined sequentially by the integral recursions
-$U_{2n+1}$ $=$ $~U_{2n}~:=U+\log{{\cal Q}(e^{-V_{2n}})}$
-$V_{2(n+1)}$ $=$ $V_{2n+1}:=V+\log{{\cal R}(e^{-U_{2n}})}.$ (18)
-When $n\to\infty$, the sequence of distributions ${\cal P}_{2n}$ converges towards the Schrödinger bridge (4) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t.
-the reference measure ${\cal P}_{0}$) while the sequence ${\cal P}_{2n+1}$ converges towards the Schrödinger bridge from $\nu_{V}$ to $\lambda_{U}$ (w.r.t.
-the reference measure ${\cal P}_{1}$).
-Under mild integrability assumptions (see for instance Section 6 in [adm-24]) the Schrödinger bridge distribution $\mathbb{P}$ solving (4) can also be computed from any Sinkhorn bridge reference distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
-$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{2n}).$ (19)
+$U_{2n+1}$ $=$ $~U_{2n}~:=U+\log{{\mathcal{Q}}(e^{-V_{2n}})}$
+(2.4) $V_{2(n+1)}$ $=$ $V_{2n+1}:=V+\log{{\mathcal{R}}(e^{-U_{2n}})}.$
+When $n\to\infty$, the sequence of distributions ${\mathcal{P}}_{2n}$ converges towards the Schrödinger bridge (1.4) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t.
+the reference measure ${\mathcal{P}}_{0}$) while the sequence ${\mathcal{P}}_{2n+1}$ converges towards the Schrödinger bridge from $\nu_{V}$ to $\lambda_{U}$ (w.r.t.
+the reference measure ${\mathcal{P}}_{1}$).
+Under mild integrability assumptions (see for instance Section 6 in [adm-24]) the Schrödinger bridge distribution $\mathbb{P}$ solving (1.4) can also be computed from any Sinkhorn bridge reference distribution ${\mathcal{P}}_{2n}$, in the sense that for any $n\geq 0$ we have
+(2.5) $\mathbb{P}=\operatorname*{arg\,min}_{{\mathcal{P}}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{2n}).$
 The Sinkhorn transition operators can also be written in terms of the potential functions $(U_{n},V_{n})$, namely,
-${\cal S}_{2n}(x,dy)=\frac{{\cal Q}(x,dy)e^{-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\cal S}_{2n+1}(y,dx)=\frac{{\cal R}(y,dx)~e^{-U_{2n+1}(x)}}{{\cal R}(e^{-U_{2n+1}})(y)},$ (20)
-while Proposition 6.3 in [adm-24] yields the formulae
-$V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$ (21)
-for the potentials $(U_{n},V_{n})$, $n\geq 0$, expressed in terms of the distribution flow of the internal states of the Gibbs-loop process.
-2.2 Sinkhorn semigroups
-For any $n\geq 0$ we have the fixed point equations
-$\lambda_{U}{\cal S}^{\circ}_{2n+1}=\lambda_{U}\quad\mbox{\rm and}\quad\nu_{V}{\cal S}^{\circ}_{2(n+1)}=\nu_{V}$ (22)
-with the forward-backward Gibbs-type Markov transitions
-${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}.$
-The above fixed point equations yield, for any $n\geq 1$, the time inhomogeneous Markov evolution equations
-$\pi_{2n}=\pi_{2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2n-1}{\cal S}^{\circ}_{2n+1}.$ (23)
-For any $0\leq p\leq n$ we have the semigroup equations
-$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
-The Sinkhorn transitions ${\cal S}^{\circ}_{2p,2n}$ and ${\cal S}^{\circ}_{2p+1,2n+1}$ are defined inductively by the time inhomogeneous semigroup equations
-${\cal S}^{\circ}_{2p,2n}={\cal S}^{\circ}_{2p,2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2p+1,2n+1}={\cal S}^{\circ}_{2p+1,2n-1}{\cal S}^{\circ}_{2n+1},$
-with ${\cal S}^{\circ}_{2n,2n}=I={\cal S}^{\circ}_{2n+1,2n+1}$ the identity integral operator.
-The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (23) sharing the same invariant measures (22) at each individual step.
+(2.6) ${\mathcal{S}}_{2n}(x,dy)=\frac{{\mathcal{Q}}(x,dy)e^{-V_{2n}(y)}}{{\mathcal{Q}}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\mathcal{S}}_{2n+1}(y,dx)=\frac{{\mathcal{R}}(y,dx)~e^{-U_{2n+1}(x)}}{{\mathcal{R}}(e^{-U_{2n+1}})(y)},$
+Moreover, Proposition 6.3 in [adm-24] shows that the potentials can be expressed through the marginal flows as
+(2.7) $V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$
+2.2. Sinkhorn semigroups
+The Sinkhorn iterations naturally induce a family of time-inhomogeneous Markov semigroups. For each $n\geq 0$ define the forward-backward Gibbs-type transitions
+${\mathcal{S}}^{\circ}_{2n+1}:={\mathcal{S}}_{2n}{\mathcal{S}}_{2n+1}\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2(n+1)}:={\mathcal{S}}_{2n+1}{\mathcal{S}}_{2(n+1)}.$
+These kernels satisfy the fixed-point relations
+(2.8) $\lambda_{U}{\mathcal{S}}^{\circ}_{2n+1}=\lambda_{U}\quad\mbox{\rm and}\quad\nu_{V}{\mathcal{S}}^{\circ}_{2(n+1)}=\nu_{V}$
+and yield the marginal evolution equations
+(2.9) $\pi_{2n}=\pi_{2(n-1)}{\mathcal{S}}^{\circ}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2n-1}{\mathcal{S}}^{\circ}_{2n+1}.$
+For $0\leq p\leq n$, we define the associated inhomogeneous semigroups by composition,
+${\mathcal{S}}^{\circ}_{2p,2n}={\mathcal{S}}^{\circ}_{2p,2(n-1)}{\mathcal{S}}^{\circ}_{2n}\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2p+1,2n+1}={\mathcal{S}}^{\circ}_{2p+1,2n-1}{\mathcal{S}}^{\circ}_{2n+1},$
+with identity operators on the diagonal. This representation highlights that Sinkhorn dynamics belong to the class of time-varying Markov chains sharing common invariant measures at each step.
 Remark 2.1.
-At every time step $n\geq 0$, the conjugate formulae (16) also show that ${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure
-${\cal P}_{2n}(d(x,y))=\lambda_{U}(dx)~{\cal S}_{2n}(x,dy)=\pi_{2n}(dy)~{\cal S}_{2n+1}(y,dx)$
-For $n=0$, the two-blocks Gibbs sampler with the forward/backward transition ${\cal S}^{\circ}_{1}:={\cal S}_{0}{\cal S}_{1}$ and target product measure ${\cal P}_{0}=\lambda_{U}\times{\cal S}_{0}$ can be interpreted as a sampling technique for $\lambda_{U}$ based on data augmentation or auxiliary variables (a.k.a.
-proximal sampling), see for instance [damlen1999gibbs, rendell2020global, vono2020asymptotically, vono2019split] and references therein.
-For a more detailed discussion on proximal samplers we also refer to the series of articles [mou, chewi-chen, chewi-phd, jiaming, jiaojiao, lee, guan].
-Similarly, ${\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure ${\cal P}_{2n+1}$.
-2.3 Some stability theorems
+Each kernel ${\mathcal{S}}^{\circ}_{2n+1}:={\mathcal{S}}_{2n}{\mathcal{S}}_{2n+1}$ corresponds to a two-block Gibbs sampler with target distribution
+${\mathcal{P}}_{2n}(d(x,y))=\lambda_{U}(dx)~{\mathcal{S}}_{2n}(x,dy)=\pi_{2n}(dy)~{\mathcal{S}}_{2n+1}(y,dx)$
+For $n=0$, this recovers classical proximal and data-augmentation samplers; see, e.g., [damlen1999gibbs, rendell2020global, vono2019split, vono2020asymptotically] and related works [chewi-chen, chewi-phd, jiaojiao, guan, lee, jiaming, mou].
+2.3. Some stability theorems
 In contrast with Kantorovich criteria or related Wasserstein-type semi-distances, even when $\mathbb{X}=\mathbb{Y}$ and the marginals coincide, $\lambda_{U}=\nu_{V}$, the identity-map deterministic coupling has infinite entropy.
-That is, for any reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ of the form (3) we have
-${\cal P}(d(x,y)):=\lambda_{U}(dx)~\delta_{x}(dy)\in\Pi_{U,V}\Longrightarrow\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=+\infty.$
-This fact, together with (5), shows that the Schrödinger bridge $\mathbb{P}$ and the convergence analysis of Sinkhorn bridges ${\cal P}_{2n}$ strongly depend on the form of the reference transition $\cal Q$} and the transition potential function $W$ discussed in (2).
-Markov transitions of the form (2) with a bounded potential $W$ typically arise on finite or compact spaces.
-For instance, Markov processes on compact manifolds, such as diffusions on bounded manifolds reflected at the boundary, have Markov transitions of the form (2) with a uniformly bounded transition potential $W$ that depends on the time parameter [dpa, aronson, nash, varopoulos].
+That is, for any reference measure ${\mathcal{P}}_{0}=(\lambda_{U}\times{\mathcal{Q}})$ of the form (1.3) we have
+${\mathcal{P}}(d(x,y)):=\lambda_{U}(dx)~\delta_{x}(dy)\in\Pi_{U,V}\Longrightarrow\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0})=+\infty.$
+This fact, together with (1.5), shows that the Schrödinger bridge $\mathbb{P}$ and the convergence analysis of Sinkhorn bridges ${\mathcal{P}}_{2n}$ crucially depend on the form of the reference transition $\mathcal{Q}$} and the transition potential function $W$ discussed in (1.2).
+Markov transitions of the form (1.2) with a bounded potential $W$ typically arise on finite or compact spaces.
+For instance, Markov processes on compact manifolds, such as diffusions on bounded manifolds reflected at the boundary, have Markov transitions of the form (1.2) with a uniformly bounded transition potential $W$ that depends on the time parameter [dpa, aronson, nash, varopoulos].
 The convergence analysis of Sinkhorn bridges in this context is often developed using Hilbert Hilbert projective metrics [chen, deligiannidis, franklin, marino].
-In the context of bounded transition potentials, the contraction principles of Markov operators presented in the article [dm-03] yield an alternative stability analysis that also applies to any generalized $\Phi$-entropies introduced in (13).
+In the context of bounded transition potentials, the contraction principles of Markov operators presented in the article [dm-03] yield an alternative stability analysis that also applies to any generalized $\Phi$-entropies introduced in (1.13).
 Theorem 2.2.
@@
 In this case, for any $\Phi$-entropy criteria $D_{\Phi}$ there is some parameters $\rho\in]0,1[$ and $c>0$ such that for any $n\geq 0$ we have the contraction inequalities
-$D_{\Phi}\left(\pi_{2n},\nu_{V}\right)\leq\rho^{n}~D_{\Phi}\left(\pi_{0},\nu_{V}\right)\quad\mbox{and}\quad\left\|\log{({d\pi_{2n}}/{d\nu_{V}})}\right\|\leq c~\rho^{n}.$ (24)
-The above inequalities remain valid when we replace $(\pi_{2n},\nu_{V})$ by $(\nu_{V},\pi_{2n})$.
-Similar estimates hold for $(\lambda_{U},\pi_{2n+1})$.
-There also exists $c>0$ such that for any $n\geq 0$ we have
-$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\rho^{n}.$ (25)
-The estimates (24) are a consequence of the sandwich-type contraction inequalities stated in Theorem 4.1.
-A more refined version of the entropy estimate (25) in terms of Dobrushin contraction coeffcients is provided in Corollary 4.2
+(2.10) $D_{\Phi}\left(\pi_{2n},\nu_{V}\right)\leq\rho^{n}~D_{\Phi}\left(\pi_{0},\nu_{V}\right)\quad\mbox{and}\quad\left\|\log{({d\pi_{2n}}/{d\nu_{V}})}\right\|\leq c~\rho^{n}.$
+The above inequalities remain valid when we replace $(\pi_{2n},\nu_{V})$ by $(\nu_{V},\pi_{2n})$. Analogous bounds hold for $(\lambda_{U},\pi_{2n+1})$. Moreover,
+(2.11) $\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})\leq c~\rho^{n}.$
+The estimates (2.10) are a consequence of the sandwich-type contraction inequalities stated in Theorem 4.1.
+A more refined version of the entropy estimate (2.11) in terms of Dobrushin contraction coefficients is provided in Corollary 4.2
 We underline that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}=]0,1[$ the transition potential function $W$ may be unbounded.
 Indeed, consider the reference Markov transition defined for any $x\in]0,1[$ by
-${\cal Q}(x,dy)=\frac{1}{q(x)}~\left(1-|x-y|\right)~1_{]0,1[}(y)~dy\quad\mbox{\rm with}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.$ (26)
+(2.12) ${\mathcal{Q}}(x,dy)=\frac{1}{q(x)}~\left(1-|x-y|\right)~1_{]0,1[}(y)~dy\quad\mbox{\rm with}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.$
 In this context, we have
@@
 This rather simple model does not satisfy the boundedness condition of Theorem 2.2. To handle locally bounded transition potentials, consider the functions
-${\cal U}_{\delta}:=\delta U-W^{V}\quad\mbox{\rm and}\quad{\cal V}_{\delta}:=\delta V-W_{U}$ (27)
+(2.13) ${\mathcal{U}}_{\delta}:=\delta U-W^{V}\quad\mbox{\rm and}\quad{\mathcal{V}}_{\delta}:=\delta V-W_{U}$
 with the integrated costs
-$W_{U}(y):=\int~\lambda_{U}(dx)~W(x,y)\quad\text{and}\quad W^{V}(x):=\int~\nu_{V}(dy)~W(x,y).$ (28)
+(2.14) $W_{U}(y):=\int~\lambda_{U}(dx)~W(x,y)\quad\text{and}\quad W^{V}(x):=\int~\nu_{V}(dy)~W(x,y).$
 Intuitively, the next theorem shows that Sinkhorn semigroups are stable as soon as the marginal-potentials $(U,V)$ dominate at infinity the (integrated) transition-potentials $(W_{U},W^{V})$.
@@
 Theorem 2.3.
-Assume that $W$ is locally bounded and bounded below by a real number and $(U,V)$ are locally bounded with compact sub-level sets.
-In addition, there exists some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta/2},{\cal V}_{\delta/2})$ are locally bounded with compact sub-level sets and we have $\lambda_{(1-\delta)U}(1)\wedge\nu_{(1-\delta)V}(1)<\infty$.
-Then, there is some $\rho\in]0,1[$, $n_{0}\geq 1$ and $c>0$ s.t.
-for any $n\geq n_{0}$ we have
+Assume that $W$ is locally bounded and bounded from below by a real number and $(U,V)$ are locally bounded with compact sub-level sets.
+In addition, there exists some $\delta\in]0,1[$ such that the functions $({\mathcal{U}}_{\delta/2},{\mathcal{V}}_{\delta/2})$ are locally bounded with compact sub-level sets and we have $\lambda_{(1-\delta)U}(1)\wedge\nu_{(1-\delta)V}(1)<\infty$.
+Then, there exist some $\rho\in]0,1[$, $n_{0}\geq 1$ and $c>0$ such that for any $n\geq n_{0}$,
 […]
-When $\psi$ satisfies the lower bound estimate (9) for some semi-metric $\varrho$ we also have the exponential Wasserstein estimates
-$\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\vee\mathbb{W}_{\varrho}(\pi_{2n},\nu_{V})\leq c~\rho^{n}.$ (30)
+When $\psi$ satisfies the lower bound estimate (1.9) for some semi-metric $\varrho$ then we also have the exponential Wasserstein estimates
+(2.16) $\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\vee\mathbb{W}_{\varrho}(\pi_{2n},\nu_{V})\leq c~\rho^{n}.$
 […]
-The estimates (29) are a consequence of Corollary 4.15 and the Wasserstein formulation (7) of weighted norms.
-The estimates (30) are a consequence of (10) and (29).
-More refined strict contraction estimates are provided in Corollary 4.14 and Corollary 4.16.
-2.4 Illustrations and comments
+The estimates (2.15) are a consequence of Corollary 4.10 and the Wasserstein formulation (1.7) of weighted norms.
+The estimates (2.16) are a consequence of (1.10) and (2.15).
+More refined strict contraction estimates are established in Corollary 4.9 and Corollary 4.11.
+2.4. Illustrations and comments
 Some illustrations
-In Section 5, the stability theorems presented in Section 2.3 and Section 4 are illustrated in a variety of situations ranging from polynomial growth potentials (Section 5.1) and heavy tailed marginals on general normed spaces (Section 5.2), to more sophisticated boundary state space models (Section 5.3), semi-compact models including semi-discrete target spaces (Section 5.7) as well as finite mixture of statistical models (Section 5.9).
-Section 5.8 is dedicated to explicit descriptions and equivalent regularity conditions in the context of linear Gaussian models.
-When $W$ is bounded, it does not appear useful to consider weighted total variation norms.
-In this context, the stability of Sinkhorn semigroups w.r.t.
-Hilbert’s metric is well known [borwein, chen, deligiannidis, franklin] and already yields exponential decays w.r.t.
-the total variation norm.
-On bounded subspaces $\mathbb{X}\subset\mathbb{R}^{d}$ the total variation norm also dominates Wasserstein distances w.r.t.
-the Euclidian norm but in general the target space $\mathbb{X}$ is non complete and then not Polish w.r.t.
-the Euclidean distance.
+Section 5 provides a range of examples illustrating the stability results established in Sections 2.3 and 4.
+These include polynomial growth potentials (Section 5.1), heavy-tailed marginals on general normed spaces (Section 5.2), boundary state-space models (Section 5.3), semi-compact and semi-discrete settings (Section 5.7), as well as finite mixtures of statistical models (Section 5.9).
+Section 5.8 is devoted to linear Gaussian models, where the regularity conditions and Lyapunov criteria admit explicit characterizations.
+When the transition potential $W$ is uniformly bounded, weighted total variation norms do not provide additional insight.
+In this regime, exponential convergence of Sinkhorn semigroups is well understood through Hilbert’s projective metric and related contraction techniques [borwein, chen, deligiannidis, franklin], which already imply exponential decay in total variation.
+On bounded subsets $\mathbb{X}\subset\mathbb{R}^{d}$, total variation also dominates Wasserstein distances associated with the Euclidean metric.
+However, the state space $\mathbb{X}$ is often non-complete and therefore non-Polish with respect to the Euclidean distance, which complicates Wasserstein convergence analysis.
 In this context, it is desirable to find a complete distance on the target space that makes the corresponding Wasserstein distance in distribution space complete [bolley2008separability].
-It is easy to construct examples of marginal models on non complete metric spaces.
-For instance, consider the marginal Beta distributions $\lambda_{U}=\nu_{V}$ on $\mathbb{X}=\mathbb{Y}=]0,1[$ equipped with the Euclidean metric $\varrho_{e}\left(x_{1},x_{2}\right)=|x_{1}-x_{2}|$ and defined by
+This issue naturally motivates the introduction of alternative complete metrics.
+As a simple illustration, consider the Beta marginal distributions $\lambda_{U}=\nu_{V}$ on $\mathbb{X}=\mathbb{Y}=]0,1[$ equipped with the Euclidean metric $\varrho_{e}\left(x_{1},x_{2}\right)=|x_{1}-x_{2}|$ and defined by
 $\nu_{V}(dx)=\frac{1}{6}~x(1-x)~\nu(dx)\quad\mbox{\rm with}\quad\nu(dx):=1_{]0,1[}(x)~dx.$
-Applying (24) to the total variation distance (see (14)) we readily check the Wasserstein exponential decays
+Applying (2.10) to the total variation distance (see (1.14)) we readily check the Wasserstein exponential decays
 $\mathbb{W}_{\varrho_{e}}(\pi_{2n},\nu_{V})\leq\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\leq\rho^{n}~\|\pi_{0}-\nu_{V}\|_{\tiny tv}.$
-Note that the open unit interval $\mathbb{X}=]0,1[$ is not complete for the Euclidean metric but it is complete for the metric defined as
+Although the open unit interval $\mathbb{X}=]0,1[$ is not complete for the Euclidean metric, it becomes complete under
 $\varrho_{c}\left(x_{1},x_{2}\right):=\left|\frac{1}{x_{1}}-\frac{1}{x_{2}}\right|+\left|\frac{1}{1-x_{1}}-\frac{1}{1-x_{2}}\right|\geq 2~{\varrho_{e}}(x_{1},x_{2}).$
@@
 $\psi(x_{1})+\psi(x_{2})\geq\varrho_{c,\delta}\left(x_{1},x_{2}\right):=c_{\delta}~\varrho_{c}\left(x_{1},x_{2}\right)^{\delta}$
-for some $c_{\delta}>0$. In this context, the l.h.s. estimate (29) yields the $\varrho_{c,\delta}$-Wasserstein estimates
+for some $c_{\delta}>0$. In this context, the l.h.s. estimate (2.15) yields the $\varrho_{c,\delta}$-Wasserstein estimates
 […]
@@
 Section 5.1).
-In this context, applying (11) to $(p,q)=(2,1)$ we have the $2$-Wasserstein estimate (30), which implies
+In this context, for $(p,q)=(2,1)$ in (1.11), we obtain exponential convergence in the the $2$-Wasserstein distance,
 […]
-Section 5.8 provides an explicit description of the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ in the context of linear-Gaussian models.
-For instance, for one-dimensional centered Gaussian marginals with unit variance and a symmetric Gaussian reference transition $\cal Q$} with variance $\tau>2$ we have
-${2}/{\tau}<\delta<1\Longrightarrow{\cal U}_{\delta/2}(x)=c_{\delta}+\frac{x^{2}}{2}~\left(\frac{\delta}{2}-\frac{1}{\tau}\right)\stackrel{{\scriptstyle|x|\to\infty}}{{\longrightarrow}}+\infty\quad\mbox{\rm for some $c_{\delta}\in\mathbb{R}$.
+Section 5.8 provides an explicit description of the functions $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ in the context of linear-Gaussian models.
+For instance, for one-dimensional centered Gaussian marginals with unit variance and a symmetric Gaussian reference transition $\mathcal{Q}$} with variance $\tau>2$ we have
+${2}/{\tau}<\delta<1\Longrightarrow{\mathcal{U}}_{\delta/2}(x)=c_{\delta}+\frac{x^{2}}{2}~\left(\frac{\delta}{2}-\frac{1}{\tau}\right)\stackrel{{\scriptstyle|x|\to\infty}}{{\longrightarrow}}+\infty\quad\mbox{\rm for some $c_{\delta}\in\mathbb{R}$.
 }$
-A symmetric Gaussian reference transition $\cal Q$} with variance $\tau$ can be interpreted as the transition probability of a Brownian motion on a time horizon $\tau$.
+A symmetric Gaussian reference transition $\mathcal{Q}$} with variance $\tau$ can be interpreted as the transition probability of a Brownian motion on a time horizon $\tau$.
 In this context, the exponential stability of Sinkhorn bridges w.r.t.
-the relative entropy for log-concave marginals at infinity, including Gaussian marginals, also holds for any time horizon $\tau>0$, see for instance [del2025stability, chiarini, durmus].
+the relative entropy for log-concave marginals at infinity, including Gaussian marginals, also holds for any time horizon $\tau>0$, see for instance [chiarini, durmus, del2025stability].
 The article [adm-24] also provides Wasserstein exponential decays for any Gaussian marginals and any linear-Gaussian reference transition.
@@
 Energy barriers
-Section 2.3 and Section 5 discuss a series of application models that can be handled using our Lyapunov approach.
-This section is concerned with some limitations of our approach and related important open research problems.
+Sections 2.3 and 5 demonstrate that Lyapunov techniques provide a flexible tool for establishing quantitative stability of Sinkhorn semigroups. Nevertheless, several important limitations remain.
 Firstly, note that Theorem 2.3 applies when $W$ is locally bounded.
@@
 Indeed, consider the reference Markov transition defined for any $x\in]0,1[$ by
-$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{q(x)}~|x-y|~\nu(dy)\\ \\ \displaystyle\mbox{\rm with}\quad\nu(dy):=1_{]0,1[}(y)~dy\quad\mbox{\rm and}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.\end{array}$
-This rather elementary example is a a slight modification of (26) but in this case the function $W$ is not defined on the diagonal and for any $x\not=y$ we have
+$\begin{array}[]{l}\displaystyle{\mathcal{Q}}(x,dy)=\frac{1}{q(x)}~|x-y|~\nu(dy)\\ \\ \displaystyle\mbox{\rm with}\quad\nu(dy):=1_{]0,1[}(y)~dy\quad\mbox{\rm and}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.\end{array}$
+This rather elementary example is a a slight modification of (2.12) but in this case the function $W$ is not defined on the diagonal and for any $x\not=y$ we have
 $\begin{array}[]{l}\displaystyle W(x,y)=-\log{|x-y|}+\log{q(x)}\Longrightarrow W(x,x)=+\infty.\end{array}$
@@
 $\begin{array}[]{l}\displaystyle\lambda_{U}(dx):=4~|x-1/2|~\lambda(dx)\quad\mbox{\rm with}\quad\lambda(dx):=1_{]0,1[}(x)~dx\\ \\ \displaystyle\text{~~yields~~}U(x)=-\log{4}-\log{|x-1/2|}\Longrightarrow U(1/2)=+\infty.\end{array}$
-In this case the potential function $U$ is not well defined on the whole state space, thus this marginal model cannot be written in terms of a Boltzmann-Gibbs measure of the form (1).
+In this case the potential function $U$ is not well defined on the whole state space, thus this marginal model cannot be written in terms of a Boltzmann-Gibbs measure of the form (1.1).
 Nevertheless for uniformly bounded transition potential functions, Theorem 2.2 applies to any class of marginal distributions.
@@
 Even for bounded transition potential functions Theorem 2.3 cannot be used (but as mentioned above, Theorem 2.2 applies to any marginal models).
-3 Contraction of Markov operators
+3. Contraction of Markov operators
 Section 3.1 is dedicated to a brief description of the Hilbert (projective) metric and integral operator norms.
@@
 These coefficients will be key to the analysis of contraction properties of Sinkhorn semigroups in Section 4.
-3.1 Metrics and divergences
-3.1.1 Hilbert projective metric
-The Hilbert (projective) metric is defined, for any comparable measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$, by the formula
-$H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~\|\nu_{1}-\nu_{2}\|_{\sf tv},$ (31)
+3.1. Metrics and divergences
+3.1.1. Hilbert projective metric
+The Hilbert (projective) metric is defined, for any comparable measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(\mathbb{X})$, by the formula
+(3.1) $H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~\|\nu_{1}-\nu_{2}\|_{\sf tv},$
 where $\|\cdot\|$ and $\|\cdot\|_{\sf tv}$ denote the supremum and total variation norms, respectively, and $H(\nu_{1},\nu_{2})=\infty$ when the measures $\nu_{1},\nu_{2}$ are not comparable.
 The proof of the r.h.s.
-estimate in (31) is provided in [atar] (cf.
+estimate in (3.1) is provided in [atar] (cf.
 Lemma 1), see also [legland].
 For a more thorough discussion on Hilbert projective metrics we refer the reader to the recent review and self contained article on Hilbert projective metrics [cohen].
-3.1.2 Integral operator norms
-Given some functions $\varphi\in{\cal B}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\cal M}_{\varphi}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\cal M}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
+3.1.2. Integral operator norms
+Given some functions $\varphi\in{\mathcal{B}}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\mathcal{M}}_{\varphi}(\mathbb{X})\subset{\mathcal{M}}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\mathcal{M}}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
 […]
 $=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation measure}\quad|\nu|:=\nu_{+}+\nu_{-}.$
-In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\cal B}(\mathbb{X})$, which is given by
+In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\mathcal{B}}(\mathbb{X})$, which is given by
 $\mbox{\rm osc}_{\varphi}(f):=\sup_{x_{1},x_{2}}\frac{|f(x_{1})-f(x_{2})|}{\varphi(x_{1})+\varphi(x_{2})}\leq\|f\|_{\varphi}.$
-For a detailed proof of the equivalent formulations in the latter definition (32), we refer to Proposition 8.2.16 in [penev].
-The weighted Wasserstein formulation (7) of the $\varphi$-norm is a consequence of Kellerer’s extension of the Kantorovich-Rubinstein theorem [edwards2011kantorovich, kellerer1984duality, kellerer1982duality], see also the recent article [delgerber2025].
+For a detailed proof of the equivalent formulations in the latter definition (3.2), we refer to Proposition 8.2.16 in [penev].
+The weighted Wasserstein formulation (1.7) of the $\varphi$-norm is a consequence of Kellerer’s extension of the Kantorovich-Rubinstein theorem [edwards2011kantorovich, kellerer1984duality, kellerer1982duality], see also the recent article [delgerber2025].
 […]
-Recall that, given an arbitrary Polish space $S$, for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
-$\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\cal M}_{1}(S)~:~\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$ (33)
-Choosing $\varphi\geq 1$ and $\kappa>0$, for any $\nu\in{\cal M}_{b}(\mathbb{X})$ we also have the norm equivalence formulae
+(3.3) $\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\mathcal{M}}_{1}(S)~:~\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$
+Choosing $\varphi\geq 1$ and $\kappa>0$, for any $\nu\in{\mathcal{M}}_{b}(\mathbb{X})$ we also have the norm equivalence formulae
 […]
@@
 for some function $b(x)$ and $a,p>0$. In this case we have
-$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho(x_{1},x_{2}):=\frac{1}{2^{(p-1)_{+}}}~\|b(x_{1})-b(x_{2})\|^{p}$ (36)
-as well as the Wasserstein estimate (10).
-3.2 Some contraction coefficients
+(3.6) $\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho(x_{1},x_{2}):=\frac{1}{2^{(p-1)_{+}}}~\|b(x_{1})-b(x_{2})\|^{p}$
+as well as the Wasserstein estimate (1.10).
+3.2. Some contraction coefficients
 Let $\delta_{x}$ denote the Dirac delta measure located at $x$.
-Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$, we denote by $\imath_{x_{1},x_{2}}({\cal K})(y)$ the Radon-Nikodym function
-$\imath_{x_{1},x_{2}}({\cal K})(y):=\frac{d\delta_{x_{1}}{\cal K}}{d\delta_{x_{2}}{\cal K}}(y).$ (37)
+Whenever $\delta_{x_{1}}{\mathcal{K}}\ll\delta_{x_{2}}{\mathcal{K}}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$, we denote by $\imath_{x_{1},x_{2}}({\mathcal{K}})(y)$ the Radon-Nikodym function
+(3.7) $\imath_{x_{1},x_{2}}({\mathcal{K}})(y):=\frac{d\delta_{x_{1}}{\mathcal{K}}}{d\delta_{x_{2}}{\mathcal{K}}}(y).$
 We also consider the $[0,1]$-valued parameters
-$\jmath({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~\inf_{y\in\mathbb{Y}}~\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
-$\hbar({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (38)
+$\jmath({\mathcal{K}})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~\inf_{y\in\mathbb{Y}}~\imath_{x_{1},x_{2}}({\mathcal{K}})(y)\quad\text{and}$
+(3.8) $\hbar({\mathcal{K}})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\mathcal{K}})(y_{1})~\imath_{x_{2},x_{1}}({\mathcal{K}})(y_{2})\right)\geq\jmath({\mathcal{K}})^{2},$
 and we note that
-$\delta_{x_{1}}{\cal K}(dy_{1})~\delta_{x_{2}}{\cal K}(dy_{2})~\geq\hbar({\cal K})~\delta_{x_{2}}{\cal K}(dy_{1})~\delta_{x_{1}}{\cal K}(dy_{2}).$
+$\delta_{x_{1}}{\mathcal{K}}(dy_{1})~\delta_{x_{2}}{\mathcal{K}}(dy_{2})~\geq\hbar({\mathcal{K}})~\delta_{x_{2}}{\mathcal{K}}(dy_{1})~\delta_{x_{1}}{\mathcal{K}}(dy_{2}).$
 Integrating the above inequality w.r.t. the variable $y_{2}$ we readily see that
-$\delta_{x_{1}}{\cal K}(dy_{1})~\geq\hbar({\cal K})~\delta_{x_{2}}{\cal K}(dy_{1}),\quad\text{which readily yields}\quad\jmath({\cal K})\geq\hbar({\cal K})\geq\jmath({\cal K})^{2}.$ (39)
-3.2.1 Birkhoff contraction coefficient
-The Birkhoff contraction coefficient of a Markov transition $\cal K$} from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
-$\raisebox{2.0pt}{$\chi$}_{H}({\cal K}):=\sup\frac{H(\nu_{1}{\cal K},\nu_{2}{\cal K})}{H(\nu_{1},\nu_{2})}=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}},$ (40)
-where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B.
-Combining (39) and (40) we readily obtain the equivalent inequalities
-${\raisebox{2.0pt}{$\chi$}_{H}({\cal K})<1}\Longleftrightarrow\hbar({\cal K})>0\Longleftrightarrow\jmath({\cal K})>0.$
-We also emphasize that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}$ there is an interesting twist because the uniform minorization condition $\jmath({\cal K})>0$ is generally not satisfied.
+(3.9) $\delta_{x_{1}}{\mathcal{K}}(dy_{1})~\geq\hbar({\mathcal{K}})~\delta_{x_{2}}{\mathcal{K}}(dy_{1}),\quad\text{which readily yields}\quad\jmath({\mathcal{K}})\geq\hbar({\mathcal{K}})\geq\jmath({\mathcal{K}})^{2}.$
+3.2.1. Birkhoff contraction coefficient
+The Birkhoff contraction coefficient of a Markov transition $\mathcal{K}$} from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
+(3.10) $\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{K}}):=\sup\frac{H(\nu_{1}{\mathcal{K}},\nu_{2}{\mathcal{K}})}{H(\nu_{1},\nu_{2})}=\frac{1-\sqrt{\hbar({\mathcal{K}})}}{1+\sqrt{\hbar({\mathcal{K}})}},$
+where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B.
+Combining (3.9) and (3.10) we readily obtain the equivalent inequalities
+${\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{K}})<1}\Longleftrightarrow\hbar({\mathcal{K}})>0\Longleftrightarrow\jmath({\mathcal{K}})>0.$
+We also emphasize that even on totally bounded state spaces there is an interesting twist because the uniform minorization condition $\jmath({\mathcal{K}})>0$ is generally not satisfied.
 We illustrate this assertion with the semi-circle transition on the unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ defined by
-${\cal K}(x,dy)=\frac{1}{q(x)}~\sqrt{1-(x-y)^{2}}~1_{]0,1[}(y)~dy$ (41)
-for some normalizing constant $1/3\leq q(x)\leq 1$. In this case, choosing $x_{1}=1/(2n)=1-y=1-x_{2}$ in (37) we check that
-$0\leq\jmath({\cal K})\leq 3~\sqrt{1-(1-1/n)^{2}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0\Longrightarrow\jmath({\cal K})=0.$
-The entropic transport model (4) on the unit interval ]0,1[]0,1[ with reference transition (41) is discussed in Section 5.4. Note that in that case, the cost function is unbounded.
-3.2.2 Dobrushin contraction coefficients
-Consider functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ for any $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.
-The $(\varphi,\psi)$-Dobrushin coefficient $\raisebox{2.0pt}{$\chi$}_{\varphi,\psi}({\cal K})$ of a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by the norm operator
+(3.11) ${\mathcal{K}}(x,dy)=\frac{1}{q(x)}~\sqrt{1-(x-y)^{2}}~1_{]0,1[}(y)~dy$
+for some normalizing constant $1/3\leq q(x)\leq 1$. In this case, choosing $x_{1}=1/(2n)=1-y=1-x_{2}$ in (3.7) we check that
+$0\leq\jmath({\mathcal{K}})\leq 3~\sqrt{1-(1-1/n)^{2}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0\Longrightarrow\jmath({\mathcal{K}})=0.$
+The entropic transport model (1.4) on the unit interval ]0,1[]0,1[ with reference transition (3.11) is discussed in Section 5.4. Note that in that case, the cost function is unbounded.
+3.2.2. Dobrushin contraction coefficients
+Consider functions $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $\psi\in{\mathcal{B}}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ for any $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.
+The $(\varphi,\psi)$-Dobrushin coefficient $\raisebox{2.0pt}{$\chi$}_{\varphi,\psi}({\mathcal{K}})$ of a Markov transition ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by the norm operator
 […]
-When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\cal K})$ instead of $\chi_{\varphi,\varphi}({\cal K})$.
-Given a Markov transition ${\cal K}_{0}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a Markov transition ${\cal K}_{1}(x,dy)$ from $\mathbb{Y}$ into $\mathbb{X}$ we have
-$\chi_{\varphi}({\cal K}_{0}{\cal K}_{1})=\chi_{\varphi,\varphi}({\cal K}_{0}{\cal K}_{1})\leq\chi_{\varphi,\psi}({\cal K}_{0})~\chi_{\psi,\varphi}({\cal K}_{1}).$ (42)
-The terminology $\varphi$-Dobrushin coefficient comes from the fact that we recover the standard Dobrushin contraction coefficient $\raisebox{2.0pt}{$\chi$}({\cal K})$ by choosing constant functions.
+When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\mathcal{K}})$ instead of $\chi_{\varphi,\varphi}({\mathcal{K}})$.
+Given a Markov transition ${\mathcal{K}}_{0}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a Markov transition ${\mathcal{K}}_{1}(x,dy)$ from $\mathbb{Y}$ into $\mathbb{X}$ we have
+(3.12) $\chi_{\varphi}({\mathcal{K}}_{0}{\mathcal{K}}_{1})=\chi_{\varphi,\varphi}({\mathcal{K}}_{0}{\mathcal{K}}_{1})\leq\chi_{\varphi,\psi}({\mathcal{K}}_{0})~\chi_{\psi,\varphi}({\mathcal{K}}_{1}).$
+The terminology $\varphi$-Dobrushin coefficient comes from the fact that we recover the standard Dobrushin contraction coefficient $\raisebox{2.0pt}{$\chi$}({\mathcal{K}})$ by choosing constant functions.
 Indeed, if $\varphi=1=\psi$, then
-$\raisebox{2.0pt}{$\chi$}_{1,1}({\cal K})=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{\|\nu_{1}{\cal K}-\nu_{2}{\cal K}\|_{\sf tv}}{\|\nu_{1}-\nu_{2}\|_{\sf tv}}=:\raisebox{2.0pt}{$\chi$}({\cal K})=1-\varepsilon({\cal K}),$ (43)
+(3.13) $\raisebox{2.0pt}{$\chi$}_{1,1}({\mathcal{K}})=\sup_{\nu_{1},\nu_{2}\,\in{\mathcal{M}}_{1}(\mathbb{X})}\frac{\|\nu_{1}{\mathcal{K}}-\nu_{2}{\mathcal{K}}\|_{\sf tv}}{\|\nu_{1}-\nu_{2}\|_{\sf tv}}=:\raisebox{2.0pt}{$\chi$}({\mathcal{K}})=1-\varepsilon({\mathcal{K}}),$
 with the parameter
-$\varepsilon({\cal K}):=1-\sup_{(x_{1},x_{2})\,\in\mathbb{X}^{2}}{\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}}.$ (44)
-Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$ using the estimate (39) as well as (33) we readily find that
-$\varepsilon({\cal K})\geq\jmath({\cal K})\geq\hbar({\cal K})\quad\mbox{\rm and}\quad\raisebox{2.0pt}{$\chi$}({\cal K})\leq 1-\jmath({\cal K})\leq 1-\hbar({\cal K}).$ (45)
-3.2.3 Entropy contraction coefficient
-Given a convex function $\Phi$ satisfying (12), the Dobrushin $\Phi$-contraction coefficient $\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})$ of some Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
-$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K}):=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{D_{\Phi}(\nu_{1}{\cal K},\nu_{2}{\cal K})}{D_{\Phi}(\nu_{1},\nu_{2})},$ (46)
-where the supremum is taken over all probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ such that $D_{\Phi}(\nu_{1},\nu_{2})>0$.
-Here, again, the terminology $\Phi$-Dobrushin coefficient comes from the fact that choosing $\Phi(u,v)=\Phi_{0}(u,v)=\frac{1}{2}|u-v|$ we recover the standard Dobrushin contraction coefficient, i.e., $\raisebox{2.0pt}{$\chi$}_{\Phi_{0}}({\cal K})=\raisebox{2.0pt}{$\chi$}({\cal K})$.
-For any convex function $\Phi$ satisfying (12) we recall that
-$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})\leq\raisebox{2.0pt}{$\chi$}({\cal K}).$ (47)
-The estimate (47) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [cohen-0] for finite state spaces and [dm-03] for general measurable spaces).
+(3.14) $\varepsilon({\mathcal{K}}):=1-\sup_{(x_{1},x_{2})\,\in\mathbb{X}^{2}}{\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{K}}\|_{\sf tv}}.$
+Whenever $\delta_{x_{1}}{\mathcal{K}}\ll\delta_{x_{2}}{\mathcal{K}}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$ using the estimate (3.9) as well as (3.3) we readily find that
+(3.15) $\varepsilon({\mathcal{K}})\geq\jmath({\mathcal{K}})\geq\hbar({\mathcal{K}})\quad\mbox{\rm and}\quad\raisebox{2.0pt}{$\chi$}({\mathcal{K}})\leq 1-\jmath({\mathcal{K}})\leq 1-\hbar({\mathcal{K}}).$
+3.2.3. Entropy contraction coefficient
+Given a convex function $\Phi$ satisfying (1.12), the Dobrushin $\Phi$-contraction coefficient $\raisebox{2.0pt}{$\chi$}_{\Phi}({\mathcal{K}})$ of some Markov transition ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
+(3.16) $\raisebox{2.0pt}{$\chi$}_{\Phi}({\mathcal{K}}):=\sup_{\nu_{1},\nu_{2}\,\in{\mathcal{M}}_{1}(\mathbb{X})}\frac{D_{\Phi}(\nu_{1}{\mathcal{K}},\nu_{2}{\mathcal{K}})}{D_{\Phi}(\nu_{1},\nu_{2})},$
+where the supremum is taken over all probability measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(S)$ such that $D_{\Phi}(\nu_{1},\nu_{2})>0$.
+Here, again, the terminology $\Phi$-Dobrushin coefficient comes from the fact that choosing $\Phi(u,v)=\Phi_{0}(u,v)=\frac{1}{2}|u-v|$ we recover the standard Dobrushin contraction coefficient, i.e., $\raisebox{2.0pt}{$\chi$}_{\Phi_{0}}({\mathcal{K}})=\raisebox{2.0pt}{$\chi$}({\mathcal{K}})$.
+For any convex function $\Phi$ satisfying (1.12) we recall that
+(3.17) $\raisebox{2.0pt}{$\chi$}_{\Phi}({\mathcal{K}})\leq\raisebox{2.0pt}{$\chi$}({\mathcal{K}}).$
+The estimate (3.17) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [cohen-0] for finite state spaces and [dm-03] for general measurable spaces).
 For further details on the Dobrushin contraction coefficient we refer the reader to the article [dob-56], see also [dg-01, dm-03] and Chapter 4 in [dm-04].
-4 Sinkhorn contraction properties
-4.1 Strong mixing models
-As underlined in [fienberg, mosteller] in the context of finite state space models, Sinkhorn iterations of the form (16) preserve the cross product Radon-Nikodym ratio in (38) and, hence, for any $n\geq 0$ we have
-$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}).$ (48)
+4. Sinkhorn contraction properties
+This section investigates the contraction and stability properties of Sinkhorn iterations and their associated Gibbs-loop semigroups.
+We first recall the strong-mixing regime, in which uniform positivity of the reference kernel yields Hilbert-metric contractions and exponential convergence.
+We then move beyond this setting by introducing Lyapunov drift and local minorization conditions tailored to the forward-backward Sinkhorn transitions.
+These conditions allow us to establish weighted-norm contraction inequalities, from which we derive exponential convergence rates in total variation, weighted norms, Wasserstein distances, and relative entropy.
+The results provide a unified framework for analyzing Sinkhorn dynamics on non-compact state spaces and for treating models that fall outside the classical strong-mixing assumptions.
+4.1. Strong mixing models
+As underlined in [fienberg, mosteller] in the context of finite state space models, Sinkhorn iterations of the form (2.2) preserve the cross product Radon-Nikodym ratio in (3.8) and, hence, for any $n\geq 0$ we have
+(4.1) $\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{S}}_{0})=\hbar({\mathcal{Q}}).$
 This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [borwein, chen, deligiannidis, franklin].
-Recall that the Hilbert contraction coefficient of a Markov transition $\cal Q$} can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq.
-(40)), hence the exponential convergence of Sinkhorn iterations w.r.t.
-the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$.
+Recall that the Hilbert contraction coefficient of a Markov transition $\mathcal{Q}$} can be explicitly computed in terms of $\hbar({\mathcal{Q}})$ (see Eq.
+(3.10)), hence the exponential convergence of Sinkhorn iterations w.r.t.
+the Hilbert metric takes place if, and only if, $\hbar({\mathcal{Q}})>0$.
 An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
-From (39), let us observe that
-$\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})\geq\jmath({\cal Q})^{2}.$
-Thus, for any given convex function $\Phi$ satisfying (12), we can use (40) and (47) to show that
-$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
+From (3.9), let us observe that
+$\jmath({\mathcal{S}}_{n})\wedge\jmath({\mathcal{Q}})\geq\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{Q}})\geq\jmath({\mathcal{Q}})^{2}.$
+Thus, for any given convex function $\Phi$ satisfying (1.12), we can use (3.10) and (3.17) to show that
+$\hbar({\mathcal{Q}})>0\quad\text{if and only if}\quad\jmath({\mathcal{Q}})>0,$
 while
@@
 Theorem 4.1.
-For any convex function $\Phi$ satisfying (12) we have the contraction estimates
-$\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$ (50)
+For any convex function $\Phi$ satisfying (1.12) we have the contraction estimates
+(4.3) $\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})~D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2}~D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})~D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2}~D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$
 The above inequalities remain valid when we replace $(\nu_{V},\pi_{2n})$ by $(\lambda_{U},\pi_{2n+1})$.
-If we assume that $\jmath({\cal Q})>0$, then $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ and, in addition, there exists some $c>0$ such that
-$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~\raisebox{2.0pt}{$\chi$}({\cal S})^{2n}\quad\text{for any $n\geq 0$.}$ (51)
+If we assume that $\jmath({\mathcal{Q}})>0$, then $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})<1$ and, in addition, there exists some $c>0$ such that
+(4.4) $\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2n}\quad\text{for any $n\geq 0$.}$
 Proof.
-The estimates (50) are a direct consequence of (47) combined with the Markov transport formulae
-$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\cal S}_{2(n+1)},\pi_{2n+1}{\cal S}_{2(n+1)}),\quad\text{and}$
-$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
-which are themselves easily obtained from expressions (16) and (22). The proof of (51) follows exactly the same arguments as the proof of the inequality (71) in Theorem 4.13, hence it is skipped.
+The estimates (4.3) are a direct consequence of (3.17) combined with the Markov transport formulae
+$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\mathcal{S}}_{2(n+1)},\pi_{2n+1}{\mathcal{S}}_{2(n+1)}),\quad\text{and}$
+$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\mathcal{S}}_{2n+1},\nu_{V}{\mathcal{S}}_{2n+1}),$
+which are themselves easily obtained from expressions (2.2) and (2.8). The proof of (4.4) follows exactly the same arguments as the proof of the inequality (4.24) in Theorem 4.8, hence it is skipped.
 The following result can be obtained from the series expansion given by Corollary 4.15.
 Corollary 4.2.
-Under the assumptions of Theorem 4.1, we have the entropy exponential decay estimates
-$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\raisebox{2.0pt}{$\chi$}({\cal S})^{2n},\quad\text{for any $n\geq 0$}.$
-Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [dg-01].
+If we assume that $\jmath({\mathcal{Q}})>0$, then $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})<1$ and, in addition, there exists some $c>0$ such that
+$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})\leq c~\raisebox{2.0pt}{$\chi$}({\mathcal{S}})^{2n},\quad\text{for any $n\geq 0$}.$
+Several examples satisfying condition $\jmath({\mathcal{Q}})>0$ are discussed in Section 2.4 in [dg-01].
 For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself.
-The choice ${\cal Q}(x,dy)~\propto~e^{-\delta\,\|y-b(x)\|}~dx$ yields
-$\jmath({\cal Q})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
-In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [aronson], Nash [nash] and Varopoulos [varopoulos] on Gaussian estimates for heat kernels on manifolds.
+The choice ${\mathcal{Q}}(x,dy)~\propto~e^{-\delta\,\|y-b(x)\|}~dx$ yields
+$\jmath({\mathcal{Q}})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
+In continuous time settings, condition $\jmath({\mathcal{Q}})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [aronson], Nash [nash] and Varopoulos [varopoulos] on Gaussian estimates for heat kernels on manifolds.
 Nevertheless, the above condition is not met for Gaussian transitions.
-4.2 Gibbs-Loop Sinkhorn semigroups
-The $r$-sub-level set of a positive function $\varphi\in{\cal B}(\mathbb{X})$ is denoted as
-$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~:~\varphi(x)\leq r\right\}.$ (52)
-When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$.
-Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [dpa] and references therein).
-In the context of Gibbs-loop semigroups (23), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
-Condition 4.3 (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
-There exist some parameters $\epsilon\in]0,1[$ some constant $c>0$ and some integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and for any $n\geq n_{0}$ we have the estimates
-${\cal S}_{2n}^{\circ}(\psi)\leq\epsilon~\psi+c\quad\mbox{and}\quad{\cal S}_{2n+1}^{\circ}(\varphi)\leq\epsilon~\varphi+c.$ (53)
-There also exists some $r_{0}\geq 1$ and some function
-[…]
-such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
-$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{2n+1}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}^{\circ}_{2n}\|_{\sf tv}\leq 1-\varepsilon(r).$ (54)
-The inequality (54) is satisfied if, and only if, there exists some ${\cal X}\in{\cal M}_{1}(\mathbb{X})$ (respectively, ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$) that may depend on $(n,r,(x_{1},x_{2}))$ (respectively, on $(n,r,(y_{1},y_{2}))$) satisfying, for any $i\in\{1,2\}$, the minorization conditions
-${\cal S}^{\circ}_{2n}(y_{i},dy)\geq\varepsilon(r)~{\cal Y}(dy)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}(x_{i},dx)\geq\varepsilon(r)~{\cal X}(dx).$ (55)
-The above assertion is rather well known and it can be checked using (33).
-Note that the drift condition (53) is trivially met for the unit functions $\varphi=1=\psi$ but in this case for any $r\geq 1$ we have $C_{\varphi}(r)=\mathbb{X}$ and $C_{\psi}(r)=\mathbb{Y}$.
-In that case, the inequalities in (55) are uniform minorization conditions over $\mathbb{X}$ and $\mathbb{Y}$.
-Let ${\cal B}_{\infty}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the subalgebra of locally bounded and uniformly positive functions $\varphi$ that grow at infinity; that is, $\sup_{C}\varphi<\infty$ for any compact set $C\subset\mathbb{X}$, and for any
-$r\geq\varphi_{\star}:=\inf_{\mathbb{X}}\varphi>0$ (56)
-the $r$-sub-level set $C_{\varphi}(r)$ is a non empty compact subset.
-We denote by ${\cal B}_{0}(\mathbb{X}):=\{1/\varphi~:~\varphi\in{\cal B}_{\infty}(\mathbb{X})\}\subset{\cal B}_{b}(\mathbb{X})$ the sub-algebra of (bounded) positive functions, locally lower bounded and that vanish at infinity.
-When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ the $r$-sub-level sets $C_{\varphi}(r)$ and $C_{\psi}(r)$ are compact and the inequality (55) yields local minorization conditions.
-The assumption $\varphi\wedge\psi\geq 1$ is only made to simplify the presentation.
-Indeed, for any function $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ satisfying (53), the rescaled functions $\varphi/\varphi_{\star}\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi/\psi_{\star}\in{\cal B}_{\infty}(\mathbb{Y})$ also satisfy (53).
-The drift conditions in (53) ensure that for any $n\geq n_{0}$ we have
-${\cal S}^{\circ}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\varphi}(\mathbb{X}).$
-Whenever Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, it can be easily shown that
-$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.\$(57)
-These estimates are direct consequences of the drift conditions (53). A detailed proof is provided in Section E on page E.
-4.3 Weighted norm contractions
-The drift conditions (53) combined with the local minorization conditions (54) provide a way to control the contraction coefficients of Sinkhorn transitions.
-More precisely, we have the following estimate for the contraction coefficients of the forward-backward Markov transitions ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$.
+4.2. Gibbs-Loop Sinkhorn semigroups
+For a positive function $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $r>0$, we define its $r$-sublevel set by
+(4.5) $C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~:~\varphi(x)\leq r\right\}.$
+When there is no ambiguity, we write $C_{\varphi}(r)=\{\varphi\leq r\}$.
+Lyapunov functions combined with local minorization conditions are classical tools in the analysis of time-inhomogeneous Markov chains; see, for instance, [dpa] and the references therein.
+In the context of Gibbs-loop semigroups (2.9), such conditions are formulated using Lyapunov-type functions $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $\psi\in{\mathcal{B}}(\mathbb{Y})$ satisfying $\varphi\wedge\psi\geq 1$, as described below.
+Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$: There exist $\epsilon\in(0,1)$, a constant $c>0$, and an integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and, for all $n\geq n_{0}$,
+(4.6) ${\mathcal{S}}_{2n}^{\circ}(\psi)\leq\epsilon,\psi+c,\qquad{\mathcal{S}}_{2n+1}^{\circ}(\varphi)\leq\epsilon,\varphi+c.$
+Moreover, there exist $r_{0}\geq 1$ and a function
+$\varepsilon:[r_{0},\infty[\rightarrow]0,1]$
+such that, for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$, and $n\geq n_{0}$,
+(4.7) $\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}^{\circ}_{2n+1}\|{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\mathcal{S}}^{\circ}_{2n}\|{\sf tv}\leq 1-\varepsilon(r).$
+Condition (4.7) holds if and only if there exist probability measures ${\mathcal{X}}\in{\mathcal{M}}_{1}(\mathbb{X})$ and ${\mathcal{Y}}\in{\mathcal{M}}_{1}(\mathbb{Y})$, possibly depending on $(n,r,(x_{1},x_{2}))$ and $(n,r,(y_{1},y_{2}))$, respectively, such that, for $i\in{1,2}$,
+(4.8) ${\mathcal{S}}^{\circ}_{2n}(y_{i},dy)\geq\varepsilon(r)~{\mathcal{Y}}(dy)\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}^{\circ}(x_{i},dx)\geq\varepsilon(r)~{\mathcal{X}}(dx).$
+The drift condition (4.6) is trivially satisfied for the constant functions $\varphi=1=\psi$.
+In this case, $C_{\varphi}(r)=\mathbb{X}$ and $C_{\psi}(r)=\mathbb{Y}$ for all $r\geq 1$, and (4.8) reduces to a uniform minorization condition over the whole state spaces.
+Let ${\mathcal{B}}_{\infty}(\mathbb{X})\subset{\mathcal{B}}(\mathbb{X})$ denote the subalgebra of locally bounded, uniformly positive functions $\varphi$ that grow at infinity, in the sense that $\sup_{C}\varphi<\infty$ for any compact set $C\subset\mathbb{X}$ and that, for any
+(4.9) $r\geq\varphi_{\star}:=\inf_{\mathbb{X}}\varphi>0,$
+the sublevel set $C_{\varphi}(r)$ is nonempty and compact. We further define
+${\mathcal{B}}_{0}(\mathbb{X}):=\{1/\varphi~:~\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})\}\subset{\mathcal{B}}_{b}(\mathbb{X})$
+the algebra of bounded, positive functions that are locally lower bounded and vanish at infinity.
+If $\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})$ and $\psi\in{\mathcal{B}}_{\infty}(\mathbb{Y})$, then the sets $C_{\varphi}(r)$ and $C_{\psi}(r)$ are compact, and (4.8) yields genuine local minorization conditions.
+The assumption $\varphi\wedge\psi\geq 1$ is made solely for notational convenience.
+Indeed, if $\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})$ and $\psi\in{\mathcal{B}}_{\infty}(\mathbb{Y})$ satisfy (4.6), then the rescaled functions $\varphi/\varphi_{\star}$ and $\psi/\psi_{\star}$ also satisfy the same drift inequalities.
+The drift conditions (4.6) imply that, for all $n\geq n_{0}$,
+${\mathcal{S}}^{\circ}_{2n}\left({\mathcal{B}}_{\psi}(\mathbb{Y})\right)\subset{\mathcal{B}}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}^{\circ}\left({\mathcal{B}}_{\varphi}(\mathbb{X})\right)\subset{\mathcal{B}}_{\varphi}(\mathbb{X}).$
+Moreover, whenever ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds,
+(4.10) $\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.\$
+These bounds follow directly from the drift inequalities (4.6); a detailed proof is given in Section E.
+4.3. Weighted norm contractions
+Combining the drift conditions (4.6) with the local minorization conditions (4.7) allows us to control the contraction coefficients of the Sinkhorn transitions.
+More precisely, we obtain the following uniform contraction estimate for the forward–backward Markov kernels ${\mathcal{S}}^{\circ}_{2n}$ and ${\mathcal{S}}^{\circ}_{2n+1}$.
 Theorem 4.3.
-If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied some integer $n_{0}\geq 1$, then there exist parameters $\kappa>0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (53) and on the function $\varepsilon(r)$ in (54) and such that
-$\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\cal S}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (58)
-Theorem 4.4 is a straightforward consequence of Theorem C.2 in Appendix C.
-The contraction estimate (58) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $n_{0}\leq p\leq n$ we readily obtain
+Assume that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds for some $n_{0}\geq 1$.
+Then there exist parameters $\kappa>0$ and $\rho\in[0,1[$ depending only on $(\epsilon,c)$ in (4.6) and on $\varepsilon(r)$ in (4.7) and such that
+(4.11) $\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\mathcal{S}}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\mathcal{S}}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
+Theorem 4.3 is an immediate consequence of Theorem C.1 in Appendix C.
+The contraction estimate (4.11) ensures the exponential decay of Sinkhorn semigroups.
+For instance, for any $\mu_{1},\mu_{2}\in{\mathcal{M}}_{1,\psi}(\mathbb{Y})$ and $n_{0}\leq p\leq n$ we readily obtain
 […]
@@
 […]
-with the parameters $\rho\in[0,1[$ and $\kappa>0$ as in (58).
+with the parameters $\rho\in[0,1[$ and $\kappa>0$ as in (4.11).
 Choosing $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [persi, saloff-zuniga, saloff-zuniga-2]
-$\|{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{2(n_{0}-1),2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{1+n-n_{0}}~\left(\psi_{\kappa}(y_{1})+\psi_{\kappa}(y_{2})\right).$
+$\|{\mathcal{S}}^{\circ}_{2(n_{0}-1),2n}(y_{1},\mbox{\LARGE.})-{\mathcal{S}}^{\circ}_{2(n_{0}-1),2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{1+n-n_{0}}~\left(\psi_{\kappa}(y_{1})+\psi_{\kappa}(y_{2})\right).$
 Remark 4.5.
-The proximal sampler discussed in Remark 2.1 has a Markov transition ${\cal S}^{\circ}_{1}$ with invariant measure $\lambda_{U}$.
+The proximal sampler discussed in Remark 2.1 has a Markov transition ${\mathcal{S}}^{\circ}_{1}$ with invariant measure $\lambda_{U}$.
 The weighted-norm contraction properties of these models are simply based on the existence of a Lyapunov function $\varphi$ satisfying the r.h.s.
-condition in (53) with $n=0$.
-As in (54) we also require a local minorization condition $\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{1}\|_{\sf tv}\leq 1-\varepsilon(r)$ on the sub-level sets $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$ of the Lyapunov function to ensure that $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{1})<1$, for some parameter $\kappa\geq 0$ (see for instance Theorem C.2).
-Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may not be easy to test – except in the case of Gaussian and linear models.
+condition in (4.6) with $n=0$.
+As in (4.7) we also require a local minorization condition $\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}^{\circ}_{1}\|_{\sf tv}\leq 1-\varepsilon(r)$ on the sub-level sets $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$ of the Lyapunov function to ensure that $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\mathcal{S}}^{\circ}_{1})<1$, for some parameter $\kappa\geq 0$ (see for instance Theorem C.1).
+Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\mathcal{S}}^{\circ}_{2n}$ and ${\mathcal{S}}^{\circ}_{2n+1}$, which may not be easy to test – except in the case of Gaussian and linear models.
 Several Lyapunov functions in this context are designed in Section 5.8.
-A stronger (but easier to test) condition can be provided in terms of the transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
-Condition 4.7 (${\cal D}_{\varphi,\psi}({\cal S})$).
-There exists some parameter $\epsilon\in]0,1[$, some constant $c>0$ and some integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$ we have the estimates
-${\cal S}_{2n}(\psi)\leq\epsilon~\varphi+c\quad\mbox{and}\quad{\cal S}_{2n+1}(\varphi)\leq\epsilon~\psi+c.$ (60)
-There also exists some $r_{0}\geq 1$ and some function
-[…]
-such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
-$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (61)
-When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, for any $n\geq n_{0}$ we readily see that
-${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y}).$
+A stronger (but easier to test) condition can be provided in terms of the transitions ${\mathcal{S}}_{2n}$ and ${\mathcal{S}}_{2n+1}$.
+Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ There exists $\epsilon\in]0,1[$, a constant $c>0$ and an integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$,
+(4.13) ${\mathcal{S}}_{2n}(\psi)\leq\epsilon~\varphi+c,\qquad{\mathcal{S}}_{2n+1}(\varphi)\leq\epsilon~\psi+c.$
+Moreover, there exists some $r_{0}\geq 1$ and a function
+$\varepsilon:[r_{0},\infty[\rightarrow]0,1]$
+such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$,
+(4.14) $\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\mathcal{S}}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$
+When Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is met, for any $n\geq n_{0}$ we readily see that
+${\mathcal{S}}_{2n}\left({\mathcal{B}}_{\psi}(\mathbb{Y})\right)\subset{\mathcal{B}}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}\left({\mathcal{B}}_{\varphi}(\mathbb{X})\right)\subset{\mathcal{B}}_{\psi}(\mathbb{Y}).$
 In addition, we have
-$\pi_{2n+1}(\varphi)=\nu_{V}{\cal S}_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c$ (62)
+(4.15) $\pi_{2n+1}(\varphi)=\nu_{V}{\mathcal{S}}_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c$
 […]
 Theorem 4.6.
-Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds as soon as condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied and the estimates (59) hold for some $n_{0}\geq 1$ and some parameters $\rho\in[0,1[$ and $\kappa>0$.
-When ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied there also exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (60) and on the function $\varepsilon(r)$ in (61) and such that, for any $n\geq n_{0}$,
-$\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa},\psi_{\kappa}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa},\varphi_{\kappa}}({\cal S}_{2n+1})<\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
+We have the following:
+  • Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds as soon as condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied and the estimates (4.12) hold for some $n_{0}\geq 1$ and some parameters $\rho\in[0,1[$ and $\kappa>0$.
+  • When ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (4.13) and on the function $\varepsilon(r)$ in (4.14) and such that, for any $n\geq n_{0}$,
+    $\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa},\psi_{\kappa}}({\mathcal{S}}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa},\varphi_{\kappa}}({\mathcal{S}}_{2n+1})<\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
 A proof of Theorem 4.6 is provided in Appendix C (on page C).
-Let us emphasize that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is stronger than Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$.
-In particular, the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ can be stable even when the Sinkhorn transitions ${\cal S}_{n}$ are unstable.
-Intuitively, Gibbs-loop transitions go forward and backward towards their invariant measures (22) while the Sinkhorn transitions in (16) transport asymptotically a target marginal measure to the other.
-For instance, by Lemma 4.9 in [adm-24] (see also Section A) the Gibbs-loop transitions ${\cal S}^{\circ}_{n}$ of the Gaussian Sinkhorn algorithm are always stable.
-However, in the context of regularized Gaussian entropic problems on a time horizon $t$ (cf.
-Section 3.4 in [adm-24]) the Sinkhorn transitions ${\cal S}_{n}$ may be unstable when the time horizon is not sufficiently large (by Remarks 3.9 and 4.6 in [adm-24]).
-4.4 A Lyapunov stability theorem
-Regularity conditions
-This section presents some sufficient conditions for ${\cal D}_{\varphi,\psi}({\cal S})$, in terms of the potential functions $(U,V,W)$.
+Let us emphasize that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is stronger than Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$.
+In particular, the Gibbs-loop transitions ${\mathcal{S}}^{\circ}_{n}$ may be stable even in situations where the Sinkhorn transitions ${\mathcal{S}}_{n}$ are unstable.
+Intuitively, Gibbs-loop transitions alternate between forward and backward steps that steer the dynamics toward their invariant measures (2.8), whereas the Sinkhorn transitions in (2.2) asymptotically transport one target marginal measure to the other.
+As an illustration, Lemma 4.9 in [adm-24] (see also Section A) shows that the Gibbs-loop transitions ${\mathcal{S}}^{\circ}_{n}$ of the Gaussian Sinkhorn algorithm are always stable.
+By contrast, in the setting of regularized Gaussian entropic problems over a finite time horizon $t$ (see Section 3.4 in [adm-24]), the Sinkhorn transitions ${\mathcal{S}}_{n}$ may become unstable when the time horizon is not sufficiently large; see Remarks 3.9 and 4.6 in [adm-24].
+4.4. A Lyapunov stability theorem
+4.4.1. Regularity conditions
+This section presents some sufficient conditions for ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$, in terms of the potential functions $(U,V,W)$.
 These simple conditions enable us to obtain geometric rates for the convergence of the Sinkhorn iterations for a variety of models, including polynomial growth models (Section 5.1) and heavy tailed models (Section 5.2), as well as boundary states models including Beta and Weibull distributions (Section 5.3 and Section 5.4).
@@
 These conditions ensure there exists some $W_{\star}\in\mathbb{R}$ such that
-$W(x,y)\geq W_{\star}\quad\mbox{\rm for $(\lambda\otimes\nu)$-almost every $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.}$ (63)
+(4.16) $W(x,y)\geq W_{\star}\quad\mbox{\rm for $(\lambda\otimes\nu)$-almost every $(x,y)\in(\mathbb{X}\times\mathbb{Y})$.}$
 In addition, $(U,V)$ are bounded below by a real number so that $e^{-U}$ and $e^{-V}$ are uniformly bounded.
@@
 Also note that for any $\delta>0$ we have
-$(e^{\delta U},e^{\delta V})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
+$(e^{\delta U},e^{\delta V})\in({\mathcal{B}}_{\infty}(\mathbb{X})\times{\mathcal{B}}_{\infty}(\mathbb{Y})).$
 Consider the following conditions:
-Condition 4.9 (${\cal H}_{\delta}(U,V)$).
-There is some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ defined in (27) are locally bounded with compact sub-level sets and we have
-$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$ (64)
-Note that (63) and (64) together imply the inequalities
+Condition ${\mathcal{H}}_{\delta}(U,V)$: There is some $\delta\in]0,1[$ such that the functions $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ defined in (2.13) are locally bounded with compact sub-level sets and we have
+(4.17) $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$
+Note that (4.16) and (4.17) together imply the inequalities
 $\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)\leq e^{-W_{\star}}~\left(\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\right)<\infty.$
 This yields the exponential moment estimates
-${\cal H}_{\delta}(U,V)\Longrightarrow\lambda_{(1-\delta)U}\left(1\right)=\lambda_{U}\left(\exp{(\delta U)}\right)<\infty\quad\mbox{\rm and}\quad\nu_{(1-\delta)V}(1)=\nu_{V}\left(\exp{(\delta V)}\right).$
-In particular all absolute moments $\lambda_{U}(|U|^{p})\vee\nu_{V}(|V|^{p})<\infty$, with $p\geq 0$ are finite. Given the reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ also have
-${\cal H}_{\delta}(U,V)\Longrightarrow\mbox{\rm Ent}((\lambda_{U}\otimes\nu_{V})~|~{\cal P}_{0})=(\lambda_{U}\otimes\nu_{V})(W)-\nu_{V}(V)<\infty.$ (65)
+${\mathcal{H}}_{\delta}(U,V)\Longrightarrow\lambda_{(1-\delta)U}\left(1\right)=\lambda_{U}\left(\exp{(\delta U)}\right)<\infty\quad\mbox{\rm and}\quad\nu_{(1-\delta)V}(1)=\nu_{V}\left(\exp{(\delta V)}\right).$
+In particular all absolute moments $\lambda_{U}(|U|^{p})\vee\nu_{V}(|V|^{p})<\infty$, with $p\geq 0$ are finite.
+Given the reference measure ${\mathcal{P}}_{0}=(\lambda_{U}\times{\mathcal{Q}})$ also have
+(4.18) ${\mathcal{H}}_{\delta}(U,V)\Longrightarrow\mbox{\rm Ent}((\lambda_{U}\otimes\nu_{V})~|~{\mathcal{P}}_{0})=(\lambda_{U}\otimes\nu_{V})(W)-\nu_{V}(V)<\infty.$
 To check this claim, note that
 $(\lambda_{U}\otimes\nu_{V})(W)$ $=$ $\lambda_{U}(W^{V})=W_{\star}+\lambda_{U}(W^{V}-W_{\star})$
-$\leq$ $W_{\star}+\lambda_{U}(e^{W^{V}-W_{\star}})=W_{\star}+e^{-W_{\star}}~\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}).$
+$\leq$ $W_{\star}+\lambda_{U}(e^{W^{V}-W_{\star}})=W_{\star}+e^{-W_{\star}}~\lambda_{(1-\delta)U}(e^{-{\mathcal{U}}_{\delta}}).$
 On the other hand, for any $r>0$ we have
-$\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}})$ $\leq$ $e^{-r}~\lambda_{(1-\delta)U}(1)+\lambda_{(1-\delta)U}(e^{-{\cal U}_{\delta}}~1_{{\cal U}_{\delta}\leq r})$
+$\lambda_{(1-\delta)U}(e^{-{\mathcal{U}}_{\delta}})$ $\leq$ $e^{-r}~\lambda_{(1-\delta)U}(1)+\lambda_{(1-\delta)U}(e^{-{\mathcal{U}}_{\delta}}~1_{{\mathcal{U}}_{\delta}\leq r})$
 […]
-Condition 4.10 (${\cal H}^{\prime}_{\delta}(U,V)$).
-There is some $\delta\in]0,1/2[$ such that ${\cal U}_{\delta}$ and ${\cal V}_{\delta}$ are locally bounded with compact sub-level sets and we have
-$\lambda_{(1-2\delta)U}\left(1\right)\vee\nu_{(1-2\delta)V}\left(1\right)<\infty.$ (66)
-We have
-${\cal H}^{\prime}_{\delta}(U,V)\Longrightarrow{\cal H}_{\delta}(U,V).$ (67)
-To verify this claim, note that the regularity condition on $({\cal U}_{\delta},{\cal V}_{\delta})$ ensures that
-$(e^{{\cal U}_{\delta}},e^{{\cal V}_{\delta}})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
-When ${\cal U}_{\delta}$ is locally bounded with compact sub-level sets $C_{{\cal U}}(r):=\{{\cal U}_{\delta}\leq r\}$ for any $0<\delta<1/2$ we have
-$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $=$ $\lambda_{(1-2\delta)U}\left(e^{-{\cal U}_{\delta}}\right)$
-$\leq$ $\left(e^{-r}+e^{c_{\delta}(r)}\right)\lambda_{(1-2\delta)U}\left(1\right)\quad\mbox{\rm with}\quad c_{\delta}(r):=\sup_{x\in C_{{\cal U}}(r)}|{\cal U}_{\delta}(x)|.$
+Condition ${\mathcal{H}}^{\prime}_{\delta}(U,V)$: There is some $\delta\in]0,1/2[$ such that ${\mathcal{U}}_{\delta}$ and ${\mathcal{V}}_{\delta}$ are locally bounded with compact sub-level sets and we have
+(4.19) $\lambda_{(1-2\delta)U}\left(1\right)\vee\nu_{(1-2\delta)V}\left(1\right)<\infty.$
+Let us remark that
+(4.20) ${\mathcal{H}}^{\prime}_{\delta}(U,V)\Longrightarrow{\mathcal{H}}_{\delta}(U,V).$
+To verify this claim, note that the regularity condition on $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ ensures that
+$(e^{{\mathcal{U}}_{\delta}},e^{{\mathcal{V}}_{\delta}})\in({\mathcal{B}}_{\infty}(\mathbb{X})\times{\mathcal{B}}_{\infty}(\mathbb{Y})).$
+When ${\mathcal{U}}_{\delta}$ is locally bounded with compact sub-level sets $C_{{\mathcal{U}}}(r):=\{{\mathcal{U}}_{\delta}\leq r\}$ for any $0<\delta<1/2$ we have
+$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $=$ $\lambda_{(1-2\delta)U}\left(e^{-{\mathcal{U}}_{\delta}}\right)$
+$\leq$ $\left(e^{-r}+e^{c_{\delta}(r)}\right)\lambda_{(1-2\delta)U}\left(1\right)\quad\mbox{\rm with}\quad c_{\delta}(r):=\sup_{x\in C_{{\mathcal{U}}}(r)}|{\mathcal{U}}_{\delta}(x)|.$
 In this context, we readily see that
 […]
-This ends the proof of (67).
+This ends the proof of (4.20).
 Remark 4.7.
-Whenever $\|W\|<\infty$ we have $\|W_{U}\|\vee\|W^{V}\|<\infty$. In this case ${\cal H}_{\delta}(U,V)$ is satisfied as soon as there is some $\delta\in]0,1[$ such that
-$\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)<\infty.$ (69)
-Remark 4.12.
-Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions and we have $\delta U-W^{V}\geq{\cal U}_{\delta}^{-}$ and $\delta V-W_{U}\geq{\cal V}_{\delta}^{-}$, for some functions $({\cal U}_{\delta}^{-},{\cal V}_{\delta}^{-})$ with compact level sets.
-In this case, the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ are locally bounded with compact level sets.
-A Lyapunov stability theorem
+Whenever $\|W\|<\infty$ we have $\|W_{U}\|\vee\|W^{V}\|<\infty$. In this case ${\mathcal{H}}_{\delta}(U,V)$ is satisfied as soon as there is some $\delta\in]0,1[$ such that
+(4.22) $\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)<\infty.$
+4.4.2. A Lyapunov stability theorem
 Theorem 4.8.
-If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions
-$(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})\quad\mbox{with}\quad(U_{\star},V_{\star}):=(\inf_{\mathbb{X}}U,\inf_{\mathbb{Y}}V)\in(\mathbb{R}\times\mathbb{R}).$ (70)
+If Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions
+(4.23) $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})\quad\mbox{with}\quad(U_{\star},V_{\star}):=(\inf_{\mathbb{X}}U,\inf_{\mathbb{Y}}V)\in(\mathbb{R}\times\mathbb{R}).$
 In addition, there exist some parameter $\rho\in[0,1[$ and some constant $c>0$ such that
-$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c~\rho^{2n}\quad\text{for any $n\geq 1$.}$ (71)
+(4.24) $\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c~\rho^{2n}\quad\text{for any $n\geq 1$.}$
 A proof of Theorem 4.8 is provided in Appendix D.4.
-The above theorem combined with the estimates (62) ensures that
+The above theorem combined with the estimates (4.15) ensures that
 $\sup_{n\geq n_{0}}\pi_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c=\epsilon~e^{-\delta V_{\star}}~\nu_{(1-\delta)V}(1)+c$
 $\sup_{n\geq n_{0}}\pi_{2n}(\psi)$ $\leq$ $\epsilon~\lambda_{U}(\varphi)+c=\epsilon~e^{-\delta U_{\star}}~\lambda_{(1-\delta)U}(1)+c$
-with the parameters $(\epsilon,c)$ as in (60).
-The rescaling of the potential functions (70) is only made to have $\varphi\wedge\psi\geq 1$.
+with the parameters $(\epsilon,c)$ as in (4.13).
+The rescaling of the potential functions (4.23) is only made to have $\varphi\wedge\psi\geq 1$.
 Combining Theorem 4.8 with Theorem 4.6 we obtain the following corollary.
 Corollary 4.9.
-Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
 In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
 with the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$.
-The r.h.s. term in (72) can be estimate with (62). Also note that
+The r.h.s. term in (4.25) can be estimate with (4.15). Also note that
 […]
@@
 Corollary 4.10.
-Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
 In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
 with the functions $(\psi,\varphi)$ as in Corollary 4.9.
-4.5 Wasserstein exponential decays
-Assume that the Lyapunov functions $(\psi,\varphi)$ defined in (70) satisfy the estimates
-$e^{\delta U(x_{1})}+e^{\delta U(x_{2})}\geq\varrho_{\varphi}(x_{1},x_{2})\quad\mbox{\rm and}\quad e^{\delta V(y_{1})}+e^{\delta V(y_{2})}\geq\varrho_{\psi}(y_{1},y_{2})$ (74)
-for some semi-metric $\varrho_{\varphi},\varrho_{\psi}$. In this case, by (10) there exists $c_{u},c_{v}>0$ such that
+Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions and we have
+$\delta U-W^{V}\geq{\mathcal{U}}_{\delta}^{-}\quad\mbox{\rm and}\quad\delta V-W_{U}\geq{\mathcal{V}}_{\delta}^{-},$
+for some functions $({\mathcal{U}}_{\delta}^{-},{\mathcal{V}}_{\delta}^{-})$ with compact level sets.
+In this case, the functions $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ are locally bounded with compact level sets.
+As underlined in [ajay-dp-25], using (C.7) and Proposition D.9, the continuity condition in the above assertion can be relaxed.
+More precisely, Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$, Theorem 4.8 as well as Corollaries  4.9 and 4.10 hold when $({\mathcal{U}}_{\delta}^{-},{\mathcal{V}}_{\delta}^{-})$ are locally bounded with compact sub-level sets.
+Assume we have
+$W(x,y)\leq a_{W}(x)+b_{W}(y)$
+for some locally bounded functions $(a_{W},b_{W})$ such that $\lambda_{U}(a_{W})\vee\nu_{V}(b_{W})<\infty$. Note that
+${\mathcal{U}}^{-}_{\delta}(x):=\delta U(x)-a_{W}(x)-\nu_{V}(b_{V})\leq{\mathcal{U}}_{\delta}(x)\leq\delta U(x)-W_{\star}.$
+with $W_{\star}$ as in (4.16).
+This shows that $({\mathcal{U}}_{\delta},{\mathcal{U}}^{-}_{\delta})$ and similarly $({\mathcal{V}}_{\delta},{\mathcal{V}}^{-}_{\delta})$ with ${\mathcal{V}}^{-}_{\delta}:=\delta V-b_{W}$ are locally bounded.
+Moreover $({\mathcal{U}}^{-}_{\delta},{\mathcal{V}}^{-}_{\delta})$ have compact compact sub-level sets if and only if $(\delta U-a_{W},\delta V-b_{W})$ have compact compact sub-level sets.
+4.5. Wasserstein exponential decays
+Assume that the Lyapunov functions $(\psi,\varphi)$ defined in (4.23) satisfy the estimates
+(4.27) $e^{\delta U(x_{1})}+e^{\delta U(x_{2})}\geq\varrho_{\varphi}(x_{1},x_{2})\quad\mbox{\rm and}\quad e^{\delta V(y_{1})}+e^{\delta V(y_{2})}\geq\varrho_{\psi}(y_{1},y_{2})$
+for some semi-metric $\varrho_{\varphi},\varrho_{\psi}$. In this case, by (1.10) there exists $c_{u},c_{v}>0$ such that
 […]
-with the Wasserstein semi-metrics $\mathbb{W}_{\varrho_{\varphi}},\mathbb{W}_{\varrho_{\psi}}$ defined in (8). This yields the following Wasserstein estimates.
+with the Wasserstein semi-metrics $\mathbb{W}_{\varrho_{\varphi}},\mathbb{W}_{\varrho_{\psi}}$ defined in (1.8). This yields the following Wasserstein estimates.
 Corollary 4.11.
-Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
-In addition, the Lyapunov functions $(e^{\delta U},e^{\delta V})$ satisfy (74) for some semi-metric $(\varrho_{\varphi},\varrho_{\psi})$.
+Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+In addition, the Lyapunov functions $(e^{\delta U},e^{\delta V})$ satisfy (4.27) for some semi-metric $(\varrho_{\varphi},\varrho_{\psi})$.
 In this situation, there exists some $n_{0}\geq 0$ and $c_{u},c_{v}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
 […]
-In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
+In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (4.25) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (4.23).
 Assume there exists some semi-metric $\varrho_{U},\varrho_{V}$ and some constants $c_{u},c_{v}\in\mathbb{R}$ such that
 $U(x_{1})+U(x_{2})$ $\geq$ $c_{u}+\varrho_{U}(x_{1},x_{2}),$
-$V(y_{1})+V(y_{2})$ $\geq$ $c_{v}+\varrho_{V}(y_{1},y_{2}).$ (76)
-In this case, by (35), for any $p\geq 1$ there exists some constant $c_{p}>0$ such that
+(4.29) $V(y_{1})+V(y_{2})$ $\geq$ $c_{v}+\varrho_{V}(y_{1},y_{2}).$
+In this case, by (3.5), for any $p\geq 1$ there exists some constant $c_{p}>0$ such that
 […]
-with the $p$-Wasserstein semi-metrics $\mathbb{W}_{p,\varrho_{U}},\mathbb{W}_{p,\varrho_{V}}$ defined in (8). This yields the following Wasserstein estimates.
+with the $p$-Wasserstein semi-metrics $\mathbb{W}_{p,\varrho_{U}},\mathbb{W}_{p,\varrho_{V}}$ defined in (1.8). This yields the following Wasserstein estimates.
 Corollary 4.12.
-Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
-In addition, the target potential functions $(U,V)$ satisfy (76) for some semi-metric $(\varrho_{U},\varrho_{V})$.
+Assume ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$.
+In addition, the target potential functions $(U,V)$ satisfy (4.29) for some semi-metric $(\varrho_{U},\varrho_{V})$.
 In this situation, for any $p\geq 1$ there exists some $c_{p}\geq 0$ such that for any $n\geq n_{0}$ we have
@@
 […]
-In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
+In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (4.25) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (4.23).
 Remark 4.13.
-Using Proposition D.3 in Appendix D and following exactly the same argument as in the proof of Theorem 4.13, the estimate (71) remains valid when the weaker condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, provided that
+Using Proposition D.3 in Appendix D and following exactly the same argument as in the proof of Theorem 4.8, the estimate (4.24) remains valid when the weaker condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ holds for some $n_{0}\geq 0$, provided that
 $\left(\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\right)\vee\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right)<\infty.$
-4.6 Entropy exponential decays
-Let us discuss some fairly straightforward consequences of the estimate (71). First note that (71) ensures the convergence of the series
-$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ $:=$ $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
-$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ $:=$ $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (78)
-where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (18) converge as $n\rightarrow\infty$ to a system of integral equations
-$\mathbb{U}=U+\log{{\cal Q}(e^{-\mathbb{V}})}\quad\mbox{\rm and}\quad\mathbb{V}=V+\log{{\cal R}(e^{-\mathbb{U}})}.$
-More precisely, using the estimates in (71) it is straightforward to prove the following corollary.
+4.6. Entropy exponential decays
+Let us discuss some fairly straightforward consequences of the estimate (4.24). First note that (4.24) ensures the convergence of the series
+$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ $:=$ $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\mathcal{B}}_{\varphi}(\mathbb{X})$,}$
+(4.31) $\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ $:=$ $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\mathcal{B}}_{\psi}(\mathbb{Y})$},$
+where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (2.4) converge as $n\rightarrow\infty$ to a system of integral equations
+$\mathbb{U}=U+\log{{\mathcal{Q}}(e^{-\mathbb{V}})}\quad\mbox{\rm and}\quad\mathbb{V}=V+\log{{\mathcal{R}}(e^{-\mathbb{U}})}.$
+More precisely, using the estimates in (4.24) it is straightforward to prove the following corollary.
 Corollary 4.14.
 Under the assumptions of Theorem 4.8 for any $n\geq 1$ we have
-$\|\mathbb{U}-U_{2n}\|_{\psi}\vee\|\mathbb{V}-V_{2n}\|_{\varphi}\leq\frac{c}{1-\rho^{2}}~\rho^{2n}$ (79)
-with the constant $c>0$ as in (71). In addition, we have the relative entropy estimates
-$\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2n+1}\right)\leq c~\lambda_{U}(\varphi)~\rho^{2n}~\quad\mbox{and}\quad\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2n}\right)\leq c~\nu_{V}(\psi)~\rho^{2n}.$ (80)
-With this notation at hand, the bridge distribution (4) has the form
-$\lim_{n\to\infty}{\cal P}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~\lambda_{\mathbb{U}}(dx)~\nu_{\mathbb{V}}(dy).$ (81)
+(4.32) $\|\mathbb{U}-U_{2n}\|_{\psi}\vee\|\mathbb{V}-V_{2n}\|_{\varphi}\leq\frac{c}{1-\rho^{2}}~\rho^{2n}$
+with the constant $c>0$ as in (4.24). In addition, we have the relative entropy estimates
+(4.33) $\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2n+1}\right)\leq c~\lambda_{U}(\varphi)~\rho^{2n}~\quad\mbox{and}\quad\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2n}\right)\leq c~\nu_{V}(\psi)~\rho^{2n}.$
+With this notation at hand, the bridge distribution (1.4) has the form
+(4.34) $\lim_{n\to\infty}{\mathcal{P}}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~\lambda_{\mathbb{U}}(dx)~\nu_{\mathbb{V}}(dy).$
 Moreover, we have
-$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
-$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right).$
-The exponential estimate (71) also allows the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (80).
+$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
+$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2q+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right).$
+The exponential estimate (4.24) also justifies the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (4.33).
 Corollary 4.15.
 Under the assumptions of Theorem 4.8 we obtain
-$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})$ $=$ $\sum_{p\geq n}\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2p+1}\right)+\sum_{p\geq n}\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2p}\right)\quad\text{for any $n\geq 1$}.$
+$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})$ $=$ $\sum_{p\geq n}\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2p+1}\right)+\sum_{p\geq n}\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2p}\right)\quad\text{for any $n\geq 1$}.$
 In addition, there exists some $c>0$ such that, for any $n\geq 1$, we have the entropy exponential decay estimates
-$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\rho^{2n}~(\lambda_{U}(\varphi)+\nu_{V}(\psi)).$
-5 Some illustrations
-5.1 Polynomial growth models
+$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})\leq c~\rho^{2n}~(\lambda_{U}(\varphi)+\nu_{V}(\psi)).$
+5. Some illustrations
+In this section, we illustrate the scope and flexibility of the theoretical results developed above through a collection of representative examples.
+We show how the Lyapunov, drift-minorization, and weighted contraction conditions can be verified in a variety of settings, ranging from polynomial and heavy-tailed models to boundary-driven, semi-compact, and Gaussian frameworks.
+These examples highlight how the abstract assumptions translate into explicit growth, integrability, and regularity conditions on the target potentials and transition kernels.
+5.1. Polynomial growth models
 Consider some finite dimensional normed vector space $\mathbb{X}=\mathbb{Y}$ equipped with measure $\lambda=\nu$ such that for any $p>0$ and $\tau>0$ we have
-$\int e^{-\tau\|x\|^{p}}~\lambda(dx)<\infty.$ (82)
+(5.1) $\int e^{-\tau\|x\|^{p}}~\lambda(dx)<\infty.$
 In this context, we have $\lambda_{U}(dx)=e^{-U(x)}~\lambda(dx)$ and $\nu_{V}(dx)=\lambda_{V}(dx)=e^{-V(x)}~\lambda(dx)$.
-Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions.
+Assume that the functions $U,V$ and $W_{U},W^{V}$ are continuous.
 In addition, there exists some parameters $c_{w},c_{u},c_{v}\in\mathbb{R}$ as well as $\varsigma_{u},\varsigma_{v},\tau_{u},\tau_{v},p_{u},p_{v}>0$ and $q_{u},q_{v}\geq 0$ such that
 $W(x,y)$ $\leq$ $c_{w}+\varsigma_{u}~\|x\|^{p_{u}}+\varsigma_{v}~\|y\|^{p_{v}},$
-$U(x)$ $\geq$ $c_{u}+\tau_{u}~\|x\|^{p_{u}+q_{u}}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\tau_{v}~\|x\|^{p_{v}+q_{v}}.$ (83)
+(5.2) $U(x)$ $\geq$ $c_{u}+\tau_{u}~\|x\|^{p_{u}+q_{u}}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\tau_{v}~\|x\|^{p_{v}+q_{v}}.$
 Proposition 5.1.
-Condition ${\cal H}^{\prime}_{\delta}(U,V)$ is satisfied for some $0<\delta<1/2$ whenever, for any parameter $\iota\in\{u,v\}$, one of the following assertions is satisfied
-$\left(q_{\iota}>0\right)\quad\mbox{or}\quad\left(q_{\iota}=0\quad\mbox{and}\quad\tau_{\iota}/2>\varsigma_{\iota}\right).$ (84)
+Condition ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ is satisfied for some $0<\delta<1/2$ whenever, for any parameter $\iota\in\{u,v\}$, one of the following assertions is satisfied
+(5.3) $\left(q_{\iota}>0\right)\quad\mbox{or}\quad\left(q_{\iota}=0\quad\mbox{and}\quad\tau_{\iota}/2>\varsigma_{\iota}\right).$
 Proof.
-The exponential moment condition (82) ensures that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments
+The exponential moment condition (5.1) ensures that (4.21) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments
 $m^{p_{u}}_{u}:=\int\|x\|^{p_{u}}\lambda_{U}(dx)<\infty\quad\mbox{\rm and}\quad m^{p_{v}}_{v}:=\int\|x\|^{p_{u}}\lambda_{V}(dx)<\infty.$
 Also note that
-${\cal U}_{\delta}(x)=\delta U(x)-W^{V}(x)$ $\geq$ $(\delta c_{u}-c_{w}-\varsigma_{v}~m^{p_{v}})+\delta\tau_{u}~\|x\|^{p_{u}+q_{u}}-\varsigma_{u}~\|x\|^{p_{u}}.$
-This shows that ${\cal U}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
+${\mathcal{U}}_{\delta}(x)=\delta U(x)-W^{V}(x)$ $\geq$ $(\delta c_{u}-c_{w}-\varsigma_{v}~m^{p_{v}})+\delta\tau_{u}~\|x\|^{p_{u}+q_{u}}-\varsigma_{u}~\|x\|^{p_{u}}.$
+This shows that ${\mathcal{U}}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
 $q_{u}>0\quad\mbox{\rm or}\quad\left(q_{u}=0\quad\mbox{\rm and}\quad\delta\tau_{u}>\varsigma_{u}\right).$
-In the same vein, condition (84) with $a=v$ ensures ${\cal V}_{\delta}$ has compact sub-level sets.
+In the same vein, condition (5.3) with $a=v$ ensures ${\mathcal{V}}_{\delta}$ has compact sub-level sets.
 For instance, for bi-Laplace transitions on $\mathbb{X}=\mathbb{R}=\mathbb{Y}$ we have
-$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{\varsigma}{2}~e^{-\varsigma|x-y|}~dy\quad\mbox{\rm with $\varsigma>0$}\\ \\ \Longrightarrow W(x,y)=c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma|x|+\varsigma|y|\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }\\ \\ \Longrightarrow|W(x,z_{1})-W(x,z_{2})|\vee|W(z_{1},y)-W(z_{2},y)|\leq\varsigma~|z_{1}-z_{2}|.\end{array}$
+$\begin{array}[]{l}\displaystyle{\mathcal{Q}}(x,dy)=\frac{\varsigma}{2}~e^{-\varsigma|x-y|}~dy\quad\mbox{\rm with $\varsigma>0$}\\ \\ \Longrightarrow W(x,y)=c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma|x|+\varsigma|y|\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }\\ \\ \Longrightarrow|W(x,z_{1})-W(x,z_{2})|\vee|W(z_{1},y)-W(z_{2},y)|\leq\varsigma~|z_{1}-z_{2}|.\end{array}$
 For linear Gaussian models, the r.h.s.
-condition in (84) is satisfied as soon as the covariance of the linear Gaussian transition $\cal Q$} is defined by (2) is sufficiently large.
-In the same context of linear and Gaussian models, we show in Proposition 5.3 that this condition is in fact equivalent to condition ${\cal H}_{\delta}(U,V)$.
-By (83) there exists some constant $\overline{\tau}_{u}>0$ such that
+condition in (5.3) is satisfied as soon as the covariance of the linear Gaussian transition $\mathcal{Q}$} is defined by (1.2) is sufficiently large.
+In the same context of linear and Gaussian models, we show in Proposition 5.3 that this condition is in fact equivalent to condition ${\mathcal{H}}_{\delta}(U,V)$.
+By (5.2) there exists some constant $\overline{\tau}_{u}>0$ such that
 $U(x_{1})+U(x_{2})$ $\geq$ $\overline{c}_{u}+\overline{\tau}_{u}~\left(\|x_{1}\|+\|x_{2}\|\right)^{p_{u}+q_{u}}\quad\mbox{\rm with}\quad\overline{c}_{u}:=2c_{u}.$
@@
 $V(x_{1})+V(x_{2})\geq\overline{c}_{v}+\varrho_{V}(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho_{V}(x_{1},x_{2}):=\overline{\tau}_{v}~\|x_{1}-x_{2}\|^{p_{v}+q_{v}}.$
-We conclude that $(U,V)$ satisfy (76) with the semi-metric $(\varrho_{U},\varrho_{V})$ so that the Wasserstein estimates (77) hold.
-5.2 Heavy tailed transitions
+We conclude that $(U,V)$ satisfy (4.29) with the semi-metric $(\varrho_{U},\varrho_{V})$ so that the Wasserstein estimates (4.30) hold.
+5.2. Heavy tailed transitions
 Consider the normed vector space $\mathbb{X}=\mathbb{Y}$ and the reference measure $\lambda=\nu$ discussed in Section 5.1.
@@
 For instance, for Cauchy transitions on $\mathbb{X}=\mathbb{R}=\mathbb{Y}$ we have
-$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{\pi a}~\frac{1}{1+((x-y)/a)^{2}}~dy\quad\mbox{\rm with}\quad a>0\\ \\ \displaystyle\Longrightarrow W(x,y)=\log{(a\pi)}+\log{\left(1+((x-y)/a)^{2}\right)}\quad\mbox{\rm and}\quad|\partial_{x}W(x,y)|\vee|\partial_{y}W(x,y)|\leq 1/a\\ \\ \displaystyle\Longrightarrow c_{w}=\log{(a\pi)},\qquad\tau_{p,w}=2/a^{2}=\tau_{q,w}\quad\mbox{\rm and}\quad p_{w}=q_{w}=2.\end{array}$
-Exponential targets
+$\begin{array}[]{l}\displaystyle{\mathcal{Q}}(x,dy)=\frac{1}{\pi a}~\frac{1}{1+((x-y)/a)^{2}}~dy\quad\mbox{\rm with}\quad a>0\\ \\ \displaystyle\Longrightarrow W(x,y)=\log{(a\pi)}+\log{\left(1+((x-y)/a)^{2}\right)}\quad\mbox{\rm and}\quad|\partial_{x}W(x,y)|\vee|\partial_{y}W(x,y)|\leq 1/a\\ \\ \displaystyle\Longrightarrow c_{w}=\log{(a\pi)},\qquad\tau_{p,w}=2/a^{2}=\tau_{q,w}\quad\mbox{\rm and}\quad p_{w}=q_{w}=2.\end{array}$
+5.2.1. Exponential targets
 Consider some continuous target potential functions $U,V$ such that
@@
 Also assume that $(W_{U},W^{V})$ are continuous functions.
-Following the proof of Proposition 5.1, the exponential moment condition (82) ensures that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\nu_{V}=\lambda_{V}$ have finite $p_{w}$ and $q_{w}$ absolute moments.
+Following the proof of Proposition 5.1, the exponential moment condition (5.1) ensures that (4.21) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\nu_{V}=\lambda_{V}$ have finite $p_{w}$ and $q_{w}$ absolute moments.
 By Jensen’s inequality, we also have
-${\cal U}_{\delta}(x)$ $\geq$ $\delta c_{u}+\delta\tau_{u}~\|x\|^{p_{u}}-c_{w}-\log\left(1+\tau_{p,w}~\|x\|^{p_{w}}+\tau_{q,w}~m_{v}^{q_{w}}\right)\stackrel{{\scriptstyle\|x\|\rightarrow\infty}}{{\longrightarrow}}\infty$
+${\mathcal{U}}_{\delta}(x)$ $\geq$ $\delta c_{u}+\delta\tau_{u}~\|x\|^{p_{u}}-c_{w}-\log\left(1+\tau_{p,w}~\|x\|^{p_{w}}+\tau_{q,w}~m_{v}^{q_{w}}\right)\stackrel{{\scriptstyle\|x\|\rightarrow\infty}}{{\longrightarrow}}\infty$
 In the same vein we check that
-$\lim_{\|x\|\rightarrow\infty}(\delta U(x)-W^{V}(x))=\infty=\lim_{\|x\|\rightarrow\infty}(\delta V(x)-W_{U}(x)).$ (85)
-We conclude that ${\cal H}^{\prime}_{\delta}(U,V)$ is met for any $0<\delta<1/2$. Also note that $(U,V)$ satisfy (76) with the semi-metrics
+(5.4) $\lim_{\|x\|\rightarrow\infty}(\delta U(x)-W^{V}(x))=\infty=\lim_{\|x\|\rightarrow\infty}(\delta V(x)-W_{U}(x)).$
+We conclude that ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ is met for any $0<\delta<1/2$. Also note that $(U,V)$ satisfy (4.29) with the semi-metrics
 $\varrho_{U}(x_{1},x_{2}):=\overline{\tau}_{u}~\|x_{1}-x_{2}\|^{p_{u}}\quad\mbox{\rm and}\quad\varrho_{V}(x_{1},x_{2}):=\overline{\tau}_{v}~\|x_{1}-x_{2}\|^{q_{v}}$
-for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (77) hold.
-Heavy tailed targets
+for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (4.30) hold.
+5.2.2. Heavy tailed targets
 Consider continuous target potential functions $U,V$ such that
@@
 $p_{u}>1+2p_{w}\quad\mbox{\rm and}\quad q_{v}>1+2q_{w}.$
-Condition (68) is met as soon as
+Condition (4.21) is met as soon as
 $(1-2\delta)p_{u}>1\quad\mbox{\rm and}\quad(1-2\delta)q_{v}>1$
@@
 $p_{u}-p_{w}>1\quad\mbox{\rm and}\quad q_{v}-q_{w}>1.$
-Finally, (85) holds when
+Finally, (5.4) holds when
 $\delta p_{u}-p_{w}>0\quad\mbox{\rm and}\quad\delta q_{v}-q_{w}>0.$
-This shows that ${\cal H}_{\delta}(U,V)$ for any $0<\delta<1/2$ such that
+This shows that ${\mathcal{H}}_{\delta}(U,V)$ for any $0<\delta<1/2$ such that
 $\frac{p_{w}}{p_{u}}<\delta<\frac{1}{2}~\left(1-\frac{1}{p_{u}}\right)\quad\mbox{\rm and}\quad\frac{q_{w}}{q_{v}}<\delta<\frac{1}{2}~\left(1-\frac{1}{q_{v}}\right).$
-Also note that the exponential functions $(e^{\delta U},e^{\delta V})$ satisfy (74) with the semi-metrics
+Also note that the exponential functions $(e^{\delta U},e^{\delta V})$ satisfy (4.27) with the semi-metrics
 $\varrho_{\varphi}(x_{1},x_{2}):=\overline{\tau}_{u}~\|x_{1}-x_{2}\|^{p_{u}}\quad\mbox{\rm and}\quad\varrho_{\psi}(x_{1},x_{2}):=\overline{\tau}_{v}~\|x_{1}-x_{2}\|^{q_{v}}$
-for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (75) hold.
-5.3 Some boundary states
+for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (4.28) hold.
+5.3. Some boundary states
 Consider a bounded Lipschitz domain $\mathbb{X}=\mathbb{Y}\subset\mathbb{R}^{d}$ equipped with the uniform measure $\lambda=\nu$.
@@
 Finally, we assume that the integrated functions $(W^{V},W_{U})$ are continuous.
-When $W(x,y)$ is bounded, Sinkhorn transitions are strongly mixing, in the sense that $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ with the uniform contraction parameter $\raisebox{2.0pt}{$\chi$}({\cal S})$ defined in (49).
+When $W(x,y)$ is bounded, Sinkhorn transitions are strongly mixing, in the sense that $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})<1$ with the uniform contraction parameter $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})$ defined in (4.2).
 The exponential decays of $\Phi$-entropies associated with these strong mixing models are discussed in Section 4.1.
@@
 $U(x)\geq-\log{c_{q,u}}+q_{u}~\log{\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})}}\Longleftrightarrow e^{\delta U(x)}\geq\frac{1}{c_{q,u}^{\delta}}~\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})^{\delta q_{u}}}.$
-By Remark 4.11 the above models satisfy ${\cal H}_{\delta}(U,V)$ for any $\delta\in]0,1[$.
-In addition, following the proof of (36) we check that the exponential function $e^{\delta U}$ satisfies (74) with the semi-metric
+By Remark 4.7 the above models satisfy ${\mathcal{H}}_{\delta}(U,V)$ for any $\delta\in]0,1[$.
+In addition, following the proof of (3.6) we check that the exponential function $e^{\delta U}$ satisfies (4.27) with the semi-metric
 $\varrho_{\varphi}(x_{1},x_{2})=\overline{c}_{u}~\left(\frac{1}{\varrho_{\partial}(x_{1},\partial\mathbb{X})}-\frac{1}{\varrho_{\partial}(x_{2},\partial\mathbb{X})}\right)^{\delta q_{u}}~1_{x_{1}\not=x_{2}}\quad\mbox{\rm for some parameter $\overline{c}_{u}>0$.}$
-Similarly, the exponential function $e^{\delta V}$ satisfies (74) with the metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(c_{u},q_{u})$ by $(c_{v},q_{v})$, so that the Wasserstein estimates (75) hold.
+Similarly, the exponential function $e^{\delta V}$ satisfies (4.27) with the metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(c_{u},q_{u})$ by $(c_{v},q_{v})$, so that the Wasserstein estimates (4.28) hold.
 When $W(x,y)=W(y,x)$ is symmetric and non necessarily bounded, assume we have an estimate of the form
 $W(x,y)\leq c_{w}+q_{w}~\log{\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})}}$
-for some $c_{w}\in\mathbb{R}$ and some parameters $0\leq q_{w}<(q_{u}\wedge q_{v})/2$. In this case condition ${\cal H}_{\delta}(U,V)$ holds for any $\frac{q_{w}}{q_{u}\wedge q_{v}}<\delta<1/2$.
-5.4 Beta marginal models
+for some $c_{w}\in\mathbb{R}$ and some parameters $0\leq q_{w}<(q_{u}\wedge q_{v})/2$. In this case condition ${\mathcal{H}}_{\delta}(U,V)$ holds for any $\frac{q_{w}}{q_{u}\wedge q_{v}}<\delta<1/2$.
+5.4. Beta marginal models
 Next we illustrate the boundary state models discussed in Section 5.3 with the open unit interval $\mathbb{X}=\mathbb{Y}=]0,1[$ with boundary $\partial\mathbb{X}=\{0,1\}$.
@@
 $\frac{1}{2}~\varrho_{\partial}(x,\partial\mathbb{X})\leq x(1-x)\leq\varrho_{\partial}(x,\partial\mathbb{X})$
-and consider the reference Markov transition $\cal Q$} defined by the semi-circle distribution (41). Let $(\lambda_{U},\lambda_{V})$ be some Beta marginal distributions given by
+and consider the reference Markov transition $\mathcal{Q}$} defined by the semi-circle distribution (3.11). Let $(\lambda_{U},\lambda_{V})$ be some Beta marginal distributions given by
 $\lambda_{U}(dx)=\frac{1}{c_{u}}~x^{a_{u}}(1-x)^{b_{u}}~\lambda(dx)\quad\mbox{\rm and}\quad\lambda_{V}(dx)=\frac{1}{c_{v}}~x^{a_{v}}(1-x)^{b_{v}}~\lambda(dx)$
@@
 $|\partial_{x}\log{(1-(x-y)^{2})}|\leq g(x)^{2}\wedge g(y)^{2}.$
-In this example, the functions $(U,V)$ and $(W_{U},W^{V})$ are differentiable functions. Condition (66) is clearly met for any $0<\delta<1/2$ and we have
+In this example, the functions $(U,V)$ and $(W_{U},W^{V})$ are differentiable functions. Condition (4.19) is clearly met for any $0<\delta<1/2$ and we have
 $\delta U(x)-W^{V}(x)\geq\delta\log{(c_{u})}+(\delta a_{u}-1/2)\log{\frac{1}{x}}+(\delta b_{u}-1/2)\log{\frac{1}{1-x}}.$
-This shows that ${\cal H}^{\prime}_{\delta}(U,V)$ is met for any $((a_{u}\wedge b_{u})\wedge(a_{v}\wedge b_{v}))^{-1}<2\delta<1$. We also have
+This shows that ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ is met for any $((a_{u}\wedge b_{u})\wedge(a_{v}\wedge b_{v}))^{-1}<2\delta<1$. We also have
 $U(x_{1})+U(x_{2})\geq\overline{c}_{u}:+\varrho_{U}(x_{1},x_{2})$
@@
 $\varrho_{U}(x_{1},x_{2}):=a_{u}~\left|\log{\frac{1}{x_{1}}}-\log{\frac{1}{x_{2}}}\right|+b_{u}\left|\log{\frac{1}{1-x_{1}}}-\log{\frac{1}{1-x_{2}}}\right|.$
-We conclude that $U$ satisfies (76). Similarly $V$ satisfies (76) with a metric $\varrho_{V}$ defined as $\varrho_{U}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$.
+We conclude that $U$ satisfies (4.29). Similarly $V$ satisfies (4.29) with a metric $\varrho_{V}$ defined as $\varrho_{U}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$.
 Recall that $uv\geq(u+v)/2$ and $u^{p}v^{q}\geq(uv)^{p\wedge q}$ for any $u,v\geq 1$ and any parameters $p,q>0$ we see that
@@
 $e^{\delta U(x)}$ $\geq$ $c_{u}^{\delta}~\left(\frac{1}{x(1-x)}\right)^{\delta(a_{u}\wedge b_{u})}\geq\overline{c}_{u}^{\delta}~\frac{1}{\varrho_{\partial}(x,\partial\mathbb{X})^{\delta(a_{u}\wedge b_{u})}}\quad\mbox{\rm with}\quad\overline{c}_{u}:=2^{(a_{u}\wedge b_{u})}c_{u}.$
-Following the proof of (36) we check that the exponential function $e^{\delta U}$ satisfies (74) with
+Following the proof of (3.6) we check that the exponential function $e^{\delta U}$ satisfies (4.27) with
 $\varrho_{\varphi}(x_{1},x_{2})=\overline{c}_{u}~\left(\frac{1}{\varrho_{\partial}(x_{1},\partial\mathbb{X})}-\frac{1}{\varrho_{\partial}(x_{2},\partial\mathbb{X})}\right)^{\delta(a_{u}\wedge b_{u})}$
@@
 for some parameter $\overline{c}_{u}>0$.
-Similarly, the exponential function $e^{\delta V}$ satisfies (74) with the complete metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$, so that the Wasserstein estimates (75) hold.
-5.5 Weibull marginal models
+Similarly, the exponential function $e^{\delta V}$ satisfies (4.27) with the complete metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$, so that the Wasserstein estimates (4.28) hold.
+5.5. Weibull marginal models
 We equip the open half-line $\mathbb{X}=\mathbb{Y}=]0,\infty[$ with the Lebesgue measure $\lambda(dx)=\nu(dx)=1_{]0,\infty[}(x)dx$ and consider the Weibull marginal measure
@@
 and $a_{u},b_{u}>0$.
-Note that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments $m^{p_{u}}_{u}$ and $m^{p_{v}}_{v}$ for any $p_{u},p_{v}$.
+Note that (4.21) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments $m^{p_{u}}_{u}$ and $m^{p_{v}}_{v}$ for any $p_{u},p_{v}$.
 We define similarly $\lambda_{V}$ with some parameters $a_{v},b_{v}>0$.
-Consider a continuous function $W(x,y)$ satisfying (83) for some parameters $c_{w}\in\mathbb{R}$ as well as $p_{u},p_{v}>0$. In this context, we have
+Consider a continuous function $W(x,y)$ satisfying (5.2) for some parameters $c_{w}\in\mathbb{R}$ as well as $p_{u},p_{v}>0$. In this context, we have
 $W^{V}(x)\leq c_{w}+\varsigma_{u}~x^{p_{u}}+\varsigma_{v}~m_{v}^{p_{v}}\quad\mbox{\rm with}\quad m^{p_{v}}_{v}:=\int_{0}^{\infty}~x^{p_{u}}\lambda_{V}(dx)<\infty.$
@@
 $\delta U(x)-W^{V}(x)\geq\overline{c}_{u}+\delta a_{u}~\log{\frac{1}{x}}+\frac{\delta}{b_{u}^{a_{u}+1}}~x^{a_{u}+1}-\varsigma_{u}~x^{p_{u}}\quad\mbox{\rm with}\quad\overline{c}_{u}:=\delta c_{u}-c_{w}-\varsigma_{v}~m_{v}^{p_{v}}$
-Assume that $(W_{U},W^{V})$ are both continuous functions. In this situation, ${\cal U}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
+Assume that $(W_{U},W^{V})$ are both continuous functions. In this situation, ${\mathcal{U}}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
 $a_{u}+1>p_{u}\quad\mbox{\rm or}\quad\left(p_{u}=a_{u}+1\quad\mbox{\rm and}\quad\varsigma_{u}~b_{u}^{a_{u}+1}<\delta<1/2\right).$
 The r.h.s. condition requires that $\varsigma_{u}<b_{u}^{-(a_{u}+1)}/2$. The analysis of these Weibull models now follows the same line of arguments as the ones we used for the Beta models.
-5.6 Exponential marginal models
+5.6. Exponential marginal models
 We equip the half-line $\mathbb{X}=\mathbb{Y}=[0,\infty[$ with the Lebesgue measure $\lambda(dx)=1_{[0,\infty[}(x)dx$ and consider the exponential marginal measures
@@
 Note that the functions $x\in[0,\infty[\mapsto\tau_{\iota}x\in[0,\infty[$ with $\iota\in\{u,v\}$ have compact sub-level sets.
-Let ${\cal Q}(x,dy)$ be the exponential transition defined for any $x\in[0,\infty[$ by
+Let ${\mathcal{Q}}(x,dy)$ be the exponential transition defined for any $x\in[0,\infty[$ by
 […]
@@
 $W(x,y)\leq c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma~x+\varsigma~y,\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }.$
-In this context, condition ${\cal H}^{\prime}_{\delta}(U,V)$ holds for any $({\varsigma}/{\tau_{u}})\vee({\varsigma}/{\tau_{v}})<\delta<{1}/{2}$.
-5.7 Semi-compact models
+In this context, condition ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ holds for any $({\varsigma}/{\tau_{u}})\vee({\varsigma}/{\tau_{v}})<\delta<{1}/{2}$.
+5.7. Semi-compact models
 In many practical situations of interest, the target state spaces are different. Next we illustrate this situation with a semi-discrete model and a semi-compact models.
-Semi-discrete model
+Semi-discrete model.
 Consider a finite or countable subset $\mathbb{X}\subset\mathbb{R}^{d}$ equipped with the counting measure $\lambda$ and the discrete topology.
@@
 $U~:~x\in\mathbb{X}\mapsto U(x)\in[0,\infty[\quad\mbox{\rm such that}\quad\sum_{x\in\mathbb{X}}e^{-U(x)}=1.$
-Let $\nu_{V}(dy)=e^{-V(y)}~dy$ be a Boltzmann-Gibbs probability measure on $\mathbb{R}^{d}$ and assume that $(U,V,W)$ are continuous functions satisfying the polynomial growth conditions (83).
+Let $\nu_{V}(dy)=e^{-V(y)}~dy$ be a Boltzmann-Gibbs probability measure on $\mathbb{R}^{d}$ and assume that $(U,V,W)$ are continuous functions satisfying the polynomial growth conditions (5.2).
 In this context, the sub-level sets of $U$ are finite, thus compact in $\mathbb{X}$.
-Following the proof of Proposition 5.1 these conditions ensure that ${\cal H}^{\prime}_{\delta}(U,V)$ holds for some $0<\delta<1/2$ as soon as condition (84) is satisfied.
+Following the proof of Proposition 5.1 these conditions ensure that ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ holds for some $0<\delta<1/2$ as soon as condition (5.3) is satisfied.
 The analysis of heavy tailed transitions follows word-for-word the same lines of arguments as in Section 5.2, this it is skipped.
-Semi-compact model
+Semi-compact model.
 Consider the uniform measure $\lambda_{U}$ on a compact space $\mathbb{X}\subset\mathbb{Y}=\mathbb{R}^{d}$ associated with a null potential $U=0$, which yields $\lambda_{U}=\lambda$.
@@
 $W(x,y)\leq w(x)+a~\|y\|^{p}\Longrightarrow W_{U}(y)\leq\lambda(w)+b~\|y\|^{p}$
-for some continuous function $w(x)$ and some parameters $a\leq 2$ and $p\leq 1$. These models satisfy ${\cal H}^{\prime}_{\delta}(U,V)$ as soon as
-$(p=2~~\mbox{\rm and}~~b<\delta/2)\quad\mbox{\rm or}\quad(p<2)$ (86)
+for some continuous function $w(x)$ and some parameters $a\leq 2$ and $p\leq 1$. These models satisfy ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ as soon as
+(5.5) $(p=2~~\mbox{\rm and}~~b<\delta/2)\quad\mbox{\rm or}\quad(p<2)$
 We check this claim using the fact that
 […]
-5.8 Linear-Gaussian models
+5.8. Linear-Gaussian models
 Consider the state spaces $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, for some integer $d\geq 1$ and the Lebesgue measure $\lambda(dz)=\nu(dz)=dz$.
 Hereafter, points $x$ in the Euclidean space $\mathbb{X}=\mathbb{R}^{d}$ are represented by $d$-dimensional column vectors (or, equivalently, by $d\times 1$ matrices).
-When there is no possible confusion, we use the notation $\|\cdot\|$ for any equivalent matrix or vector norm.
-Let ${\cal G}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\cal S}^{+}_{d}\subset{\cal G}l_{d}$ the subset of positive definite matrices.
+Let ${\mathcal{G}}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\mathcal{S}}^{+}_{d}\subset{\mathcal{G}}l_{d}$ the subset of positive definite matrices.
 We sometimes use the Löwner partial ordering notation $v_{1}\geq v_{2}$ to mean that a symmetric matrix $v_{1}-v_{2}$ is positive semi-definite (equivalently, $v_{2}-v_{1}$ is negative semi-definite), and $v_{1}>v_{2}$ when $v_{1}-v_{2}$ is positive definite (equivalently, $v_{2}-v_{1}$ is negative definite).
-Given $v\in{\cal S}_{d}^{+}$ we denote by $v^{1/2}$ the principal (unique) symmetric square root.
-Denote by $\nu_{m,\sigma}$ the Gaussian distribution on $\mathbb{R}^{d}$ with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\sigma\in{\cal S}^{+}_{d}$.
-In addition, let $g_{\sigma}$ denote the probability density function of the distribution $\nu_{0,\sigma}$, with covariance matrix $\sigma\in{\cal S}^{+}_{d}$.
+Given $v\in{\mathcal{S}}_{d}^{+}$ we denote by $v^{1/2}$ the principal (unique) symmetric square root.
+Denote by $\nu_{m,\sigma}$ the Gaussian distribution on $\mathbb{R}^{d}$ with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\sigma\in{\mathcal{S}}^{+}_{d}$.
+In addition, let $g_{\sigma}$ denote the probability density function of the distribution $\nu_{0,\sigma}$, with covariance matrix $\sigma\in{\mathcal{S}}^{+}_{d}$.
 Consider the target probability measures $(\lambda_{U},\nu_{V})=(\nu_{m,\sigma},\nu_{\overline{m},\overline{\sigma}})$ for some given parameters
@@
 Equivalently, the log-density functions $(U,V)$ are defined by
-$U(x)=-\log{g_{\sigma}(x-m)}\quad\mbox{\rm and}\quad V(y):=-\log{g_{\overline{\sigma}}(y-\overline{m})}.$ (87)
-The linear Gaussian transition $\cal Q$} is defined by (2), with the potential
-$W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$ (88)
+(5.6) $U(x)=-\log{g_{\sigma}(x-m)}\quad\mbox{\rm and}\quad V(y):=-\log{g_{\overline{\sigma}}(y-\overline{m})}.$
+The linear Gaussian transition $\mathcal{Q}$} is defined by (1.2), with the potential
+(5.7) $W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$
 […]
-$\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$ (89)
-The matrix $\Delta_{\varpi}$ in expression (89) above is defined as
-$\Delta_{\varpi}:=\frac{\varpi}{2}+\left(\varpi+\left(\frac{\varpi}{2}\right)^{2}\right)^{1/2}\geq\varpi+(I+\varpi^{-1})^{-1}$ (90)
+(5.8) $\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$
+The matrix $\Delta_{\varpi}$ in expression (5.8) above is defined as
+(5.9) $\Delta_{\varpi}:=\frac{\varpi}{2}+\left(\varpi+\left(\frac{\varpi}{2}\right)^{2}\right)^{1/2}\geq\varpi+(I+\varpi^{-1})^{-1}$
 with
@@
 Proposition 5.2.
-Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied with the Lyapunov functions
-$\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$ (91)
+Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is satisfied with the Lyapunov functions
+(5.10) $\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$
 A proof is provided in the Appendix A (on page A).
-The choice of the Lyapunov functions is not unique. For instance, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
+The choice of the Lyapunov functions is not unique. For instance, Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
 $\varphi(x)$ $:=$ $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
-$\psi(y)$ $:=$ $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (92)
-There also exists some $\delta_{0}>0$ such that ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
+(5.11) $\psi(y)$ $:=$ $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$
+There also exists some $\delta_{0}>0$ such that ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
 $\varphi(x)$ $:=$ $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
-$\psi(y)$ $:=$ $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (93)
+(5.12) $\psi(y)$ $:=$ $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$
 A detailed proof of the above assertions is provided in Appendix A (on page A).
-We end this section with an equivalent formulation of condition ${\cal H}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (87) and (88).
+We end this section with an equivalent formulation of condition ${\mathcal{H}}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (5.6) and (5.7).
 Proposition 5.3.
-Condition ${\cal H}_{\delta}(U,V)$ is satisfied for the Lyapunov functions in (93) if, and only if,
-$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma^{-1}\quad\text{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}^{-1}.$ (94)
+Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for the Lyapunov functions in (5.12) if, and only if,
+(5.13) $\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma^{-1}\quad\text{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}^{-1}.$
 A proof is provided in Appendix E (on page E).
-5.9 Finite mixture models
+5.9. Finite mixture models
 Consider some (fully supported) discrete probability measures $\xi_{0}$ and $\xi_{1}$ on some some finite states $E_{0}$ and $E_{1}$ and some collection of probability measures $(\lambda_{U_{i}})_{i\in E_{0}}$ and $(\nu_{V_{j}})_{j\in E_{1}}$ for some continuous potential functions $U_{i}:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V_{j}:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$ indexed by $i\in E_{0}$ and $j\in E_{1}$.
 The mixture distributions $\lambda_{U}$ and $\nu_{V}$ associated with these potential functions take the form
-$U=-\log{\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}}}\quad\mbox{\rm and}\quad V=-\log{\sum_{j\in E_{1}}\xi_{1}(j)~e^{-V_{j}}}.$ (95)
+(5.14) $U=-\log{\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}}}\quad\mbox{\rm and}\quad V=-\log{\sum_{j\in E_{1}}\xi_{1}(j)~e^{-V_{j}}}.$
 Proposition 5.4.
-Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1[$ if conditions ${\cal H}_{\delta}(U_{i},V_{j})$ are met for the same $\delta\in]0,1[$ and almost every $(i,j)\in(E_{0}\times E_{1})$.
+Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1[$ if conditions ${\mathcal{H}}_{\delta}(U_{i},V_{j})$ are met for the same $\delta\in]0,1[$ and almost every $(i,j)\in(E_{0}\times E_{1})$.
 A proof is provided in Appendix E (on page E).
-Let us now consider the mixture model (95) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (88) and the log-densities
-$U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (96)
-where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters.
+Let us now consider the mixture model (5.14) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (5.7) and the log-densities
+(5.15) $U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$
+where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\mathcal{S}}^{+}_{d}\times{\mathcal{S}}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters.
 Combining Proposition 5.4 with Proposition 5.3 we readily obtain the following result.
 Proposition 5.5.
-The Gaussian mixture model defined in (95) with the Gaussian log-densities of (96) defined on $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$ satisfies Condition ${\cal H}_{\delta}(U,V)$ for some $\delta\in]0,1[$ provided that
-$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma_{i}^{-1}\quad\mbox{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}_{j}^{-1}$ (97)
+The Gaussian mixture model defined in (5.14) with the Gaussian log-densities of (5.15) defined on $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$ satisfies Condition ${\mathcal{H}}_{\delta}(U,V)$ for some $\delta\in]0,1[$ provided that
+(5.16) $\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma_{i}^{-1}\quad\mbox{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}_{j}^{-1}$
 for almost every $(i,j)\in(E_{0}\times E_{1})$.
@@
 The Gaussian-kernel estimation model defined by
-$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~dx$ (98)
+(5.17) $\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~dx$
 is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$).
 The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing.
-Note that (98) coincides with (95) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$.
+Note that (5.17) coincides with (5.14) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$.
 For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [parzen, rosenblatt] as well as the books [chacon, simonov, wand].
-The Gaussian mixture distribution (98) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [doucet-bortoli] and references therein.
+The Gaussian mixture distribution (5.17) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [doucet-bortoli] and references therein.
 Remark 5.6.
 Schrödinger bridges and the Sinkhorn algorithm are often used in practice to solve high dimensional optimal transport problems.
-In our context, choosing $\tau=tI$ and $(\alpha,\beta)=(0,I)$ in (88), the Schrödinger bridge in the limit $t\rightarrow 0$ of vanishing regularization solves the Monge-Kantorovich problem and converge towards the optimal $2$-Wasserstein coupling.
-In this case, ${\cal S}_{0}=Q$ coincides with the heat kernel on a time horizon $t$ (see also Remark 2.3 in [adm-24]), namely,
-${\cal Q}(x,dy)=(2\pi t)^{-d/2}~\exp{\left(-\frac{1}{2t}~\|y-x\|^{2}\right)}~dy.$
-In this context, the l.h.s. condition in (97) applied to (98) becomes
-$\frac{1}{t}~I<(\delta\wedge(1-\delta))~\sigma^{-1}.$ (99)
-In many cases, the Markov transition ${\cal Q}(x,dy)$ of interest is generated by a stochastic diffusion between time 0 and a time horizon $t$ –see, e.g., [leonard, stromme2023sampling, kato2024large].
-In our context, if we choose $(\alpha,\beta)=(0,e^{tA})$ and $\tau=\tau_{t}:=\int_{0}^{t}e^{sA}e^{sA^{\prime}}ds$ for some Hurwitz matrix $A$ and some $t>0$, ${\cal Q}(x,dy)$ coincides with the transition semigroup of the Ornstein-Uhlenbeck diffusion (see Remark 2.3 in [adm-24]).
+In our context, choosing $\tau=tI$ and $(\alpha,\beta)=(0,I)$ in (5.7), the Schrödinger bridge in the limit $t\rightarrow 0$ of vanishing regularization solves the Monge-Kantorovich problem and converge towards the optimal $2$-Wasserstein coupling.
+In this case, ${\mathcal{S}}_{0}=Q$ coincides with the heat kernel on a time horizon $t$ (see also Remark 2.3 in [adm-24]), namely,
+${\mathcal{Q}}(x,dy)=(2\pi t)^{-d/2}~\exp{\left(-\frac{1}{2t}~\|y-x\|^{2}\right)}~dy.$
+with the Euclidian norm $\|z\|^{2}=z^{\prime}z$. In this context, the l.h.s. condition in (5.16) applied to (5.17) becomes
+(5.18) $\frac{1}{t}~I<(\delta\wedge(1-\delta))~\sigma^{-1}.$
+In many cases, the Markov transition ${\mathcal{Q}}(x,dy)$ of interest is generated by a stochastic diffusion between time 0 and a time horizon $t$ –see, e.g., [leonard, stromme2023sampling, kato2024large].
+In our context, if we choose $(\alpha,\beta)=(0,e^{tA})$ and $\tau=\tau_{t}:=\int_{0}^{t}e^{sA}e^{sA^{\prime}}ds$ for some Hurwitz matrix $A$ and some $t>0$, ${\mathcal{Q}}(x,dy)$ coincides with the transition semigroup of the Ornstein-Uhlenbeck diffusion (see Remark 2.3 in [adm-24]).
 In this case, the l.h.s.
-condition in (97) applied to (98) is met for any $t\geq t_{0}$ provided that
-$e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~\sigma^{-1}.$ (100)
-Both transitions gradually transform the data distribution $\lambda_{U}$ into a simpler distribution $\lambda_{U}{\cal Q}$ by injecting noise into data.
-Also note that (99) and (100) are met when the time horizon is chosen sufficiently large.
+condition in (5.16) applied to (5.17) is met for any $t\geq t_{0}$ provided that
+(5.19) $e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~\sigma^{-1}.$
+Both transitions gradually transform the data distribution $\lambda_{U}$ into a simpler distribution $\lambda_{U}{\mathcal{Q}}$ by injecting noise into the data.
+Also note that (5.18) and (5.19) are met when the time horizon is chosen sufficiently large.
 The choice of the prior distribution $\nu_{V}$ is rather flexible.
 For instance, we can choose a Gaussian distribution, say $\nu_{0,\overline{\sigma}}$, satisfying the r.h.s.
-condition in (94).
-Starting from such Gaussian random states, the backward Sinkhorn transitions (indexed by odd indices) eventually generates samples back into the initial distribution by slowly removing the noise distortion.
-Conditions (99) and (100) reflect the fact that the forward noise injection as well as the reverse denoising procedure in diffusion-based generative models often requires very large time horizons.
+condition in (5.13).
+Starting from such Gaussian random states, the backward Sinkhorn transitions (indexed by odd indices) eventually generate samples back into the initial distribution by slowly removing the noise distortion.
+Conditions (5.18) and (5.19) reflect the fact that the forward noise injection as well as the reverse denoising procedure in diffusion-based generative models often require very large time horizons.
 Acknowledgements
+We would like to thank the anonymous reviewers for their excellent suggestions for improving the paper. Their detailed comments greatly improved the presentation of the article.
 This work is supported by the Innovation and Talent Base for Digital Technology and Finance (B21038) and ‘the Fundamental Research Funds for the Central Universities’, Zhongnan University of Economics and Law (2722023EJ002).
-J. Miguez acknowledges the financial support of the Office of Naval Research (award no. N00014-22-1-2647) and MICIU/AEI/10.13039/501100011033/FEDER UE (grant no. PID2021-125159NB-I00 TYCHE).
+J.
+Miguez acknowledges the financial support of the Office of Naval Research (award no.
+N00014-22-1-2647), MICIU/AEI/10.13039/501100011033/FEDER UE (grants no.
+PID2021-125159NB-I00 TYCHE and PID2024-158181NB-I00 NISA), and Community of Madrid, under grant IDEA-CM (TEC- 2024/COM-89).
+Appendix
+This appendix compiles technical material and supplementary results that bolster the paper’s central arguments.
+Appendix A focuses on Gaussian Sinkhorn algorithms.
+This material is taken from [adm-24].
+We review some properties of the Riccati matrix differential equations and the random maps associated with Sinkhorn transitions.
+We also provide equivalent formulations of the drift-minorization conditions for the Lyapunov functions presented in Proposition 5.2.
+Appendix B is devoted to basic properties of Hilbert’s projective metric.
+Appendix C recalls several analytical tools related to weighted Banach spaces, operator norms, and contraction coefficients that are used throughout the stability analysis.
+For a more thorough discussion on weighted norms and contraction coefficients, we refer the reader to the articles [dg-01, dm-03, horton-dp, delgerber2025] as well as the book [penev].
+Appendix D reviews Sinkhorn potentials.
+The first part, Appendix D.1, presents some commutation formulae.
+Building on this framework, several quantitative estimates are provided in Appendix D.2.
+Section D.3 is dedicated to Lyapunov inequalities and minorization conditions.
+The Proof of Theorem 4.8 is provided in Appendix D.4 and Appendix E contains further technical arguments.
 Appendix A The Gaussian Sinkhorn algorithm
@@
 Riccati difference equations
-Let us associate with some $\varpi\in{\cal S}^{+}_{d}$ the increasing map $\mbox{\rm Ricc}_{\varpi}$ from ${\cal S}^{0}_{d}$ into ${\cal S}^{+}_{d}$ defined by
-$\begin{array}[]{lccl}\mbox{\rm Ricc}_{\varpi}:&{\cal S}^{0}_{d}&\mapsto&{\cal S}^{+}_{d}\\ &v&\leadsto&\mbox{\rm Ricc}_{\varpi}(v):=(I+(\varpi+v)^{-1})^{-1}\\ \end{array}.$ (101)
+Let us associate with some $\varpi\in{\mathcal{S}}^{+}_{d}$ the increasing map $\mbox{\rm Ricc}_{\varpi}$ from ${\mathcal{S}}^{0}_{d}$ into ${\mathcal{S}}^{+}_{d}$ defined by
+(A.1) $\begin{array}[]{lccl}\mbox{\rm Ricc}_{\varpi}:&{\mathcal{S}}^{0}_{d}&\mapsto&{\mathcal{S}}^{+}_{d}\\ &v&\leadsto&\mbox{\rm Ricc}_{\varpi}(v):=(I+(\varpi+v)^{-1})^{-1}\\ \end{array}.$
 If we define the matrices
-$\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\cal S}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~\tau^{-1}\beta~\sigma^{1/2},$
+$\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\mathcal{S}}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~\tau^{-1}\beta~\sigma^{1/2},$
 and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [adm-24]) that the recursions
-$\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~\upsilon_{2n}~\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~\upsilon_{2n+1}~\gamma^{\prime},$ (102)
+(A.2) $\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~\upsilon_{2n}~\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~\upsilon_{2n+1}~\gamma^{\prime},$
 satisfy the Riccati difference equations
-$\upsilon_{2(n+1)}=\mbox{\rm Ricc}_{\varpi}\left(\upsilon_{2n}\right)\quad\mbox{and}\quad\upsilon_{2n+1}=\mbox{\rm Ricc}_{\overline{\varpi}}\left(\upsilon_{2n-1}\right),\quad\text{for $n\geq 0$}.$ (103)
+(A.3) $\upsilon_{2(n+1)}=\mbox{\rm Ricc}_{\varpi}\left(\upsilon_{2n}\right)\quad\mbox{and}\quad\upsilon_{2n+1}=\mbox{\rm Ricc}_{\overline{\varpi}}\left(\upsilon_{2n-1}\right),\quad\text{for $n\geq 0$}.$
 Let us also define
@@
 Sinkhorn transitions
-We also denote by $G$ a $d$-dimensional centred Gaussian random variable (r.v.) with unit covariance. The Sinkhorn transitions in (16) can be expressed as
-${\cal S}_{2n}(x,dy)=\mathbb{P}\left({\cal Z}_{2n}(x)\in dy\right)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}(y,dx)=\mathbb{P}\left({\cal Z}_{2n+1}(y)\in dx\right)$
+We also denote by $G$ a $d$-dimensional centred Gaussian random variable (r.v.) with unit covariance. The Sinkhorn transitions in (2.2) can be expressed as
+${\mathcal{S}}_{2n}(x,dy)=\mathbb{P}\left({\mathcal{Z}}_{2n}(x)\in dy\right)\quad\mbox{\rm and}\quad{\mathcal{S}}_{2n+1}(y,dx)=\mathbb{P}\left({\mathcal{Z}}_{2n+1}(y)\in dx\right)$
 with random maps defined, for any $n\geq 0$, by
-${\cal Z}_{2n}(x)$ $=$ $m_{2n}+\beta_{2n}(x-m)+\tau_{2n}^{1/2}~G\quad\text{and}$
-${\cal Z}_{2n+1}(y)$ $=$ $m_{2n+1}+\beta_{2n+1}(y-\overline{m})+\tau_{2n+1}^{1/2}~G.$
+${\mathcal{Z}}_{2n}(x)$ $=$ $m_{2n}+\beta_{2n}(x-m)+\tau_{2n}^{1/2}~G\quad\text{and}$
+${\mathcal{Z}}_{2n+1}(y)$ $=$ $m_{2n+1}+\beta_{2n+1}(y-\overline{m})+\tau_{2n+1}^{1/2}~G.$
 The parameters $m_{n}$ in the latter expressions are defined, for any $n\geq 0$, by the recursion
@@
 The transitions of the Gibbs-loop process introduced in Section 2.2 can also be rewritten as
-${\cal S}^{\circ}_{2n}(y_{1},dy_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n}(y_{1})\in dy_{2}\right)\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1}(x_{1},dx_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n+1}(x_{1})\in dx_{2}\right)$
+${\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2}):=\mathbb{P}\left({\mathcal{Z}}^{\circ}_{2n}(y_{1})\in dy_{2}\right)\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2n+1}(x_{1},dx_{2}):=\mathbb{P}\left({\mathcal{Z}}^{\circ}_{2n+1}(x_{1})\in dx_{2}\right)$
 with the random maps
-${\cal Z}^{\circ}_{2n}(y)$ $:=$ $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G$
-${\cal Z}^{\circ}_{2n+1}(x)$ $:=$ $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~G$
+${\mathcal{Z}}^{\circ}_{2n}(y)$ $:=$ $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G$
+${\mathcal{Z}}^{\circ}_{2n+1}(x)$ $:=$ $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~G$
 defined in terms of the parameters
@@
 Lyapunov functions
-From the fixed point equations (22) we have
-$\overline{\sigma}=\beta^{\circ}_{2n}\overline{\sigma}\left(\beta^{\circ}_{2n}\right)^{\prime}+\tau^{\circ}_{2n}\geq\tau^{\circ}_{2n}\quad\mbox{and}\quad\sigma=\beta^{\circ}_{2n+1}\sigma\left(\beta^{\circ}_{2n+1}\right)^{\prime}+\tau^{\circ}_{2n+1}\geq\tau^{\circ}_{2n+1},$ (104)
+From the fixed point equations (2.8) we have
+(A.4) $\overline{\sigma}=\beta^{\circ}_{2n}\overline{\sigma}\left(\beta^{\circ}_{2n}\right)^{\prime}+\tau^{\circ}_{2n}\geq\tau^{\circ}_{2n}\quad\mbox{and}\quad\sigma=\beta^{\circ}_{2n+1}\sigma\left(\beta^{\circ}_{2n+1}\right)^{\prime}+\tau^{\circ}_{2n+1}\geq\tau^{\circ}_{2n+1},$
 which, in turn, imply that
-$\overline{\sigma}_{\varpi}:=\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}^{\circ}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma_{\overline{\varpi}}:=\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}^{\circ}\leq\sigma.$ (105)
+(A.5) $\overline{\sigma}_{\varpi}:=\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}^{\circ}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma_{\overline{\varpi}}:=\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}^{\circ}\leq\sigma.$
 Also note that if $\overline{\sigma}^{-1}\leq(\tau_{2n}^{\circ})^{-1}$ then $(\tau_{2n}^{\circ})^{1/2}\overline{\sigma}^{-1}(\tau_{2n}^{\circ})^{1/2}\leq I$. Moreover, we also readily see that
-$\begin{array}[]{l}\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\\ \\ =(y-\overline{m})^{\prime}\left(\beta^{\circ}_{2n}\right)^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)\\ \\ =(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)^{\prime}\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)\right]\overline{\sigma}^{-1/2}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right),\end{array}$
+$\begin{array}[]{l}\mathbb{E}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]\\ \\ =(y-\overline{m})^{\prime}\left(\beta^{\circ}_{2n}\right)^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)\\ \\ =(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)^{\prime}\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)\right]\overline{\sigma}^{-1/2}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right),\end{array}$
 which implies
-$\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]$ $=$ $(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[I-\upsilon_{2n}\right]^{2}\overline{\sigma}^{-1/2}(y-\overline{m})$ (106)
+(A.6) $\mathbb{E}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]$ $=$ $(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[I-\upsilon_{2n}\right]^{2}\overline{\sigma}^{-1/2}(y-\overline{m})$
 $+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)$
@@
 $\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$
-The estimate in (106) ensures that
-${\cal S}_{2n}^{\circ}(\psi)(y)=\mathbb{E}\left[\psi({\cal Z}^{\circ}_{2n}(y))\right]\leq(1+\lambda_{\sf min}(\varpi))^{-2}~\psi(y)+\left(\left(1-(1+\lambda_{\sf min}(\varpi))^{-2}\right)+d\right).$
+The estimate in (A.6) ensures that
+${\mathcal{S}}_{2n}^{\circ}(\psi)(y)=\mathbb{E}\left[\psi({\mathcal{Z}}^{\circ}_{2n}(y))\right]\leq(1+\lambda_{\sf min}(\varpi))^{-2}~\psi(y)+\left(\left(1-(1+\lambda_{\sf min}(\varpi))^{-2}\right)+d\right).$
 Next, let us observe that
-$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})=\frac{1}{\sqrt{\mbox{\rm det}(2\pi\tau_{2n}^{\circ})}}\\ \\ \displaystyle\times~\exp{\left(-\frac{1}{2}~\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\right)}~dy_{2}\end{array}$
+$\begin{array}[]{l}\displaystyle{\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2})=\frac{1}{\sqrt{\mbox{\rm det}(2\pi\tau_{2n}^{\circ})}}\\ \\ \displaystyle\times~\exp{\left(-\frac{1}{2}~\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\right)}~dy_{2}\end{array}$
 and
 $\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left(y_{2}-\overline{m}\right)+2\left(\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left(\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left((y_{2}-\overline{m})\right)+2~(1+\lambda_{\sf min}(\varpi^{-1}))~\left(y_{1}-\overline{m}\right)^{\prime}(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\left(y_{1}-\overline{m}\right),\end{array}$
-where the second inequality follows from (105). Moreover, arguing again as above we arrive at
-$\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)+2~(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\end{array}$ (107)
-The inequality (107) above together with (104) readily yields
-$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\frac{1}{\sqrt{\mbox{\rm det}(2\pi\overline{\sigma})}}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\right)}\\ \\ \displaystyle\times~\exp{\left(-\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)\right)}~dy_{2}.\end{array}$
+where the second inequality follows from (A.5). Moreover, arguing again as above we arrive at
+(A.7) $\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)+2~(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right).\end{array}$
+The inequality (A.7) above together with (A.4) readily yields
+$\begin{array}[]{l}\displaystyle{\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2})\geq\frac{1}{\sqrt{\mbox{\rm det}(2\pi\overline{\sigma})}}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\right)}\\ \\ \displaystyle\times~\exp{\left(-\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)\right)}~dy_{2}.\end{array}$
 Thus for any $y_{1}$ such that
@@
 we have
-$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\varepsilon(r)~\nu_{\overline{m},\overline{\sigma}_{\varpi}/2}(dy_{2})\end{array}$
+$\begin{array}[]{l}\displaystyle{\mathcal{S}}^{\circ}_{2n}(y_{1},dy_{2})\geq\varepsilon(r)~\nu_{\overline{m},\overline{\sigma}_{\varpi}/2}(dy_{2}),\end{array}$
 where
 $\varepsilon(r)=\frac{1}{\sqrt{\mbox{\rm det}(2\overline{\sigma}\overline{\sigma}_{\varpi}^{-1})}}~\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~(r-1)\right)}.$
-This shows that the function $\psi$, and similarly the function $\varphi$ defined in (91) satisfy the drift-minorization conditions stated in (53) and (54).
-Lyapunov functions in (92) and (93)
+This shows that the function $\psi$, and similarly the function $\varphi$ defined in (5.10) satisfy the drift-minorization conditions stated in (4.6) and (4.7).
+Lyapunov functions in (5.11) and (5.12)
 Using the triangle inequality for Mahalanobis distances we check that
-$\begin{array}[]{l}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\\ \\ \leq\left[(\beta^{\circ}_{2n}(y-\overline{m}))^{\prime}\overline{\sigma}^{-1}(\beta^{\circ}_{2n}(y-\overline{m}))\right]^{1/2}+\left[(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)^{\prime}\overline{\sigma}^{-1}(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)\right]^{1/2}\\ \\ \leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2},\end{array}$
+$\begin{array}[]{l}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\\ \\ \leq\left[(\beta^{\circ}_{2n}(y-\overline{m}))^{\prime}\overline{\sigma}^{-1}(\beta^{\circ}_{2n}(y-\overline{m}))\right]^{1/2}+\left[(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)^{\prime}\overline{\sigma}^{-1}(\left(\tau^{\circ}_{2n}\right)^{1/2}~G)\right]^{1/2}\\ \\ \leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2},\end{array}$
 which yields the almost sure estimate
-$\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2}.$
-This implies that the function $\psi$, and similarly the function $\varphi$ defined in (92) satisfy the drift-minorization conditions stated in (53) and (54).
-The analysis of the Lyapunov functions (93) is based on the following technical lemma.
+$\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\leq(1+\lambda_{\sf min}(\varpi))^{-1}~((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2}.$
+This implies that the function $\psi$, and similarly the function $\varphi$ defined in (5.11) satisfy the drift-minorization conditions stated in (4.6) and (4.7).
+The analysis of the Lyapunov functions (5.12) is based on the following technical lemma.
 Lemma A.1.
@@
 Using Lemma A.1 we check that
-$\begin{array}[]{l}\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\right)}\right]\\ \\ =\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)^{\prime}\overline{\sigma}^{-1}\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)\right]\right)}\right]\\ \\ =\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2})}}~\exp{\left(\frac{\delta}{2}~(y-\overline{m})^{\prime}~(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}~(y-\overline{m})\right)}.\end{array}$
+$\begin{array}[]{l}\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\mathcal{Z}}^{\circ}_{2n}(y)-\overline{m})\right]\right)}\right]\\ \\ =\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)^{\prime}\overline{\sigma}^{-1}\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~G\right)\right]\right)}\right]\\ \\ =\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2})}}~\exp{\left(\frac{\delta}{2}~(y-\overline{m})^{\prime}~(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}~(y-\overline{m})\right)}.\end{array}$
 On the other hand, we have
@@
 and arguing as above we conclude that
-$(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}$ $=$ $(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n})^{\prime}\left(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\right)^{-1}\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}$ (108)
-$\leq$ $\frac{1}{1-\delta}~(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\leq\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}~\overline{\sigma}^{-1}.$ (109)
+(A.8) $(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}$ $=$ $(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n})^{\prime}\left(I-\delta~\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\right)^{-1}\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}$
+(A.9) $\leq$ $\frac{1}{1-\delta}~(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\leq\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}~\overline{\sigma}^{-1}.$
 If we now consider the function
@@
 then we have
-$\frac{1}{\psi(y)}~\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)$ $\leq$ $\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~I)}}$$\times$
+$\frac{1}{\psi(y)}~\mathbb{E}\left(\psi\left({\mathcal{Z}}^{\circ}_{2n}(y)\right)\right)$ $\leq$ $\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~I)}}$$\times$
 $\times~\exp{\left(-\frac{\delta}{2}~\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~(y-\overline{m})^{\prime}~\overline{\sigma}^{-1}~(y-\overline{m})\right)}.$
@@
 we check that
-$\frac{1}{(1-\delta)(1+\lambda_{\sf min}(\varpi))^{2}}<1\quad\text{and}\quad\lim_{|y|\to\infty}\frac{1}{\psi(y)}~{\cal S}^{\circ}_{2n}(\psi)(y)=0.$
-This implies that the function $\psi$, and similarly the function $\varphi$, defined in (93) satisfy the drift-minorization conditions stated in (53) and (54).
+$\frac{1}{(1-\delta)(1+\lambda_{\sf min}(\varpi))^{2}}<1\quad\text{and}\quad\lim_{|y|\to\infty}\frac{1}{\psi(y)}~{\mathcal{S}}^{\circ}_{2n}(\psi)(y)=0.$
+This implies that the function $\psi$, and similarly the function $\varphi$, defined in (5.12) satisfy the drift-minorization conditions stated in (4.6) and (4.7).
 Appendix B Hilbert metric contractions
 We recall the Hilbert contraction coefficient formula ([bauer, birkhoff, birkhoff-2, hopf], see also [atar, cohen, legland])
-$\raisebox{2.0pt}{$\chi$}_{H}({\cal K})=\tanh{\left(\frac{1}{4}\log\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})\right)}.$ (110)
+(B.1) $\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{K}})=\tanh{\left(\frac{1}{4}\log\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}H(\delta_{x_{1}}{\mathcal{K}},\delta_{x_{2}}{\mathcal{K}})\right)}.$
 Also note that
-$H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})=\sup_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)$
-with the Radon-Nikodym function $\imath_{x_{1},x_{2}}({\cal K})(y)$ defined in (37). We also have
-$\sup_{(x_{1},x_{2})\in S^{2}}H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})=1/\hbar({\cal K})$
-with the parameter $\hbar({\cal K})$ defined in (38). Using the fact that
+$H(\delta_{x_{1}}{\mathcal{K}},\delta_{x_{2}}{\mathcal{K}})=\sup_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\mathcal{K}})(y_{1})~\imath_{x_{2},x_{1}}({\mathcal{K}})(y_{2})\right)$
+with the Radon-Nikodym function $\imath_{x_{1},x_{2}}({\mathcal{K}})(y)$ defined in (3.7). We also have
+$\sup_{(x_{1},x_{2})\in S^{2}}H(\delta_{x_{1}}{\mathcal{K}},\delta_{x_{2}}{\mathcal{K}})=1/\hbar({\mathcal{K}})$
+with the parameter $\hbar({\mathcal{K}})$ defined in (3.8). Using the fact that
 $\tanh\left(\frac{1}{4}~\log{u}\right)=\frac{u^{1/4}-u^{-1/4}}{u^{1/4}-u^{-1/4}}=\frac{u^{1/2}-1}{u^{1/2}+1}=\frac{1-u^{-1/2}}{1+u^{-1/2}}$
 we conclude that
-$\raisebox{2.0pt}{$\chi$}_{H}({\cal K})=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}}.$ (111)
+(B.2) $\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{K}})=\frac{1-\sqrt{\hbar({\mathcal{K}})}}{1+\sqrt{\hbar({\mathcal{K}})}}.$
 Proposition B.1.
-Consider the collection of Sinkhorn transitions ${\cal S}_{n}$ defined in (15). For any $n\geq 0$,
-$\hbar({\cal S}_{n})=\hbar({\cal Q})\quad\mbox{and}\quad\raisebox{2.0pt}{$\chi$}_{H}({\cal S}_{n})=\frac{1-\sqrt{\hbar({\cal Q})}}{1+\sqrt{\hbar({\cal Q})}}.$
+Consider the collection of Sinkhorn transitions ${\mathcal{S}}_{n}$ defined in (2.1). For any $n\geq 0$,
+$\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{Q}})\quad\mbox{and}\quad\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{S}}_{n})=\frac{1-\sqrt{\hbar({\mathcal{Q}})}}{1+\sqrt{\hbar({\mathcal{Q}})}}.$
 Proof.
 For any $n\geq 0$, we have
-$\imath_{y_{1},y_{2}}({\cal S}_{2n+1})(x_{1})~\imath_{y_{2},y_{1}}({\cal S}_{2n+1})(x_{2})$ $=$ $\frac{(d\delta_{x_{1}}{\cal S}_{2n}/d\pi_{2n})(y_{1})}{(d\delta_{x_{1}}{\cal S}_{2n}/d\pi_{2n})(y_{2})}\frac{(d\delta_{x_{2}}{\cal S}_{2n}/d\pi_{2n})(y_{2})}{(d\delta_{x_{2}}{\cal S}_{2n}/d\pi_{2n})(y_{1})}$
-$=$ $\imath_{x_{2},x_{1}}({\cal S}_{2n})(y_{2})~\imath_{x_{1},x_{2}}({\cal S}_{2n})(y_{1})$
-which, together with (38), yields
-$\hbar({\cal S}_{2n+1})=\hbar({\cal S}_{2n}).$
+$\imath_{y_{1},y_{2}}({\mathcal{S}}_{2n+1})(x_{1})~\imath_{y_{2},y_{1}}({\mathcal{S}}_{2n+1})(x_{2})$ $=$ $\frac{(d\delta_{x_{1}}{\mathcal{S}}_{2n}/d\pi_{2n})(y_{1})}{(d\delta_{x_{1}}{\mathcal{S}}_{2n}/d\pi_{2n})(y_{2})}\frac{(d\delta_{x_{2}}{\mathcal{S}}_{2n}/d\pi_{2n})(y_{2})}{(d\delta_{x_{2}}{\mathcal{S}}_{2n}/d\pi_{2n})(y_{1})}$
+$=$ $\imath_{x_{2},x_{1}}({\mathcal{S}}_{2n})(y_{2})~\imath_{x_{1},x_{2}}({\mathcal{S}}_{2n})(y_{1})$
+which, together with (3.8), yields
+$\hbar({\mathcal{S}}_{2n+1})=\hbar({\mathcal{S}}_{2n}).$
 In the same vein, for any $n\geq 1$ we have
-$\imath_{x_{1},x_{2}}({\cal S}_{2n})(y_{1})~\imath_{x_{2},x_{1}}({\cal S}_{2n})(y_{2})$ $=$ $\frac{(d\delta_{y_{1}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{1})}{(d\delta_{y_{1}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{2})}~\frac{(d\delta_{y_{2}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{2})}{(d\delta_{y_{2}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{1})}$
-$=$ $\imath_{y_{2},y_{1}}({\cal S}_{2n-1})(x_{2})~\imath_{y_{1},y_{2}}({\cal S}_{2n-1})(x_{1}),$
+$\imath_{x_{1},x_{2}}({\mathcal{S}}_{2n})(y_{1})~\imath_{x_{2},x_{1}}({\mathcal{S}}_{2n})(y_{2})$ $=$ $\frac{(d\delta_{y_{1}}{\mathcal{S}}_{2n-1}/d\pi_{2n-1})(x_{1})}{(d\delta_{y_{1}}{\mathcal{S}}_{2n-1}/d\pi_{2n-1})(x_{2})}~\frac{(d\delta_{y_{2}}{\mathcal{S}}_{2n-1}/d\pi_{2n-1})(x_{2})}{(d\delta_{y_{2}}{\mathcal{S}}_{2n-1}/d\pi_{2n-1})(x_{1})}$
+$=$ $\imath_{y_{2},y_{1}}({\mathcal{S}}_{2n-1})(x_{2})~\imath_{y_{1},y_{2}}({\mathcal{S}}_{2n-1})(x_{1}),$
 hence
-$\hbar({\cal S}_{2n})=\hbar({\cal S}_{2n-1}).$
-Recalling that ${\cal S}_{0}={\cal Q}$, this ends the proof of the proposition.
+$\hbar({\mathcal{S}}_{2n})=\hbar({\mathcal{S}}_{2n-1}).$
+Recalling that ${\mathcal{S}}_{0}={\mathcal{Q}}$, this ends the proof of the proposition.
 Appendix C Weighted norm contractions
-We associate with some $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ the following condition.
-Condition C.1 (${\cal L}_{\varphi,\psi}({\cal K})$).
-There exists some parameters $\epsilon\in]0,1[$ and $c<\infty$ such that
-${\cal K}(\psi)\leq\epsilon~\varphi+c.$ (112)
+We associate with some $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $\psi\in{\mathcal{B}}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ the following condition.
+${\mathcal{L}}_{\varphi,\psi}({\mathcal{K}})$ There exists some parameters $\epsilon\in]0,1[$ and $c<\infty$ such that
+(C.1) ${\mathcal{K}}(\psi)\leq\epsilon~\varphi+c.$
 In addition, there exist some constant $r_{0}\geq 1$ and some function
@@
 such that, for any $r\geq r_{0}$ and $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$,
-$\chi_{\varphi}(r,{\cal K}):=\sup_{(x_{1},x_{2})\in C_{\varphi}(r)^{2}}\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}<1-\varepsilon(r).$ (113)
-By (33), condition (113) is equivalent to the existence of some ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$ (that may depends on $(r,x_{1},x_{2})$) satisfying for $i\in\{1,2\}$ the local minorization condition
-${\cal K}(x_{i},dy)\geq\varepsilon(r)~{\cal Y}(dy).$ (114)
-For instance, assume that $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $x_{1},x_{2}\in\mathbb{X}$ and set
-$\jmath_{\varphi,\psi}(r,{\cal K}):=\inf_{(x_{1},x_{2})\in C_{\varphi}(r)}~\inf_{y\in C_{\psi}(r)}\imath_{x_{1},x_{2}}({\cal K})(y).$ (115)
-In this notation, we see that if $\jmath_{\varphi,\psi}(r,{\cal K})\geq\varepsilon(r)$ then (114) holds.
-Note that the strength of conditions (112), (113) depends on the strength of the functions $(\varphi,\psi)$.
-When the functions are bounded, the geometric drift condition (112) is trivially met, but in this case (114) is a uniform contraction condition on the space $\mathbb{X}$.
-When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact and condition (114) is met as soon as we have
-${\cal K}(x,dy)\geq q(x,y)~\nu(dy)$ (116)
+(C.2) $\chi_{\varphi}(r,{\mathcal{K}}):=\sup_{(x_{1},x_{2})\in C_{\varphi}(r)^{2}}\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{K}}\|_{\sf tv}<1-\varepsilon(r).$
+By (3.3), condition (C.2) is equivalent to the existence of some ${\mathcal{Y}}\in{\mathcal{M}}_{1}(\mathbb{Y})$ (that may depends on $(r,x_{1},x_{2})$) satisfying for $i\in\{1,2\}$ the local minorization condition
+(C.3) ${\mathcal{K}}(x_{i},dy)\geq\varepsilon(r)~{\mathcal{Y}}(dy).$
+For instance, assume that $\delta_{x_{1}}{\mathcal{K}}\ll\delta_{x_{2}}{\mathcal{K}}$ for any $x_{1},x_{2}\in\mathbb{X}$ and set
+(C.4) $\jmath_{\varphi,\psi}(r,{\mathcal{K}}):=\inf_{(x_{1},x_{2})\in C_{\varphi}(r)}~\inf_{y\in C_{\psi}(r)}\imath_{x_{1},x_{2}}({\mathcal{K}})(y).$
+In this notation, we see that if $\jmath_{\varphi,\psi}(r,{\mathcal{K}})\geq\varepsilon(r)$ then (C.3) holds.
+Note that the strength of conditions (C.1), (C.2) depends on the strength of the functions $(\varphi,\psi)$.
+When the functions are bounded, the geometric drift condition (C.1) is trivially met, but in this case (C.3) is a uniform contraction condition on the space $\mathbb{X}$.
+When $\varphi\in{\mathcal{B}}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact and condition (C.3) is met as soon as we have
+(C.5) ${\mathcal{K}}(x,dy)\geq q(x,y)~\nu(dy)$
 for some Radon positive measure $\nu$ on $\mathbb{Y}$ and some density function $q$, satisfying the local minorization condition
-$\inf_{C_{\varphi}(r)\times C_{\psi}(r)}q>0\quad\mbox{\rm and}\quad 0<\nu(C_{\psi}(r))<\infty$ (117)
+(C.6) $\inf_{C_{\varphi}(r)\times C_{\psi}(r)}q>0\quad\mbox{\rm and}\quad 0<\nu(C_{\psi}(r))<\infty$
 for any $r\geq r_{0}$.
@@
 Also note that the l.h.s.
-minorization condition (117) is satisfied when $(x,y)\in(\mathbb{X}\times\mathbb{Y})^{\circ}\mapsto q(x,y)$ is a continuous positive function on the interior $(\mathbb{X}\times\mathbb{Y})^{\circ}:=(\mathbb{X}^{\circ}\times\mathbb{Y}^{\circ})$ of $(\mathbb{X}\times\mathbb{Y})$.
+minorization condition (C.6) is satisfied when $(x,y)\in(\mathbb{X}\times\mathbb{Y})^{\circ}\mapsto q(x,y)$ is a continuous positive function on the interior $(\mathbb{X}\times\mathbb{Y})^{\circ}:=(\mathbb{X}^{\circ}\times\mathbb{Y}^{\circ})$ of $(\mathbb{X}\times\mathbb{Y})$.
 Also note that
@@
 Indeed, if the inequality above holds then for any $0<\epsilon<\|\Theta\|$, the set $A_{\epsilon}:=\{\Theta\geq\epsilon\}=C_{1/\Theta}(1/\epsilon)$ is a non empty compact subset and we have
-${\cal K}(\psi)\leq\epsilon~1_{S-A_{\epsilon}}\varphi+1_{A_{\epsilon}}(\Theta~\varphi)\leq\epsilon~\varphi+c_{\epsilon}\quad\mbox{\rm with}\quad c_{\epsilon}:=\|\Theta\|\sup_{A_{\epsilon}}\varphi.$ (118)
+(C.8) ${\mathcal{K}}(\psi)\leq\epsilon~1_{S-A_{\epsilon}}\varphi+1_{A_{\epsilon}}(\Theta~\varphi)\leq\epsilon~\varphi+c_{\epsilon}\quad\mbox{\rm with}\quad c_{\epsilon}:=\|\Theta\|\sup_{A_{\epsilon}}\varphi.$
 Theorem C.1.
-Assume Condition ${\cal L}_{\varphi,\psi}({\cal K})$ is satisfied for some uniformly positive functions $(\varphi,\psi)\in({\cal B}(\mathbb{X})\times{\cal B}(\mathbb{Y}))$.
-Then, there exist constants $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (112) and on the function $\varepsilon(r)$ in (113) and such that
-$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\cal K})\leq\rho\quad\mbox{with}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (119)
-Theorem C.2 is a consequence of Lemma 2.3 in [horton-dp], see also Theorem 8.2.21 in [penev] and Lemma 2.1 in [dpa].
-Several examples satisfying the Lyapunov condition (112) and the local minorization conditions (119) are discussed in [dpa, penev].
+Assume Condition ${\mathcal{L}}_{\varphi,\psi}({\mathcal{K}})$ is satisfied for some uniformly positive functions $(\varphi,\psi)\in({\mathcal{B}}(\mathbb{X})\times{\mathcal{B}}(\mathbb{Y}))$.
+Then, there exist constants $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (C.1) and on the function $\varepsilon(r)$ in (C.2) and such that
+(C.9) $\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\mathcal{K}})\leq\rho\quad\mbox{with}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$
+Theorem C.1 is a consequence of Lemma 2.3 in [horton-dp], see also [delgerber2025], as well as Theorem 8.2.21 in [penev] and Lemma 2.1 in [dpa].
+Several examples satisfying the Lyapunov condition (C.1) and the local minorization conditions (C.9) are discussed in [dpa, penev].
 Proof of Theorem 4.6
-We check that if Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (53) and (54) also hold for any $n>n_{0}$.
-Assume that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$. From (60), for any $n>n_{0}$ we have the estimates
-${\cal S}^{\circ}_{2n}(\psi)$ $=$ ${\cal S}_{2n-1}({\cal S}_{2n}(\psi))\leq\epsilon~{\cal S}_{2n-1}(\varphi)+c\leq\epsilon^{2}~\psi+(1+\epsilon)~c\quad\text{and}$
-${\cal S}^{\circ}_{2n+1}(\varphi)$ $=$ ${\cal S}_{2n}({\cal S}_{2n+1}(\varphi))\leq\epsilon~{\cal S}_{2n}(\psi)+c\leq\epsilon^{2}~\varphi+(1+\epsilon)~c.$ (120)
-Also note that for any $n>n_{0}$ the inequality (61) in Condition ${\cal D}_{\varphi,\psi}({\cal S})$ implies that
-$\|((\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n-1}){\cal S}_{2n}\|_{\sf tv}\vee\|((\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (121)
-The estimates in (120) imply that (53) holds, while the inequality (121) implies that (54) is satisfied.
-Moreover, combining the fixed point equations in (22) with (120) we arrive at
-$\sup_{n\geq n_{0}}\left(\|{\cal S}_{2n}(\psi)\|_{\varphi}\vee\|{\cal S}_{2n+1}(\varphi)\|_{\psi}\right)<\infty\quad\mbox{\rm and}\quad(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))\leq\frac{c}{1-\epsilon},$ (122)
-and (60) together with (122) yields
+We check that if Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied for some integer $n_{0}\geq 0$, then (4.6) and (4.7) also hold for any $n>n_{0}$.
+Assume that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ holds for some $n_{0}\geq 0$. From (4.13), for any $n>n_{0}$ we have the estimates
+${\mathcal{S}}^{\circ}_{2n}(\psi)$ $=$ ${\mathcal{S}}_{2n-1}({\mathcal{S}}_{2n}(\psi))\leq\epsilon~{\mathcal{S}}_{2n-1}(\varphi)+c\leq\epsilon^{2}~\psi+(1+\epsilon)~c\quad\text{and}$
+(C.10) ${\mathcal{S}}^{\circ}_{2n+1}(\varphi)$ $=$ ${\mathcal{S}}_{2n}({\mathcal{S}}_{2n+1}(\varphi))\leq\epsilon~{\mathcal{S}}_{2n}(\psi)+c\leq\epsilon^{2}~\varphi+(1+\epsilon)~c.$
+Also note that for any $n>n_{0}$ the inequality (4.14) in Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ implies that
+(C.11) $\|((\delta_{y_{1}}-\delta_{y_{2}}){\mathcal{S}}_{2n-1}){\mathcal{S}}_{2n}\|_{\sf tv}\vee\|((\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{S}}_{2n}){\mathcal{S}}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$
+The estimates in (C.10) imply that (4.6) holds, while the inequality (C.11) implies that (4.7) is satisfied.
+Moreover, combining the fixed point equations in (2.8) with (C.10) we arrive at
+(C.12) $\sup_{n\geq n_{0}}\left(\|{\mathcal{S}}_{2n}(\psi)\|_{\varphi}\vee\|{\mathcal{S}}_{2n+1}(\varphi)\|_{\psi}\right)<\infty\quad\mbox{\rm and}\quad(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))\leq\frac{c}{1-\epsilon},$
+and (4.13) together with (C.12) yields
 $\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))\leq c_{1}~(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))$
-for some $c_{1}>0$. Therefore, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied.
-Finally, we observe that (60) ensures that inequality (112) is satisfied for the Markov transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$, while (61) yields
-$\chi_{\varphi}(r,{\cal S}_{2n})\vee\chi_{\psi}(r,{\cal S}_{2n+1})\leq 1-\varepsilon(r)$
-and, hence, the inequality (113) also holds for ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
+for some $c_{1}>0$. Therefore, Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is satisfied.
+Finally, we observe that (4.13) ensures that inequality (C.1) is satisfied for the Markov transitions ${\mathcal{S}}_{2n}$ and ${\mathcal{S}}_{2n+1}$, while (4.14) yields
+$\chi_{\varphi}(r,{\mathcal{S}}_{2n})\vee\chi_{\psi}(r,{\mathcal{S}}_{2n+1})\leq 1-\varepsilon(r)$
+and, hence, the inequality (C.2) also holds for ${\mathcal{S}}_{2n}$ and ${\mathcal{S}}_{2n+1}$.
 As a consequence, Theorem C.1 can be applied to show that there are parameters $\kappa\geq 0$ and $\rho\in[0,1[$ such that
-$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\cal S}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{{\psi_{\kappa},\varphi_{\kappa}}}({\cal S}_{2n+1})\leq\rho.$
-Appendix D Schrödinger potential functions
-D.1 Commutation formulae
+$\raisebox{2.0pt}{$\chi$}_{{\varphi_{\kappa},\psi_{\kappa}}}({\mathcal{S}}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{{\psi_{\kappa},\varphi_{\kappa}}}({\mathcal{S}}_{2n+1})\leq\rho.$
+Appendix D Sinkhorn potential functions
+This appendix collects a number of technical results concerning the Sinkhorn potential functions associated with the alternating Sinkhorn dynamics.
+We first establish a set of commutation formulae linking the successive marginals and Radon–Nikodym derivatives along the iteration.
+These identities are then used to derive quantitative bounds on the Sinkhorn potentials and on the corresponding density ratios, both in uniform and weighted norms.
+Under suitable integrability and Lyapunov-type assumptions, we further obtain refined estimates leading to drift and local minorization inequalities for the Sinkhorn transition kernels.
+Together, these results provide the analytical backbone needed to verify the contraction and stability properties used in the proofs of the main theorems.
+D.1. Commutation formulae
 For any $n\geq 0$ we have the commutation formulae
-${\cal S}_{2n}\left(\frac{d\nu_{V}}{d\pi_{2n}}\right)=\frac{d\pi_{2n+1}}{d\lambda_{U}}\quad\mbox{and}\quad{\cal S}_{2(n+1)}\left(\frac{d\pi_{2n}}{d\nu_{V}}\right)=\frac{d\lambda_{U}}{d\pi_{2n+1}}.$ (123)
+(D.1) ${\mathcal{S}}_{2n}\left(\frac{d\nu_{V}}{d\pi_{2n}}\right)=\frac{d\pi_{2n+1}}{d\lambda_{U}}\quad\mbox{and}\quad{\mathcal{S}}_{2(n+1)}\left(\frac{d\pi_{2n}}{d\nu_{V}}\right)=\frac{d\lambda_{U}}{d\pi_{2n+1}}.$
 In the same vein, for any $n\geq 1$ we have
-${\cal S}_{2n-1}\left(\frac{d\lambda_{U}}{d\pi_{2n-1}}\right)=\frac{d\pi_{2n}}{d\nu_{V}}\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)=\frac{d\nu_{V}}{d\pi_{2n}}.$ (124)
-A proof for (123) and (124) is provided in Appendix E (on page E).
-Using the commutation properties (123) and (124) for any $n\geq p$ we readily check that
-$\frac{d\pi_{2n}}{d\nu_{V}}$ $=$ ${\cal S}^{\circ}_{2n}\left(\frac{d\pi_{2(n-1)}}{d\nu_{V}}\right)={\cal S}^{\circ}_{2n,2p}\left(\frac{d\pi_{2p}}{d\nu_{V}}\right)\quad\text{and}$
-$\frac{d\pi_{2n+1}}{d\lambda_{U}}$ $=$ ${\cal S}_{2n+1}^{\circ}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)={\cal S}^{\circ}_{2n+1,2p-1}\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right),$ (125)
+(D.2) ${\mathcal{S}}_{2n-1}\left(\frac{d\lambda_{U}}{d\pi_{2n-1}}\right)=\frac{d\pi_{2n}}{d\nu_{V}}\quad\mbox{and}\quad{\mathcal{S}}_{2n+1}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)=\frac{d\nu_{V}}{d\pi_{2n}}.$
+A proof for (D.1) and (D.2) is provided in Appendix E (on page E).
+Using the commutation properties (D.1) and (D.2) for any $n\geq p$ we readily check that
+$\frac{d\pi_{2n}}{d\nu_{V}}$ $=$ ${\mathcal{S}}^{\circ}_{2n}\left(\frac{d\pi_{2(n-1)}}{d\nu_{V}}\right)={\mathcal{S}}^{\circ}_{2n,2p}\left(\frac{d\pi_{2p}}{d\nu_{V}}\right)\quad\text{and}$
+(D.3) $\frac{d\pi_{2n+1}}{d\lambda_{U}}$ $=$ ${\mathcal{S}}_{2n+1}^{\circ}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)={\mathcal{S}}^{\circ}_{2n+1,2p-1}\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right),$
 with the backward semigroups
-${\cal S}^{\circ}_{2n,2p}:={\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1,2p-1}:={\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1}.$
-By (123) and (124) we also obtain
-$\frac{d\lambda_{U}}{d\pi_{2n+1}}$ $=$ $\left({\cal S}_{2(n+1)}\,{\cal S}^{\circ}_{2n,2p}\,{\cal S}_{2p-1}\right)\left(\frac{d\lambda_{U}}{d\pi_{2p-1}}\right)\quad\text{and}$
-$\frac{d\nu_{V}}{d\pi_{2n}}$ $=$ $\left({\cal S}_{2n+1}\,{\cal S}^{\circ}_{2n-1,2p-1}\,{\cal S}_{2(p-1)}\right)\left(\frac{d\nu_{V}}{d\pi_{2(p-1)}}\right).$ (126)
-D.2 Some quantitative estimates
+${\mathcal{S}}^{\circ}_{2n,2p}:={\mathcal{S}}^{\circ}_{2n}\ldots{\mathcal{S}}^{\circ}_{2(p+1)}\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2n+1,2p-1}:={\mathcal{S}}^{\circ}_{2n+1}\ldots{\mathcal{S}}^{\circ}_{2p+1}.$
+By (D.1) and (D.2) we also obtain
+$\frac{d\lambda_{U}}{d\pi_{2n+1}}$ $=$ $\left({\mathcal{S}}_{2(n+1)}\,{\mathcal{S}}^{\circ}_{2n,2p}\,{\mathcal{S}}_{2p-1}\right)\left(\frac{d\lambda_{U}}{d\pi_{2p-1}}\right)\quad\text{and}$
+(D.4) $\frac{d\nu_{V}}{d\pi_{2n}}$ $=$ $\left({\mathcal{S}}_{2n+1}\,{\mathcal{S}}^{\circ}_{2n-1,2p-1}\,{\mathcal{S}}_{2(p-1)}\right)\left(\frac{d\nu_{V}}{d\pi_{2(p-1)}}\right).$
+D.2. Some quantitative estimates
 Proposition D.1.
-When $\jmath({\cal Q})>0$, for any $n\geq 0$ we have the uniform estimates
-$\|\frac{d\pi_{2n}}{d\nu_{V}}\|\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|\leq\hbar({\cal Q})^{-2}\quad\mbox{and}\quad\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|\leq\hbar({\cal Q})^{-2}.$ (127)
+When $\jmath({\mathcal{Q}})>0$, for any $n\geq 0$ we have the uniform estimates
+(D.5) $\|\frac{d\pi_{2n}}{d\nu_{V}}\|\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|\leq\hbar({\mathcal{Q}})^{-2}\quad\mbox{and}\quad\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|\leq\hbar({\mathcal{Q}})^{-2}.$
 Proof.
-Recalling that $\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})$, for any given $x\in\mathbb{X}$ we have
-$\hbar({\cal Q})~{\cal S}_{2n}(x,dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})=({\cal S}_{2n-1}{\cal S}_{2n})(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})}~{\cal S}_{2n}(x,dy_{2}).$
+Recalling that $\jmath({\mathcal{S}}_{n})\wedge\jmath({\mathcal{Q}})\geq\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{Q}})$, for any given $x\in\mathbb{X}$ we have
+$\hbar({\mathcal{Q}})~{\mathcal{S}}_{2n}(x,dy_{2})\leq{\mathcal{S}}_{2n}^{\circ}(y_{1},dy_{2})=({\mathcal{S}}_{2n-1}{\mathcal{S}}_{2n})(y_{1},dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})}~{\mathcal{S}}_{2n}(x,dy_{2}).$
 This yields, for any $(y_{1},\overline{y}_{1})\in\mathbb{Y}^{2}$, the estimate
-$\hbar({\cal Q})^{2}~{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})$
+$\hbar({\mathcal{Q}})^{2}~{\mathcal{S}}_{2n}^{\circ}(\overline{y}_{1},dy_{2})\leq{\mathcal{S}}_{2n}^{\circ}(y_{1},dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})^{2}}~{\mathcal{S}}_{2n}^{\circ}(\overline{y}_{1},dy_{2})$
 and integrating w.r.t. $\nu_{V}(d\overline{y}_{1})$ and $\pi_{2(n-1)}(dy_{1})$ we find that
-$\hbar({\cal Q})^{2}~\nu_{V}(dy_{2})\leq\pi_{2n}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~\nu_{V}(dy_{2}).$
+$\hbar({\mathcal{Q}})^{2}~\nu_{V}(dy_{2})\leq\pi_{2n}(dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})^{2}}~\nu_{V}(dy_{2}).$
 Similarly, integrating w.r.t. $\nu_{V}(dy_{1})$ and $\pi_{2(n-1)}(d\overline{y}_{1})$ we also obtain
-$\hbar({\cal Q})^{2}~\pi_{2n}(dy_{2})\leq\nu_{V}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~\pi_{2n}(dy_{2}).$
+$\hbar({\mathcal{Q}})^{2}~\pi_{2n}(dy_{2})\leq\nu_{V}(dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})^{2}}~\pi_{2n}(dy_{2}).$
 From the inequalities above we conclude that
-$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq\hbar({\cal Q})^{-2}\quad\text{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq\hbar({\cal Q})^{-2}.$
+$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq\hbar({\mathcal{Q}})^{-2}\quad\text{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq\hbar({\mathcal{Q}})^{-2}.$
 More generally these Radon Nikodym functions are unbounded and we need to carry a more refined analysis.
 Proposition D.2.
-If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds for some $n_{0}\geq 1$ then there exists $c_{\epsilon}>0$ such that
-$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$ (128)
+If Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds for some $n_{0}\geq 1$ then there exists $c_{\epsilon}>0$ such that
+(D.6) $\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$
 Proof.
-Using (125) and (53), we see that, for any $n\geq n_{0}$,
-$\frac{d\pi_{2n}}{d\nu_{V}}$ $\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2n_{0}}\right)\left(\psi\right)$
+Using (D.3) and (4.6), we see that, for any $n\geq n_{0}$,
+$\frac{d\pi_{2n}}{d\nu_{V}}$ $\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~\left({\mathcal{S}}^{\circ}_{2n}\ldots{\mathcal{S}}^{\circ}_{2n_{0}}\right)\left(\psi\right)$
 $\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\left(\epsilon^{n-n_{0}}\psi+\frac{c}{1-\epsilon}\right)\leq c_{\epsilon}~\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~\psi,\quad\text{with}~c_{\epsilon}:=\left(1+\frac{c}{1-\epsilon}\right).$
@@
 Proposition D.3.
-In addition, when Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, there exists $c_{\epsilon}>0$ such that
+In addition, when Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ holds for some $n_{0}\geq 0$, there exists $c_{\epsilon}>0$ such that
 $\sup_{n\geq n_{0}}\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)\leq c_{\epsilon}~\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right),$
@@
 Proof.
-If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$ then, by Theorem 4.8, the drift condition (53) also holds for any $n>n_{0}$, which implies that
+If Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied for some integer $n_{0}\geq 0$ then, by Theorem 4.6, the drift condition (4.6) also holds for any $n>n_{0}$, which implies that
 $\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\quad\text{and, similarly,}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}.$
-Arguing as in the proof of Proposition D.2, now combining the drift conditions (60) with (126), we find that
+Arguing as in the proof of Proposition D.2, now combining the drift conditions (4.13) with (D.4), we find that
 $\sup_{n\geq n_{0}}\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\leq c_{\epsilon}~\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\leq c_{\epsilon}~\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}.$
-Under the stronger Condition ${\cal H}_{\delta}(U,V)$, more refined estimates can be obtained. With that end, consider the functions
-${\cal W}_{V,U}(x):=\log{{\cal Q}_{V}(e^{W_{U}})(x)}\quad\mbox{\rm and}\quad{\cal W}^{U,V}(y):=\log{{\cal R}_{U}(e^{W^{V}})(y)},$ (129)
-with the integrated cost functions $(W_{U},W^{V})$ defined in (28) and the integral operators
-${\cal Q}_{V}(x,dy)=e^{-W(x,y)}~\nu_{V}(dy)\quad\mbox{\rm and}\quad{\cal R}_{U}(y,dx)=e^{-W(x,y)}~\lambda_{U}(dx).$
-Next technical lemma is pivotal.
+Under the stronger Condition ${\mathcal{H}}_{\delta}(U,V)$, more refined estimates can be obtained. With that end, consider the functions
+(D.7) ${\mathcal{W}}_{V,U}(x):=\log{{\mathcal{Q}}_{V}(e^{W_{U}})(x)}\quad\mbox{\rm and}\quad{\mathcal{W}}^{U,V}(y):=\log{{\mathcal{R}}_{U}(e^{W^{V}})(y)},$
+with the integrated cost functions $(W_{U},W^{V})$ defined in (2.14) and the integral operators
+${\mathcal{Q}}_{V}(x,dy)=e^{-W(x,y)}~\nu_{V}(dy)\quad\mbox{\rm and}\quad{\mathcal{R}}_{U}(y,dx)=e^{-W(x,y)}~\lambda_{U}(dx).$
+The next technical lemma is pivotal.
 Lemma D.4.
 For any $n\geq 0$ we have
-$\begin{array}[]{rcl}\log{{\cal Q}(\exp{(-V_{2n})})}&\geq&-W^{V}+\nu_{V}(V)\quad\text{and}\\ &&\\ -\nu_{V}(V)+{\cal W}^{U,V}\geq\log{{\cal R}(\exp{(-U_{2n})})}&\geq&-W_{U},\end{array}$ (130)
+(D.8) $\begin{array}[]{rcl}\log{{\mathcal{Q}}(\exp{(-V_{2n})})}&\geq&-W^{V}+\nu_{V}(V)\quad\text{and}\\ &&\\ -\nu_{V}(V)+{\mathcal{W}}^{U,V}\geq\log{{\mathcal{R}}(\exp{(-U_{2n})})}&\geq&-W_{U},\end{array}$
 as well as
@@
 In addition, for any $n\geq 1$ we have
-${\cal W}_{V,U}\geq\log{{\cal Q}(\exp{(-V_{2n})})}\geq-W^{V}+\nu_{V}(V)$ (131)
+(D.9) ${\mathcal{W}}_{V,U}\geq\log{{\mathcal{Q}}(\exp{(-V_{2n})})}\geq-W^{V}+\nu_{V}(V)$
 as well as
-$\begin{array}[]{rcccl}U-W^{V}+\nu_{V}(V)&\leq&U_{2n}&\leq&U+{\cal W}_{V,U}\quad\text{and}\\ &&&&\\ V-W_{U}&\leq&V_{2n}&\leq&V+{\cal W}^{U,V}-\nu_{V}(V).\end{array}$ (132)
+(D.10) $\begin{array}[]{rcccl}U-W^{V}+\nu_{V}(V)&\leq&U_{2n}&\leq&U+{\mathcal{W}}_{V,U}\quad\text{and}\\ &&&&\\ V-W_{U}&\leq&V_{2n}&\leq&V+{\mathcal{W}}^{U,V}-\nu_{V}(V).\end{array}$
 Proof.
-Formula (21) yields the following monotone properties
-$\begin{array}[]{rcccccl}\nu_{V}(V_{2(n+1)})&\leq&\nu_{V}(V_{2n})&=&\nu_{V}(V_{2(n+1)})+\mbox{\rm Ent}(\nu_{V}~|~\pi_{2n})&\leq&\nu_{V}(V_{0})=0,\\ &&&&&&\\ \lambda_{U}(U_{2(n+1)})&\leq&\lambda_{U}(U_{2n})&=&\lambda_{U}(U_{2(n+1)})+\mbox{\rm Ent}(\lambda_{U}~|~\pi_{2n+1})&\leq&\lambda_{U}(U_{0}).\end{array}$ (133)
+Formula (2.7) yields the following monotone properties
+(D.11) $\begin{array}[]{rcccccl}\nu_{V}(V_{2(n+1)})&\leq&\nu_{V}(V_{2n})&=&\nu_{V}(V_{2(n+1)})+\mbox{\rm Ent}(\nu_{V}~|~\pi_{2n})&\leq&\nu_{V}(V_{0})=0,\\ &&&&&&\\ \lambda_{U}(U_{2(n+1)})&\leq&\lambda_{U}(U_{2n})&=&\lambda_{U}(U_{2(n+1)})+\mbox{\rm Ent}(\lambda_{U}~|~\pi_{2n+1})&\leq&\lambda_{U}(U_{0}).\end{array}$
 Applying Jensen’s inequality, for any $n\geq 0$ we have
-$\log{{\cal Q}(\exp{(-V_{2n})})(x)}$ $=$ $\log{\int~\nu_{V}(dy)~\exp{(-W(x,y)+V(y)-V_{2n}(y))}}$
+$\log{{\mathcal{Q}}(\exp{(-V_{2n})})(x)}$ $=$ $\log{\int~\nu_{V}(dy)~\exp{(-W(x,y)+V(y)-V_{2n}(y))}}$
 $\geq$ $-W^{V}(x)+\nu_{V}(V)-\nu_{V}(V_{2n})\geq-W^{V}(x)+\nu_{V}(V).$
-By (133) we also have $\lambda_{U}(U)\geq\lambda_{U}(U_{2n})$ and, therefore, for any $n\geq 0$ we obtain
-$\log{{\cal R}(\exp{(-U_{2n})})(y)}$ $=$ $\log{\int~\lambda_{U}(dx)~\exp{(-W(x,y)+U(x)-U_{2n}(x))}}$
+By (D.11) we also have $\lambda_{U}(U)\geq\lambda_{U}(U_{2n})$ and, therefore, for any $n\geq 0$ we obtain
+$\log{{\mathcal{R}}(\exp{(-U_{2n})})(y)}$ $=$ $\log{\int~\lambda_{U}(dx)~\exp{(-W(x,y)+U(x)-U_{2n}(x))}}$
 $\geq$ $-W_{U}(y)+\lambda_{U}(U)-\lambda_{U}(U_{2n})\geq-W_{U}(y)$
-using Jensen’s inequality again. This ends the proof of (130).
-The integral equations in (18) imply that for any $n\geq 0$ we have
+using Jensen’s inequality again. This ends the proof of (D.8).
+The integral equations in (2.4) imply that for any $n\geq 0$ we have
 $U_{2n}\geq U-W^{V}+\nu_{V}(V)\quad\mbox{\rm and}\quad V_{2(n+1)}\geq V-W_{U}.$
-Using again (18) we see that
-$U_{2(n+1)}-U=\log{{\cal Q}(e^{-V_{2(n+1)}})}$ $\leq$ $\log{{\cal Q}_{V}(e^{W_{U}})}={\cal W}_{V,U}\quad\text{and}$
-$V_{2(n+1)}-V=\log{{\cal R}(e^{-U_{2n}})}$ $\leq$ $-\nu_{V}(V)+\log{{\cal R}_{U}(e^{W^{V}})}=-\nu_{V}(V)+{\cal W}^{U,V}.$
+Using again (2.4) we see that
+$U_{2(n+1)}-U=\log{{\mathcal{Q}}(e^{-V_{2(n+1)}})}$ $\leq$ $\log{{\mathcal{Q}}_{V}(e^{W_{U}})}={\mathcal{W}}_{V,U}\quad\text{and}$
+$V_{2(n+1)}-V=\log{{\mathcal{R}}(e^{-U_{2n}})}$ $\leq$ $-\nu_{V}(V)+\log{{\mathcal{R}}_{U}(e^{W^{V}})}=-\nu_{V}(V)+{\mathcal{W}}^{U,V}.$
 Proposition D.5.
-If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then
-$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}\vee\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty.$ (134)
+If Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then
+(D.12) $\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}\vee\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty.$
 In addition, the exists some constant $c>0$ such that, for any $n\geq 1$,
-$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq c~e^{W_{U}}\quad\mbox{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq c~e^{W^{V}}.$ (135)
+(D.13) $\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq c~e^{W_{U}}\quad\mbox{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq c~e^{W^{V}}.$
 Proof.
@@
 and conclude that
-$\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}$ $=$ $\|{\cal Q}_{V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}$
+$\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}$ $=$ $\|{\mathcal{Q}}_{V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}$
 $\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{V}\left(e^{W_{U}}\right)$
@@
 as well as
-$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{U}(e^{W^{V}})\|\leq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
-where $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$ by Condition ${\cal H}_{\delta}(U,V)$. This ends the proof of (134).
-Using (132) we see that, for any $n\geq 1$,
-${\cal W}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\cal W}_{V,U}+W^{V}-\nu_{V}(V)$
+$\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\mathcal{R}}_{U}(e^{W^{V}})\|\leq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
+where $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$ by Condition ${\mathcal{H}}_{\delta}(U,V)$. This ends the proof of (D.12).
+Using (D.10) we see that, for any $n\geq 1$,
+${\mathcal{W}}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\mathcal{W}}_{V,U}+W^{V}-\nu_{V}(V)$
 as well as
-${\cal W}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\cal W}^{U,V}+W_{U}-\nu_{V}(V).$
-Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (140) and (141) in the proof of the commutation formulae (123), we readily obtain the estimates
-$e^{\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W_{U}}\leq\frac{d\pi_{2n}}{d\nu_{V}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W_{U}}$
+${\mathcal{W}}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\mathcal{W}}^{U,V}+W_{U}-\nu_{V}(V).$
+Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (E.1) and (E.2) in the proof of the commutation formulae (D.1), we readily obtain the estimates
+$e^{\nu_{V}(V)}\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W_{U}}\leq\frac{d\pi_{2n}}{d\nu_{V}}\leq e^{-\nu_{V}(V)}\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W_{U}}$
 as well as
-$e^{\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W^{V}}\leq\frac{d\pi_{2n+1}}{d\lambda_{U}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W^{V}}$
+$e^{\nu_{V}(V)}\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W^{V}}\leq\frac{d\pi_{2n+1}}{d\lambda_{U}}\leq e^{-\nu_{V}(V)}\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W^{V}}$
 Remark D.6.
-The functions $({\cal W}_{V,U},{\cal W}^{U,V})$ defined in (129) can be computed explicitly for the the Gaussian model defined in (87) and (88).
+The functions $({\mathcal{W}}_{V,U},{\mathcal{W}}^{U,V})$ defined in (D.7) can be computed explicitly for the the Gaussian model defined in (5.6) and (5.7).
 We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [adm-24]) we have
-${\cal W}_{V,U}(x)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~(\beta_{\tau}~(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
+${\mathcal{W}}_{V,U}(x)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~(\beta_{\tau}~(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
 and
-${\cal W}^{U,V}(y)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\frac{1}{2}(y-\overline{m})^{\prime}\left(\beta_{\tau}\,\sigma\,\beta_{\tau}^{\prime}-\tau^{-1}\right)(y-\overline{m})-(y-\overline{m})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$ (136)
+(D.14) ${\mathcal{W}}^{U,V}(y)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\frac{1}{2}(y-\overline{m})^{\prime}\left(\beta_{\tau}\,\sigma\,\beta_{\tau}^{\prime}-\tau^{-1}\right)(y-\overline{m})-(y-\overline{m})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$
 Note that one readily obtain the relationships
 […]
-Using (135), Condition ${\cal H}_{\delta}(U,V)$ ensures that for any $n\geq 1$ we have
-$e^{-\delta V}~\left(\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\right)\leq c~e^{-{\cal V}_{\delta}}\quad\mbox{\rm and}\quad e^{-\delta U}~~\left(\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\right)\leq c~e^{-{\cal U}_{\delta}}$
+Using (D.13), Condition ${\mathcal{H}}_{\delta}(U,V)$ ensures that for any $n\geq 1$ we have
+$e^{-\delta V}~\left(\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\right)\leq c~e^{-{\mathcal{V}}_{\delta}}\quad\mbox{\rm and}\quad e^{-\delta U}~~\left(\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\right)\leq c~e^{-{\mathcal{U}}_{\delta}}$
 for some constant $c<\infty$. These estimates immediately yield the following result.
@@
 Under the assumptions of Proposition D.5, for any $n\geq 1$ we have
-$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\in{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\in{\cal B}_{\varphi}(\mathbb{X}).$
+$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\in{\mathcal{B}}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\in{\mathcal{B}}_{\varphi}(\mathbb{X}).$
 In addition, we have the uniform estimates
 $\sup_{n\geq 1}~\left(\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\right)<\infty\quad\mbox{and}\quad\sup_{n\geq 1}~\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)<\infty.$
-D.3 Lyapunov inequalities
+D.3. Lyapunov inequalities
 The result below is a fairly direct consequence of the estimates presented in Lemma D.4.
@@
 For any $n\geq 1$ we have the estimates
-$e^{\nu_{V}(V)}~e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{x}{\cal S}_{2n}}{d\delta_{x}{\cal Q}_{V}}(y)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}$
+$e^{\nu_{V}(V)}~e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}\leq\frac{d\delta_{x}{\mathcal{S}}_{2n}}{d\delta_{x}{\mathcal{Q}}_{V}}(y)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}$
 and
-$e^{\nu_{V}(V)}~e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{y}{\cal S}_{2n+1}}{d\delta_{y}{\cal R}_{U}}(x)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}.$
+$e^{\nu_{V}(V)}~e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}\leq\frac{d\delta_{y}{\mathcal{S}}_{2n+1}}{d\delta_{y}{\mathcal{R}}_{U}}(x)\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)+W_{U}(y)}.$
 Proof.
-Using the Sinkhorn transition operators as expressed in (20) together with (131) and (132), we find that, for any $n\geq 1$,
-${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{(V-V_{2n})(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\geq e^{\nu_{V}(V)}e^{-({\cal W}_{V,U}(x)+{\cal W}^{V,U}(y))}{\cal Q}_{V}(x,dy)$~
+Using the Sinkhorn transition operators as expressed in (2.6) together with (D.9) and (D.10), we find that, for any $n\geq 1$,
+${\mathcal{S}}_{2n}(x,dy)$ $=$ $\frac{{\mathcal{Q}}_{V}(x,dy)e^{(V-V_{2n})(y)}}{{\mathcal{Q}}(e^{-V_{2n}})(x)}\geq e^{\nu_{V}(V)}e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{V,U}(y))}{\mathcal{Q}}_{V}(x,dy)$~
 and, in the same vein,
-${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\geq e^{\nu_{V}(V)}~e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{\cal R}_{U}(y,dx).$
-Similarly, one can resort to (131) and (132) again to obtain the lower bounds
-${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{V-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)}~{\cal Q}_{V}(x,dy)~e^{W_{U}(y)}$
+${\mathcal{S}}_{2n+1}(y,dx)$ $=$ $\frac{{\mathcal{R}}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\mathcal{R}}(e^{-U_{2n}})(y)}\geq e^{\nu_{V}(V)}~e^{-({\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}~{\mathcal{R}}_{U}(y,dx).$
+Similarly, one can resort to (D.9) and (D.10) again to obtain the lower bounds
+${\mathcal{S}}_{2n}(x,dy)$ $=$ $\frac{{\mathcal{Q}}_{V}(x,dy)e^{V-V_{2n}(y)}}{{\mathcal{Q}}(e^{-V_{2n}})(x)}\leq e^{-\nu_{V}(V)}~e^{W^{V}(x)}~{\mathcal{Q}}_{V}(x,dy)~e^{W_{U}(y)}$
 and, in the same way,
-${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\leq e^{-\nu_{V}(V)}~e^{W_{U}(y)}~{\cal R}_{U}(y,dx)~e^{W^{V}(x)}.$
+${\mathcal{S}}_{2n+1}(y,dx)$ $=$ $\frac{{\mathcal{R}}_{U}(y,dx)~e^{(U-U_{2n})(x)}}{{\mathcal{R}}(e^{-U_{2n}})(y)}\leq e^{-\nu_{V}(V)}~e^{W_{U}(y)}~{\mathcal{R}}_{U}(y,dx)~e^{W^{V}(x)}.$
 With these estimates at hand, we can now state and prove the following Lyapunov inequalities.
 Proposition D.9.
-Assume that Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
-${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~e^{-{\cal U}_{\delta}}~\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n-1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}~\in{\cal B}_{0}(\mathbb{Y})$ (137)
+Assume that Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
+(D.15) ${\mathcal{S}}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~e^{-{\mathcal{U}}_{\delta}}~\in{\mathcal{B}}_{0}(\mathbb{X})\quad\mbox{and}\quad{\mathcal{S}}_{2n-1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\mathcal{V}}_{\delta}}~\in{\mathcal{B}}_{0}(\mathbb{Y})$
 for every $n\geq 1$ and some finite constant $c_{\delta}<\infty$.
@@
 By Lemma D.8 we have
-${\cal S}_{2n}(x,dy)~\psi(y)/\varphi(x)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal U}_{\delta}(x)}~{\cal Q}_{(1-\delta)V}(x,dy)~e^{W_{U}(y)},\quad\text{and}$
-${\cal S}_{2n+1}(y,dx)~\varphi(x)/\psi(y)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal V}_{\delta}(y)}~{\cal R}_{(1-\delta)U}(y,dx)~e^{W^{V}(x)},$
-which (via (129)) imply the inequalities
-${\cal S}_{2n}\left(\psi\right)/\varphi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal U}_{\delta}}~e^{{\cal W}_{(1-\delta)V,U}},\quad\text{and}$
-${\cal S}_{2n+1}\left(\varphi\right)/\psi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\cal V}_{\delta}}~e^{{\cal W}^{(1-\delta)U,V}}.$
+${\mathcal{S}}_{2n}(x,dy)~\psi(y)/\varphi(x)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{U}}_{\delta}(x)}~{\mathcal{Q}}_{(1-\delta)V}(x,dy)~e^{W_{U}(y)},\quad\text{and}$
+${\mathcal{S}}_{2n+1}(y,dx)~\varphi(x)/\psi(y)$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{V}}_{\delta}(y)}~{\mathcal{R}}_{(1-\delta)U}(y,dx)~e^{W^{V}(x)},$
+which (via (D.7)) imply the inequalities
+${\mathcal{S}}_{2n}\left(\psi\right)/\varphi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{U}}_{\delta}}~e^{{\mathcal{W}}_{(1-\delta)V,U}},\quad\text{and}$
+${\mathcal{S}}_{2n+1}\left(\varphi\right)/\psi$ $\leq$ $e^{-\nu_{V}(V)}~e^{-{\mathcal{V}}_{\delta}}~e^{{\mathcal{W}}^{(1-\delta)U,V}}.$
 On the other hand, we have the estimates
-$\|e^{{\cal W}_{(1-\delta)V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}=\|{\cal Q}_{(1-\delta)V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
+$\|e^{{\mathcal{W}}_{(1-\delta)V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}=\|{\mathcal{Q}}_{(1-\delta)V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
 and
-$\|e^{{\cal W}^{(1-\delta)U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{(1-\delta)U}(e^{W^{V}})\|_{\mathbb{L}_{\infty}(\nu)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
-where $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ when Condition ${\cal H}_{\delta}(U,V)$ holds. This yields (137) with the constant
+$\|e^{{\mathcal{W}}^{(1-\delta)U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\mathcal{R}}_{(1-\delta)U}(e^{W^{V}})\|_{\mathbb{L}_{\infty}(\nu)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
+where $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ when Condition ${\mathcal{H}}_{\delta}(U,V)$ holds. This yields (D.15) with the constant
 $c_{\delta}:=e^{-\nu_{V}(V)}~\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\left(\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\right).$
-By (130) we have
-${\cal S}_{1}(y,dx)$ $=$ $\frac{{\cal R}(y,dx)~e^{-U(x)}}{{\cal R}(e^{-U})(y)}\leq e^{W_{U}(y)}~{\cal R}(y,dx)~e^{-U(x)}$
+By (D.8) we have
+${\mathcal{S}}_{1}(y,dx)$ $=$ $\frac{{\mathcal{R}}(y,dx)~e^{-U(x)}}{{\mathcal{R}}(e^{-U})(y)}\leq e^{W_{U}(y)}~{\mathcal{R}}(y,dx)~e^{-U(x)}$
 This implies that
-$e^{-\delta V(y)}{\cal S}_{1}\left(e^{\delta U}\right)(y)\leq e^{-(\delta V(y)-W_{U}(y))}~\|e^{-W}\|~\lambda_{(1-\delta)U}\left(1\right)$
+$e^{-\delta V(y)}{\mathcal{S}}_{1}\left(e^{\delta U}\right)(y)\leq e^{-(\delta V(y)-W_{U}(y))}~\|e^{-W}\|~\lambda_{(1-\delta)U}\left(1\right)$
 Equivalently, some finite constant $c_{\delta}<\infty$ we have
-${\cal S}_{1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\cal V}_{\delta}}$~
+${\mathcal{S}}_{1}\left(\varphi\right)/\psi\leq c_{\delta}~e^{-{\mathcal{V}}_{\delta}}$~
 This ends the proof of the proposition.
-D.4 Proof of Theorem 4.13
-The Lyapunov inequalities (137) in Proposition D.9 ensure that the drift conditions stated in (60) hold with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
+D.4. Proof of Theorem 4.8
+By (C.7), the Lyapunov inequalities (D.15) in Proposition D.9 ensure that the drift conditions stated in (4.13) hold with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
 Recall from the proof of Proposition D.5 that $e^{-\delta U}$ and $e^{-\delta V}$ are uniformly bounded for any $\delta\geq 0$.
 Since $\mathbb{X}$ and $\mathbb{Y}$ can be exhausted respectively by the compact sub-levels sets $C_{U}(r)$ and $C_{V}(r)$ of $U$ and $V$, there exists some $r_{1}>1$ such that $\lambda(C_{U}(r))\wedge\nu(C_{V}(r))>0$ for $r=r_{1}$ and thus for any $r\geq r_{1}$.
-This ensure there exists some $r_{0}>0$ such that for any $r\geq r_{0}$ we have
+This ensures there exists some $r_{0}>0$ such that for any $r\geq r_{0}$ we have
 $\lambda(C_{\varphi}(r))\wedge\nu(C_{\psi}(r))>0.$
 Following the proof of Proposition D.5, for $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
-$e^{{\cal W}_{V,U}(x)}={\cal Q}_{V}(e^{W_{U}})(x)$ $=$ $\int~e^{W_{U}(y)-W(x,y)}~\nu_{V}(dy)$
+$e^{{\mathcal{W}}_{V,U}(x)}={\mathcal{Q}}_{V}(e^{W_{U}})(x)$ $=$ $\int~e^{W_{U}(y)-W(x,y)}~\nu_{V}(dy)$
 $\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\nu_{V}(e^{W_{U}})$
@@
 and we conclude that
-$w_{V,U}(r):=\sup_{C_{\varphi}(r)}{\cal W}_{V,U}<\infty.$
+$w_{V,U}(r):=\sup_{C_{\varphi}(r)}{\mathcal{W}}_{V,U}<\infty.$
 Similarly, for $\nu$-almost any $y\in C_{\psi}(r)$ we have
-$e^{{\cal W}^{U,V}(y)}={\cal R}_{U}(e^{W^{V}})(y)$ $=$ $\int~e^{-W(x,y)+W^{V}(x)}~\lambda_{U}(dx)$
+$e^{{\mathcal{W}}^{U,V}(y)}={\mathcal{R}}_{U}(e^{W^{V}})(y)$ $=$ $\int~e^{-W(x,y)+W^{V}(x)}~\lambda_{U}(dx)$
 $\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{U}(e^{W^{V}})$
@@
 which yields
-$w^{U,V}(r):=\sup_{C_{\psi}(r)}{\cal W}^{U,V}<\infty.$
+$w^{U,V}(r):=\sup_{C_{\psi}(r)}{\mathcal{W}}^{U,V}<\infty.$
 On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ and $r\geq r_{0}$ we arrive at
-${\cal S}_{2n}(x,dy)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~\nu_{V}(dy)~1_{C_{\psi}(r)}(y)$
-$\geq$ $e^{-w(r)}~\nu_{V}\left(C_{\psi}(r)\right)~{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad{\cal Y}_{r}(dy):=\frac{\nu_{V}(dy)~1_{C_{\psi}(r)}(y)}{\nu_{V}\left(C_{\psi}(r)\right)}$
+${\mathcal{S}}_{2n}(x,dy)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}~\nu_{V}(dy)~1_{C_{\psi}(r)}(y)$
+$\geq$ $e^{-w(r)}~\nu_{V}\left(C_{\psi}(r)\right)~{\mathcal{Y}}_{r}(dy)\quad\mbox{\rm with}\quad{\mathcal{Y}}_{r}(dy):=\frac{\nu_{V}(dy)~1_{C_{\psi}(r)}(y)}{\nu_{V}\left(C_{\psi}(r)\right)}$
 and the parameter
@@
 and this yields, for any $x\in C_{\varphi}(r)$ and $r\geq r_{0}$, the estimate
-${\cal S}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~e^{-w(r)}.$
+${\mathcal{S}}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{\mathcal{Y}}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~e^{-w(r)}.$
 In the same vein, for any $y\in C_{\psi}(r)$ we have
-${\cal S}_{2n+1}(y,dx)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)$
+${\mathcal{S}}_{2n+1}(y,dx)$ $\geq$ $e^{\nu_{V}(V)}~e^{-(W(x,y)+{\mathcal{W}}_{V,U}(x)+{\mathcal{W}}^{U,V}(y))}~\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)$
 and, arguing as above, we arrive at
@@
 Thus, for any $y\in C_{\psi}(r)$ we have
-${\cal S}_{2n+1}(y,dx)\geq\varepsilon_{1}(r)~{\cal X}_{r}(dx)$
+${\mathcal{S}}_{2n+1}(y,dx)\geq\varepsilon_{1}(r)~{\mathcal{X}}_{r}(dx)$
 where
-$\varepsilon_{1}(r):=w_{1}(r)~e^{-w(r)}\quad\mbox{\rm and}\quad{\cal X}_{r}(dx):=\frac{\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)}{\lambda_{U}\left(C_{\varphi}(r)\right)}.$
-The above estimates ensure that the local minorization condition (61) is met with $n_{0}=1$ and the function
+$\varepsilon_{1}(r):=w_{1}(r)~e^{-w(r)}\quad\mbox{\rm and}\quad{\mathcal{X}}_{r}(dx):=\frac{\lambda_{U}(dx)~1_{C_{\varphi}(r)}(x)}{\lambda_{U}\left(C_{\varphi}(r)\right)}.$
+The above estimates ensure that the local minorization condition (4.14) is met with $n_{0}=1$ and the function
 $\varepsilon(r):=\varepsilon_{0}(r)\wedge\varepsilon_{1}(r)>0.$
-Hence, we conclude that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied with $n_{0}=1$.
-Using (22) and (125), for any $n\geq p\geq 1$ we verify that
-$\frac{d\pi_{2n}}{d\nu_{V}}(y)-1=\left[(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right]\left(\frac{d\pi_{2p}}{d\nu_{V}}\right),$
+Hence, we conclude that Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied with $n_{0}=1$.
+Using (2.8) and (D.3), for any $n\geq p\geq 1$ we verify that
+$\frac{d\pi_{2n}}{d\nu_{V}}(y)-1=\left[(\delta_{y}-\nu_{V})\left({\mathcal{S}}^{\circ}_{2n}\ldots{\mathcal{S}}^{\circ}_{2(p+1)}\right)\right]\left(\frac{d\pi_{2p}}{d\nu_{V}}\right),$
 which implies the inequality
 […]
-On the other hand, combining Proposition D.7 with the norm equivalence (34) and Theorem 4.8, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
+On the other hand, combining Proposition D.7 with the norm equivalence (3.4) and Theorem 4.6, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
 […]
@@
 Therefore, for any $n\geq 1$, we obtain
-$\left\|\frac{d\pi_{2n}}{d\nu_{V}}-1\right\|_{\psi}\leq c_{\kappa,3}~\rho^{2n}~\quad\text{and, similarly,}\quad\left\|\frac{d\nu_{V}}{d\pi_{2n}}-1\right\|_{\psi}\leq c_{\kappa,4}~\rho^{2n},$ (138)
-for some constants $c_{\kappa,i}$, $i=3,4$. By the same argument, using (22) we have, for any $n\geq p\geq 1$,
-$\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)-1=\left[(\delta_{x}-\lambda_{U})({\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1})\right]\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right)$
+(D.16) $\left\|\frac{d\pi_{2n}}{d\nu_{V}}-1\right\|_{\psi}\leq c_{\kappa,3}~\rho^{2n}~\quad\text{and, similarly,}\quad\left\|\frac{d\nu_{V}}{d\pi_{2n}}-1\right\|_{\psi}\leq c_{\kappa,4}~\rho^{2n},$
+for some constants $c_{\kappa,i}$, $i=3,4$. By the same argument, using (2.8) we have, for any $n\geq p\geq 1$,
+$\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)-1=\left[(\delta_{x}-\lambda_{U})({\mathcal{S}}^{\circ}_{2n+1}\ldots{\mathcal{S}}^{\circ}_{2p+1})\right]\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right)$
 and arrive at
-$\left\|\frac{d\pi_{2n+1}}{d\lambda_{U}}-1\right\|_{\varphi}\vee\left\|\frac{d\lambda_{U}}{d\pi_{2n+1}}-1\right\|_{\varphi}\leq c_{\kappa,5}~\rho^{2n}$~ (139)
+(D.17) $\left\|\frac{d\pi_{2n+1}}{d\lambda_{U}}-1\right\|_{\varphi}\vee\left\|\frac{d\lambda_{U}}{d\pi_{2n+1}}-1\right\|_{\varphi}\leq c_{\kappa,5}~\rho^{2n}$~
 for any $n\geq 1$ and some constant $c_{\kappa,5}$.
@@
 where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $1+u>0$. This implies that
-$\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}$
-The inequality above combined with (138) and (139) yields the estimates
+$\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}.$
+The inequality above combined with (D.16) and (D.17) yields the estimates
 $\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c_{6,\kappa}~\rho^{2n},$
-for some finite constant $c_{\kappa,6}$. This completes the proof of (71).
+for some finite constant $c_{\kappa,6}$. This completes the proof of (4.24).
 Appendix E Some technical proofs
-Proof of (123) and (124)
-We first note, using (16) and (17), that
-$\frac{d{\cal P}_{2n}}{d{\cal P}_{2n+1}}(x,y)=\frac{d\pi_{2n}}{d\nu_{V}}(y)=\exp{\left(V_{2n+1}(y)-V_{2n}(y)\right)}$ (140)
+Proof of (D.1) and (D.2)
+We first note, using (2.2) and (2.3), that
+(E.1) $\frac{d{\mathcal{P}}_{2n}}{d{\mathcal{P}}_{2n+1}}(x,y)=\frac{d\pi_{2n}}{d\nu_{V}}(y)=\exp{\left(V_{2n+1}(y)-V_{2n}(y)\right)}$
 and, in the same vein,
-$\frac{d{\cal P}_{2n+1}}{d{\cal P}_{2(n+1)}}(x,y)=\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)=\exp{\left(U_{2(n+1)}(x)-U_{2n+1}(x)\right)}.$ (141)
+(E.2) $\frac{d{\mathcal{P}}_{2n+1}}{d{\mathcal{P}}_{2(n+1)}}(x,y)=\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)=\exp{\left(U_{2(n+1)}(x)-U_{2n+1}(x)\right)}.$
 If we additionally observe that
-${\cal P}_{2(n+1)}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\cal P}_{2n}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
-${\cal P}_{2n+1}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n-1}}(x)~{\cal P}_{2n-1}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)$
-then, using (15), we obtain the formulae
-${\cal S}_{2(n+1)}(x,dy)$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\cal S}_{2n}(x,dy)~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
-${\cal S}_{2n+1}(y,dx)$ $=$ $~\frac{d\nu_{V}}{d\pi_{2n}}(y)~{\cal S}_{2n-1}(y,dx)~\frac{d\lambda_{U}}{d\pi_{2n-1}}(x),$
+${\mathcal{P}}_{2(n+1)}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\mathcal{P}}_{2n}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
+${\mathcal{P}}_{2n+1}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n-1}}(x)~{\mathcal{P}}_{2n-1}(d(x,y))~\frac{d\nu_{V}}{d\pi_{2n}}(y)$
+then, using (2.1), we obtain the formulae
+${\mathcal{S}}_{2(n+1)}(x,dy)$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{\mathcal{S}}_{2n}(x,dy)~\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
+${\mathcal{S}}_{2n+1}(y,dx)$ $=$ $~\frac{d\nu_{V}}{d\pi_{2n}}(y)~{\mathcal{S}}_{2n-1}(y,dx)~\frac{d\lambda_{U}}{d\pi_{2n-1}}(x),$
 and the commutation properties now follow from elementary manipulations.
-Proof of (57)
-Assume Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied. Combining (53) with (22) we arrive at
-$\nu_{V}(\psi)=\nu_{V}({\cal S}_{2n}^{\circ}(\psi))\leq\epsilon~\nu_{V}(\psi)+c\quad\text{which implies}\quad\nu_{V}(\psi)\leq\frac{c}{1-\epsilon}.$
+Proof of (4.10)
+Assume Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is satisfied. Combining (4.6) with (2.8) we arrive at
+$\nu_{V}(\psi)=\nu_{V}({\mathcal{S}}_{2n}^{\circ}(\psi))\leq\epsilon~\nu_{V}(\psi)+c\quad\text{which implies}\quad\nu_{V}(\psi)\leq\frac{c}{1-\epsilon}.$
 In a similar way, one can also show that
 $\lambda_{U}(\varphi)\leq\frac{c}{1-\epsilon}.$
-Using the Markov evolution equations in (23) together with (53) we find that, for any $n\geq n_{0}$,
-$\pi_{2n}(\psi)$ $=$ $\pi_{2(n-1)}({\cal S}_{2n}^{\circ}(\psi))$
+Using the Markov evolution equations in (2.9) together with (4.6) we find that, for any $n\geq n_{0}$,
+$\pi_{2n}(\psi)$ $=$ $\pi_{2(n-1)}({\mathcal{S}}_{2n}^{\circ}(\psi))$
 $\leq$ $\epsilon~\pi_{2(n-1)}(\psi)+c\leq\epsilon^{n-n_{0}}~\pi_{2n_{0}}(\psi)+\frac{c}{1-\epsilon}$
 and, by the same argument,
-$\pi_{2n+1}(\varphi)$ $=$ $\pi_{2n-1}({\cal S}_{2n+1}^{\circ}(\varphi))$
+$\pi_{2n+1}(\varphi)$ $=$ $\pi_{2n-1}({\mathcal{S}}_{2n+1}^{\circ}(\varphi))$
 $\leq$ $\epsilon~\pi_{2n-1}(\varphi)+c\leq\epsilon^{n-n_{0}}~\pi_{2n_{0}+1}(\varphi)+\frac{c}{1-\epsilon}.$
@@
 Therefore,
-${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{R}^{d})\quad\mbox{\rm and}\quad{\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{R}^{d})$
+${\mathcal{V}}_{\delta}=\delta V-W_{U}\in{\mathcal{B}}_{\infty}(\mathbb{R}^{d})\quad\mbox{\rm and}\quad{\mathcal{U}}_{\delta}=\delta U-W^{V}\in{\mathcal{B}}_{\infty}(\mathbb{R}^{d})$
 if, and only if,
@@
 which implies that
-$W_{U}=\sum_{i\in E_{0}}~\xi_{0}(i)~W_{U_{i}}\quad\mbox{\rm and}\quad W^{V}=\sum_{j\in E_{1}}~\xi_{1}(j)~W^{V_{j}}$ (142)
+(E.3) $W_{U}=\sum_{i\in E_{0}}~\xi_{0}(i)~W_{U_{i}}\quad\mbox{\rm and}\quad W^{V}=\sum_{j\in E_{1}}~\xi_{1}(j)~W^{V_{j}}$
 and ensures that the functions $W_{U}$ and $W^{V}$ are locally bounded.
@@
 $=$ $\bigcup_{(i,j)\in(E_{0}\times E_{1})}\left\{x~:~(\delta~U_{i}(x)-~W^{V_{j}}(x))\leq r\right\}$
-and this ensures that ${\cal U}_{\delta}=\delta U-W^{V}$ have compact sub-level sets. In the same vein, one can also prove that ${\cal V}_{\delta}=\delta V-W_{U}$ have compact sub-level sets.
+and this ensures that ${\mathcal{U}}_{\delta}=\delta U-W^{V}$ have compact sub-level sets.
+In the same vein, one can also prove that ${\mathcal{V}}_{\delta}=\delta V-W_{U}$ have compact sub-level sets.
 On the other hand, using the fact that $(a+b)^{1-\delta}\leq a^{1-\delta}+b^{1-\delta}$ for any $a,b\geq 0$, we have
-$\lambda_{(1-\delta)U}(dx)$ $=$ $\left(\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}(x)}\right)^{1-\delta}~\lambda(dx)$ (143)
+(E.4) $\lambda_{(1-\delta)U}(dx)$ $=$ $\left(\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}(x)}\right)^{1-\delta}~\lambda(dx)$
 $\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~e^{-(1-\delta)U_{i}(x)}~\lambda(dx)=\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~\lambda_{(1-\delta)U_{i}}(dx)$
 and, by way of Hölder’s inequality, we also have
-$\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~e^{\xi_{1}(j)~W^{V_{j}}}\right)\leq\prod_{j\in E_{1}}~\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}.$ (144)
-Combining (142), (143) and (144) we obtain the estimates
+(E.5) $\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~e^{\xi_{1}(j)~W^{V_{j}}}\right)\leq\prod_{j\in E_{1}}~\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}.$
+Combining (E.3), (E.4) and (E.5) we obtain the estimates
 $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~e^{\xi_{1}(j)~W^{V_{j}}}\right)$